### v1

# Flexible Deep Neural Networks for Partially Linear Survival Data

Abstract
We propose a flexible deep neural network (DNN) framework for modeling survival data within a partially linear regression structure. The approach preserves interpretability through a parametric linear component for covariates of primary interest, while a nonparametric DNN component captures complex time–covariate interactions among nuisance variables. We refer to the method as FLEXI–Haz, a FLEXIble Hazard model with a partially linear structure. In contrast to existing DNN approaches for partially linear Cox models, FLEXI–Haz does not rely on the proportional hazards assumption. We establish theoretical guarantees: the neural network component attains minimax-optimal convergence rates that depends on composite Hölder classes, and the linear estimator is $\sqrt{n}$–consistent, asymptotically normal, and semiparametrically efficient. Extensive simulations and real-data analyses demonstrate that FLEXI–Haz provides accurate estimation of the linear effect, offering a principled and interpretable alternative to modern methods based on proportional hazards. Code for implementing FLEXI-Haz, as well as scripts for reproducing data analyses and simulations is available at GitHub site https://github.com/AsafBanana/FLEXI-Haz.
keywords: Deep neural networks; Interpretability; Partial linear regression; Survival analysis; Time to event data.
1 Introduction
Deep neural networks (DNNs) are widely used in modern data analysis because of their ability to capture complex nonlinear patterns, handle high-dimensional inputs, and scale to large datasets. They have achieved strong performance in diverse domains such as image recognition, speech processing, and natural language understanding, often improving on traditional statistical and machine learning methods (lecun2015deep). Continued progress in architectures, optimization, and computational resources has further expanded their range of applications in both scientific and industrial domains.
Deep neural networks for survival analysis.
Survival analysis aims to model the distribution of event times in the presence of censoring, with applications in medicine, public health, epidemiology, engineering, and finance (kalbfleisch2011; klein2006survival). In recent years, DNNs have become increasingly popular in survival analysis thanks to their ability to model complex nonlinear relationships between covariates and event times and to scale efficiently to high-dimensional data. Many existing approaches extend the classical semi-parametric Cox proportional hazards (PH) model (cox1972regression) by replacing the parametric linear risk score with a neural network, while using the negative partial likelihood as the network’s loss function. Early work by faraggi1995neural employed a single hidden-layer network in place of the linear component, and katzman2018deepsurv extended this to deeper architectures (DeepSurv), demonstrating improved concordance in several benchmark datasets. Variants of this approach have been proposed for genomics (yousefi2017predicting; ching2018coxnnet), medical imaging (haarburger2019image; li2019deep), and recurrent or discrete-time formulations (giunchiglia2018rnnsurv; gensheimer2019nnet). While these PH-based DNN methods can yield strong empirical performance, they inherit the PH restriction, which may limit their ability to capture rich time-dependent effects.
Non-PH modeling with deep learning.
To address the limitations of the PH assumption, a growing body of work has focused on DNN-based models that allow covariate effects to vary with time. One strategy augments the network input with time itself, enabling non PH while retaining an unspecified baseline hazard, as in Cox-Time of kvamme2019time. Other approaches abandon the Cox structure entirely in favor of modeling the event time distribution or hazard function directly. In particular, hu2023conditional proposed a framework for estimating the full survival function, using a likelihood function derived from the continuous-time hazard without introducing a baseline hazard function. Their formulation naturally accommodates time-varying covariates through a data-expansion scheme and applies optimization to the log-hazard. By sidestepping the baseline hazard and proportionality constraints, such approaches offer greater modeling flexibility and can better reflect complex temporal dynamics in the hazard. The utility of this approach is illustrated through simulations and real-data analyses, but no asymptotic results are provided.
Semiparametric models and deep learning.
Alongside fully nonparametric formulations, there is considerable interest in semiparametric survival models that combine a parametric component for covariates of primary interest with a flexible, nonparametric component for high-dimensional or nuisance covariates. This structure preserves interpretability and enables formal inference for the parametric component, while leveraging modern machine learning to estimate the nuisance function. Theoretical foundations for this paradigm have been established in the non-survival setting by schmidthieber2020nonparametric, who derived minimax-optimal rates for ReLU-activated DNNs in nonparametric regression, and extended to survival analysis by zhong2022deep in the partially linear Cox model. In the latter, the log-hazard at time $t$ is modeled as $\log h_{0}(t)+\mbox{${\zeta}$}^{\top}{\bf Z}+f({\bf X})$, where $h_{0}(t)$ is an unspecified baseline hazard function, ${\bf Z}\in\mathbb{R}^{p}$ is the vector of covariates of primary interest, ${\bf X}\in\mathbb{R}^{d}$ is the vectors of nuisance covariates, and $\mbox{${\zeta}$}\in\mathbb{R}^{p}$ is the main parameter of interest, estimated at the semiparametric efficiency bound via partial likelihood. The nuisance function $f$ is approximated by a DNN. This approach achieves both optimal convergence for $g$ and $\sqrt{n}$-consistency for $\zeta$} under the PH assumption. The work of zhong2022deep demonstrates how semiparametric theory and deep learning can be integrated to yield interpretable, efficient estimators in survival analysis, but like other Cox-based methods, it remains tied to PH and baseline hazard estimation.
Contributions.
In this work, we develop a semiparametric survival model that extends the partially linear framework of zhong2022deep to the non-PH setting. We model the log-hazard function at time $t$ as $\mbox{${\theta}$}_{o}^{\sf T}{\bf Z}+g_{o}(t,{\bf X})$, where $\mbox{${\theta}$}_{o}\in\mathbb{R}^{p}$ is the parameter of primary interest and $g_{o}$ is an unknown nuisance function approximated by a DNN. This formulation allows arbitrary interactions between the nuisance covariates $\bf X$} and time $t$, naturally accommodates time-dependent effects and time-dependent covariates within $\bf X$}, and does not require specifying or estimating a baseline hazard function.
A central methodological difference from existing semiparametric approaches is that we work with the full likelihood, rather than the partial likelihood commonly used in Cox-type models. Despite removing the PH structure, we establish that the estimator of the parametric component $\mbox{${\theta}$}_{o}$ remains $\sqrt{n}$-consistent, asymptotically normal, and semiparametric efficient. In particular, we show that the theoretical guarantees of zhong2022deep, developed for the partially linear Cox model, continue to hold in the substantially more general non-PH setting, provided the hazard is almost surely bounded away from zero.
While parts of our analysis build on the approximation theory of schmidthieber2020nonparametric and the high-level strategy of zhong2022deep, several key components of our asymptotic theory are fundamentally different. The core technical challenges arise from the fact that our neural network estimator is a function of both $t$ and $\bf X$}, rather than $\bf X$} alone, and the fact that we use the full likelihood rather than a profiling argument. These differences require new empirical process tools, a new characterization of the nuisance tangent space, and a new argument for controlling the bias induced by the DNN approximation.
Finally, when the linear effect $\mbox{${\theta}$}_{o}^{\top}{\bf Z}$ is removed and $\bf Z$} is absorbed into $g_{o}$, our model collapses to the conditional hazard model of hu2023conditional. Our analysis therefore fills an important theoretical gap by providing the first asymptotic guarantees for that model as well.
Code for implementing FLEXI-Haz, as well as scripts for reproducing all data analyses and simulations, is available at GitHub site https://github.com/AsafBanana/FLEXI-Haz.
2 Model Formulation
We consider right-censored survival data, where for each individual the true event time is denoted by $U$ and the censoring time by $C$. We observe only $T=\min(U,C)$ and $\Delta=\mathbf{1}(U\leq C)$, indicating the observed time and whether the event occurred. Let the covariate vector be $({{\bf Z}}^{\sf T},{{\bf X}}^{\sf T})^{\sf T}$, where ${{\bf Z}}\in\mathbb{R}^{p}$ contains the variables of primary scientific interest and ${{\bf X}}\in\mathbb{R}^{d}$ denotes nuisance covariates. Our goal is to characterize the effect of $\bf Z$} on the event time while appropriately adjusting for $\bf X$}. For example, $\mathbf{X}$ may represent a high-dimensional set of genomic or imaging features, whereas $\mathbf{Z}$ may contain key clinical variables such as treatment assignment or targeted biomarkers. Understanding the effect of $\bf Z$} while properly adjusting for $\bf X$} is essential for identifying clinically actionable signals. For simplicity, we focus here on time-independent covariates; the extension to time-dependent covariates will be discussed in Section 7. We assume that $U$ and $C$ are conditionally independent given $(\mathbf{X},\mathbf{Z})$. The observed data consist of $n$ independent replicates, $\{(T_{i},\Delta_{i},\mathbf{X}_{i},\mathbf{Z}_{i})\}_{i=1}^{n}$.
Under the above setting, the Cox PH model (cox1972regression) specifies the instantaneous hazard function at time $t>0$ as
$h_{1}(t\mid{{\bf X}},{{\bf Z}})=h_{0}(t)\exp({\mbox{${\alpha}$}}_{1}^{\top}{{\bf X}}+{\mbox{${\alpha}$}}_{2}^{\top}{{\bf Z}})\,$,
where $h_{0}$ is an unspecified baseline hazard function and ${\mbox{${\alpha}$}}_{1}\in\mathbb{R}^{d}$ and ${\mbox{${\alpha}$}}_{2}\in\mathbb{R}^{p}$ are unknown parameters. The model remains widely used due to its interpretability, semiparametric structure, and the availability of efficient estimation procedures. DeepSurv (katzman2018deepsurv) extends the Cox model by replacing the linear risk score $\exp({\mbox{${\alpha}$}}_{1}^{\top}\mathbf{X}+{\mbox{${\alpha}$}}_{2}^{\top}\mathbf{Z})$ with a flexible nonlinear function $f_{1}({{\bf X}},{{\bf Z}})$ estimated by a DNN. Cox-Time (kvamme2019time) further generalized this formulation by replacing $f_{1}$ with a time-depend function, $f_{2}(t,{{\bf X}},{{\bf Z}})$, thereby relaxing the PH assumption and permitting time-covariate interactions. Because $f_{2}$ is time-dependent while the baseline hazard is left unspecified, identifiability becomes nontrivial; kvamme2019time address this challenge through a numerical normalization strategy. More recently, hu2023conditional proposed a baseline-free hazard model of the form
$h_{3}(t\mid{{\bf X}},{{\bf Z}})=\exp\{f_{3}(t,{{\bf X}},{{\bf Z}})\}\,$,
which provides full flexibility for modeling non-proportional, time-varying effects via a DNN. Their estimation procedure is evaluated only through simulations, with no accompanying asymptotic guarantees. Although these DNN-based approaches are highly flexible, they do not provide interpretable estimates for the effect of $\bf Z$}}. To address this limitation, zhong2022deep introduced the deep partially linear Cox model
$h_{4}(t\mid\mathbf{X},\mathbf{Z})=h_{0}(t)\exp\{\mbox{${\zeta}$}^{\top}\mathbf{Z}+f(\mathbf{X})\}\,$,
which augments the Cox framework by separating time component from the covariates and retaining a parametric term $\mbox{${\zeta}$}^{\top}\mathbf{Z}$ for interpretability, while modeling nuisance effects through a DNN. This structure enables valid inference for $\zeta$} while still allowing complex, nonlinear effect of $\bf X$}}.
In this work, we extend the deep partially linear framework of zhong2022deep to the non-PH setting by considering a baseline-free model of the form
$h_{o}(t\mid\mathbf{X},{{\bf Z}})=\exp\{\mbox{${\theta}$}_{o}^{\top}\mathbf{Z}+g_{o}(t,\mathbf{X})\},$ (1)
where $g_{o}(t,\mathbf{X})$ is estimated by a DNN. Our formulation retains the interpretability and inference guarantees for the parametric component ${\mbox{${\theta}$}}_{o}$, but generalizes the nuisance component from $f(\mathbf{X})$ to $g_{o}(t,\mathbf{X})$, thereby allowing fully flexible, time-varying effects and removing the reliance on a baseline hazard function. This extension overcomes the PH restriction inherent in zhong2022deep, resolves the non-identifiability issues that arise when time is included in the DNN, and greatly expands the class of survival models that can be handled within a semiparametric, partially linear framework.
As we demonstrate in Section 6, restricting the nuisance function to depend only on only on $\mathbf{X}$, as in the partially linear Cox model, can induce substantial bias in the estimation of the linear effect whenever the nuisance covariates exhibit time-varying effects. Allowing $g_{o}$ to depend on both $t$ and $\mathbf{X}$ eliminates this source of bias but introduces significant methodological challenges. First, the negative partial likelihood can no longer be used as the DNN loss function. Instead, the full-likelihood is adopted which includes the predictable cumulative hazard that must be numerically approximated at every training step, increasing the computational burden. Second, the presence of this integral markedly complicates the asymptotic analysis: the time dimension introduces additional stochastic terms that must be controlled, requiring new empirical process arguments beyond those used for time-invariant nuisance functions. Our estimation framework in Section 3 addresses these computational difficulties, and the technical developments leading to the asymptotic results are presented in Section 4.3.
2.1 Composite Hölder Class of $g_{o}$
To formalize the complexity assumptions on the nuisance component $g_{o}$, we follow the compositional function framework of juditsky2009 and schmidthieber2020nonparametric, and use the notation of zhong2022deep. We assume that the nuisance function $g_{o}$ admits a hierarchical, low-dimensional representation of the form
$g_{o}=g_{q}\circ g_{q-1}\circ\cdots\circ g_{1}\circ g_{0},$
with the convention that $(g\circ f)(x)=g(f(x))$. This composition is associate with a dimension vector ${\bf d}^{*}=(d^{*}_{0},\ldots,d^{*}_{q+1})\in\mathbb{N}_{+}^{q+2}$ and an intrinsic dimension vector $\tilde{\bf d}=(\tilde{d}_{0},\ldots,\tilde{d}_{q})\in\mathbb{N}_{+}^{q+1}$, $\tilde{d}_{\ell}\leq d^{*}_{\ell}$ for each $\ell$. The outermost input is $d^{*}_{0}=r=1+d$, corresponding to observed times and covariates $\mathbf{X}$, and the final output is scalar $d^{*}_{q+1}=1$. For each layer $l$, $l=0,\ldots,q$, the vector-valued mapping $g_{\ell}=(g_{\ell 1},\ldots,g_{\ell\,d^{*}_{\ell+1}})^{\top}$ maps the bounded set $a_{\ell},b_{\ell}]^{\,d^{*}_{\ell}}$ into $a_{\ell+1},b_{\ell+1}]^{\,d^{*}_{\ell+1}}$, where $a_{l},b_{l}\in\mathbb{R}$ are fixed finite constants ensuring each intermediate representation remains bounded. Additionally, each component $g_{\ell j}$ depending on at most $\tilde{d}_{\ell}$ coordinates of its input.
For illustration, consider $f:\mathbb{R}^{6}\to\mathbb{R}$ defined by
$f(x_{1},\ldots,x_{6})\;=\;\sin(x_{1}+x_{2})\;+\;\log(1+x_{3})\;+\;x_{4}^{2}.$
which depends on low-dimensional groupings of the input coordinates. A valid three-layer decomposition is $g\;=\;g_{2}\circ g_{1}\circ g_{0}$ where
$g_{0}:\mathbb{R}^{6}\to\mathbb{R}^{3}\,,\quad g_{0}(x)\;=\;\big(\,x_{1}+x_{2},\;\;x_{3},\;\;x_{4}\,\big)\,$,
$g_{1}:\mathbb{R}^{3}\to\mathbb{R}^{3}\,,\quad g_{1}(y_{1},y_{2},y_{3})\;=\;\big(\,\sin(y_{1}),\;\;\log(1+y_{2}),\;\;y_{3}^{2}\,\big)\,$,
$g_{2}:\mathbb{R}^{3}\to\mathbb{R}\,,\quad g_{2}(z_{1},z_{2},z_{3})\;=\;z_{1}+z_{2}+z_{3}\,.$
For this representation, ${\bf d}^{*}=({d^{*}_{0}},{d^{*}_{1}},{d^{*}_{2}},{d^{*}_{3}})=(6,3,3,1)$, and the intrinsic dimension vector is $\tilde{\bf d}=(\tilde{d}_{0},\tilde{d}_{1},\tilde{d}_{2})=(2,1,3)$. That is, components of $g_{0}$ depend on at most two inputs, those of $g_{1}$ are univariate, and $g_{2}$ combines all three outputs. Other valid decompositions may exist. In general, convergence rates depend on the choice of the optimal decomposition used in the asymptotic analysis.
We also assume that each component function satisfies a standard Hölder–smoothness condition. Let $\alpha>0$ and $M>0$. A function is said to have Hölder smoothness index $\alpha$ if all mixed partial derivatives up to order $\lfloor\alpha\rfloor$ exist and are uniformly bounded, and if the derivatives of order exactly $\lfloor\alpha\rfloor$ are $(\alpha-\lfloor\alpha\rfloor)$–Hölder continuous. Here, $\lfloor\alpha\rfloor$ denotes the greatest integer strictly smaller than $\alpha$. The ball of $\alpha$-Hölder functions with radius $M$ consists of all such functions whose derivatives and Hölder constants are bounded by $M$. Formally, for a domain $\mathbb{D}\subset\mathbb{R}^{m}$, the Hölder ball is defined as
$\mathcal{H}^{\alpha}_{m}(\mathbb{D},M)=\Big\{g:\mathbb{D}\to\mathbb{R}:\sum_{\boldsymbol{\beta}:|\boldsymbol{\beta}|<\alpha}\|\partial^{\boldsymbol{\beta}}g\|_{\infty}+\sum_{\boldsymbol{\beta}:|\boldsymbol{\beta}|=\lfloor\alpha\rfloor}\sup_{u,v\in\mathbb{D},u\neq v}\frac{|\partial^{\boldsymbol{\beta}}g(u)-\partial^{\boldsymbol{\beta}}g(v)|}{\|u-v\|^{\alpha-\lfloor\alpha\rfloor}}\leq M\Big\},$
where $\boldsymbol{\beta}^{\sf T}=(\beta_{1},\ldots,\beta_{m})\in\mathbb{N}^{m}$ is a multi-index $\partial^{\boldsymbol{\beta}}=\partial^{\beta_{1}}\cdots\partial^{\beta_{m}}$, $|\boldsymbol{\beta}|=\sum_{j=1}^{m}\beta_{j}$, and $\|g\|_{\infty}$ is the sup–norm of a function over its domain.
To formalize the function class used for the nuisance component, let $\boldsymbol{\alpha}^{\sf T}=(\alpha_{0},\ldots,\alpha_{q})\in\mathbb{R}_{+}^{q+1}$ denote the layerwise smoothness parameters, $\bf d$} the dimension vector, and $\tilde{{\bf d}}$ the intrinsic dimension vector. The composite Hölder class is defined as
$\mathcal{H}(q,\boldsymbol{\alpha},{\bf d},\tilde{{\bf d}},M)=\Big\{g=g_{q}\circ\cdots\circ g_{0}:g_{\ell j}\in\mathcal{H}^{\alpha_{\ell}}_{\tilde{d}_{\ell}}([a_{\ell},b_{\ell}]^{\tilde{d}_{\ell}},M),\ |a_{\ell}|,|b_{\ell}|\leq M\Big\}.$
This class accomodates a broad range of structured functions, such as additive models, single- and multiple-index models, and hierarchical interactions, through the intrinsic dimensions $\tilde{d}_{\ell}$. Finally, we assume that the true nuisance function $g_{o}$ belongs to the composite class, $g_{o}\in\mathcal{H}(q,\boldsymbol{\alpha},{\bf d},\tilde{{\bf d}},M)$.
In the classical (non-composite) Hölder class $\mathcal{H}^{\alpha}_{d}$, the minimax-optimal rate is $n^{-\alpha/(2\alpha+d)}$ (tsybakov2009). For multilayer composite Hölder classes, schmidthieber2020nonparametric established matching minimax lower and upper bounds (up to polylogarithmic factors), showing that the optimal rate is governed by $\gamma_{n}\;=\;\max_{\ell=0,\ldots,q}\,n^{-\tilde{\alpha}_{\ell}/(2\tilde{\alpha}_{\ell}+\tilde{d}_{\ell})}$, where the effective smoothness indices are $\tilde{\alpha}_{\ell}\;=\;\alpha_{\ell}\prod_{k=\ell+1}^{q}(\alpha_{k}\wedge 1),\ \ a\wedge b=\min\{a,b\}$. Thus, each layer contributes to the overall difficulty of the estimation problem through its intrinsic dimension $\tilde{d}_{\ell}$ and effective smoothness $\tilde{\alpha}_{\ell}$. In Theorem 2 we show that our estimator achieves this optimal composite rate $\gamma_{n}$, up to polylogarithmic factors.
2.2 Neural Network Class
To approximate the nuisance function $g_{o}(t,\mathbf{X})$, we use the popular class of sparse deep ReLU neural networks (schmidthieber2020nonparametric, and reference therein). This choice is motivated by two considerations: (i) ReLU networks are universal approximators capable of representing complex, high-dimensional functions; and (ii) their structure lends itself to sharp theoretical control through entropy bounds and sparsity constraints. We define a general $(K+1)$-layer feedforward ReLU network with depth $K\in\mathbb{N}_{+}$ and width vector $\boldsymbol{p}^{\top}=(p_{0},\ldots,p_{K+1})$, where $p_{0}=r$ denotes the input dimension and $p_{K+1}=1$ the scalar output. Specifically, the network is defined by
$g(\mathbf{u})\;=\;\mathbf{W}_{K}\,\sigma\!\big(\mathbf{W}_{K-1}\,\sigma(\cdots\sigma(\mathbf{W}_{0}\mathbf{u}+\mathbf{v}_{0})\cdots)+\mathbf{v}_{K-1}\big)+\mathbf{v}_{K},$
where $\mathbf{u}\in\mathbb{R}^{r}$, $\sigma(x)=\max\{x,0\}$ is applied componentwise, $\mathbf{W}_{\ell}\in\mathbb{R}^{p_{\ell+1}\times p_{\ell}}$ are weight matrices, and $\mathbf{v}_{\ell}\in\mathbb{R}^{p_{\ell+1}}$ are bias vectors, $\ell=0,\ldots,K$.
For theoretical analysis we first restrict attention to networks with uniformly bounded parameters,
$\mathcal{G}(K,\boldsymbol{p})\;=\;\Big\{\,g:\ \max_{\ell=0,\ldots,K}\big(\|\mathbf{W}_{\ell}\|_{\infty}\vee\|\mathbf{v}_{\ell}\|_{\infty}\big)\leq 1\,\Big\}\,$,
which ensures controlled Lipschitz behavior of the class, where for a matrix $\bf A$} and a vector $\bf a$} the entrywise sup–norms is adopted, $\|\mathbf{A}\|_{\infty}=\max_{i,j}|A_{ij}|$, $\|\mathbf{a}\|_{\infty}=\max_{j}|a_{j}|$,. Following schmidthieber2020nonparametric we further impose sparsity. For sparsity level $s\in\mathbb{N}_{+}$ and range bound $D>0$, the sparse deep ReLU class is define as
$\mathcal{G}(K,s,\boldsymbol{p},D)\;=\;\Big\{g\in\mathcal{G}(K,\boldsymbol{p})\ :\ \sum_{\ell=0}^{K}\big(\|\mathbf{W}_{\ell}\|_{0}+\|\mathbf{v}_{\ell}\|_{0}\big)\leq s,\ \ \|g\|_{\infty}\leq D\Big\},$
where $\|\cdot\|_{0}$ counts nonzero entries. Sparsity serves to control network complexity and is essential for obtaining optimal convergence rates. In the subsequent analysis we set the depth, width, and sparsity levels to balance bias-variance tradeoff. The bracketing entropy (vdvwellner1996, Chs. 2–3) of this class is then bounded using results of schmidthieber2020nonparametric, and Dudley’s entropy integral (vdvwellner1996, Thm. 2.14.1) is used to establish the required uniform empirical-process bounds.
3 FLEXI-Haz Procedure
Estimation is based on maximizing the full likelihood induced by the hazard model (1). Under noninformative and conditionally independent censoring given $({\bf X},{\bf Z})$, the corresponding full log–likelihood is proportional to
$\ell_{n}(\mbox{${\theta}$},g)\;=\;\frac{1}{n}\sum_{i=1}^{n}\Big[\Delta_{i}\big\{g(T_{i},\mathbf{X}_{i})+\mbox{${\theta}$}^{\top}\mathbf{Z}_{i}\big\}-\int_{0}^{\tau}Y_{i}(t)\,\exp\!\big\{g(t,\mathbf{X}_{i})+\mbox{${\theta}$}^{\top}\mathbf{Z}_{i}\big\}\,dt\Big].$ (2)
where $Y_{i}(t)\;=\;\mathbf{1}(T_{i}\geq t)$, $t\in[0,\tau]$, $i=1,\ldots,n$, are at-risk processes, and $\tau$ is a prespecified constant. Because $g$ is modeled by a DNN, the predictable cumulative hazard in (2) contains a time-dependent integral that lacks a closed-form expression and must therefore be approximated numerically during optimization. We therefore introduce a grid $0=t_{0}<t_{1}<\cdots<t_{m_{n}}=\tau$ and approximate the integral by the Riemann–type sum
$\int_{0}^{\tau}Y_{i}(t)\,\exp\!\{g(t,\mathbf{X}_{i})+\mbox{${\theta}$}^{\top}\mathbf{Z}_{i}\}\,dt\;\approx\;\sum_{j:\,Y_{i}(t_{j-1})=1}\big(\min\{T_{i},t_{j}\}-t_{j-1}\big)\,\exp\!\{g(t_{j},\mathbf{X}_{i})+\mbox{${\theta}$}^{\top}\mathbf{Z}_{i}\}.$
Defining $\Delta_{ij}=\mathbf{1}\{T_{i}\in(t_{j-1},t_{j}],\Delta_{i}=1\}$, we obtain the discretized likelihood
$\widetilde{\ell}_{n}(\mbox{${\theta}$},g)\;=\;\frac{1}{n}\sum_{i=1}^{n}\sum_{j:\,Y_{i}(t_{j-1})=1}\Big[\Delta_{ij}\,\big\{g(t_{j},\mathbf{X}_{i})+\mbox{${\theta}$}^{\top}\mathbf{Z}_{i}\big\}-\big\{\min(T_{i},t_{j})-t_{j-1}\big\}\,\exp\!\{g(t_{j},\mathbf{X}_{i})+\mbox{${\theta}$}^{\top}\mathbf{Z}_{i}\}\Big].$ (3)
In practice, discretizing the time axis on a fixed fine grid and augmenting it with all observed times (event and censoring times) provides an accurate and computationally efficient approximation, as detailed in Section 5.1. This construction automatically produces a dense grid in regions where failures occur, precisely where the integrand contributes most to the likelihood.
We consider two estimators. The oracle maximum–likelihood estimator is defined as
$(\widehat{}\mbox{${\theta}$},\widehat{g})\;=\;\arg\max_{\,g\in\mathcal{G}(K,s,\boldsymbol{p},D),\;\theta\in\Theta\subseteq\mathbb{R}^{p}}\;\ell_{n}(\mbox{${\theta}$},g)\,.$ (4)
The estimator that is actually implemented replaces the at-risk integral with its discretized version $\widetilde{\ell}_{n}$. We therefore define the numerically integrated likelihood estimator as
$(\widetilde{}\mbox{${\theta}$},\widetilde{g})\;=\;\arg\max_{\,g\in\mathcal{G}(K,s,\boldsymbol{p},D),\;\theta\in\Theta}\;\widetilde{\ell}_{n}(\mbox{${\theta}$},g)\,.$ (5)
Theorems 1–4 establish the asymptotic properties of (4), while Theorem 5 shows that under the chosen grid these properties carry over to the estimator (5) implemented in practice.
4 Asymptotic Analysis
4.1 Notation
Some additional notation will be used throughout. We write $a_{n}\lesssim b_{n}$ if there exists a constant $c>0$ such that $a_{n}\leq c\,b_{n}$ for all $n$, and $a_{n}\asymp b_{n}$ if both $a_{n}\lesssim b_{n}$ and $b_{n}\lesssim a_{n}$ hold. For a column vector $\bf a$}, ${\bf a}^{\otimes 2}={\bf a}{\bf a}^{\sf T}$. Let $P_{0}$ denote the distribution of a generic observation $(T,\Delta,{\bf X},{\bf Z})$. For any measurable function $h=h(T,{\bf X},{\bf Z})$ we define the unweighted $L^{2}_{P_{0}}$ norm $\|h\|^{2}_{L^{2}({P_{0}})}=\mathbb{E}_{0}\{h(T,{\bf X},{\bf Z})^{2}\}$ and the event–weighted norm $\|h\|^{2}_{L^{2}(P_{0},\Delta)}=\mathbb{E}_{0}\{\Delta\,h(T,{\bf X},{\bf Z})^{2}\big\}$.
Let $\eta=(\mbox{${\theta}$},g)$, write
$\chi_{\eta}(t,{\bf X},{\bf Z})=g(t,{\bf X})\ +\ \theta^{\top}{\bf Z}\,$,
and denote the true parameter by $\eta_{o}=(\mbox{${\theta}$}_{o},g_{o})$. For two parameters $\eta_{1},\eta_{2}$, we measure their respective event-weighted score distance by
$d(\eta_{1},\eta_{2})\ =\ \Big(\mathbb{E}_{0}\big[\Delta\big\{\chi_{\eta_{1}}(T,{\bf X},{\bf Z})-\chi_{\eta_{2}}(T,{\bf X},{\bf Z})\big\}^{2}\big]\Big)^{1/2}\ =\ \|\chi_{\eta_{1}}-\chi_{\eta_{2}}\|_{L^{2}({P_{0},\Delta})}\,.$
Finally, the population loss is defined as
$\ell_{0}(\eta)\;=\;\mathbb{E}_{0}\Big[\Delta\,\chi_{\eta}(T,{\bf X},{\bf Z})-\int_{0}^{\tau}Y(t)\,\exp\!\big\{\chi_{\eta}(t,{\bf X},{\bf Z})\big\}\,dt\Big]\,.$ (6)
4.2 Assumptions
A1.
The failure time $U$ and censoring time $C$ are conditionally independent given $({\bf X}^{\top},{\bf Z}^{\top})$, and the censoring mechanism is noninformative in the sense that the censoring process is noninformative for $(\mbox{${\theta}$},g)$. The true nuisance function satisfies $g_{o}\in\mathcal{H}(q,\boldsymbol{\alpha},{\bf d},\tilde{{\bf d}},M)$.
A2.
The covariate vector $({\bf X},\mathbf{Z})$ takes values in a bounded subset of $\mathbb{R}^{p+q}$, denoted $\mathcal{X}\times\mathcal{Z}$, where both sets are compact. Without loss of generality, assume $\mathcal{X}\subset[0,1]^{d}$ and that $\|\mathbf{Z}\|_{2}\leq B_{{\bf Z}}$ almost surely. The joint density of $({\bf X},{\bf Z})$ is bounded away from zero on $\mathcal{X}\times\mathcal{Z}$. The parameter space for the finite-dimensional component is restricted to the compact set $\Theta\;=\;\{\vartheta\in\mathbb{R}^{p}:\ \|\vartheta\|\leq M_{\Theta}\}$. Finally, the study follows participants over a fixed, finite time horizon $\tau>0$.
A3.
The network architecture satisfies $K=O(\log n)$, $s=O\!\big(n\gamma_{n}^{2}\log n\big)$, and $n\gamma_{n}^{2}\;\min_{0\leq\ell\leq K}p_{\ell}\lesssim\max_{0\leq\ell\leq K}p_{\ell}\lesssim n$. In addition, the sieve is uniformly bounded, $\|g\|_{\infty}\leq D_{n}$ for a known sieve radius $D_{n}=O(1)$.
A4.
There exist $\delta_{1},\delta_{2}\in(0,1]$ such that, almost surely, $\Pr(\Delta=1\mid T,{\bf X},{\bf Z})\geq\delta_{1}$ and $\Pr(U\geq\tau\mid{\bf X},{\bf Z})\geq\delta_{2}$.
A5.
The covariate vector $\mathbf{Z}$ does not include a constant (intercept) term. Furthermore, the residual covariance matrix of $\mathbf{Z}$ given $\mathbf{X}$ is positive definite, that is,
$\Sigma_{\mathbf{Z}}=\mathbb{E}\!\left[\big\{\mathbf{Z}-\mathbb{E}(\mathbf{Z}\mid\mathbf{X})\big\}\big\{\mathbf{Z}-\mathbb{E}(\mathbf{Z}\mid\mathbf{X})\big\}^{\top}\right]\succ 0\,.$
Assumptions A1A3, A5 used also in zhong2022deep under the partial linear Cox regression model. A3 determines the structure of the neural network and provides a trade-off between the approximation error and estimation error (schmidthieber2020nonparametric). Assumption A2 is a standard assumption for semi- or non-parametric regression. Moreover, A1 and A2 imply that there exist constants $0<c<C<\infty$ (that depend on $D_{n},B_{{\bf Z}}$ and $M_{\Theta}$) such that for all $t\in(0,\tau]$ and $({\bf X},{\bf Z})\in\mathcal{X}\times\mathcal{Z}$, $c_{h}\leq h_{o}(t|{\bf X},{\bf Z})\leq C_{h}$. Assumption A5 is a standard idenfiability assumption in survival analysis. Assumption A4 is required since in the hazard function (1) time is one of the components in the function $g_{o}$ estimated by the DNN and therefore the effects of the covariates $\bf X$}} and time are not disatengled, in contrast to the partial linear Cox model of zhong2022deep. Hence we need a non-vanishing chance to observe events over the entire study horizon for each covariate pattern.
4.3 Asymptotic Results
The proofs of all lemmas and theorems in this section appear in the appendix. We begin by establishing several basic properties of the model. A key step is to understand how the counting–process framework allows us to translate expressions involving at-risk process into simpler quantities expressed directly in terms of the event indicator $\Delta$. The next lemma formalizes this relationship, showing that these two representations are equivalent up to a multiplicative constant. This equivalence will be used repeatedly throughout the analysis.
Lemma 1.
Under Assumptions A1A2, for any nonnegative, predictable function $f(t,{\bf X},{\bf Z})$ with $\mathbb{E}\!\left[\int_{0}^{\tau}Y(t)\,f(t,{\bf X},{\bf Z})\,dt\right]<\infty$,
$\mathbb{E}\!\left\{\int_{0}^{\tau}Y(t)\,f(t,{\bf X},{\bf Z})\,dt\right\}\;\asymp\;\mathbb{E}\!\left\{\Delta\,f(T,{\bf X},{\bf Z})\right\}.$
The proof uses only the positivity of the hazard function together with the compensator identity for counting processes, and therefore applies uniformly to all functions that arise in our later analysis.
Lemma 2.
Let $\chi_{\eta_{j}}(t,{\bf X},{\bf Z})$, $j=1,2$, be bounded predictable processes satisfying
$\mathbb{E}\!\left\{\int_{0}^{\tau}\chi_{\eta_{j}}(t,{\bf X},{\bf Z})^{2}\,dt\right\}<\infty\qquad j=1,2\,.$
Define $\phi(t,{\bf X},{\bf Z})=\chi_{\eta_{1}}(t,{\bf X},{\bf Z})-\chi_{\eta_{2}}(t,{\bf X},{\bf Z})$. Under Assumptions A1, A2 and A4, there exist finite constants $0<c_{1}\leq c_{2}<\infty$ and $0<c_{3}\leq c_{4}<\infty$, depending only on $(\tau,\delta_{1},\delta_{2},c_{h},C_{h})$, such that
$c_{1}\,\mathbb{E}\!\big\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\big\}$ $\;\leq\;\mathbb{E}\!\big\{\phi(T,{\bf X},{\bf Z})^{2}\big\}\;\leq\;c_{2}\,\mathbb{E}\!\big\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\big\}\,$, (7)
$c_{3}\,\mathbb{E}\!\big\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\big\}$ $\;\leq\;\mathbb{E}_{{\bf X},{\bf Z}}\!\Big\{\int_{0}^{\tau}\phi(t,{\bf X},{\bf Z})^{2}\,dt\Big\}\;\leq\;c_{4}\,\mathbb{E}\!\big\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\big\}\,.$ (8)
In particular, if $\mathbb{E}\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\}=0$, then $\phi(T,{\bf X},{\bf Z})=0$ $P_{0}$–a.s. and $\phi(t,{\bf X},{\bf Z})=0$ for Lebesgue-a.e. $t\in[0,\tau]$ and $P_{{\bf X},{\bf Z}}$–a.e. $({\bf X},{\bf Z})$, i.e. $\chi_{\eta_{1}}(t,{\bf X},{\bf Z})=\chi_{\eta_{2}}(t,{\bf X},{\bf Z})$ for $dt\otimes P_{{\bf X},{\bf Z}}$–almost all $(t,{\bf X},{\bf Z})$.
Lemma 2 shows that the distance $d(\cdot,\cdot)$ is a valid metric on the parameter space: $d(\eta_{1},\eta_{2})=0$ if and only if the induced linear score coincide almost everywhere. In addition, the lemma shows that the event-weighted norm and the corresponding unweighted norms are equivalent up to multiplicative constants. As a result, any bound or convergence statement expressed in terms of the metric $d$ can be translated into an equivalent statement under the usual $L^{2}$ norm and vice versa. This equivalence is used repeatedly throughout the theoretical development.
Next we study the shape of the population loss around $\eta_{o}=(g_{o},\mbox{${\theta}$}_{o})$.
Lemma 3.
Assume A1A2. For any $\eta\in\Theta\times\mathcal{H}(q,\boldsymbol{\alpha},{\bf d},\tilde{{\bf d}},M)$,
$\ell_{0}(\eta)-\ell_{0}(\eta_{o})\;\asymp\;-\,d(\eta,\eta_{o})^{2}.$
Lemma 3 shows that the population loss is well behaved in a neighbourhood of the truth: any deviation from $\eta_{o}$ incurs a loss of order $d(\eta,\eta_{o})^{2}$. Combined with Lemma 2, this implies that $\eta_{o}$ is the unique minimizer of the population loss over the square–integrable class under consideration. This population curvature is a key ingredient in the analysis that will later be paired with empirical process bounds to obtain convergence rates for $\widehat{g}$ and to derive the asymptotic distribution of $\widehat{\mbox{${\theta}$}}$.
We now move to the main asymptotic results. Our ultimate goal is to derive the asymptotic distribution of $\widehat{}\mbox{${\theta}$}$ and to show that it attains the semiparametric information bound. To this end, we first establish the convergence rate of the nuisance estimator $\widehat{g}$, show that this rate is essentially optimal, characterize the efficient score and information for $\mbox{${\theta}$}_{o}$, and then combine these ingredients to obtain an asymptotic distribution of $\widehat{}\mbox{${\theta}$}$. Finally, we verify that the numerical approximation used in practice does not alter the limit law.
We begin by quantifying how well the estimator $\widehat{g}$ recovers the unknown function $g_{o}$. The next theorem gives a nonparametric convergence rate for $\widehat{g}$ in the event–weighted $L^{2}_{P_{0}}$ norm that naturally appears in the loss and in the distance function $d$. The rate is expressed in terms of the complexity parameter $\gamma_{n}$ and will later be used to control remainder terms in the expansion of the efficient score.
Theorem 1.
[…]
$\left(\mathbb{E}\!\Bigl[\Delta\bigl\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\bigr\}^{2}\Bigr]\right)^{1/2}=O_{p}(\tau_{n})$ (9)
where the expectation is taken with respect to an independent draw $(T,\Delta,{{\bf X}},{{\bf Z}})\sim P_{0}$ from the true distribution, and $\tau_{n}=\gamma_{n}\log^{2}n$. The quantity in (9) is random through its dependence on the training sample used to construct $\widehat{g}$.
The conclusion of Theorem 1 is strengthened by Lemma 2, which shows that $d(\eta,\eta_{0})^{2}$ is equivalent to the unweighted norm $\|\chi_{\eta}-\chi_{\eta_{0}}\|^{2}_{L^{2}(dt\otimes P_{\mathbf{X},\mathbf{Z}})}$. Consequently, the estimator converges in $L^{2}$ over the entire time domain $(0,\tau)$ with respect to Lebesgue measure, and not merely at observed event times or in a way that depends on the censoring pattern.
The next result shows that the rate obtained in Theorem 1 is essentially optimal, up to logarithmic factors. It establishes a minimax lower bound for estimation of $g_{o}$ under the same event–weighted $L^{2}_{P_{0}}$ loss, taken uniformly over the parameter space $\Theta$ and the nuisance function class $\mathcal{H}(q,\boldsymbol{\alpha},{\bf d},\tilde{{\bf d}},M)$.
Theorem 2.
Under A1A2 and A5 there exists $c>0$ such that
$\inf_{\widehat{g}}\sup_{\theta_{o}\in\Theta,\,g_{o}\in\mathcal{H}(q,\boldsymbol{\alpha},{\bf d},\tilde{{\bf d}},M)}\mathbb{E}_{P_{g_{o},\theta_{o}}}\!\left[\Delta\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\}^{2}\right]\;\;\geq\;c\,\gamma_{n}^{2}\,.$
Next, we turn to the parametric component $\mbox{${\theta}$}_{o}$ and characterize the efficient score and information bound in our semiparametric model. To characterize the nuisance tangent space associated with $g_{o}$, we consider differentiable one-dimensional submodels that pass through $g_{o}$. Let $\mathcal{H}_{g_{o}}$ denote the collection of all functions $g\in L^{2}([0,\tau]\times[0,1]^{r})$ for which there exist a submodel $\{g_{b}:b\in(-1,1)\}\subset\mathcal{H}(q,\boldsymbol{\alpha},{\bf d},\tilde{{\bf d}},M)$ satisfying
$\lim_{b\to 0}\left\|b^{-1}(g_{b}-g_{o})-g\right\|_{L^{2}([0,\tau]\times[0,1]^{r})}=0\,.$
Here, the scalar $b$ is a perturbation parameter indexing the submodel. The set of all attainable first–order directions at $g_{o}$ is then defined as
$\mathcal{T}_{g_{o}}=\Big\{g\in L^{2}([0,\tau]\times[0,1]^{r}):\exists\,\{g_{b}\}\in\mathcal{H}_{g_{o}}\text{ s.t. }\lim_{b\to 0}\left\|b^{-1}(g_{b}-g_{o})-g\right\|=0\Big\}\,.$
Finally, semiparametric efficiency theory requires a closed linear subspace, so we take the closure of the linear span of $\mathcal{T}{g_{o}}$, denoted $\overline{\mathcal{T}}{g_{o}}$, which forms the nuisance tangent space.
[…]
$\mathbb{E}\!\left\{\Delta\,\|{\bf Z}-{\bf g}^{*}(T,{\bf X})\|_{c}^{2}\right\},\quad\|v\|_{c}^{2}=(v_{1}^{2},\ldots,v_{p}^{2})^{\top}\,$,
so ${{\bf g}}^{*}(t,{\bf X})$ projects the covariates $\bf Z$} onto the tangent space of the nuisance function. Finally, under the counting-process representation of the model (andersen1982cox), the counting process $N(t)=\Delta I(T\leq t)$ admits the Doob-Meyer decomposition where the associated martingale is
$M(t)=N(t)-\int_{0}^{t}Y(s)\exp\{\mbox{${\theta}$}_{o}^{\sf T}{{\bf Z}}+g_{o}(s,{{\bf X}})\}\,ds\,.$
Theorem 3.
Under Assumptions A1A2 and A5, the efficient score for $\mbox{${\theta}$}_{o}$ in the semiparametric model (1) is
$\ell_{\theta_{o}}^{*}(T,{\bf X},{\bf Z},\Delta)\;=\;\int_{0}^{\tau}\big\{{\bf Z}-{\bf g}^{*}(t,{\bf X})\big\}\,dM(t)\,$,
where ${\bf g}^{*}(t,{\bf X})$ is the projection of $\bf Z$} onto the nuisance tangent space and $M(t)$ is the associated counting–process martingale. The semiparametric information bound for $\mbox{${\theta}$}_{0}$ is
$I(\mbox{${\theta}$}_{0})\;=\;\mathbb{E}\!\left\{\,\ell_{\theta_{o}}^{*}(T,{\bf X},{\bf Z},\Delta)^{\otimes 2}\right\}=\mathbb{E}\!\left[\,\Delta\,\{{\bf Z}-{\bf g}^{*}(T,{\bf X})\}^{\otimes 2}\right]\,.$
Consequently, the asymptotic variance of any regular asymptotically linear (RAL) estimator of $\mbox{${\theta}$}_{o}$ is bounded below by $I(\mbox{${\theta}$}_{o})^{-1}$ in the Loewner sense.
With the curvature and metric properties of the population loss, the convergence rate from Theorem 1, and the efficient score representation of Theorem 3, we can now establish an asymptotic linear expansion for $\widehat{\mbox{${\theta}$}}$. This, in turn, yields the asymptotic distribution of the estimator.
Theorem 4.
Under Assumptions A1A5, suppose that the efficient information matrix $I(\mbox{${\theta}$}_{o})$ is nonsingular and that $\sqrt{n}\,\tau_{n}^{2}\;\to\;0$ as $n\to\infty$, where $\tau_{n}=\gamma_{n}\log^{2}n$. Then,
$\sqrt{n}(\widehat{\mbox{${\theta}$}}-\mbox{${\theta}$}_{o})=I(\mbox{${\theta}$}_{o})^{-1}\,\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\ell^{*}_{\theta_{o}}(T_{i},{\bf X}_{i},{\bf Z}_{i},\Delta_{i})+o_{p}(1)\,$,
and hence
$\sqrt{n}(\widehat{\mbox{${\theta}$}}-\mbox{${\theta}$}_{o})\;\overset{D}{\longrightarrow}\;N({\bf 0},I(\mbox{${\theta}$}_{o})^{-1})\,.$
The key technical observation is that, in the Taylor expansion of the population score around $g_{o}$, the first–order derivative in $\widehat{g}-g_{o}$ direction vanishes because the efficient score is orthogonal to the nuisance tangent space. The remaining second–order term is of smaller order, specifically $o_{p}(n^{-1/2})$, provided that $n\gamma_{n}^{4}\to 0$ as $n\to\infty$.
Finally, we justify the numerical approximation used to compute the estimator in practice. The next theorem shows that, under mild conditions the estimator $\widetilde{}\mbox{${\theta}$}$ has the same asymptotic distribution as the oracle estimator $\widehat{}\mbox{${\theta}$}$.
Theorem 5.
[…]
$\delta_{n}\;=\;\max_{1\leq j\leq m_{n}}(t_{j}-t_{j-1})$
denote the mesh size of this grid. Assume that $\delta_{n}=o(\tau_{n}^{2})$ and $\sqrt{n}\,\delta_{n}\;\rightarrow\;0$, as $n\rightarrow\infty$, and that the conditions of Theorems 1 and 4 hold. Then,
  1. 1.
    The numerically integrated likelihood estimator attains the same convergence rate as the oracle estimator:
    $d(\widetilde{\eta},\eta_{o})=O_{p}(\tau_{n})\,,\qquad\|\widetilde{}\mbox{${\theta}$}-\mbox{${\theta}$}_{o}\|=O_{p}(\tau_{n})\,.$
  2. 2.
    The parametric estimators are asymptotically equivalent:
    $\sqrt{n}(\widetilde{}\mbox{${\theta}$}-\widehat{}\mbox{${\theta}$})\;\xrightarrow{P}\;{\bf 0}\,.$
    Consequently,
[…]
    where $I(\mbox{${\theta}$}_{o})$ is the efficient information matrix from Theorem 3.
Theorem 5 shows that the difference between the exact and the numerically integrated likelihood estimators is $o_{p}(n^{-1/2})$, so that both share the same asymptotic normal distribution. Part 2 verifies that the required error bounds hold for a simple equally spaced grid with $n$ time points, where the per–observation quadrature error is of order $1/n$. This ensures that the asymptotic efficiency result for $\widehat{}\mbox{${\theta}$}$ carries over directly to the implementable estimator $\widetilde{}\mbox{${\theta}$}$ based on numerical integration.
5 Pratical Implementation
5.1 Applying DNN
In practice, we approximate $g_{o}$ by a fully connected ReLU network and model the linear term $\mbox{${\theta}$}^{\top}{\bf Z}$ using a parallel linear layer. The network computes $\widetilde{g}(t,{\bf X})$ and $\widetilde{\mbox{${\theta}$}}^{\top}{\bf Z}$ and their sum is inserted into the loss based on (3). This architecture jointly learns the nuisance function $g$ and a finite–dimensional linear effect of $\bf Z$} in a way that mirrors the model structure.
To evaluate the approximated log-likelihood (3) in practice, we discretize the time axis on a fixed grid (e.g., $n$) and augment it with all observed times, yielding $0=t_{0}<t_{1}<\cdots<t_{m_{n}}=\tau$. The dataset is then expanded in the usual counting‐process manner. Each subject $i$ with data $(T_{i},\Delta_{i},\mathbf{X}_{i},\mathbf{Z}_{i})$ is represented by one dummy-observation for every interval $t_{k},t_{k+1})$ such that $t_{k}\leq T_{i}$. All dummy-observations inherit $(\mathbf{X}_{i},\mathbf{Z}_{i})$; the event indicator $\Delta_{ik}=\mathbf{1}\{T_{i}\in(t_{k-1},t_{k}],\Delta=1\}$ is set to 0 except for the final interval containing $T_{i}$, where it takes the original value $\Delta_{i}$. This choice of grid and data expansion enabling efficient computation within standard deep-learning frameworks. The construction parallels the standard expanded-data representation used for Cox models with time-dependent covariates (therneau2000cox; hu2023conditional).
In theory we allow for sparsity, but in practice we do not explicitly enforce it. Network hyperparameters (number of layers, hidden units, activation functions, learning rate, batch size) are selected via tuning on the training set. Training maximizes the empirical likelihood using the ADAM optimizer in TensorFlow. The data are split into training and validation sets, with 33% reserved for early stopping. Training continues until the validation loss stops improving, and the weights achieving the best validation performance are retained. The linear $\mathbf{Z}$-branch is initialized using the coefficients from a standard Cox model fit on $(\mathbf{X},\mathbf{Z})$, providing a good starting point for the linear component.
After training, we extract the weights from the $\mathbf{Z}$-branch as an initial estimate of $\theta$}. To mitigate the effect of early stopping on the optimization of $\theta$}, we then perform a separate numerical maximization of the likelihood with respect to $\boldsymbol{\theta}$, holding $\widetilde{g}$ fixed at its network estimate. Thus, we treat $\widetilde{g}$ as known and solve a low-dimensional optimization problem for $\boldsymbol{\theta}$ using SciPy’s optimization routines. This post-hoc refinement consistently improved the finite-sample behavior of $\widetilde{\mbox{${\theta}$}}$ in our experiments. For the theoretical results, the critical requirement is that $\widehat{g}$ achieves the convergence rate needed for Theorem 4.
5.2 Estimation of the Covariance Matrix $I(\mbox{${\theta}$}_{o})^{-1}$
Inference on $\mbox{${\theta}$}_{0}$ relies an estimating the efficient information matrix $I(\mbox{${\theta}$}_{o})$ from Theorem 4, whose inverse provides the asymptotic covariance of the estimator $\widetilde{\mbox{${\theta}$}}$. Recall that the efficient information is
$I(\mbox{${\theta}$}_{o})=\mathbb{E}\!\Big[\Delta\{{\bf Z}-{{\bf g}}^{\ast}(T,{\bf X})\}\{{\bf Z}-{\bf g}^{\ast}(T,{\bf X})\}^{\top}\Big]\,$,
where $\bf g$$^{\ast}$ is the event–weighted projection of $\bf Z$} onto the nuisance tangent space. In practice we estimate $\bf g$$^{\ast}$ using an additional set of neural networks and then construct a plug–in estimate of $I(\mbox{${\theta}$}_{o})$ from the corresponding residuals.
Specifically, we approximate $\bf g$$^{\ast}$ coordinatewise. For each $j=1,\ldots,p$ we fit a separate fully connected network $\widehat{g}^{\ast}_{j}(T,{\bf X})$ by least squares using only uncensored observations, that is by
$\min\sum_{i:\,\Delta_{i}=1}\{Z_{ij}-\widehat{g}^{\ast}_{j}(T_{i},{{\bf X}}_{i})\}^{2}\,.$
Each network $j$, $j=1,\ldots,p$, takes $(T_{i},\mathbf{X}_{i})$ as input, $Z_{ij}$ as target, and uses the same architecture and hyperparameters as the main model; empirically, these choices produce stable estimates and are not sensitive to tuning.
To avoid overfitting, we employ 55–fold cross–fitting. The data are split into five folds, for each fold and each coordinate $j$, the corresponding network is trained on the remaining four folds (restricting to $\Delta=1$) and predictions $\widehat{g}^{\ast}_{j}(T_{i},X_{i})$ are obtained for the held-out fold. for all $i$ in the held–out fold. Aggregating across folds yields, for each observation $i$, the estimated projection vector
$\widehat{\bf g}^{\ast}(T_{i},{{\bf X}}_{i})=\bigl(\widehat{g}^{\ast}_{1}(T_{i},{{\bf X}}_{i}),\ldots,\widehat{g}^{\ast}_{p}(T_{i},{{\bf X}}_{i})\bigr)^{\top}\,.$
We then compute the residuals ${\bf R}_{i}={{\bf Z}}_{i}-\hat{\bf g}^{\ast}(T_{i},{{\bf X}}_{i})\in\mathbb{R}^{p}$, $i=1,\ldots,n$, and stack them into an $n\times p$ matrix $\mathcal{R}$. Our estimator of the efficient information is $\widehat{I}=\mathcal{R}^{\top}\mathcal{R}/n$, and the asymptotic covariance is $\widehat{I}^{-1}$. In practice this procedure is numerically stable, requires no tuning beyond the main network fit, and produces reliable covariance estimates for $\widetilde{\mbox{${\theta}$}}$.
6 Simulation Study
In this section, we compare the finite-sample performance of the proposed FLEXI-Haz method with the partially linear neural PH estimator of zhong2022deep. Our evaluation focuses on the estimation accuracy of the finite-dimensional parameter $\theta$} and on the coverage properties of the corresponding Wald confidence intervals.
We consider a setting with both nuisance and primary covariates. Let ${{\bf X}}=(X_{1},X_{2},X_{3})^{\top}$ and ${{\bf Z}}=(Z_{1},Z_{2})^{\top}$, where each covariate is independently sampled from a uniform distributed over [1,1][-1,1]. The true parameter of interest is $\mbox{${\theta}$}_{o}=(b_{1},b_{2})^{\top}=(2,-1)^{\top}$. For the nuisance component, we define $f({\bf X})=0.2(X_{1}+X_{2})+0.5X_{1}X_{2}+X_{3}^{2}$, which induces nonlinear and interaction effects and interactions among the nuisance covariates. The event time $U$ is generated from a hazard function of the form
$h_{o}(t\mid{\bf X},{\bf Z})=0.1\exp\!\left\{(0.1+f({\bf X})^{2})t+\mbox{${\theta}$}_{o}^{\top}{\bf Z}\right\},\qquad t\in[0,\tau]\,.$
with $\tau=30$ chosen as the fixed study horizon. Right censoring is introduced through an independent censoring time $C$ sampled from an exponential distribution with expectation $1/30$, combined with administration censoring at time 30. The parameters of the data-generating mechanism are chosen so that the censoring rate is approximately 30%.
For both FLEXI-Haz and the estimator of zhong2022deep, the nuisance function is estimated using a fully connected feedforward network with ReLU activations. The only and difference between the two estimators lies in the underlying hazard model, and the training pipeline and tuning strategy are identical. Hyperparameters are selected once, using only one simulated dataset. We construct the expanded dataset using a fixed time grid of $m_{n}=n$ points augmented with all observed times, and perform 5-fold cross-validation over the following search space:
  • number of hidden layers: $\{1,2,3,4,5,6\}$
  • width (nodes per hidden layer): $\{20,40,80,160,320\}$
  • learning rate: $\{10^{-3},\,5\times 10^{-3},\,10^{-2},\,5\times 10^{-2}\}$
Training uses early stopping with a patience of 35 epochs and a batch size of $10^{5}$ on the expanded dataset. For each hyperparameter combination we train the network once and record the average validation loss across the five folds; the configuration achieving the smallest average loss is selected. Cross-validation selects, for FLEXI-Haz, a network with five hidden layers, width 20, and learning rate $10^{-3}$. For the estimator of zhong2022deep, the selected architecture has four hidden layers, width 80, and the same learning rate $10^{-3}$.
Table 1 presents the results for $n=8000$ based on 200 replications. The proposed estimator, FLEXI-Haz, is essentially unbiased and its empirical variability closely matches the theoretical standard errors, yielding coverage probabilities near their nominal levels. In contrast, the partially linear PH estimator of zhong2022deep exhibits substantial bias in both components of $\boldsymbol{\theta}_{o}$, leading to near-zero empirical coverage.
Table 1: Simulation results for $n=8000$ observations, based on 200 replications. The true parameter is $\mbox{${\theta}$}_{o}=(2,-1)^{\top}$. “Est” denotes the mean of the estimated coefficients; “Emp SD” the empirical standard deviation; “Est SE” the estimated theoretical standard error; and “90% , 95%” the empirical coverage rates of the corresponding Wald-type confidence intervals.
FLEXI-Haz zhong2022deep
Est Emp SD Est SE 90% 95% Est Emp SD Est SE 90% 95%
$\theta_{o1}$ 1.997 0.040 0.035 0.878 0.935 1.764 0.039 0.034 0.000 0.000
$\theta_{o2}$ -0.994 0.027 0.029 0.937 0.959 -0.886 0.030 0.028 0.015 0.035
7 Discussion
In this work we proposed FLEXI–Haz, a deep semiparametric survival model that combines a fully non-PH formulation with a partially linear structure. By modeling the log-hazard as $\boldsymbol{\theta}_{o}^{\top}\mathbf{Z}+g_{o}(t,\mathbf{X})$ and estimating $g_{o}$ with a deep ReLU network, we retain interpretability and valid inference for the low-dimensional parameter $\boldsymbol{\theta}_{o}$ while allowing highly flexible, time-varying effects of the nuisance covariates. Our main theoretical result is that FLEXI–Haz enjoys the same key properties that zhong2022deep established for partially linear Cox models under PH: the neural estimator of $g_{o}$ attains the optimal composite minimax rate, and the estimator of $\boldsymbol{\theta}_{o}$ is $\sqrt{n}$–consistent, asymptotically normal, and semiparametrically efficient. Crucially, these properties continue to hold in our more general, baseline-free non-PH setting, despite the presence of time in the nuisance component and the use of the full likelihood rather than a partial likelihood.
The simulation study supports these theoretical findings and illustrates their practical relevance. In a setting with nonlinear time–covariate interactions, FLEXI–Haz produces essentially unbiased estimates of $\boldsymbol{\theta}_{o}$, with empirical standard deviations that align closely with the estimated standard errors and nominal coverage for Wald-type confidence intervals. In contrast, the partially linear PH estimator of zhong2022deep, which forces the nuisance component to depend only on $\mathbf{X}$, exhibits substantial bias and severe undercoverage when the true hazard is non-proportional. This behavior is exactly what the theory predicts: when time-varying effects are incorrectly constrained to be time-invariant, the nuisance approximation error leaks into the finite-dimensional parameter. Allowing $g_{o}$ to depend on both $t$ and $\mathbf{X}$ removes this source of misspecification, and the empirical results confirm that the resulting estimator behaves in accordance with the asymptotic theory.
Although FLEXI–Haz provides consistent estimators of the hazard and enables estimation of the conditional survival function $S(t\mid\mathbf{X},\mathbf{Z})$, the present work does not derive the asymptotic distribution of the resulting survival-curve estimator. Nevertheless, valid pointwise confidence intervals and simultaneous confidence bands can still be constructed by applying the ensemble–bootstrap procedure developed by arie2024confidence. That method is fully nonparametric, accommodates right censoring, and can be used as a post–hoc procedure on top of FLEXI–Haz without modifying the training algorithm.
Our implementation focuses on fully connected ReLU networks, which fit naturally into the composite Hölder framework and for which sharp approximation and entropy bounds are available. This choice enables us to obtain clean minimax rates and empirical-process control. At the same time, fully connected architectures are not always the most natural representation for complex structured data such as images. Convolutional neural networks (CNNs) are often preferable in imaging applications, and several deep survival methods in the literature rely on convolutional encoders. While the FLEXI–Haz likelihood and estimation strategy can be used with convolutional backbones in practice, our current theory does not cover convolutional architectures, and the existing results on sparse ReLU networks cannot be directly transferred to CNNs. Extending the approximation and efficiency analysis to convolutional or other structured architectures is therefore an important direction for future work.
A further practical advantage of our formulation is that it meshes naturally with the standard “pseudo-observation” representation used to handle time-dependent covariates. The full-likelihood derivation already requires us to expand each subject into multiple pseudo-observations over risk intervals. This is precisely the data-expansion scheme commonly used to encode time-dependent covariates in Cox-type models. As a result, extending FLEXI–Haz to time-dependent covariates is conceptually straightforward: one augments $\mathbf{X}$ to include time-varying components and constructs the expanded dataset in the usual way. The theoretical development is also compatible with this extension, as the key arguments are expressed in terms of predictable processes and integrals over the at-risk time, and do not rely on time-independence of the covariates beyond notational convenience.
The main current limitation of FLEXI–Haz is computational rather than statistical. Because we work with the full likelihood and approximate the cumulative hazard via a Riemann-type sum on a fine time grid, the expanded dataset contains $O(n\,m_{n})$ pseudo-observations, where $m_{n}$ is the number of grid points. In the default choice $m_{n}\asymp n$, this leads to an effective $O(n^{2})$ cost per epoch, which can become prohibitive for very large studies or dense event processes. Our experiments show that this is manageable for moderate $n$, but improving scalability is crucial for applications with hundreds of thousands of subjects or very fine time resolution.
Several strategies could mitigate this computational burden. One natural approach is to sub-sample pseudo-observations corresponding to non-events in the risk sets, e.g., by keeping all intervals containing failures and a carefully chosen random subset of non-event intervals, with importance weights to preserve unbiasedness. Since the information about $\boldsymbol{\theta}_{o}$ and $g_{o}$ is concentrated near failures and in regions with high hazard, aggressive sub-sampling of “uninformative” background intervals should reduce cost with limited loss in efficiency (Keret02102023). Another line of work is to exploit structured quadrature rules or other powerful approximation methods.
Our theoretical treatment of numerical integration in Theorem 5 is deliberately abstract and suggests a broad class of admissible approximations. The only requirement is that the approximation error in the empirical loss is $O_{p}(n^{-1})$ uniformly over the parameter space, and that the corresponding error in the $\boldsymbol{\theta}$-score is $O_{p}(n^{-1/2})$. Any numerical scheme—whether based on fewer grid points, higher-order quadrature, adaptive time meshes, or stochastic approximations—that satisfies these two error bounds yields the same asymptotic distribution of $\widetilde{\mbox{${\theta}$}}$. In particular, this shows that our choice of a simple Riemann sum with $m_{n}\asymp n$ is sufficient but not necessary: one can potentially use substantially fewer time points while preserving the validity of the asymptotic theory, provided the approximation quality is high enough.
8 Appendix - Proofs
See 1
Proof of Lemma 1.
Let $N(t)=\Delta\,\mathbf{1}(T\leq t)$ be the counting process associated with the event time, and let $M(t)=N(t)-\int_{0}^{t}Y(s)\,h_{o}(s\,|\,{{\bf X}},{{\bf Z}})\,ds$ denote its martingale decomposition under the true hazard (1). Because $f(t,{\bf X},{\bf Z})$ is predictable, the stochastic integral $\int_{0}^{\tau}f(t,{{\bf X}},{{\bf Z}})\,dM(t)$ has expectation zero. Consequently,
$\mathbb{E}\!\left\{\int_{0}^{\tau}f(t,{\bf X},{\bf Z})\,dN(t)\right\}=\mathbb{E}\!\left\{\int_{0}^{\tau}f(t,{\bf X},{\bf Z})\,Y(t)\,h_{0}(t\,|\,{\bf X},{\bf Z})\,dt\right\}.$
Since $\int_{0}^{\tau}f(t,{\bf X},{\bf Z})\,dN(t)=\Delta\,f(T,{\bf X},{\bf Z})$, we obtain the identity
$\mathbb{E}\{\Delta\,f(T,{\bf X},{\bf Z})\}=\mathbb{E}\!\left\{\int_{0}^{\tau}f(t,{\bf X},{\bf Z})\,Y(t)\,h_{0}(t\,|\,{\bf X},{\bf Z})\,dt\right\}\,.$ (10)
Under Assumptions A1–A2, the hazard is uniformly bounded and bounded away from zero,
$0<c_{\lambda}\;\leq\;h_{0}(t\,|\,{\bf X},{\bf Z})\;\leq\;C_{\lambda}<\infty.$
Applying these bounds to the right-hand side of (10) yields
$c_{\lambda}\,\mathbb{E}\!\left\{\int_{0}^{\tau}f(t,{\bf X},{\bf Z})\,Y(t)\,dt\right\}\;\leq\;\mathbb{E}\{\Delta\,f(T,{\bf X},{\bf Z})\}\;\leq\;C_{\lambda}\,\mathbb{E}\!\left\{\int_{0}^{\tau}f(t,{\bf X},{\bf Z})\,Y(t)\,dt\right\}\,.$
This implies the desired equivalence:
$\mathbb{E}\!\left\{\int_{0}^{\tau}f(t,{\bf X},{\bf Z})\,Y(t)\,dt\right\}\asymp\mathbb{E}\!\{\Delta\,f(T,{\bf X},{\bf Z})\}\,.$
See 2
Proof of Lemma 2.
By the tower property of conditional expectation,
$\mathbb{E}\!\big\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\big\}=\mathbb{E}\Big\{\phi(T,{\bf X},{\bf Z})^{2}\,\Pr(\Delta=1\mid T,{\bf X},{\bf Z})\Big\}\,.$
Assumption A4 states that
$0<\delta_{1}\;\leq\;\Pr(\Delta=1\mid T,{\bf X},{\bf Z})\;\leq\;1\qquad\text{a.s}.$
Since $\phi(T,{\bf X},{\bf Z})^{2}\geq 0$, we obtain
$\delta_{1}\,\mathbb{E}\!\big\{\phi(T,{\bf X},{\bf Z})^{2}\big\}\;\leq\;\mathbb{E}\!\big\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\big\}\;\leq\;\,\mathbb{E}\!\big\{\phi(T,{\bf X},{\bf Z})^{2}\big\}\,.$
Rearranging the inequalities gives
$\mathbb{E}\!\big\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\big\}\;\leq\;\mathbb{E}\!\big\{\phi(T,{\bf X},{\bf Z})^{2}\big\}\;\leq\;\frac{1}{\delta_{1}}\,\mathbb{E}\!\big\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\big\}\,.$
Thus (7) holds with $c_{1}\leq 1$ and $c_{2}=1/\delta_{1}$.
Next, we relate the event–weighted quantity $\mathbb{E}\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\}$ to the time–integrated $L^{2}$ norm of $\phi$. Write,
$\mathbb{E}\!\big\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\big\}=\mathbb{E}_{{\bf X},{\bf Z}}\!\Big[\mathbb{E}\!\big\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\,\big|\,{\bf X},{\bf Z}\big\}\Big]\,$, (11)
and let $f_{T,\Delta=1\mid{\bf X},{\bf Z}}(t\mid{\bf X},{\bf Z})$ denote the conditional density of $T$ given $({\bf X},{\bf Z})$ on the event $\{\Delta=1\}$. Then,
$\mathbb{E}\!\big\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\,\big|\,{\bf X},{\bf Z}\big\}=\int_{0}^{\tau}\phi(t,{\bf X},{\bf Z})^{2}\,f_{T,\Delta=1\mid{\bf X},{\bf Z}}(t\mid{\bf X},{\bf Z})\,dt\,.$
Under Assumptions A1 and A2, the event-time hazard $h_{0}(t,{\bf X},{\bf Z})$ is bounded away from zero and infinity. Similarly, Assumption A4 implies the censoring mechanism is well behaved. Therefore,
$\Pr(T>t\mid{\bf X},{\bf Z})=\exp\left[-\int_{0}^{t}\{h_{o}(s\mid{\bf X},{\bf Z})+\mu(s\mid{\bf X},{\bf Z})ds\}\right]$
where $\mu(t\mid{\bf X}.{\bf Z})$ is the censoring hazard bounded away from zero and infinity, and
$f_{T,\Delta=1\mid{\bf X},{\bf Z}}(t\mid{\bf X},{\bf Z})=h_{o}(s\mid{\bf X},{\bf Z})\exp\left[-\int_{0}^{t}\{h_{o}(s\mid{\bf X},{\bf Z})+\mu(s\mid{\bf X},{\bf Z})ds\}\right]$
is bounded above and below by positive constants that depend only on $(\tau,\delta_{1},\delta_{2},c_{h},C_{h})$. Specifically, there exist $0<c_{f}\leq C_{f}<\infty$ such that, almost surely,
$c_{f}\;\leq\;f_{T,\Delta=1\mid{\bf X},{\bf Z}}(t\mid{\bf X},{\bf Z})\;\leq\;C_{f},\qquad t\in[0,\tau]\,.$
Using these bounds inside the conditional expectation (11), we obtain, that given $({\bf X},{\bf Z})$,
$c_{f}\int_{0}^{\tau}\phi(t,{\bf X},{\bf Z})^{2}\,dt\;\leq\;\mathbb{E}\!\big\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\,\big|\,{\bf X},{\bf Z}\big\}\;\leq\;C_{f}\int_{0}^{\tau}\phi(t,{\bf X},{\bf Z})^{2}\,dt\,.$
Taking expectation over $({\bf X},{\bf Z})$ yields
$c_{f}\,\mathbb{E}_{{\bf X},{\bf Z}}\!\Big\{\int_{0}^{\tau}\phi(t,{\bf X},{\bf Z})^{2}\,dt\Big\}\;\leq\;\mathbb{E}\!\big\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\big\}\;\leq\;C_{f}\,\mathbb{E}_{{\bf X},{\bf Z}}\!\Big\{\int_{0}^{\tau}\phi(t,{\bf X},{\bf Z})^{2}\,dt\Big\}.$
Thus (8) holds with $c_{3}=1/C_{f}$ and $c_{4}=1/c_{f}$ (or equivalently with $c_{3},c_{4}$ chosen directly as above).
Finally, suppose $\mathbb{E}\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\}=0$. Since $\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\geq 0$, this implies $\Delta\,\phi(T,{\bf X},{\bf Z})^{2}=0$ almost surely, hence $\phi(T,{\bf X},{\bf Z})=0$ $P_{0}$–a.s. on $\{\Delta=1\}$. From (7) we also have
$0\;\leq\;\mathbb{E}\!\big\{\phi(T,{\bf X},{\bf Z})^{2}\big\}\;\leq\;c_{2}\,\mathbb{E}\!\big\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\big\}=0\,$,
so $\mathbb{E}\{\phi(T,{\bf X},{\bf Z})^{2}\}=0$, which forces $\phi(T,{\bf X},{\bf Z})=0$ $P_{0}$–almost surely. Likewise, from (8),
$0\;\leq\;\mathbb{E}_{{\bf X},{\bf Z}}\!\Big\{\int_{0}^{\tau}\phi(t,{\bf X},{\bf Z})^{2}\,dt\Big\}\;\leq\;c_{4}\,\mathbb{E}\!\big\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\big\}=0\,$,
so
$\mathbb{E}_{{\bf X},{\bf Z}}\!\Big\{\int_{0}^{\tau}\phi(t,{\bf X},{\bf Z})^{2}\,dt\Big\}=0\,.$
Since the integrand $\phi(t,{\bf X},{\bf Z})^{2}$ is nonnegative, this implies
$\int_{0}^{\tau}\phi(t,{\bf X},{\bf Z})^{2}\,dt=0\quad\text{for $P_{{\bf X},{\bf Z}}$–a.e. }({\bf X},{\bf Z}),$
and hence $\phi(t,{\bf X},{\bf Z})=0$ for Lebesgue–a.e. $t\in[0,\tau]$ and $P_{{\bf X},{\bf Z}}$–a.e. $({\bf X},{\bf Z})$. Equivalently,
$\chi_{\eta_{1}}(t,{\bf X},{\bf Z})=\chi_{\eta_{2}}(t,{\bf X},{\bf Z})\quad\text{for }dt\otimes P_{{\bf X},{\bf Z}}\text{–almost all }(t,{\bf X},{\bf Z})\,.$
See 3
Proof of Lemma 3.
Recall that $\ell_{0}(\eta)$ denotes the population (expected) log-likelihood,
$\ell_{0}(\eta)=\mathbb{E}\Big\{\,\Delta\,\chi_{\eta}(T,{\bf X},{\bf Z})-\int_{0}^{\tau}Y(t)\,h_{\eta}(t,{\bf X},{\bf Z})\,dt\Big\},$
where $\chi_{\eta}$ is the linear predictor and $h_{\eta}(t,{\bf X},{\bf Z})$ is the corresponding instantaneous hazard process. Let $\eta_{0}=(g_{o},\theta_{o})$ denote the true parameter, with true hazard $h_{o}(t,{\bf X},{\bf Z})=h_{\eta_{0}}(t,{\bf X},{\bf Z})$.
We study the behavior of $\ell_{0}$ along a one–dimensional path through $\eta_{0}$. Fix an arbitrary direction $w=(w_{g},w_{\theta})$ and define
$\eta_{\varepsilon}=\eta_{0}+\varepsilon w\,,\qquad\varepsilon\in\mathbb{R}\,.$
Assume that $\varepsilon$ is restricted to a small neighbourhood of 0 so that $\eta_{\varepsilon}\in\mathcal{\eta}\in\Theta\times\mathcal{H}(q,\boldsymbol{\alpha},{\bf d},\tilde{{\bf d}},M)$. Because $\chi_{\eta}$ is linear in $(g,\mbox{${\theta}$})$, the Gateaux derivative of $\chi_{\eta}$ in direction $w$ at $\eta_{0}$ is simply
$\dot{\chi}_{h}(t,{\bf X},{\bf Z})=\frac{d}{d\varepsilon}\,\chi_{\eta_{\varepsilon}}(t,{\bf X},{\bf Z})\Big|_{\varepsilon=0}\,$,
and we can write
$\chi_{\eta_{\varepsilon}}(t,{\bf X},{\bf Z})=\chi_{\eta_{0}}(t,{\bf X},{\bf Z})+\varepsilon\,\dot{\chi}_{w}(t,{\bf X},{\bf Z})\,.$
For notational simplicity, when we fix a particular $\eta$, we also write
$q(t,{\bf X},{\bf Z})=\chi_{\eta}(t,{\bf X},{\bf Z})-\chi_{\eta_{o}}(t,{\bf X},{\bf Z})\,$,
so $q$ is the difference in linear predictors between $\eta$ and $\eta_{o}$. Along the path $\eta_{\varepsilon}$, $q$ is proportional to the direction $\dot{\chi}_{w}$.
Next, note that
$h_{\eta_{\varepsilon}}(t,{\bf X},{\bf Z})=h_{o}(t,{\bf X},{\bf Z})\,\exp\{\chi_{\eta_{\varepsilon}}(t,{\bf X},{\bf Z})-\chi_{\eta_{0}}(t,{\bf X},{\bf Z})\}=h_{o}(t,{\bf X},{\bf Z})\,\exp\{\varepsilon\,\dot{\chi}_{w}(t,{\bf X},{\bf Z})\}\,$,
so at $\varepsilon=0$ we have $h_{\eta_{0}}=h_{o}$ and
$\frac{d}{d\varepsilon}h_{\eta_{\varepsilon}}(t,{\bf X},{\bf Z})\Big|_{\varepsilon=0}=h_{o}(t,{\bf X},{\bf Z})\,\dot{\chi}_{w}(t,{\bf X},{\bf Z})\,$,
$\frac{d^{2}}{d\varepsilon^{2}}h_{\eta_{\varepsilon}}(t,{\bf X},{\bf Z})\Big|_{\varepsilon=0}=h_{o}(t,{\bf X},{\bf Z})\,\dot{\chi}_{w}(t,{\bf X},{\bf Z})^{2}\,.$
By the definition of the Gateaux derivative,
$\ell_{0}^{\prime}(\eta_{o})[w]=\frac{d}{d\varepsilon}\,\ell_{0}(\eta_{\varepsilon})\Big|_{\varepsilon=0}\,.$
Using the explicit form of $\ell_{0}$ and differentiating under the expectation,
$\ell_{0}^{\prime}(\eta_{0})[w]$ $=\frac{d}{d\varepsilon}\mathbb{E}\Big\{\Delta\,\chi_{\eta_{\varepsilon}}(T,{\bf X},{\bf Z})-\int_{0}^{\tau}Y(t)\,h_{\eta_{\varepsilon}}(t,{\bf X},{\bf Z})\,dt\Big\}\Big|_{\varepsilon=0}$
$=\mathbb{E}\Big\{\Delta\,\dot{\chi}_{w}(T,{\bf X},{\bf Z})-\int_{0}^{\tau}Y(t)\,\frac{d}{d\varepsilon}h_{\eta_{\varepsilon}}(t,{\bf X},{\bf Z})\Big|_{\varepsilon=0}\,dt\Big\}$
$=\mathbb{E}\Big\{\Delta\,\dot{\chi}_{w}(T,{\bf X},{\bf Z})-\int_{0}^{\tau}Y(t)\,h_{o}(t,{\bf X},{\bf Z})\,\dot{\chi}_{w}(t,{\bf X},{\bf Z})\,dt\Big\}\,.$
Since
$\Delta\,\dot{\chi}_{w}(T,{\bf X},{\bf Z})=\int_{0}^{\tau}\dot{\chi}_{w}(t,{\bf X},{\bf Z})\,dN(t)\,$,
the Doob–Meyer decomposition for the true model (1) gives
$dN(t)=Y(t)h_{o}(t,{\bf X},{\bf Z})\,dt+dM(t)\,$,
where $M(t)$ is a martingale. Therefore,
$\Delta\,\dot{\chi}_{w}(T,{\bf X},{\bf Z})$ $=\int_{0}^{\tau}\dot{\chi}_{w}(t,{\bf X},{\bf Z})\,Y(t)h_{o}(t,{\bf X},{\bf Z})\,dt+\int_{0}^{\tau}\dot{\chi}_{w}(t,{\bf X},{\bf Z})\,dM(t)\,.$
Substituting this into the expression for $\ell_{0}^{\prime}(\eta_{o})[w]$ and cancelling the compensator term,
$\ell_{0}^{\prime}(\eta_{0})[w]$ $=\mathbb{E}\Big\{\int_{0}^{\tau}\dot{\chi}_{h}(t,{\bf X},{\bf Z})\,dM(t)\Big\}\,.$
The stochastic integral $\int_{0}^{\tau}\dot{\chi}_{w}(t,{\bf X},{\bf Z})\,dM(t)$ is a martingale with mean zero. Hence,
$\ell_{0}^{\prime}(\eta_{0})[w]=0\qquad\text{for every direction }w\,.$
In particular, when we fix $\eta$ and set $q(t,{\bf X},{\bf Z})=\chi_{\eta}(t,{\bf X},{\bf Z})-\chi_{\eta_{0}}(t,{\bf X},{\bf Z})$, we can summarize this step as
$\ell_{0}^{\prime}(\eta_{0})[w]=\mathbb{E}\Big\{\Delta\,q(T,{\bf X},{\bf Z})-\int_{0}^{\tau}Y(t)h_{o}(t,{\bf X},{\bf Z})\,q(t,{\bf X},{\bf Z})\,dt\Big\}=0.$
To control $\ell_{0}(\eta)-\ell_{0}(\eta_{0})$, we apply a one–dimensional Taylor expansion of $\ell_{0}(\eta_{\varepsilon})$ around $\varepsilon=0$. Write $\eta=\eta_{1}$ and define the line segment $\eta_{v}=\eta_{0}+v(\eta-\eta_{0})$ for $v\in[0,1]$. The mean value theorem yields
$\ell_{0}(\eta)=\ell_{0}(\eta_{0})+\ell_{0}^{\prime}(\eta_{0})[w]+\frac{1}{2}\,\ell_{0}^{\prime\prime}(\eta_{\bar{v}})[w,w]$
for some $\bar{v}\in(0,1)$, where $w=\eta-\eta_{0}$. Since we have already shown that $\ell_{0}^{\prime}(\eta_{0})[w]=0$, it remains to analyse the second derivative term.
For a fixed $v\in[0,1]$, consider the path $\varepsilon\mapsto\eta_{v}+\varepsilon w$ and write
$q_{v}(t,{\bf X},{\bf Z})=\frac{d}{d\varepsilon}\,\chi_{\eta_{v}+\varepsilon w}(t,{\bf X},{\bf Z})\Big|_{\varepsilon=0}\,$,
which is again linear in $w$. Repeating the differentiation steps above but now starting from $\eta_{v}$ instead of $\eta_{0}$ gives
$\ell_{0}^{\prime\prime}(\eta_{v})[w,w]$ $=\frac{d^{2}}{d\varepsilon^{2}}\,\ell_{0}(\eta_{v}+\varepsilon w)\Big|_{\varepsilon=0}$
$=\mathbb{E}\Big\{\Delta\,\frac{d^{2}}{d\varepsilon^{2}}\chi_{\eta_{v}+\varepsilon h}(T,{\bf X},{\bf Z})\Big|_{\varepsilon=0}-\int_{0}^{\tau}Y(t)\,\frac{d^{2}}{d\varepsilon^{2}}h_{\eta_{v}+\varepsilon w}(t,{\bf X},{\bf Z})\Big|_{\varepsilon=0}\,dt\Big\}\,.$
Since $\chi_{\eta}$ is linear in $\eta$, its second derivative in $\varepsilon$ vanishes, while the second derivative of the hazard is
$\frac{d^{2}}{d\varepsilon^{2}}h_{\eta_{v}+\varepsilon w}(t,{\bf X},{\bf Z})\Big|_{\varepsilon=0}=h_{\eta_{v}}(t,{\bf X},{\bf Z})\,q_{v}(t,{\bf X},{\bf Z})^{2}\,.$
Hence
$\ell_{0}^{\prime\prime}(\eta_{v})[w,w]=-\mathbb{E}\Big\{\int_{0}^{\tau}Y(t)\,q_{v}(t,{\bf X},{\bf Z})^{2}\,h_{\eta_{v}}(t,{\bf X},{\bf Z})\,dt\Big\}\,.$
For notational simplicity (since we will ultimately evaluate this at $\varepsilon=1$), we drop the subscript $v$ and simply write $q(t,{\bf X},{\bf Z})=\chi_{\eta}(t,{\bf X},{\bf Z})-\chi_{\eta_{0}}(t,{\bf X},{\bf Z})$ along the line, so that
$\ell_{0}^{\prime\prime}(\eta_{v})[w,w]=-\mathbb{E}\Big\{\int_{0}^{\tau}Y(t)\,q(t,{\bf X},{\bf Z})^{2}\,h_{\eta_{v}}(t,{\bf X},{\bf Z})\,dt\Big\}\,.$
Next, for relating the second derivative to $d(\eta,\eta_{0})^{2}$, we apply Lemma 1 for
$f_{v}(t,{\bf X},{\bf Z})=q(t,{\bf X},{\bf Z})^{2}\,\frac{h_{\eta_{v}}(t,{\bf X},{\bf Z})}{h_{o}(t,{\bf X},{\bf Z})}\,.$
Indeed, Lemma 1 gives
$\mathbb{E}\Big\{\int_{0}^{\tau}Y(t)\,q(t,{\bf X},{\bf Z})^{2}\,h_{\eta_{v}}(t,{\bf X},{\bf Z})\,dt\Big\}\asymp\mathbb{E}\Big\{\Delta\,q(T,{\bf X},{\bf Z})^{2}\,\frac{h_{\eta_{v}}(T,{\bf X},{\bf Z})}{h_{o}(T,{\bf X},{\bf Z})}\Big\}\,.$
Under Assumptions A1A2, the processes $\chi_{\eta}$ are uniformly bounded, and hence
$0<c_{\star}\;\leq\;\frac{h_{\eta_{v}}(t,{\bf X},{\bf Z})}{h_{o}(t,{\bf X},{\bf Z})}\;\leq\;C_{\star}<\infty,\qquad\text{for all }(t,{\bf X},{\bf Z}),\;v\in[0,1]\,.$
Therefore,
$c_{\star}\,\mathbb{E}\big[\Delta\,q(T,{\bf X},{\bf Z})^{2}\big]\;\leq\;\mathbb{E}\Big\{\Delta\,q(T,{\bf X},{\bf Z})^{2}\,\frac{h_{\eta_{v}}(T,{\bf X},{\bf Z})}{h_{o}(T,{\bf X},{\bf Z})}\Big]\;\leq\;C_{\star}\,\mathbb{E}\big[\Delta\,q(T,{\bf X},{\bf Z})^{2}\big\}\,.$
Combining this with the expression for $\ell_{0}^{\prime\prime}(\eta_{v})[w,w]$ yields the two-sided bound
$C_{\star}\,\mathbb{E}\big\{\Delta\,q(T,{\bf X},{\bf Z})^{2}\big\}\;\leq\;\ell_{0}^{\prime\prime}(\eta_{v})[w,w]\;\leq\;-c_{\star}\,\mathbb{E}\big\{\Delta\,q(T,{\bf X},{\bf Z})^{2}\big\}.$
Recall that the distance $d(\eta,\eta_{0})$ is defined by
$d(\eta,\eta_{0})^{2}=\mathbb{E}\Big\{\Delta\big\{\chi_{\eta}(T,{\bf X},{\bf Z})-\chi_{\eta_{0}}(T,{\bf X},{\bf Z})\big\}^{2}\Big\}=\mathbb{E}\big\{\Delta\,q(T,{\bf X},{\bf Z})^{2}\big\}\,.$
Hence
$C_{\star}\,d(\eta,\eta_{0})^{2}\;\leq\;\ell_{0}^{\prime\prime}(\eta_{v})[w,w]\;\leq\;-c_{\star}\,d(\eta,\eta_{0})^{2}\,.$
Therefore, by the Taylor expansion with the mean value theorem,
$\ell_{0}(\eta)-\ell_{0}(\eta_{0})=\frac{1}{2}\,\ell_{0}^{\prime\prime}(\eta_{\bar{v}})[w,w]$
for some $\bar{v}\in(0,1)$, and we have just shown that $\ell_{0}^{\prime\prime}(\eta_{\bar{v}})[w,w]$ is comparable to $d(\eta,\eta_{0})^{2}$. Thus,
$\ell_{0}(\eta)-\ell_{0}(\eta_{0})\;\asymp\;-\,d(\eta,\eta_{0})^{2},$
which completes the proof. ∎
In the following, $\mathbb{P}_{n}$ and $\mathbb{P}$ denote the empirical and probability measures of $(T,\Delta,{\bf X},{\bf Z})$, and $\mathbb{G}_{n}=\sqrt{n}(\mathbb{P}_{n}-\mathbb{P})$ is the empirical process. The symbol $\mathbb{E}^{\ast}$ stands for outer expectation, used to avoid measurability issues in the supremum over an uncountable class of functions.
Lemma 4.
Consider the following classes of functions indexed by the parameter $\eta$:
  • The log-likelihood increment
    $f_{\eta}(T,\mathbf{X},\mathbf{Z},\Delta)=\Delta\,\chi_{\eta}(T,\mathbf{X},\mathbf{Z})-\int_{0}^{T}Y(t)\,\exp\{\chi_{\eta}(t,\mathbf{X},\mathbf{Z})\}\,dt,$
  • The efficient score-type term
    $\psi_{\eta}(T,\mathbf{X},\mathbf{Z},\Delta)=\int_{0}^{T}\bigl\{\mathbf{Z}-g^{\ast}(t,\mathbf{X})\bigr\}\,dM_{\eta}(t),$
    where $g^{\ast}$ denotes the efficient projection defined in Theorem 3.
For $0<\delta\leq 1$, define the localized classes
$\mathcal{F}_{\delta}=\{\,f_{\eta}-f_{\eta_{0}}:d(\eta,\eta_{0})\leq\delta\,\}\,,\qquad\mathcal{G}_{\delta}=\{\,\psi_{\eta}-\psi_{\eta_{0}}:d(\eta,\eta_{0})\leq\delta\,\}\,.$
Under Assumptions A1A3 and A5, there exists a constant $C>0$ such that for every $\delta\in(0,1]$,
$\mathbb{E}^{\ast}\Bigl\{\sup_{\mu\in\mathcal{F}_{\delta}\cup\mathcal{G}_{\delta}}\bigl|\sqrt{n}\,(\mathbb{P}_{n}-\mathbb{P})\,\mu\bigr|\Bigr\}\;\leq\;C\Bigl\{\delta\,\sqrt{s\log(\mathcal{U}/\delta)}+n^{-1/2}\,s\log(\mathcal{U}/\delta)\Bigr\},$
where $s$ is the sparsity of the neural network class, and
$\mathcal{U}=K\prod_{k=0}^{K}(p_{k}+1)\sum_{k=0}^{K}p_{k}p_{k+1}.$
Proof of Lemma 4.
Fix $\eta_{1},\eta_{2}$ and denote $q(t,{\bf X},{\bf Z})=\chi_{\eta_{1}}(t,{\bf X},{\bf Z})-\chi_{\eta_{2}}(t,{\bf X},{\bf Z})$. For notational convenience, we write $\chi_{\eta}(t)=\chi_{\eta}(t,{\bf X},{\bf Z})$ and $q(t)=\chi_{\eta_{1}}(t)-\chi_{\eta_{2}}(t)$. Because $\chi_{\eta}$ is uniformly bounded under A1–A2 and depends on the network parameters in a Lipschitz way, we have
$\bigl|\exp\{\chi_{\eta_{1}}(t,{\bf X},{\bf Z})\}-\exp\{\chi_{\eta_{2}}(t,{\bf X},{\bf Z})\}\bigr|\;\lesssim\;|q(t,{\bf X},{\bf Z})|\,.$
Using the definition of $f_{\eta}$ and the triangle and Cauchy–Schwarz inequalities,
$|f_{\eta_{1}}-f_{\eta_{2}}|\leq\Delta\,|q(T)|+\int_{0}^{\tau}Y(t)\,|\exp\{\chi_{\eta_{1}}(t)\}-\exp\{\chi_{\eta_{2}}(t)\}|\,dt\,.$
Plugging the bound, squaring and applying the inequality $(a+b)^{2}\leq 2a^{2}+2b^{2}$ gives
$(f_{\eta_{1}}-f_{\eta_{2}})^{2}\;\lesssim\;2\Delta\,q(T)^{2}+2\int_{0}^{\tau}Y(t)\,q(t)^{2}\,dt,$
where all quantities are uniform in $({\bf X},{\bf Z})$ by A1A2. By Lemma 1 we get
$E\!\left[\int_{0}^{\tau}Y(t)\,q(t)^{2}\,dt\right]\asymp E\bigl[\Delta\,q(T)^{2}\bigr]\,.$
The right-hand side is exactly the squared metric $d^{2}(\eta_{1},\eta_{2})$. Thus,
$\|f_{\eta_{1}}-f_{\eta_{2}}\|_{L^{2}(P)}\;\lesssim\;d(\eta_{1},\eta_{2})\,.$
Repeating the argument for $\psi_{\eta}$ requires only replacing the exponential term with the stochastic integral. Assumption A5 ensures boundedness of the predictable integrand ${\bf Z}-g^{\ast}(t,{\bf X})$ and square-integrability of the martingale increments; A1A2 control the compensator. The same calculation shows
$\|\psi_{\eta_{1}}-\psi_{\eta_{2}}\|_{L^{2}(P)}\;\lesssim\;d(\eta_{1},\eta_{2})\,.$
Hence both localized classes $\mathcal{F}_{\delta}$ and $\mathcal{G}_{\delta}$ lie in a common $L^{2}(P)$-ball of radius $O(\delta)$ and admit a common envelope of order $O(\delta)$. The parametrization $\eta=(\mbox{${\theta}$},g)$ is $s$-sparse in the neural-network coordinates, and Assumption A4 ensures that the weights are uniformly bounded and that the architecture has fixed depth and layer widths. In particular, the map $\eta\mapsto\chi_{\eta}$ belongs to the class of sparse feedforward neural networks considered in zhong2022deep. Moreover, the maps $\eta\longmapsto f_{\eta}$ and $\eta\longmapsto\psi_{\eta}$ are Lipschitz transformations of $\chi_{\eta}$ in the $L^{2}(P)$-metric, with Lipschitz constant bounded uniformly under Assumptions A1A3 and A5. Since Lipschitz images of a function class cannot increase its covering number by more than a multiplicative constant, the metric entropy of the localized classes $\mathcal{F}_{\delta}$ and $\mathcal{G}_{\delta}$ is therefore controlled by the entropy of the underlying sparse neural-network ball
$\mathcal{H}_{\delta}=\{\chi_{\eta}-\chi_{\eta_{0}}:d(\eta,\eta_{0})\leq\delta\}.$
Lemma 6 of zhong2022deep establishes that for such a sparse network class,
$\log N_{[]}\!\left(\varepsilon,\,\mathcal{H}_{\delta},\,L^{2}(P)\right)\;\lesssim\;s\log\!\left(\frac{\mathcal{U}\delta}{\varepsilon}\right),$
where $\mathcal{U}$ is the architectural complexity factor. Because $\mathcal{F}_{\delta}$ and $\mathcal{G}_{\delta}$ lie in a common $L^{2}(P)$-ball of radius $O(\delta)$ and share a common envelope of order $O(\delta)$, the same entropy bound applies to their union. Consequently,
$\log N_{[]}\!\left(\varepsilon,\,\mathcal{F}_{\delta}\cup\mathcal{G}_{\delta},\,L^{2}(P)\right)\;\lesssim\;s\log\!\left(\frac{\mathcal{U}\delta}{\varepsilon}\right)\,.$
Next, we bound the bracketing entropy integral. Using the entropy estimate derived above, the associated Dudley integral satisfies
$J_{[]}(\delta)=\int_{0}^{\delta}\!\sqrt{\log N_{[]}(\varepsilon)}\;d\varepsilon\;\leq\;C_{2}\,\delta\,\sqrt{s\log\!\left(\frac{\mathcal{U}}{\delta}\right)}\,.$
Finally, applying Lemma 3.4.2 of vdvwellner1996 to the empirical process $\mathbb{G}_{n}=\sqrt{n}(\mathbb{P}_{n}-\mathbb{P})$, with envelope $F_{\delta}=O(\delta)$, gives
$\mathbb{E}^{\ast}\!\left\{\sup_{\mu\in\mathcal{F}_{\delta}\cup\mathcal{G}_{\delta}}|\mathbb{G}_{n}\mu|\right\}\;\leq\;C_{3}\Bigl\{J_{[]}(\delta)+n^{-1/2}\,\|F_{\delta}\|_{L^{2}(P)}\,\log N_{[]}(\delta)\Bigr\}\,.$
Substituting the bound on $J_{[]}(\delta)$ and noting that $\|F_{\delta}\|_{L^{2}(P)}=O(\delta)$ and $\log N_{[]}(\delta)\lesssim s\log(\mathcal{U}/\delta)$ yields the asserted maximal inequality. ∎
The following lemma establishes equivalence of expectations under two models whose hazard functions differ only through the nuisance function $g$. Because the hazards are uniformly bounded above and below, the distributions under $g_{0}$ and $g_{1}$ are mutually absolutely continuous with Radon–Nikodym derivatives uniformly bounded. This equivalence allows us to control expectations under one model using expectations under another.
Lemma 5.
Suppose Assumptions A1A2 and A4 hold. Let $g_{0},g_{1}\in\mathcal{H}(q,\boldsymbol{\alpha},{\bf d},\tilde{{\bf d}},M)$ and fix $\mbox{${\theta}$}\in\Theta$. For $j=0,1$, let $P_{g_{j}}$ denote the distribution of $(T,{\bf X},{\bf Z},\Delta)$ under event hazard $h_{g_{j}}(t\mid{\bf X},{\bf Z})=\exp\{g_{j}(t,{\bf X})+\mbox{${\theta}$}^{\top}{\bf Z}\}$ and a common censoring hazard $\mu(t\mid{\bf X},{\bf Z})$. As a consequence of Assumptions A1A2, there exists a constant $0<D<\infty$ such that, almost surely,
$e^{-D}\;\leq\;h_{g_{j}}(t\mid\boldsymbol{X},\boldsymbol{Z})\;\leq\;e^{D}\,,\qquad\mbox{for all}\,\,t\in[0,\tau]\,,\ j=0,1\,.$
Define $c=e^{-D}$ and $C=e^{D}$. Then, for any nonnegative measurable function $q:[0,\tau]\times\mathcal{X}\to[0,\infty)$ with $\mathbb{E}_{P_{g_{0}}}[q(T,{\bf X})]<\infty$,
$K_{-}\,\mathbb{E}_{P_{g_{0}}}\!\big\{q(T,{\bf X})\big\}\;\leq\;\mathbb{E}_{P_{g_{1}}}\!\big\{q(T,{\bf X})\big\}\;\leq\;K_{+}\,\mathbb{E}_{P_{g_{0}}}\!\big\{q(T,{\bf X})\big\}\,$,
where
$K_{-}=\frac{c}{C}\,e^{-\tau(C-c)}\,,\qquad K_{+}=\frac{C}{c}\,e^{\,\tau(C-c)}\,.$
In particular, $\mathbb{E}_{P_{g_{1}}}[q(T,{\bf X})]\asymp\mathbb{E}_{P_{g_{0}}}[q(T,{\bf X})]$ with constants depending only on $(c,C,\tau)$.
Proof of Lemma 5.
We work conditional on the covariates $({\bf x},{\bf z})$ and then integrate over their distribution. Under model $g_{j}$, the event time $U_{j}$ has density and survival
$f^{U}_{j}(t\mid{\bf x},{\bf z})=h_{j}(t\mid{\bf x},{\bf z})\,S^{U}_{j}(t\mid{\bf x},{\bf z}),\qquad S^{U}_{j}(t\mid{\bf x},{\bf z})=\exp\!\Big\{-\!\int_{0}^{t}h_{j}(s\mid{\bf x},{\bf z})\,ds\Big\}.$
The censoring time $C$ has density $f^{C}(t\mid{\bf x},{\bf z})=\mu(t\mid{\bf x},{\bf z})S^{C}(t\mid{\bf x},{\bf z})$ and survival $S^{C}(t\mid{\bf x},{\bf z})$, common to both models. The observed time and event indicator are $T_{j}=\min(U_{j},C)$ and $\Delta_{j}=\mathbf{1}\{U_{j}\leq C\}$. Given $({\bf x},{\bf z})$, their joint density under $g_{j}$ is
$f_{j}(t,\delta\mid{\bf x},{\bf z})=\begin{cases}h_{j}(t\mid{\bf x},{\bf z})\,S^{U}_{j}(t\mid{\bf x},{\bf z})\,S^{C}(t\mid{\bf x},{\bf z})\,,&\delta=1\\[3.99994pt] \mu(t\mid{\bf x},{\bf z})\,S^{C}(t\mid{\bf x},{\bf z})\,S^{U}_{j}(t\mid{\bf x},{\bf z})\,,&\delta=0\,.\end{cases}$
By Assumptions A1A2,
$ct\leq\int_{0}^{t}h_{j}(s\mid{\bf x},{\bf z})\,ds\leq Ct,\qquad e^{-Ct}\leq S^{U}_{j}(t\mid{\bf x},{\bf z})\leq e^{-ct}.$
The likelihood ratio of $(T_{j},\Delta_{j})$ under $g_{1}$ versus $g_{0}$, conditional on $({\bf x},{\bf z})$, is
$L(t,\delta\mid{\bf x},{\bf z})=\frac{f_{1}(t,\delta\mid{\bf x},{\bf z})}{f_{0}(t,\delta\mid{\bf x},{\bf z})}=\left\{\frac{h_{1}(t\mid{\bf x},{\bf z})}{h_{0}(t\mid{\bf x},{\bf z})}\right\}^{\delta}\frac{S^{U}_{1}(t\mid{\bf x},{\bf z})}{S^{U}_{0}(t\mid{\bf x},{\bf z})}.$
Since $c\leq h_{j}\leq C$,
$\frac{c}{C}\leq\frac{h_{1}(t\mid{\bf x},{\bf z})}{h_{0}(t\mid{\bf x},{\bf z})}\leq\frac{C}{c}.$
Moreover,
$\frac{S^{U}_{1}(t\mid{\bf x},{\bf z})}{S^{U}_{0}(t\mid{\bf x},{\bf z})}=\exp\!\Big[-\!\int_{0}^{t}\{h_{1}(s\mid{\bf x},{\bf z})-h_{0}(s\mid{\bf x},{\bf z})\}\,ds\Big],$
and because $|h_{1}-h_{0}|\leq C-c$,
$e^{-(C-c)t}\leq\frac{S^{U}_{1}(t\mid{\bf x},{\bf z})}{S^{U}_{0}(t\mid{\bf x},{\bf z})}\leq e^{(C-c)t}.$
For $t\leq\tau$ this yields
$\frac{c}{C}\,e^{-(C-c)\tau}\;\leq\;L(t,\delta\mid{\bf x},{\bf z})\;\leq\;\frac{C}{c}\,e^{(C-c)\tau}.$
Thus the Radon–Nikodým derivative satisfies
$K_{-}\;\leq\;\frac{dP_{g_{1}}}{dP_{g_{0}}}(T,\Delta,{\bf X},{\bf Z})\;\leq\;K_{+},\qquad P_{g_{0}}\text{--a.s.},$
where $K_{-}=\frac{c}{C}e^{-\tau(C-c)}$ and $K_{+}=\frac{C}{c}e^{\tau(C-c)}$. For any nonnegative measurable function $q$ with finite expectation under $P_{g_{0}}$,
$\mathbb{E}_{P_{g_{1}}}\{q(T,{\bf X})\}=\mathbb{E}_{P_{g_{0}}}\!\left\{q(T,{\bf X})\,\frac{dP_{g_{1}}}{dP_{g_{0}}}(T,\Delta,{\bf X},{\bf Z})\right\}$
and the bounds on the likelihood ratio give
$K_{-}\,\mathbb{E}_{P_{g_{0}}}\{q(T,{\bf X})\}\;\leq\;\mathbb{E}_{P_{g_{1}}}\{q(T,{\bf X})\}\;\leq\;K_{+}\,\mathbb{E}_{P_{g_{0}}}\{q(T,{\bf X})\}\,.$
Lemma 6.
Assume A2. Let $\eta_{j}=(\mbox{${\theta}$}_{j},g_{j})$, $j=1,2$, be two parameter values in $(\Theta\times\mathcal{H}(q,\boldsymbol{\alpha},{\bf d},\tilde{{\bf d}},M))\cup(\Theta\times\mathcal{G}(K,s,\boldsymbol{p},D))$. Define
$\chi_{\eta_{j}}(t,{\bf X},{\bf Z})=g_{j}(t,{\bf X})+\mbox{${\theta}$}_{j}^{\top}{\bf Z},\qquad j=1,2.$
Assume that $\|{\bf Z}\|\leq B_{\bf Z}$ a.s. for some finite constant $B_{\bf Z}$. Then
$\mathbb{E}\Bigl[\Delta\bigl\{g_{1}(T,{\bf X})-g_{2}(T,{\bf X})\bigr\}^{2}\Bigr]\;\leq\;2\,d(\eta_{1},\eta_{2})^{2}\;+\;2B_{\bf Z}^{2}\|\mbox{${\theta}$}_{1}-\mbox{${\theta}$}_{2}\|^{2}\,.$
Proof of Lemma 6.
Fix $(T,\Delta,{\bf X},{\bf Z})$ and set $a=g_{1}(T,{\bf X})-g_{2}(T,{\bf X})$, $b=(\mbox{${\theta}$}_{1}-\mbox{${\theta}$}_{2})^{\top}{\bf Z}$. Then $\chi_{\eta_{1}}(T,{\bf X},{\bf Z})-\chi_{\eta_{2}}(T,{\bf X},{\bf Z})=a+b$. Use the elementary inequality
$(u-v)^{2}\;\leq\;2u^{2}+2v^{2}\qquad\text{for all }u,v\in\mathbb{R},$
with $u=a+b$ and $v=b$. This gives
$a^{2}=(a+b-b)^{2}\;\leq\;2(a+b)^{2}+2b^{2}.$
Multiplying by $\Delta\in\{0,1\}$ preserves the inequality
$\Delta\,a^{2}\;\leq\;2\Delta\,(a+b)^{2}+2\Delta\,b^{2}\,.$
Taking expectations and using the definition of $d(\eta_{1},\eta_{2})$,
$\mathbb{E}\Bigl[\Delta\{g_{1}(T,{\bf X})-g_{2}(T,{\bf X})\}^{2}\Bigr]$ $=\mathbb{E}(\Delta a^{2})$
$\leq 2\,\mathbb{E}\{\Delta(a+b)^{2}\}+2\,\mathbb{E}(\Delta b^{2})$
$=2\,d(\eta_{1},\eta_{2})^{2}+2\,\mathbb{E}(\Delta b^{2}).$
For the second term, by boundedness of $\bf Z$},
$|b|=|(\mbox{${\theta}$}_{1}-\mbox{${\theta}$}_{2})^{\top}{\bf Z}|\leq\|\mbox{${\theta}$}_{1}-\mbox{${\theta}$}_{2}\|\,\|{\bf Z}\|\leq B_{\bf Z}\,\|\mbox{${\theta}$}_{1}-\mbox{${\theta}$}_{2}\|,$
and hence
$b^{2}\;\leq\;B_{\bf Z}^{2}\,\|\mbox{${\theta}$}_{1}-\mbox{${\theta}$}_{2}\|^{2}.$
Since $\Delta\leq 1$,
$\mathbb{E}(\Delta b^{2})\leq\mathbb{E}(b^{2})\leq B_{\bf Z}^{2}\,\|\mbox{${\theta}$}_{1}-\mbox{${\theta}$}_{2}\|^{2}\,.$
Putting everything together,
$\mathbb{E}\Bigl[\Delta\{g_{1}(T,{\bf X})-g_{2}(T,{\bf X})\}^{2}\Bigr]\leq 2\,d(\eta_{1},\eta_{2})^{2}+2B_{\bf Z}^{2}\,\|\mbox{${\theta}$}_{1}-\mbox{${\theta}$}_{2}\|^{2}\,.$
Lemma 7.
If Assumptions A1A5 hold, then there exists a constant $c_{\theta}>0$ depending only on the constants in A1A4 such that, for all parameters in a neighborhood of $\eta_{o}$,
$d(\eta_{1},\eta_{2})^{2}\;\geq\;c_{\theta}\,\|\mbox{${\theta}$}_{1}-\mbox{${\theta}$}_{2}\|^{2}\,.$
Proof of Lemma 7.
Let $\mbox{${\beta}$}=\mbox{${\theta}$}_{1}-\mbox{${\theta}$}_{2}$, $a(t,{\bf X})=g_{1}(t,{\bf X})-g_{2}(t,{\bf X})$, $b=\mbox{${\beta}$}^{\top}{\bf Z}$. At a failure time, $\chi_{\eta_{1}}-\chi_{\eta_{2}}=a+b$. Thus
$d(\eta_{1},\eta_{2})^{2}=\mathbb{E}_{0}\{\Delta(a+b)^{2}\}\,.$
To find the smallest possible value of this expression over all possible nuisance differences $g_{1}-g_{2}$, minimize $(a+b)^{2}$ over scalar functions $a(T,{\bf X})$. The minimizing choice is
$a^{\ast}(T,{\bf X})=-\mathbb{E}_{0}\{b\mid T,{\bf X},\Delta=1\}\,.$
Substituting this optimal choice yields
$\inf_{g_{1}-g_{2}}d(\eta_{1},\eta_{2})^{2}=\mathbb{E}_{0}\!\Big\{\Delta\,Var(b\mid T,{\bf X},\Delta=1)\Big\}\,.$
This is exactly the standard least–squares argument. Assumption A4 ensures
$\Pr(\Delta=1\mid T,{\bf X})\;\geq\;c_{\Delta}\quad\text{a.s.}$
Therefore
$\mathbb{E}_{0}\Big[\Pr(\Delta=1\mid T,{\bf X})\,Var(b\mid T,{\bf X},\Delta=1)\Big]\;\geq\;c_{\Delta}\,\mathbb{E}_{0}\bigl[Var(b\mid T,{\bf X},\Delta=1)\bigr]\,.$
Assumption A5 implies that for some $c_{Z}>0$,
$Var({\bf Z}\mid T,{\bf X},\Delta=1)\succeq c_{Z}I_{p}\,$,
thus for any $\beta$},
$Var(b\mid T,{\bf X},\Delta=1)=Var(\mbox{${\beta}$}^{\top}{\bf Z}\mid T,{\bf X},\Delta=1)\geq c_{\Delta}c_{Z}\|\mbox{${\beta}$}\|^{2}\,.$
Combining the above, we get
$d(\eta_{1},\eta_{2})^{2}\geq c_{Z}\|\mbox{${\theta}$}_{1}-\mbox{${\theta}$}_{2}\|^{2}.$
Setting $c_{\theta}=c_{\Delta}c_{Z}>0$ completes the proof. ∎
See 1
Proof of Theorem 1.
Let $\ell(\eta;{\bf W})$, ${\bf W}=(T,\Delta,{\bf X}^{\top},{\bf Z}^{\top})^{\top}$, denote the per–observation log–likelihood contribution, so that $\ell_{n}(\eta)=\mathbb{P}_{n}\ell(\eta;\cdot)$ is the empirical loss and $\ell_{0}(\eta)=\mathbb{P}_{0}\ell(\eta;\cdot)$ is the population loss. Let $\mathbb{G}_{n}=\sqrt{n}(\mathbb{P}_{n}-\mathbb{P}_{0})$ be the empirical process. Recall the distance
$d(\eta_{1},\eta_{2})\;=\;\Big(\mathbb{E}_{0}\big[\Delta\{\chi_{\eta_{1}}(T,{\bf X},{\bf Z})-\chi_{\eta_{2}}(T,{\bf X},{\bf Z})\}^{2}\big]\Big)^{1/2},\qquad\eta=(\mbox{${\theta}$},g),$
and let $\eta_{o}=(\mbox{${\theta}$}_{o},g_{o})$ denote the true parameter. By Lemma 4, there exists a constant $C>0$ such that for every $0<\delta\leq 1$,
$\mathbb{E}^{*}\!\sup_{d(\eta,\eta_{o})\leq\delta}\bigl||\mathbb{G}_{n}\big\{\ell(\eta;\cdot)-\ell(\eta_{o};\cdot)\big\}\bigr|\;\leq\;\Psi_{n}(\delta)\equiv\delta\,\sqrt{s\log\frac{\mathcal{U}}{\delta}}+\frac{s}{\sqrt{n}}\log\frac{\mathcal{U}}{\delta}\,$, (12)
and by Markov’s inequality, this implies
$\sup_{d(\eta,\eta_{o})\leq\delta}\bigl|\mathbb{G}_{n}\big\{\ell(\eta;\cdot)-\ell(\eta_{o};\cdot)\big\}\bigr|=O_{p}\bigl(\Psi_{n}(\delta)\bigr)\,.$
Set $\tau_{n}=\gamma_{n}\log^{2}n$ as in the theorem. By Assumption A3 and the specific choice of $(K,s,\boldsymbol{p})$ (cf. the construction in schmidthieber2020nonparametric and zhong2022deep), we have
$s\log\!\Bigl(\frac{\mathcal{U}}{\tau_{n}}\Bigr)=O\!\bigl(n\,\gamma_{n}^{2}\log n\bigr)\,$,
so that
[…]
for $n$ large enough. Hence, there exists a constant $c>0$ such that
$\Psi_{n}(\tau_{n})\;\leq\;\frac{c}{2}\,\sqrt{n}\,\tau_{n}^{2}\quad\mbox{and}\quad\frac{1}{\sqrt{n}}\Psi_{n}(\tau_{n})\;\leq\;\frac{c}{2}\,\tau_{n}^{2}\,.$ (13)
Consequently,
$\sup_{d(\eta,\eta_{o})\leq\tau_{n}}\bigl|(\ell_{n}-\ell_{0})(\eta)-(\ell_{n}-\ell_{0})(\eta_{o})\bigr|=\sup_{d(\eta,\eta_{o})\leq\tau_{n}}\frac{1}{\sqrt{n}}\bigl|\mathbb{G}_{n}\big\{\ell(\eta;\cdot)-\ell(\eta_{o};\cdot)\big\}\bigr|=O_{p}\bigl(\tfrac{1}{\sqrt{n}}\Psi_{n}(\tau_{n})\bigr)=O_{p}(\tau_{n}^{2}),$ (14)
where we used (12) and (13).
By Theorem 1 of schmidthieber2020nonparametric, applied to the composite Hölder class $\mathcal{H}(q,\boldsymbol{\alpha},{\bf d},\tilde{\bf d},M)$ with the choice of $(K,s,\boldsymbol{p},D)$ given in Assumption A3, there exists $\tilde{g}_{1}\in\mathcal{G}(K,s,\boldsymbol{p},D)$ such that
$\mathbb{E}\bigl\{\tilde{g}_{1}(T,{\bf X})-g_{o}(T,{\bf X})\bigr\}^{2}\leq C_{1}\,\tau_{n}^{2}\,,\qquad\mathbb{E}\bigl\{\Delta\big(\tilde{g}_{1}(T,{\bf X})-g_{o}(T,{\bf X})\big)^{2}\bigr\}\leq C_{1}\,\tau_{n}^{2}$
for some constant $C_{1}>0$ independent of $n$. Define $\tilde{\eta}_{1}=(\mbox{${\theta}$}_{o},\tilde{g}_{1})$. Then,
$d^{2}(\tilde{\eta}_{1},\eta_{o})=\mathbb{E}\bigl[\Delta\big\{\chi_{\tilde{\eta}_{1}}(T,{\bf X},{\bf Z})-\chi_{\eta_{o}}(T,{\bf X},{\bf Z})\big\}^{2}\bigr]=\mathbb{E}\bigl[\Delta\big\{\tilde{g}_{1}(T,{\bf X})-g_{o}(T,{\bf X})\big\}^{2}\bigr]\leq C_{1}\,\tau_{n}^{2}\,$,
since the $\theta$}–component is the same in both $\tilde{\eta}_{1}$ and $\eta_{o}$. Therefore,
$d(\tilde{\eta}_{1},\eta_{o})\;\leq\;\sqrt{C_{1}}\,\tau_{n}\,$, (15)
and in particular $\tilde{\eta}_{1}$ lies inside a ball of radius proportional to $\tau_{n}$ around $\eta_{o}$.
Lemma 3 states that, for any $\eta\in\mathcal{F}_{D}$,
$\ell_{0}(\eta)-\ell_{0}(\eta_{o})\;\asymp\;-\,d(\eta,\eta_{o})^{2}\,$,
that is, there exist constants $0<c_{-},c_{+}<\infty$ such that
$c_{+}\,d(\eta,\eta_{o})^{2}\;\leq\;\ell_{0}(\eta)-\ell_{0}(\eta_{o})\;\leq\;-\,c_{-}\,d(\eta,\eta_{o})^{2}\,.$
Applying this to $\eta=\tilde{\eta}_{1}$ and using (15), we obtain
$\ell_{0}(\tilde{\eta}_{1})-\ell_{0}(\eta_{o})\;\geq\;-\,c_{+}\,d(\tilde{\eta}_{1},\eta_{o})^{2}\;\geq\;-\,c_{+}C_{1}\,\tau_{n}^{2}\;=\;-\,C_{2}\,\tau_{n}^{2}\,$, (16)
for some constant $C_{2}>0$.
By combining (14) and (16) we get
$\ell_{n}(\tilde{\eta}_{1})-\ell_{n}(\eta_{o})$ $=\bigl\{\ell_{n}(\tilde{\eta}_{1})-\ell_{0}(\tilde{\eta}_{1})\bigr\}-\bigl\{\ell_{n}(\eta_{o})-\ell_{0}(\eta_{o})\bigr\}+\bigl\{\ell_{0}(\tilde{\eta}_{1})-\ell_{0}(\eta_{o})\bigr\}$
$\geq-\,\sup_{d(\eta,\eta_{o})\leq\tau_{n}}\bigl|(\ell_{n}-\ell_{0})(\eta)-(\ell_{n}-\ell_{0})(\eta_{o})\bigr|\;-\;C_{2}\,\tau_{n}^{2}$
$=-\,O_{p}(\tau_{n}^{2}).$ (17)
[…]
$\ell_{n}(\widehat{\eta})=\max_{\eta}\ell_{n}(\eta)\;\geq\;\ell_{n}(\tilde{\eta}_{1})\;\geq\;\ell_{n}(\eta_{o})-O_{p}(\tau_{n}^{2})\,.$ (18)
On the other hand, by Lemma 3, the population loss satisfies
$\ell_{0}(\eta)-\ell_{0}(\eta_{o})\;\leq\;-\,c_{-}\,d(\eta,\eta_{o})^{2}\quad\text{for all }\eta\in\mathcal{F}_{D}\,.$
The empirical-process bound (14) and the likelihood chain (18) verify the conditions (i)–(ii) of Theorem 3.4.1 in vdvwellner1996 with radius of order $\tau_{n}$. Therefore,
$d(\widehat{\eta},\eta_{o})=O_{p}(\tau_{n})\,.$
Finally, by writing
$\chi_{\widehat{\eta}}(t,{\bf X},{\bf Z})-\chi_{\eta_{o}}(t,{\bf X},{\bf Z})=\{\widehat{g}(t,{\bf X})-g_{o}(t,{\bf X})\}+(\widehat{}\mbox{${\theta}$}-\mbox{${\theta}$}_{o})^{\top}{\bf Z}\,$,
[…]
$\mathbb{E}\Bigl[\Delta\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\}^{2}\Bigr]\;\leq\;C\Bigl(d(\widehat{\eta},\eta_{o})^{2}+\|\widehat{}\mbox{${\theta}$}-\mbox{${\theta}$}_{o}\|^{2}\Bigr)\,.$ (19)
Applying Lemma 7 with $\eta_{1}=\widehat{\eta}$ and $\eta_{2}=\eta_{o}$ yields
$\|\widehat{}\mbox{${\theta}$}-\mbox{${\theta}$}_{o}\|^{2}\leq c_{\theta}^{-1}d(\widehat{\eta},\eta_{o})^{2}=O_{p}(\tau_{n}^{2})\,.$
Hence, $\|\widehat{}\mbox{${\theta}$}-\mbox{${\theta}$}_{o}\|=O_{p}(\tau_{n})$, and the right-hand side of (19) is $O_{p}(\tau_{n}^{2})$, which is exactly the claim of Theorem 1. The expectation is taken with respect to a new observation $(T,\Delta,{\bf X},{\bf Z})\sim P_{0}$, while the $O_{p}(\cdot)$ is with respect to the randomness in the training sample used to construct $\widehat{g}$. ∎
See 2
Proof of Theorem 2.
For $g\in\mathcal{H}(q,\boldsymbol{\alpha},{\bf d},\tilde{{\bf d}},M)$ and and $\mbox{${\theta}$}_{o}\in\Theta$, let $P_{g,\theta_{o}}$ denote the joint distribution of $(T,\Delta,{\bf X},{\bf Z})$ induced by the hazard $h_{g,\theta_{o}}(t\mid{\bf X},{\bf Z})=\exp\{g(t,{\bf X})+\mbox{${\theta}$}_{o}^{\top}{\bf Z}\}$ together with a censoring hazard that is held fixed across all $g$. By the standard Kullback-Leibler (KL) distance for bounded hazards (Assumptions A1A2), there exists $C>0$ such that for any $g_{0},g_{1}\in\mathcal{\mathcal{}}{H}(q,\boldsymbol{\alpha},{\bf d},\tilde{{\bf d}},M)$,
$\mbox{KL}(P_{g_{1},\theta_{o}}\,\|\,P_{g_{0},\theta_{o}})\;\leq\;C\,n\,\mathbb{E}_{P_{g_{1},\theta_{o}}}\!\left[\Delta\{g_{1}(T,{\bf X})-g_{0}(T,{\bf X})\}^{2}\right]\,.$
Since $\Delta\leq 1$,
$\mbox{KL}(P_{g_{1},\theta_{o}}\,\|\,P_{g_{0},\theta_{o}})\;\leq\;C\,n\,\mathbb{E}_{P_{g_{1},\theta_{o}}}\!\bigl[\{g_{1}(T,{\bf X})-g_{0}(T,{\bf X})\}^{2}\bigr].$
Lemma 5 yields the equivalence
$\mbox{KL}(P_{g_{1},\theta_{o}}\,\|\,P_{g_{0},\theta_{o}})\;\leq\;C^{\prime}\,n\,\mathbb{E}_{P_{g_{0},\theta_{o}}}\!\bigl[\{g_{1}(T,{\bf X})-g_{0}(T,{\bf X})\}^{2}\bigr]$
for a constant $C^{\prime}$ depending only on A1A2.
By the packing construction for composite Hölder networks (schmidthieber2020nonparametric), there exist functions $g^{(0)},\dots,g^{(N)}\in\mathcal{H}(q,\boldsymbol{\alpha},{\bf d},\tilde{{\bf d}},M)$ and constants $c_{1},c_{2}>0$ such that
$\|g^{(j)}-g^{(k)}\|_{L^{2}(P_{g^{(0)},\theta_{o}})}\;\geq\;2c_{1}\,\gamma_{n}\,,\qquad j\neq k\,$, (20)
and
$\frac{1}{N}\sum_{j=1}^{N}\|g^{(j)}-g^{(0)}\|_{L^{2}(P_{g^{(0)},\theta_{o}})}^{2}\;\leq\;\frac{c_{2}\,\log N}{n}\,.$ (21)
Applying the KL bound with $g_{0}=g^{(0)}$ and $g_{1}=g^{(j)}$ and the using (21),
$\frac{1}{N}\sum_{j=1}^{N}\mbox{KL}\!\big(P_{g^{(j)},\theta_{o}}\,\|\,P_{g^{(0)},\theta_{o}}\big)\;\leq\;C^{\prime}\,c_{2}\,\log N\,.$
Choose $\alpha\in(0,1/8)$ so that $C^{\prime}c_{2}\leq\alpha$. Then,
$\frac{1}{N}\sum_{j=1}^{N}\mbox{KL}(P_{g^{(j)},\theta_{o}}\,\|\,P_{g^{(0)},\theta_{o}})\;\leq\;\alpha\log N\,.$
Applying Tsybakov’s Fano inequality to $\{g^{(0)},\dots,g^{(N)}\}$ with separation (20), we obtain that there exists $c>0$ (depending only on $c_{1}$ and $\alpha$) such that for any estimator $\widehat{g}$,
$\sup_{j=0,\dots,N}\ \mathbb{E}_{P_{g^{(j)},\theta_{o}}}\!\Big(\|\widehat{g}-g^{(j)}\|_{L^{2}(P_{g^{(0)},\theta_{o}})}^{2}\Big)\;\geq\;c\,\gamma_{n}^{2}\,.$
Since the supremum over $g_{o}\in\mathcal{H}$ dominates this finite set,
$\inf_{\widehat{g}}\ \sup_{g_{o}\in\mathcal{H}}\mathbb{E}_{P_{g_{o},\theta_{o}}}\!\Bigl(\|\widehat{g}-g_{o}\|_{L^{2}(P_{g^{(0)},\theta_{o}})}^{2}\Bigr)\;\geq\;c\,\gamma_{n}^{2}\,.$
By Lemma 5, there exists a constant $C_{\mathrm{eq}}>0$ (depending only on A1A2) such that, uniformly over $g_{o}\in\mathcal{H}$ and all sieve estimators $\widehat{g}$,
$\mathbb{E}_{P_{g^{(0)},\theta_{o}}}\!\big[\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\}^{2}\big]\;\leq\;C_{\mathrm{eq}}\,\mathbb{E}_{P_{g_{o},\theta_{o}}}\!\big[\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\}^{2}\big].$
Combining this with the previous display yields
$\inf_{\widehat{g}}\ \sup_{g_{o}\in\mathcal{H}}\mathbb{E}_{P_{g_{o},\theta_{o}}}\!\big[\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\}^{2}\big]\;\geq\;\frac{c}{C_{\mathrm{eq}}}\,\gamma_{n}^{2}.$
Finally, since the supremum in the theorem is also taken over $\mbox{${\theta}$}_{o}\in\Theta$, and the above bound holds for each fixed $\theta_{o}$, we obtain
$\inf_{\widehat{g}}\ \sup_{\theta_{o}\in\Theta,\ g_{o}\in\mathcal{H}}\mathbb{E}_{P_{g_{o},\theta_{o}}}\!\big[\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\}^{2}\big]\;\geq\;c_{1}\,\gamma_{n}^{2}$ (22)
for some $c_{1}>0$. By Lemma 2 and Assumptions A1, A2, A4, there exists a constant $C_{\Delta}>0$ such that, for all $g_{o}\in\mathcal{H}$ and all $\widehat{g}$,
$\mathbb{E}_{P_{g_{o},\theta_{o}}}\!\big[\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\}^{2}\big]\;\leq\;C_{\Delta}\,\mathbb{E}_{P_{g_{o},\theta_{o}}}\!\big[\Delta\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\}^{2}\big]\,.$
Combining this with (22) gives
$\inf_{\widehat{g}}\ \sup_{\theta_{o}\in\Theta,\ g_{o}\in\mathcal{H}}\mathbb{E}_{P_{g_{o},\theta_{o}}}\!\big[\Delta\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\}^{2}\big]\;\geq\;\frac{c_{1}}{C_{\Delta}}\,\gamma_{n}^{2}\,.$
Renaming the constant $c=c_{1}/C_{\Delta}$ completes the proof. ∎
See 3
Proof of Theorem 3.
Let $\mbox{${\theta}$}\in\mathbb{R}^{p}$ and let $\{g_{s}:s\in(-1,1)\}\subset\mathcal{H}_{g_{o}}$ be a one–dimensional regular submodel of $g$, with
$g_{s}=g_{o}+sg,\qquad g=\left.\frac{\partial g_{s}}{\partial s}\right|_{s=0}\in\mathcal{T}_{g_{o}}\,.$
Consider the semiparametric submodel
$\left\{P_{\theta_{o}+s\theta,\;g_{s}}:s\in(-1,1)\right\}\,.$
For a single observation $(T,\Delta,{\bf X},{\bf Z})$ the log-likelihood is
$\ell(\mbox{${\theta}$},g)=\Delta\{g(T,{\bf X})+\mbox{${\theta}$}^{\top}{\bf Z}\}-\int_{0}^{\tau}Y(t)\exp\{g(t,{\bf X})+\mbox{${\theta}$}^{\top}{\bf Z}\}\,dt.$
Differentiating $\ell(\mbox{${\theta}$}_{o}+s\mbox{${\theta}$},g_{s})$ at $s=0$ gives
$\left.\frac{d}{ds}\ell(\mbox{${\theta}$}_{o}+s\mbox{${\theta}$},g_{s})\right|_{s=0}=\mbox{${\theta}$}^{\top}\dot{\ell}_{\theta_{o}}+\dot{\ell}_{g_{o},g},$
where
$\dot{\ell}_{\theta_{o}}=\Delta{\bf Z}-\int_{0}^{\tau}{\bf Z}Y(t)\exp\{g_{o}(t,{\bf X})+\mbox{${\theta}$}_{o}^{\top}{\bf Z}\}\,dt=\int_{0}^{\tau}{\bf Z}\,dM(t),$
and
$\dot{\ell}_{g_{o},g}=\Delta g(T,{\bf X})-\int_{0}^{\tau}g(t,{\bf X})\,Y(t)\exp\{g_{o}(t,{\bf X})+\mbox{${\theta}$}_{o}^{\top}{\bf Z}\}\,dt=\int_{0}^{\tau}g(t,{\bf X})\,dM(t)\,.$
Thus the (naive) score for $\mbox{${\theta}$}_{o}$ is $\dot{\ell}_{\theta_{o}}$. To obtain the efficient score, we project this onto the orthocomplement of the nuisance tangent space
$\mathcal{T}_{g_{o}}^{\mathrm{nuis}}=\left\{\dot{\ell}_{g_{o},g}:g\in\mathcal{T}_{g_{o}}\right\},$
but the projection must be taken onto the closure $\overline{\mathcal{T}_{g_{o}}^{\mathrm{nuis}}}$ in $L^{2}(P_{0})$. Hence the efficient score is
$\ell^{*}_{\theta_{o}}=\dot{\ell}_{\theta_{o}}-\Pi_{g_{o}}(\dot{\ell}_{\theta_{o}})\,$,
where $\Pi_{g_{o}}$ denotes the $L^{2}(P_{0})$–projection onto $\overline{\mathcal{T}_{g_{o}}^{\mathrm{nuis}}}$. The projection $\Pi_{g_{o}}(\dot{\ell}_{\theta_{o}})$ has the representation $\dot{\ell}_{g_{o},{\bf g}^{*}}$ for a pp–vector ${\bf g}^{*}=(g_{1}^{*},\dots,g_{p}^{*})^{\top}\in(\overline{\mathcal{T}_{g_{o}}})^{p}$ satisfying the orthogonality condition
$\mathbb{E}\!\Big\{(\dot{\ell}_{\theta_{o}}-\dot{\ell}_{g_{o},{\bf g}^{*}})\,\dot{\ell}_{g_{o},g}\Big\}={\bf 0}\qquad\text{for all }g\in\mathcal{T}_{g_{o}}\,.$
Using Lemma 1 of sasieni1992information, this orthogonality is equivalent to
$\mathbb{E}\!\left[\Delta\{{\bf Z}-{\bf g}^{*}(T,{\bf X})\}\,g(T,{\bf X})\right]={\bf 0}\qquad\text{for all }g\in\mathcal{T}_{g_{o}}\,.$
Thus, componentwise, $g_{j}^{*}$ is the (unique) $\Delta$–weighted $L^{2}(P_{0})$ projection of $Z_{j}$ onto $\overline{\mathcal{T}_{g_{o}}}$. Equivalently, $\bf g$$^{*}$ minimizes
$\mathbb{E}\!\left(\Delta\|{\bf Z}-{\bf g}(T,{\bf X})\|_{c}^{2}\right)\,,\qquad{\bf g}\in(\overline{\mathcal{T}_{g_{o}}})^{p}\,.$
Substituting $\bf g$$^{*}$, the efficient score becomes
$\ell^{*}_{\theta_{o}}(T,{\bf X},{\bf Z},\Delta)=\int_{0}^{\tau}\{{\bf Z}-{\bf g}^{*}(t,{\bf X})\}\,dM(t).$
By the martingale isometry (again Lemma 1 of sasieni1992information), the efficient information matrix is
$I(\theta_{o})=\mathbb{E}\left[\Delta\,\{{\bf Z}-{\bf g}^{*}(T,{\bf X})\}^{\otimes 2}\right].$
Finally, under the regularity conditions in Assumptions (A1)–(A4), existence and uniqueness of the weighted projection $\bf g$$^{*}$ in $\overline{\mathcal{T}_{g_{o}}}^{p}$ follow from the approximation theorems of stone1985additive and the general semiparametric projection theory in Appendix A.4 of bickel1993efficient. Therefore, both the efficient score and the efficient information matrix are well-defined and given by the formulas above. ∎
See 4
Proof of Theorem 4.
For any $(\mbox{${\theta}$},g)$, define the per–observation score
$r_{\theta,g}(T,{\bf X},{\bf Z},\Delta)=\Delta\{{\bf Z}-{\bf g}^{*}(T,{\bf X})\}-\int_{0}^{\tau}\{{\bf Z}-{\bf g}^{*}(t,{\bf X})\}Y(t)\exp\{g(t,{\bf X})+\mbox{${\theta}$}^{\top}{\bf Z}\}\,dt\,$,
where $\bf g$$^{*}$ is the pp–dimensional minimizer from Theorem 3, which depends only on $(\mbox{${\theta}$}_{o},g_{o})$. By Theorem 3, at the true parameter $(\mbox{${\theta}$}_{o},g_{o})$ this coincides with the efficient score,
$r_{\theta_{o},g_{o}}(T,{\bf X},{\bf Z},\Delta)=\ell^{*}_{\theta_{o}}(T,{\bf X},{\bf Z},\Delta),\qquad\mathbb{P}\,r_{\theta_{o},g_{o}}={\bf 0},$
and the efficient information matrix is
$I(\mbox{${\theta}$}_{o})=\mathbb{P}\bigl\{r_{\theta_{o},g_{o}}^{\otimes 2}\bigr\}=\mathbb{E}\!\bigl[\Delta\{{\bf Z}-{\bf g}^{*}(T,{\bf X})\}^{\otimes 2}\bigr]\,.$
The empirical efficient score at $(\mbox{${\theta}$},g)$ is $\mathbb{P}_{n}r_{\theta,g}$. By the first–order condition for the (sieve) maximum–likelihood estimator, the estimated parameters satisfy
$\mathbb{P}_{n}r_{\widehat{\theta},\widehat{g}}={\bf 0}\,.$
Write
$\mathbb{P}_{n}\bigl\{r_{\widehat{\theta},\widehat{g}}-r_{\theta_{o},g_{o}}\bigr\}=\mathbb{P}\bigl\{r_{\widehat{\theta},\widehat{g}}-r_{\theta_{o},g_{o}}\bigr\}+(\mathbb{P}_{n}-\mathbb{P})\bigl\{r_{\widehat{\theta},\widehat{g}}-r_{\theta_{o},g_{o}}\bigr\}\,.$
By Theorem 1, we have $d(\widehat{\eta},\eta_{o})=O_{p}(\tau_{n})$, where $\eta=(\mbox{${\theta}$},g)$ and $\eta_{o}=(\mbox{${\theta}$}_{o},g_{o})$. Using the same entropy bound and maximal inequality as in Lemma 4, one shows that the class
$\mathcal{R}_{n}=\bigl\{r_{\theta,g}:d(\eta,\eta_{o})\leq C\tau_{n}\bigr\}$
is stochastically equicontinuous, and in particular
$\sup_{r\in\mathcal{R}_{n}}|(\mathbb{P}_{n}-\mathbb{P})r|=o_{p}(n^{-1/2})\,.$
[…]
$(\mathbb{P}_{n}-\mathbb{P})\bigl\{r_{\widehat{\theta},\widehat{g}}-r_{\theta_{o},g_{o}}\bigr\}=o_{p}(n^{-1/2})\,.$
Using $\mathbb{P}_{n}r_{\widehat{\theta},\widehat{g}}={\bf 0}$, this identity can be written as
$\mathbb{P}\bigl\{r_{\widehat{\theta},\widehat{g}}-r_{\theta_{o},g_{o}}\bigr\}=-\,P_{n}r_{\theta_{o},g_{o}}+o_{p}(n^{-1/2}).$ (23)
Consider the map
$m(\mbox{${\theta}$},g)=\mathbb{P}\,r_{\theta,g}.$
Since $m(\mbox{${\theta}$}_{o},g_{o})={\bf 0}$ and $r_{\theta,g}$ is smooth in $(\mbox{${\theta}$},g)$ in a neighbourhood of $(\mbox{${\theta}$}_{o},g_{o})$, a first–order expansion gives
$m(\widehat{}\mbox{${\theta}$},\widehat{g})-m(\mbox{${\theta}$}_{o},g_{o})=\dot{m}_{\theta}(\mbox{${\theta}$}_{o},g_{o})(\widehat{}\mbox{${\theta}$}-\mbox{${\theta}$}_{o})+\dot{m}_{g}(\mbox{${\theta}$}_{o},g_{o})[\widehat{g}-g_{o}]+R_{n},$
where $\dot{m}_{\theta}(\mbox{${\theta}$}_{o},g_{o})$ is the derivative of $m$ with respect to $\theta$} at $(\mbox{${\theta}$}_{o},g_{o})$, $\dot{m}_{g}(\mbox{${\theta}$}_{o},g_{o})[\cdot]$ is the Gateaux derivative in the gg–direction, and $R_{n}$ is a second–order remainder.
Differentiating $r_{\theta,g_{o}}$ with respect to $\theta$} and using the representation of the efficient score in Theorem 3 together with the martingale isometry, one obtains
$\dot{m}_{\theta}(\mbox{${\theta}$}_{o},g_{o})=\left.\frac{\partial}{\partial\mbox{${\theta}$}}\right|_{\mbox{${\theta}$}=\mbox{${\theta}$}_{o}}\mathbb{P}\,r_{\theta,g_{o}}=-\,I(\mbox{${\theta}$}_{o})\,.$
Hence the contribution of $\widehat{}\mbox{${\theta}$}-\mbox{${\theta}$}_{o}$ in the expansion is $I(\mbox{${\theta}$}_{o})(\widehat{}\mbox{${\theta}$}-\mbox{${\theta}$}_{o})$.
Fix a deterministic direction $g$ and consider $g_{s}=g_{o}+sg$. Differentiating under the integral sign and applying the martingale compensation identity yields
$\left.\frac{d}{ds}\right|_{s=0}m(\mbox{${\theta}$}_{o},g_{s})=-\,\mathbb{E}\bigl[\Delta\{{\bf Z}-{\bf g}^{*}(T,{\bf X})\}g(T,{\bf X})\bigr]\,.$
By Theorem 3, $\bf g$$^{*}$ is the $\Delta$–weighted $L^{2}(P)$–projection of $\bf Z$} onto the nuisance tangent space $(\overline{\mathcal{T}}_{g_{o}})^{p}$, so
$\mathbb{E}\bigl[\Delta\{{\bf Z}-{\bf g}^{*}(T,{\bf X})\}g(T,{\bf X})\bigr]={\bf 0}\quad\text{for all }h\in\overline{\mathcal{T}}_{g_{o}}.$
By Assumptions A1A4, $\widehat{g}-g_{o}\in\overline{\mathcal{T}}_{g_{o}}$, and therefore
$\dot{m}_{g}(\mbox{${\theta}$}_{o},g_{o})[\widehat{g}-g_{o}]=0.$
The map $g\mapsto m(\mbox{${\theta}$}_{o},g)$ is twice Fréchet differentiable, and its second derivative is bounded by a constant multiple of $\mathbb{E}[\Delta\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\}^{2}]$. Consequently,
$|R_{n}|\;\lesssim\;\mathbb{E}\Bigl[\Delta\bigl\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\bigr\}^{2}\Bigr]=O_{p}(\tau_{n}^{2}),$
where Theorem 1 gives $\|\widehat{g}-g_{o}\|_{L^{2}(\Delta)}=O_{p}(\tau_{n})$. Our assumption $\sqrt{n}\,\tau_{n}^{2}\to 0$ implies $\tau_{n}^{2}=o(n^{-1/2})$, so $R_{n}=o_{p}(n^{-1/2})$.
[…]
$\mathbb{P}\bigl\{r_{\widehat{\theta},\widehat{g}}-r_{\theta_{o},g_{o}}\bigr\}=-\,I(_{o})(\widehat{}\mbox{${\theta}$}-\mbox{${\theta}$}_{o})+o_{p}(n^{-1/2})\,.$
Substituting this into (23) yields
$I(\mbox{${\theta}$}_{o})(\widehat{}\mbox{${\theta}$}-\mbox{${\theta}$}_{o})=-\,\mathbb{P}_{n}r_{\theta_{o},g_{o}}+o_{p}(n^{-1/2})\,$,
or, equivalently,
$\widehat{}\mbox{${\theta}$}-\mbox{${\theta}$}_{o}=I(\mbox{${\theta}$}_{o})^{-1}\,\mathbb{P}_{n}r_{\theta_{o},g_{o}}+o_{p}(n^{-1/2})\,.$
Recalling that $r_{\theta_{o},g_{o}}=\ell^{*}_{\theta_{o}}$ from Theorem 3, we obtain the stochastic expansion
$\sqrt{n}(\widehat{}\mbox{${\theta}$}-\mbox{${\theta}$}_{o})=I(\mbox{${\theta}$}_{o})^{-1}\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\ell^{*}_{\theta_{o}}(T_{i},{\bf X}_{i},{\bf Z}_{i},\Delta_{i})+o_{p}(1)\,.$
The efficient score $\ell^{*}_{\theta_{o}}$ has mean zero and covariance $I(\mbox{${\theta}$}_{o})$, again by Theorem 3. By the classical central limit theorem,
$\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\ell^{*}_{\theta_{o}}(T_{i},{\bf X}_{i},{\bf Z}_{i},\Delta_{i})\;\overset{D}{\longrightarrow}\;N\bigl({\bf 0},I(\mbox{${\theta}$}_{o})\bigr)\,.$
Slutsky’s theorem then gives
$\sqrt{n}(\widehat{}\mbox{${\theta}$}-\mbox{${\theta}$}_{o})\;\overset{D}{\longrightarrow}\;N\bigl({\bf 0},I(\mbox{${\theta}$}_{o})^{-1}\bigr)\,$,
which completes the proof. ∎
Lemma 8.
Let $\delta_{n}=\max_{j}(t_{j}-t_{j-1})$ be the mesh size of the time grid. Under Assumptions A1A4, there exist constants $C_{1},C_{2}>0$ such that, for all $n$ and all $\eta$ in the neighborhood $\mathcal{N}_{n}=\{\eta:d(\eta,\eta_{o})\leq M\tau_{n}\}$,
$\sup_{\eta\in\mathcal{N}_{n}}\bigl|\widetilde{\ell}_{n}(\eta)-\ell_{n}(\eta)\bigr|\leq C_{1}\,\delta_{n}$ (24)
and
$\sup_{\eta\in\mathcal{N}_{n}}\Bigl\|\nabla_{\theta}\widetilde{\ell}_{n}(\eta)-\nabla_{\theta}\ell_{n}(\eta)\Bigr\|\leq C_{2}\,\delta_{n}\,$, (25)
where $\nabla_{\theta}\widetilde{\ell}_{n}(\eta)$ and $\nabla_{\theta}\ell_{n}(\eta)$ are the $\theta$}-score functions of $\widetilde{\ell}_{n}(\eta)$ and $\ell_{n}(\eta)$, respectively.
The proof of Lemma 8 follows directly from classical approximation results for Riemann sums. Under Assumptions A1–A4, $g(t,{\bf x})$ is uniformly Hölder continuous in $t$, uniformly over $\bf x$} and over $\eta\in\mathcal{N}_{n}$. Because the hazard is Lipschitz in $g(t,{\bf x})$, the integrand $Y_{i}(t)\exp\{g(t,{\bf X}_{i})+\mbox{${\theta}$}^{\top}{\bf Z}_{i}\}$ is Lipschitz on each interval where $Y_{i}(t)=1$. Therefore the Riemann approximation error is $O(\delta_{n})$. Applying this bound uniformly over all observations and all $\eta\in\mathcal{N}_{n}$ yields inequalities (24) and (25). Because the argument involves no stochastic elements and is fully deterministic, we omit the proof.
Lemma 9.
Let $S_{n}(\eta)=\nabla_{\theta}\ell_{n}(\eta)$, and $\dot{S}_{n}(\eta)=\nabla_{\theta}S_{n}(\eta)$ is the Hessian of the empirical log–likelihood with respect to $\theta$}. For a fixed $M>0$, define the shrinking neighbourhood $\mathcal{N}_{n}=\bigl\{\eta:d(\eta,\eta_{o})\leq M\tau_{n}\bigr\}$, $\tau_{n}=\gamma_{n}\log^{2}n$. Assume A1A4. Then,
$\sup_{\eta\in\mathcal{N}_{n}}\bigl\|\dot{S}_{n}(\eta)+I(\mbox{${\theta}$}_{o})\bigr\|\;\xrightarrow{P}\;0\,$,
where $I(\mbox{${\theta}$}_{o})$ is the efficient information matrix defined in Theorem 3, and here $\|\cdot\|$ may be any fixed matrix norm; for concreteness, and because it provides the appropriate control on the inverse Hessian, we take it to be the operator norm induced by the Euclidean norm.
Proof of Lemma 9.
For a single observation $(T,\Delta,{\bf X},{\bf Z})$, the per–observation log–likelihood is
$\ell(\mbox{${\theta}$},g)=\Delta\{\mbox{${\theta}$}^{\top}{\bf Z}+g(T,{\bf X})\}-\int_{0}^{\tau}Y(t)\exp\{\mbox{${\theta}$}^{\top}{\bf Z}+g(t,{\bf X})\}\,dt\,.$
The corresponding $\theta$}-score is
$s(\mbox{${\theta}$},g)=\nabla_{\theta}\ell(\mbox{${\theta}$},g)=\Delta{\bf Z}-\int_{0}^{\tau}{\bf Z}\,Y(t)\exp\{\mbox{${\theta}$}^{\top}{\bf Z}+g(t,{\bf X})\}\,dt\,$,
and the $\theta$}-Hessian is
$\dot{s}(\mbox{${\theta}$},g)=\nabla_{\theta}s(\mbox{${\theta}$},g)=-\int_{0}^{\tau}{\bf Z}{\bf Z}^{\top}Y(t)\exp\{\mbox{${\theta}$}^{\top}{\bf Z}+g(t,{\bf X})\}\,dt\,.$
Therefore,
$\dot{S}_{n}(\eta)=\mathbb{P}_{n}\dot{s}(\mbox{${\theta}$},g)$
and
$\mathbb{P}\,\dot{s}(\mbox{${\theta}$},g)=E_{0}\!\left[-\int_{0}^{\tau}{\bf Z}{\bf Z}^{\top}Y(t)\exp\{\mbox{${\theta}$}^{\top}{\bf Z}+g(t,{\bf X})\}\,dt\right],$
where $\mathbb{P}_{n}$ is the empirical measure and $\mathbb{P}=P_{0}$ denotes expectation under the true law.
Define the matrix-valued per-observation information integrand by
$J(\eta;{\bf W})=\int_{0}^{\tau}{\bf Z}{\bf Z}^{\top}Y(t)\exp\{\mbox{${\theta}$}^{\top}{\bf Z}+g(t,{\bf X})\}\,dt,$
where ${\bf W}=(T,\Delta,{\bf X},{\bf Z})$ so that $\dot{S}_{n}(\eta)=-\,\mathbb{P}_{n}J(\eta;\cdot)$ and $\mathbb{P}\dot{s}(\mbox{${\theta}$},g)=-\,\mathbb{P}J(\eta;\cdot)$.
By Assumptions A1A4, the covariates are bounded, $\|{\bf Z}\|\leq B_{{\bf Z}}$ a.s., and the hazards are uniformly bounded. Then, there exist $0<c_{h}<C_{h}<\infty$ such that
$c_{h}\;\leq\;\exp\{\mbox{${\theta}$}^{\top}{\bf Z}+g(t,{\bf X})\}\;\leq\;C_{h}$
for all $\eta$ in the sieve and all $(t,{\bf X},{\bf Z})$. Hence the integrand in $J(\eta;{\bf W})$ is uniformly bounded by
$\|{\bf Z}{\bf Z}^{\top}Y(t)\exp\{\mbox{${\theta}$}^{\top}{\bf Z}+g(t,{\bf X})\}\|\;\leq\;B_{{\bf Z}}^{2}C_{h}$
for all $\eta\in\mathcal{N}_{n}$ and all $\bf W$}. Thus the class $\mathcal{J}_{n}=\bigl\{J(\eta;\cdot):\eta\in\mathcal{N}_{n}\bigr\}$ has a finite envelope and inherits the same entropy bounds as the likelihood class used in Lemma 4. In particular, $\mathcal{J}_{n}$ is $\mathbb{P}$–Glivenko–Cantelli, so
$\sup_{\eta\in\mathcal{N}_{n}}\bigl\|\mathbb{P}_{n}J(\eta;\cdot)-\mathbb{P}J(\eta;\cdot)\bigr\|\;\xrightarrow{P}\;0\,.$ (26)
Fix $\eta_{1}=(\mbox{${\theta}$}_{1},g_{1})$ and $\eta_{2}=(\mbox{${\theta}$}_{2},g_{2})$ in $\mathcal{N}_{n}$. For each observation $\bf W$},
$\bigl\|J(\eta_{1};{\bf W})-J(\eta_{2};{\bf W})\bigr\|$ $=\Biggl\|\int_{0}^{\tau}{\bf Z}{\bf Z}^{\top}Y(t)\Bigl[\exp\{\chi_{\eta_{1}}(t,{\bf X},{\bf Z})\}-\exp\{\chi_{\eta_{2}}(t,{\bf X},{\bf Z})\}\Bigr]\,dt\Biggr\|$
$\leq\int_{0}^{\tau}\|{\bf Z}{\bf Z}^{\top}\|\,Y(t)\Bigl|\exp\{\chi_{\eta_{1}}(t,{\bf X},{\bf Z})\}-\exp\{\chi_{\eta_{2}}(t,{\bf X},{\bf Z})\}\Bigr|\,dt\,.$
Using the mean-value theorem for the exponential and boundedness of $\bf Z$} and $g$, there exists a constant $C>0$, depending only on model constants, such that
$\Bigl|\exp\{\mbox{${\theta}$}_{1}^{\top}{\bf Z}+g_{1}(t,{\bf X})\}-\exp\{\mbox{${\theta}$}_{2}^{\top}{\bf Z}+g_{2}(t,{\bf X})\}\Bigr|\;\leq\;C\Bigl\{\|\mbox{${\theta}$}_{1}-\mbox{${\theta}$}_{2}\|+|g_{1}(t,{\bf X})-g_{2}(t,{\bf X})|\Bigr\}\,.$
Therefore,
$\bigl\|J(\eta_{1};{\bf W})-J(\eta_{2};{\bf W})\bigr\|\;\leq\;C^{\prime}\int_{0}^{\tau}Y(t)\Bigl\{\|\mbox{${\theta}$}_{1}-\mbox{${\theta}$}_{2}\|+|g_{1}(t,{\bf X})-g_{2}(t,{\bf X})|\Bigr\}\,dt$
for some $C^{\prime}>0$. Taking expectations and using $\int_{0}^{\tau}Y(t)\,dt\leq\tau$,
$\bigl\|\mathbb{P}J(\eta_{1};\cdot)-\mathbb{P}J(\eta_{2};\cdot)\bigr\|$ $\leq C^{\prime}\tau\|\mbox{${\theta}$}_{1}-\mbox{${\theta}$}_{2}\|+C^{\prime}\,E_{0}\!\left\{\int_{0}^{\tau}Y(t)\,|g_{1}(t,{\bf X})-g_{2}(t,{\bf X})|\,dt\right\}.$
By the Hölder regularity and boundedness assumptions on $g$, together with the fact that $d(\eta_{1},\eta_{2})$ controls the $L^{2}$–distance between $\chi_{\eta_{1}}$ and $\chi_{\eta_{2}}$, one can bound the right-hand side by $C^{\prime\prime}d(\eta_{1},\eta_{2})$ for some $C^{\prime\prime}>0$. In particular, for all $\eta\in\mathcal{N}_{n}$,
$\bigl\|\mathbb{P}J(\eta;\cdot)-\mathbb{P}J(\eta_{o};\cdot)\bigr\|\;\leq\;C^{\prime\prime}d(\eta,\eta_{o})\;\leq\;C^{\prime\prime}M\tau_{n}\;\longrightarrow\;0\,$,
since $\tau_{n}\to 0$. Hence
$\sup_{\eta\in\mathcal{N}_{n}}\bigl\|\mathbb{P}J(\eta;\cdot)-\mathbb{P}J(\eta_{o};\cdot)\bigr\|\;\longrightarrow\;0\,.$ (27)
By Theorem 3, the efficient score for $\mbox{${\theta}$}_{o}$ is
$\ell^{*}_{\theta_{o}}(T,{\bf X},{\bf Z},\Delta)=\int_{0}^{\tau}\{{\bf Z}-{\bf g}^{*}(t,{\bf X})\}\,dM(t)\,$,
and the efficient information matrix is
$I(\mbox{${\theta}$}_{o})=E_{0}\bigl[\Delta\{{\bf Z}-{\bf g}^{*}(T,{\bf X})\}^{\otimes 2}\bigr]\,.$
By semiparametric information theory, the efficient score satisfies the usual information identity along the least favorable submodel. In particular, the efficient Fisher information equals minus the expectation of the derivative of the efficient score,
$\mathbb{P}\dot{s}(\mbox{${\theta}$}_{o},g_{o})=I(\mbox{${\theta}$}_{o})\,.$
Recalling that $\dot{s}(\mbox{${\theta}$}_{o},g_{o})=-J(\eta_{o};{\bf W})$, this gives
$\mathbb{P}J(\eta_{o};\cdot)=I(\mbox{${\theta}$}_{o})\,.$ (28)
Finally, from (26), (27), and (28),
$\sup_{\eta\in\mathcal{N}_{n}}\bigl\|\dot{S}_{n}(\eta)+I(\mbox{${\theta}$}_{o})\bigr\|$ $=\sup_{\eta\in\mathcal{N}_{n}}\bigl\|-\mathbb{P}_{n}J(\eta;\cdot)+I(\mbox{${\theta}$}_{o})\bigr\|$
$\leq\sup_{\eta\in\mathcal{N}_{n}}\bigl\|\mathbb{P}_{n}J(\eta;\cdot)-\mathbb{P}J(\eta;\cdot)\bigr\|+\sup_{\eta\in\mathcal{N}_{n}}\bigl\|\mathbb{P}J(\eta;\cdot)-\mathbb{P}J(\eta_{o};\cdot)\bigr\|$
$\xrightarrow{P}0+0\;=\;0.$
This proves the lemma. ∎
See 5
Proof of Theorem 5.
Recall that $\ell_{n}(\eta)$ and $\widetilde{\ell}_{n}(\eta)$ denote the exact and numerically integrated log-likelihoods, respectively, and $S_{n}(\eta)=\nabla_{\theta}\ell_{n}(\eta)$ and $\widetilde{S}_{n}(\eta)=\nabla_{\theta}\widetilde{\ell}_{n}(\eta)$ their $\theta$}-scores. By Lemma 8, for all $\eta$ in the shrinking neighborhood $\mathcal{N}_{n}=\{\eta:\ d(\eta,\eta_{o})\leq M\tau_{n}\}$ for some fixed $M\geq 1$, we have the uniform bounds
$\sup_{\eta\in\mathcal{N}_{n}}|\widetilde{\ell}_{n}(\eta)-\ell_{n}(\eta)|\;\leq\;C_{1}\delta_{n}\,$,
and
$\sup_{\eta\in\mathcal{N}_{n}}\|\widetilde{S}_{n}(\eta)-S_{n}(\eta)\|\;\leq\;C_{2}\delta_{n}\,.$
Since $\eta_{o}\in\mathcal{N}_{n}$ for all $n$ large enough, subtracting the values at $\eta_{o}$ gives
$\sup_{d(\eta,\eta_{o})\leq M\tau_{n}}\bigl|\ell_{n}(\eta)-\widetilde{\ell}_{n}(\eta)-\{\ell_{n}(\eta_{o})-\widetilde{\ell}_{n}(\eta_{o})\}\bigr|=O_{p}(\delta_{n})\,$, (29)
and
$\sup_{d(\eta,\eta_{o})\leq M\tau_{n}}\bigl\|\widetilde{S}_{n}(\eta)-S_{n}(\eta)\bigr\|=O_{p}(\delta_{n})\,.$ (30)
These bounds reflect that the per-subject approximation error is $O(\delta_{n})$ and the network outputs are uniformly bounded because the sieve $\mathcal{G}(K,s,\boldsymbol{p},D)$ is constructed with bounded weights.
By (29) and Theorem 1, on the ball $\{d(\eta,\eta_{o})\leq\tau_{n}\}$ we have
$\sup_{d(\eta,\eta_{o})\leq\tau_{n}}\bigl|(\widetilde{\ell}_{n}-\ell_{0})(\eta)-(\widetilde{\ell}_{n}-\ell_{0})(\eta_{o})\bigr|$ $\leq\sup_{d(\eta,\eta_{o})\leq\tau_{n}}\bigl|(\ell_{n}-\ell_{0})(\eta)-(\ell_{n}-\ell_{0})(\eta_{o})\bigr|$
$\qquad+\sup_{d(\eta,\eta_{o})\leq\tau_{n}}\bigl|\widetilde{\ell}_{n}(\eta)-\ell_{n}(\eta)-\{\widetilde{\ell}_{n}(\eta_{o})-\ell_{n}(\eta_{o})\}\bigr|$
$=O_{p}(\tau_{n}^{2})+O_{p}(\delta_{n})=O_{p}(\tau_{n}^{2}),$
[…]
By the score equations, $S_{n}(\widehat{\eta})={\bf 0}$ and $\widetilde{S}_{n}(\widetilde{\eta})={\bf 0}$. Therefore,
[…]
Since $d(\widehat{\eta},\eta_{o})=O_{p}(\tau_{n})$ and $d(\widetilde{\eta},\eta_{o})=O_{p}(\tau_{n})$, there exists a constant $C>0$ such that, with probability tending to one, $\{d(\widehat{\eta},\eta_{o})\leq C\tau_{n},\;d(\widetilde{\eta},\eta_{o})\leq C\tau_{n}\}\subseteq\mathcal{N}_{n}$, where $M\geq C$. On this event, since the two estimators satisfy the score equations
$S_{n}(\widehat{\eta})={\bf 0}\,,\qquad\widetilde{S}_{n}(\widetilde{\eta})={\bf 0}\,$,
we also get
[…]
where $R_{n}(\eta)=\widetilde{S}_{n}(\eta)-S_{n}(\eta)$. Subtracting $S_{n}(\widehat{\eta})=0$ yields
[…]
In addition,
$S_{n}(\widetilde{\eta})-S_{n}(\widehat{\eta})=\underbrace{S_{n}(\widetilde{g},\widetilde{}\mbox{${\theta}$})-S_{n}(\widetilde{g},\widehat{}\mbox{${\theta}$})}_{A}+\underbrace{S_{n}(\widetilde{g},\widehat{}\mbox{${\theta}$})-S_{n}(\widehat{g},\widehat{}\mbox{${\theta}$})}_{B}.$
and (31) becomes
[…]
By the mean value theorem in $\theta$}, there exists $\mbox{${\theta}$}^{\dagger}$ on the segment between $\widehat{}\mbox{${\theta}$}$ and $\widetilde{}\mbox{${\theta}$}$ such that
$A=\dot{S}_{n}(\widetilde{g},\mbox{${\theta}$}^{\dagger})\,(\widetilde{}\mbox{${\theta}$}-\widehat{}\mbox{${\theta}$})\,.$
By Lemma 9, $\dot{S}_{n}(\widetilde{g},\mbox{${\theta}$}^{\dagger})\xrightarrow{P}-I(\mbox{${\theta}$}_{o})$ and $\|\dot{S}_{n}(\widetilde{g},\mbox{${\theta}$}^{\dagger})^{-1}\|=O_{p}(1)$. Substituting this into (31), we obtain
[…]
and
[…]
Lemma 8 gives uniformly over $\mathcal{N}_{n}$,
$\|R_{n}(\eta)\|\leq C\delta_{n},$
so
$\sqrt{n}\,R_{n}(\widetilde{\eta})=O_{p}(\sqrt{n}\delta_{n})=o_{p}(1),$
because $\sqrt{n}\delta_{n}\to 0$ by assumption.
We decompose
$B=S_{n}(\widetilde{g},\widehat{}\mbox{${\theta}$})-S_{n}(\widehat{g},\widehat{}\mbox{${\theta}$})=\{\mathbb{P}S(\widetilde{g},\widehat{}\mbox{${\theta}$})-\mathbb{P}S(\widehat{g},\widehat{}\mbox{${\theta}$})\}+\{(\mathbb{P}_{n}-\mathbb{P})S(\widetilde{g},\widehat{}\mbox{${\theta}$})-(\mathbb{P}_{n}-\mathbb{P})S(\widehat{g},\widehat{}\mbox{${\theta}$})\},$
where $S(\eta;{\bf W})$ is the per–observation score in $\theta$} and $S_{n}(\eta)=\mathbb{P}_{n}S(\eta;\cdot)$. The second–order expansion of the population score (see the proof of Theorem 4) yields
$\|\mathbb{P}S(\widetilde{g},\widehat{}\mbox{${\theta}$})-\mathbb{P}S(\widehat{g},\widehat{}\mbox{${\theta}$})\|\lesssim\|\widetilde{g}-g_{o}\|_{L^{2}(\Delta)}^{2}+\|\widehat{g}-g_{o}\|_{L^{2}(\Delta)}^{2}=O_{p}(\tau_{n}^{2})\,.$
Moreover, Lemma 4 applied to the localized score class $\{S(g,\widehat{}\mbox{${\theta}$};\cdot):d((g,\widehat{}\mbox{${\theta}$}),\eta_{o})\leq C\tau_{n}\}$ implies
$\sup_{d((g,\widehat{\theta}),\eta_{o})\leq C\tau_{n}}|(\mathbb{P}_{n}-\mathbb{P})S(g,\widehat{}\mbox{${\theta}$};\cdot)|=O_{p}(\tau_{n}^{2})\,$,
and since both $(\widetilde{g},\widehat{}\mbox{${\theta}$})$ and $(\widehat{g},\widehat{}\mbox{${\theta}$})$ lie in this ball with probability 1o(1)1-o(1), we obtain
$\|B\|=O_{p}(\tau_{n}^{2}),\qquad\sqrt{n}B=o_{p}(1),$
because $\sqrt{n}\tau_{n}^{2}\to 0$. Using (32) together with $\dot{S}_{n}(\widetilde{g},\mbox{${\theta}$}^{\dagger})^{-1}=O_{p}(1)$, $\sqrt{n}B=o_{p}(1)$, and $\sqrt{n}R_{n}(\widetilde{\eta})=o_{p}(1)$, we obtain
$\sqrt{n}(\widetilde{}\mbox{${\theta}$}-\widehat{}\mbox{${\theta}$})=o_{p}(1)\,.$
Finally, combining this with Theorem 4, which states that $\sqrt{n}(\widehat{}\mbox{${\theta}$}-\mbox{${\theta}$}_{o})$ is asymptotically normal $N\bigl(\mathbf{0},I(\mbox{${\theta}$}_{o})^{-1}\bigr)$, and using Slutsky’s theorem yields
$\sqrt{n}(\widetilde{}\mbox{${\theta}$}-\mbox{${\theta}$}_{o})=\sqrt{n}(\widehat{}\mbox{${\theta}$}-\mbox{${\theta}$}_{o})+\sqrt{n}(\widetilde{}\mbox{${\theta}$}-\widehat{}\mbox{${\theta}$})\;\xRightarrow{d}\;N\bigl(\mathbf{0},I(\mbox{${\theta}$}_{o})^{-1}\bigr)\,$,
as claimed. ∎

### v2

@@
-Flexible Deep Neural Networks for Partially Linear Survival Data
+Flexible Deep Neural Networks for Partially Linear Survival Data: Estimation and Survival Inference
 Abstract
@@
 In contrast to existing DNN approaches for partially linear Cox models, FLEXI–Haz does not rely on the proportional hazards assumption.
-We establish theoretical guarantees: the neural network component attains minimax-optimal convergence rates that depends on composite Hölder classes, and the linear estimator is $\sqrt{n}$–consistent, asymptotically normal, and semiparametrically efficient.
-Extensive simulations and real-data analyses demonstrate that FLEXI–Haz provides accurate estimation of the linear effect, offering a principled and interpretable alternative to modern methods based on proportional hazards.
-Code for implementing FLEXI-Haz, as well as scripts for reproducing data analyses and simulations is available at GitHub site https://github.com/AsafBanana/FLEXI-Haz.
-keywords: Deep neural networks; Interpretability; Partial linear regression; Survival analysis; Time to event data.
+We establish theoretical guarantees: the neural network component attains minimax-optimal convergence rates over composite Hölder classes, the linear estimator is $\sqrt{n}$-consistent, asymptotically normal, and semiparametrically efficient, and we develop a cross-fitted one-step estimator of the cumulative hazard and survival function for a new subject, together with pointwise asymptotic confidence intervals.
+To the best of our knowledge, this is the first frequentist asymptotic pointwise inference result for a survival function in a DNN survival model, with or without a linear component.
+Simulations and real-data analyses demonstrate the utility of FLEXI–Haz as a principled and interpretable alternative to methods based on proportional hazards.
+Time to event data,
+keywords:
+[class=MSC]
+keywords:
+[…]
 1 Introduction
-Deep neural networks (DNNs) are widely used in modern data analysis because of their ability to capture complex nonlinear patterns, handle high-dimensional inputs, and scale to large datasets.
-They have achieved strong performance in diverse domains such as image recognition, speech processing, and natural language understanding, often improving on traditional statistical and machine learning methods (lecun2015deep).
-Continued progress in architectures, optimization, and computational resources has further expanded their range of applications in both scientific and industrial domains.
-Deep neural networks for survival analysis.
-Survival analysis aims to model the distribution of event times in the presence of censoring, with applications in medicine, public health, epidemiology, engineering, and finance (kalbfleisch2011; klein2006survival).
-In recent years, DNNs have become increasingly popular in survival analysis thanks to their ability to model complex nonlinear relationships between covariates and event times and to scale efficiently to high-dimensional data.
-Many existing approaches extend the classical semi-parametric Cox proportional hazards (PH) model (cox1972regression) by replacing the parametric linear risk score with a neural network, while using the negative partial likelihood as the network’s loss function.
-Early work by faraggi1995neural employed a single hidden-layer network in place of the linear component, and katzman2018deepsurv extended this to deeper architectures (DeepSurv), demonstrating improved concordance in several benchmark datasets.
-Variants of this approach have been proposed for genomics (yousefi2017predicting; ching2018coxnnet), medical imaging (haarburger2019image; li2019deep), and recurrent or discrete-time formulations (giunchiglia2018rnnsurv; gensheimer2019nnet).
-While these PH-based DNN methods can yield strong empirical performance, they inherit the PH restriction, which may limit their ability to capture rich time-dependent effects.
-Non-PH modeling with deep learning.
-To address the limitations of the PH assumption, a growing body of work has focused on DNN-based models that allow covariate effects to vary with time.
-One strategy augments the network input with time itself, enabling non PH while retaining an unspecified baseline hazard, as in Cox-Time of kvamme2019time.
-Other approaches abandon the Cox structure entirely in favor of modeling the event time distribution or hazard function directly.
-In particular, hu2023conditional proposed a framework for estimating the full survival function, using a likelihood function derived from the continuous-time hazard without introducing a baseline hazard function.
-Their formulation naturally accommodates time-varying covariates through a data-expansion scheme and applies optimization to the log-hazard.
-By sidestepping the baseline hazard and proportionality constraints, such approaches offer greater modeling flexibility and can better reflect complex temporal dynamics in the hazard.
-The utility of this approach is illustrated through simulations and real-data analyses, but no asymptotic results are provided.
-Semiparametric models and deep learning.
-Alongside fully nonparametric formulations, there is considerable interest in semiparametric survival models that combine a parametric component for covariates of primary interest with a flexible, nonparametric component for high-dimensional or nuisance covariates.
-This structure preserves interpretability and enables formal inference for the parametric component, while leveraging modern machine learning to estimate the nuisance function.
-Theoretical foundations for this paradigm have been established in the non-survival setting by schmidthieber2020nonparametric, who derived minimax-optimal rates for ReLU-activated DNNs in nonparametric regression, and extended to survival analysis by zhong2022deep in the partially linear Cox model.
-In the latter, the log-hazard at time $t$ is modeled as $\log h_{0}(t)+\mbox{${\zeta}$}^{\top}{\bf Z}+f({\bf X})$, where $h_{0}(t)$ is an unspecified baseline hazard function, ${\bf Z}\in\mathbb{R}^{p}$ is the vector of covariates of primary interest, ${\bf X}\in\mathbb{R}^{d}$ is the vectors of nuisance covariates, and $\mbox{${\zeta}$}\in\mathbb{R}^{p}$ is the main parameter of interest, estimated at the semiparametric efficiency bound via partial likelihood.
-The nuisance function $f$ is approximated by a DNN.
-This approach achieves both optimal convergence for $g$ and $\sqrt{n}$-consistency for $\zeta$} under the PH assumption.
-The work of zhong2022deep demonstrates how semiparametric theory and deep learning can be integrated to yield interpretable, efficient estimators in survival analysis, but like other Cox-based methods, it remains tied to PH and baseline hazard estimation.
-Contributions.
-In this work, we develop a semiparametric survival model that extends the partially linear framework of zhong2022deep to the non-PH setting.
-We model the log-hazard function at time $t$ as $\mbox{${\theta}$}_{o}^{\sf T}{\bf Z}+g_{o}(t,{\bf X})$, where $\mbox{${\theta}$}_{o}\in\mathbb{R}^{p}$ is the parameter of primary interest and $g_{o}$ is an unknown nuisance function approximated by a DNN.
-This formulation allows arbitrary interactions between the nuisance covariates $\bf X$} and time $t$, naturally accommodates time-dependent effects and time-dependent covariates within $\bf X$}, and does not require specifying or estimating a baseline hazard function.
-A central methodological difference from existing semiparametric approaches is that we work with the full likelihood, rather than the partial likelihood commonly used in Cox-type models.
-Despite removing the PH structure, we establish that the estimator of the parametric component $\mbox{${\theta}$}_{o}$ remains $\sqrt{n}$-consistent, asymptotically normal, and semiparametric efficient.
-In particular, we show that the theoretical guarantees of zhong2022deep, developed for the partially linear Cox model, continue to hold in the substantially more general non-PH setting, provided the hazard is almost surely bounded away from zero.
-While parts of our analysis build on the approximation theory of schmidthieber2020nonparametric and the high-level strategy of zhong2022deep, several key components of our asymptotic theory are fundamentally different.
-The core technical challenges arise from the fact that our neural network estimator is a function of both $t$ and $\bf X$}, rather than $\bf X$} alone, and the fact that we use the full likelihood rather than a profiling argument.
-These differences require new empirical process tools, a new characterization of the nuisance tangent space, and a new argument for controlling the bias induced by the DNN approximation.
-Finally, when the linear effect $\mbox{${\theta}$}_{o}^{\top}{\bf Z}$ is removed and $\bf Z$} is absorbed into $g_{o}$, our model collapses to the conditional hazard model of hu2023conditional.
-Our analysis therefore fills an important theoretical gap by providing the first asymptotic guarantees for that model as well.
-Code for implementing FLEXI-Haz, as well as scripts for reproducing all data analyses and simulations, is available at GitHub site https://github.com/AsafBanana/FLEXI-Haz.
+Deep neural networks (DNNs) have become a central tool in modern data analysis because of their ability to capture complex nonlinear structure, accommodate high-dimensional inputs, and scale to large data sets.
+They have achieved strong empirical performance across a wide range of domains, including image recognition, speech processing, and natural language understanding (LeCun et al., 2015).
+These developments have also stimulated substantial interest in using DNNs for survival analysis.
+Survival analysis concerns the distribution of event times in the presence of censoring and arises in many fields, including medicine, public health, epidemiology, engineering, and finance (Kalbfleisch and Prentice, 2011; Klein and Moeschberger, 2006).
+In recent years, DNN-based survival methods have become increasingly popular because they can model complex nonlinear relationships between covariates and event times while remaining computationally feasible in high dimensions.
+A large part of this literature extends the Cox proportional hazards (PH) model (Cox, 1972) by replacing the linear risk score with a neural network and training by the negative partial likelihood.
+Early work by Faraggi and Simon (1995) used a single hidden-layer network, and Katzman et al.
+(2018) extended this idea to deeper architectures in DeepSurv.
+Related PH-based DNN methods have been developed for genomics (Yousefi et al., 2017; Ching et al., 2018), medical imaging (Haarburger et al., 2019; Li et al., 2019), recurrent and discrete-time settings (Giunchiglia et al., 2018; Gensheimer and Narasimhan, 2019), and hypothesis testing for covariate effects on survival (Zhong et al., 2026).
+While these approaches can perform well empirically, they inherit the PH restriction, which may be too rigid when covariate effects vary over time.
+To move beyond the PH assumption, several DNN-based methods allow time-varying effects.
+One approach augments the network input with time while retaining an unspecified baseline hazard, as in Cox-Time (Kvamme et al., 2019).
+Other approaches use baseline-free formulations based on the continuous-time hazard; see, for example, Hu and Nan (2023); Su et al.
+(2026).
+These methods naturally accommodate time-varying effects and avoid the PH constraint, but Hu and Nan (2023) do not provide asymptotic theory, and Su et al.
+(2026) focus on one- and two-sample hypothesis testing.
+At the same time, there is strong interest in semiparametric survival models that retain a low-dimensional interpretable component for covariates of primary interest while modeling high-dimensional or nuisance covariates flexibly.
+This partly linear structure is attractive because it combines interpretability with statistical efficiency.
+In the non-survival setting, theoretical foundations for DNN estimation over structured nonparametric classes were developed by Schmidt-Hieber (2020).
+In survival analysis, Zhong et al.
+(2022) extended this paradigm to the partially linear Cox model $\log h_{0}(t)+\zeta^{\top}{\bf Z}+f({\bf X})$, where $\bf Z$} is the low-dimensional covariate vector of primary interest, $\bf X$} is a nuisance covariate vector, and $f$ is approximated by a DNN.
+Their approach yields optimal convergence of the nuisance estimator and $\sqrt{n}$-consistent estimation of $\zeta$, but it remains tied to the PH assumption and to baseline hazard estimation.
+In this paper we develop a flexible semiparametric survival model that extends the partially linear framework of Zhong et al.
+(2022) to a substantially more general non-PH setting.
+Specifically, we model the log-hazard as $\mbox{${\theta}$}_{o}^{\top}{\bf Z}+g_{o}(t,{\bf X})$, where $\mbox{${\theta}$}_{o}\in\mathbb{R}^{p}$ is the finite-dimensional parameter of interest and $g_{o}(t,{\bf X})$ is an unknown nuisance function approximated by a DNN.
+This formulation is baseline-free: all time dependence is absorbed into $g_{o}(t,{\bf X})$, rather than decomposed into a baseline hazard and covariate effects.
+As a result, the model allows arbitrary interactions between time and the nuisance covariates, naturally accommodates time-varying effects, and strictly enlarges the class of partially linear Cox-type models.
+This extension is methodologically nontrivial.
+Once the nuisance component depends on both $t$ and $\bf X$}, the Cox partial likelihood is no longer available as a loss function, and the full likelihood must be used instead.
+This introduces a predictable cumulative-hazard integral that must be numerically approximated during training and leads to a substantially different asymptotic analysis.
+In particular, the time-dependent nuisance component requires new empirical-process arguments, a new characterization of the nuisance tangent space, and a new control of the bias induced by DNN approximation.
+We show, nevertheless, that the estimator of $\mbox{${\theta}$}_{o}$ remains $\sqrt{n}$-consistent, asymptotically normal, and semiparametrically efficient.
+A second and particularly important contribution of the paper is inference for the survival curve of a new subject.
+For an independent draw $({\bf X}_{0},{\bf Z}_{0})$ from the covariate distribution, we construct a cross-fitted one-step estimator of the cumulative hazard $H_{o}(t\mid{\bf X}_{0},{\bf Z}_{0})$ and the survival function $S_{o}(t\mid{\bf X}_{0},{\bf Z}_{0})$, together with Wald-type pointwise confidence intervals.
+To the best of our knowledge, this is the first frequentist asymptotic pointwise inference result for a survival curve in a DNN survival model, with or without a separate linear component.
+This result is established under an additional assumption that the nuisance covariate vector $\bf X$} has discrete finite support, which yields an explicit closed-form Riesz representer and therefore a tractable one-step correction.
+From a practical perspective, this restriction is often acceptable because nuisance covariates frequently include categorical variables or grouped continuous features, and because the finite-support assumption is required only for the explicit survival-inference theory rather than for estimation and efficiency of $\mbox{${\theta}$}_{o}$.
+The broader survival-inference problem can also be studied under general covariate distributions through a high-level semiparametric characterization of the efficient influence function.
+Since that formulation is more technical and does not yield a comparably explicit estimator without additional approximation arguments, we present it in the appendix.
+Our framework is also relevant beyond the partially linear setting.
+If the linear component is absorbed into the nuisance function, then the one-step construction and efficient influence function derived here remain valid, with the parametric term dropping out.
+In this sense, the paper also contributes asymptotic survival-inference tools for baseline-free DNN hazard models more broadly.
+Compared with Su et al.
+(2026), who derive functional asymptotic normality for smooth hazard-function functionals and use it for testing, we study a different and sharper target: semiparametric efficiency for an interpretable linear effect and explicit Wald-type pointwise inference for the survival curve of a new subject.
+The remainder of the paper is organized as follows.
+Sections 2 and 3 introduce the model, the DNN class, and the FLEXI–Haz estimation and survival-inference procedures.
+Sections 4 and 5 develop the asymptotic theory for estimation and pointwise survival inference.
+Simulation studies and real-data analyses are presented later.
+Code for FLEXI–Haz and for reproducing the simulations and data analyses is available at https://github.com/AsafBanana/FLEXI--Haz.
 2 Model Formulation
@@
 The observed data consist of $n$ independent replicates, $\{(T_{i},\Delta_{i},\mathbf{X}_{i},\mathbf{Z}_{i})\}_{i=1}^{n}$.
-Under the above setting, the Cox PH model (cox1972regression) specifies the instantaneous hazard function at time $t>0$ as
+Under the above setting, the Cox PH model (Cox, 1972) specifies the instantaneous hazard function at time $t>0$ as
 $h_{1}(t\mid{{\bf X}},{{\bf Z}})=h_{0}(t)\exp({\mbox{${\alpha}$}}_{1}^{\top}{{\bf X}}+{\mbox{${\alpha}$}}_{2}^{\top}{{\bf Z}})\,$,
@@
 The model remains widely used due to its interpretability, semiparametric structure, and the availability of efficient estimation procedures.
-DeepSurv (katzman2018deepsurv) extends the Cox model by replacing the linear risk score $\exp({\mbox{${\alpha}$}}_{1}^{\top}\mathbf{X}+{\mbox{${\alpha}$}}_{2}^{\top}\mathbf{Z})$ with a flexible nonlinear function $f_{1}({{\bf X}},{{\bf Z}})$ estimated by a DNN.
-Cox-Time (kvamme2019time) further generalized this formulation by replacing $f_{1}$ with a time-depend function, $f_{2}(t,{{\bf X}},{{\bf Z}})$, thereby relaxing the PH assumption and permitting time-covariate interactions.
-Because $f_{2}$ is time-dependent while the baseline hazard is left unspecified, identifiability becomes nontrivial; kvamme2019time address this challenge through a numerical normalization strategy.
-More recently, hu2023conditional proposed a baseline-free hazard model of the form
+DeepSurv (Katzman et al., 2018) extends the Cox model by replacing the linear risk score $\exp({\mbox{${\alpha}$}}_{1}^{\top}\mathbf{X}+{\mbox{${\alpha}$}}_{2}^{\top}\mathbf{Z})$ with a flexible nonlinear function $f_{1}({{\bf X}},{{\bf Z}})$ estimated by a DNN.
+Cox-Time (Kvamme et al., 2019) further generalized this formulation by replacing $f_{1}$ with a time-dependent function, $f_{2}(t,{{\bf X}},{{\bf Z}})$, thereby relaxing the PH assumption and permitting time-covariate interactions.
+Because $f_{2}$ is time-dependent while the baseline hazard is left unspecified, identifiability becomes nontrivial; Kvamme et al.
+(2019) address this challenge through a numerical normalization strategy.
+More recently, Hu and Nan (2023) proposed a baseline-free hazard model of the form
 $h_{3}(t\mid{{\bf X}},{{\bf Z}})=\exp\{f_{3}(t,{{\bf X}},{{\bf Z}})\}\,$,
@@
 Although these DNN-based approaches are highly flexible, they do not provide interpretable estimates for the effect of $\bf Z$}}.
-To address this limitation, zhong2022deep introduced the deep partially linear Cox model
+To address this limitation, Zhong et al.
+(2022) introduced the deep partially linear Cox model
 $h_{4}(t\mid\mathbf{X},\mathbf{Z})=h_{0}(t)\exp\{\mbox{${\zeta}$}^{\top}\mathbf{Z}+f(\mathbf{X})\}\,$,
-which augments the Cox framework by separating time component from the covariates and retaining a parametric term $\mbox{${\zeta}$}^{\top}\mathbf{Z}$ for interpretability, while modeling nuisance effects through a DNN.
+which augments the Cox framework by separating the time component from the covariates and retaining a parametric term $\mbox{${\zeta}$}^{\top}\mathbf{Z}$ for interpretability, while modeling nuisance effects through a DNN.
 This structure enables valid inference for $\zeta$} while still allowing complex, nonlinear effect of $\bf X$}}.
-In this work, we extend the deep partially linear framework of zhong2022deep to the non-PH setting by considering a baseline-free model of the form
-$h_{o}(t\mid\mathbf{X},{{\bf Z}})=\exp\{\mbox{${\theta}$}_{o}^{\top}\mathbf{Z}+g_{o}(t,\mathbf{X})\},$ (1)
+In this work, we extend the deep partially linear framework of Zhong et al. (2022) to the non-PH setting by considering a baseline-free model of the form
+$h_{o}(t\mid\mathbf{X},{{\bf Z}})=\exp\{\mbox{${\theta}$}_{o}^{\top}\mathbf{Z}+g_{o}(t,\mathbf{X})\},$ (2.1)
 where $g_{o}(t,\mathbf{X})$ is estimated by a DNN.
-Our formulation retains the interpretability and inference guarantees for the parametric component ${\mbox{${\theta}$}}_{o}$, but generalizes the nuisance component from $f(\mathbf{X})$ to $g_{o}(t,\mathbf{X})$, thereby allowing fully flexible, time-varying effects and removing the reliance on a baseline hazard function.
-This extension overcomes the PH restriction inherent in zhong2022deep, resolves the non-identifiability issues that arise when time is included in the DNN, and greatly expands the class of survival models that can be handled within a semiparametric, partially linear framework.
-As we demonstrate in Section 6, restricting the nuisance function to depend only on only on $\mathbf{X}$, as in the partially linear Cox model, can induce substantial bias in the estimation of the linear effect whenever the nuisance covariates exhibit time-varying effects.
+Model (2.1) is baseline-free in the sense that the entire time dependence of the hazard is absorbed into the nuisance function $g_{o}(t,{\bf X})$, rather than being decomposed into a baseline hazard and covariate effects.
+This is not merely a reparameterization of a Cox-type model with time-varying coefficients: allowing $g_{o}(t,{\bf X})$ to be an unrestricted function of $(t,{\bf X})$ strictly enlarges the model class relative to representations of the form $h_{0}(t)\exp\{\zeta^{\top}{\bf Z}+f(X)\}$.
+This extension overcomes the PH restriction inherent in Zhong et al.
+(2022), resolves the non-identifiability issues that arise when time is included in the DNN, and greatly expands the class of survival models that can be handled within a semiparametric, partially linear framework.
+Identifiability of $\mbox{${\theta}$}_{o}$ in this setting follows from the partially linear structure together with Assumption A5 (see Section 4.2).
+These conditions ensure that no nontrivial linear function of $\bf Z$} can be represented almost surely as a function of $(t,{\bf X})$, and therefore prevent $g_{o}(t,{\bf X})$ from absorbing components of the linear effect.
+This is formulated in Lemma 3 (Section 4.3).
+As we demonstrate in Section 7, restricting the nuisance function to depend only on $\mathbf{X}$, as in the partially linear Cox model, can induce substantial bias in the estimation of the linear effect whenever the nuisance covariates exhibit time-varying effects.
 Allowing $g_{o}$ to depend on both $t$ and $\mathbf{X}$ eliminates this source of bias but introduces significant methodological challenges.
 First, the negative partial likelihood can no longer be used as the DNN loss function.
-Instead, the full-likelihood is adopted which includes the predictable cumulative hazard that must be numerically approximated at every training step, increasing the computational burden.
+Instead, the full likelihood is adopted, which includes the predictable cumulative hazard that must be numerically approximated at every training step, increasing the computational burden.
 Second, the presence of this integral markedly complicates the asymptotic analysis: the time dimension introduces additional stochastic terms that must be controlled, requiring new empirical process arguments beyond those used for time-invariant nuisance functions.
-Our estimation framework in Section 3 addresses these computational difficulties, and the technical developments leading to the asymptotic results are presented in Section 4.3.
+Our estimation framework in Section 3 addresses these computational difficulties, and the technical developments leading to the asymptotic results are presented in Sections 4 and 5.
 2.1 Composite Hölder Class of $g_{o}$
-To formalize the complexity assumptions on the nuisance component $g_{o}$, we follow the compositional function framework of juditsky2009 and schmidthieber2020nonparametric, and use the notation of zhong2022deep.
+To formalize the complexity assumptions on the nuisance component $g_{o}$, we follow the compositional function framework of Juditsky et al.
+(2009) and Schmidt-Hieber (2020), and use the notation of Zhong et al.
+(2022).
 We assume that the nuisance function $g_{o}$ admits a hierarchical, low-dimensional representation of the form
@@
 with the convention that $(g\circ f)(x)=g(f(x))$.
-This composition is associate with a dimension vector ${\bf d}^{*}=(d^{*}_{0},\ldots,d^{*}_{q+1})\in\mathbb{N}_{+}^{q+2}$ and an intrinsic dimension vector $\tilde{\bf d}=(\tilde{d}_{0},\ldots,\tilde{d}_{q})\in\mathbb{N}_{+}^{q+1}$, $\tilde{d}_{\ell}\leq d^{*}_{\ell}$ for each $\ell$.
+This composition is associated with a dimension vector ${\bf d}^{*}=(d^{*}_{0},\ldots,d^{*}_{q+1})\in\mathbb{N}_{+}^{q+2}$ and an intrinsic dimension vector $\tilde{\bf d}=(\tilde{d}_{0},\ldots,\tilde{d}_{q})\in\mathbb{N}_{+}^{q+1}$, $\tilde{d}_{\ell}\leq d^{*}_{\ell}$ for each $\ell$.
 The outermost input is $d^{*}_{0}=r=1+d$, corresponding to observed times and covariates $\mathbf{X}$, and the final output is scalar $d^{*}_{q+1}=1$.
 For each layer $l$, $l=0,\ldots,q$, the vector-valued mapping $g_{\ell}=(g_{\ell 1},\ldots,g_{\ell\,d^{*}_{\ell+1}})^{\top}$ maps the bounded set $a_{\ell},b_{\ell}]^{\,d^{*}_{\ell}}$ into $a_{\ell+1},b_{\ell+1}]^{\,d^{*}_{\ell+1}}$, where $a_{l},b_{l}\in\mathbb{R}$ are fixed finite constants ensuring each intermediate representation remains bounded.
-Additionally, each component $g_{\ell j}$ depending on at most $\tilde{d}_{\ell}$ coordinates of its input.
+Additionally, each component $g_{\ell j}$ depends on at most $\tilde{d}_{\ell}$ coordinates of its input.
 For illustration, consider $f:\mathbb{R}^{6}\to\mathbb{R}$ defined by
@@
 The ball of $\alpha$-Hölder functions with radius $M$ consists of all such functions whose derivatives and Hölder constants are bounded by $M$.
-Formally, for a domain $\mathbb{D}\subset\mathbb{R}^{m}$, the Hölder ball is defined as
-$\mathcal{H}^{\alpha}_{m}(\mathbb{D},M)=\Big\{g:\mathbb{D}\to\mathbb{R}:\sum_{\boldsymbol{\beta}:|\boldsymbol{\beta}|<\alpha}\|\partial^{\boldsymbol{\beta}}g\|_{\infty}+\sum_{\boldsymbol{\beta}:|\boldsymbol{\beta}|=\lfloor\alpha\rfloor}\sup_{u,v\in\mathbb{D},u\neq v}\frac{|\partial^{\boldsymbol{\beta}}g(u)-\partial^{\boldsymbol{\beta}}g(v)|}{\|u-v\|^{\alpha-\lfloor\alpha\rfloor}}\leq M\Big\},$
-where $\boldsymbol{\beta}^{\sf T}=(\beta_{1},\ldots,\beta_{m})\in\mathbb{N}^{m}$ is a multi-index $\partial^{\boldsymbol{\beta}}=\partial^{\beta_{1}}\cdots\partial^{\beta_{m}}$, $|\boldsymbol{\beta}|=\sum_{j=1}^{m}\beta_{j}$, and $\|g\|_{\infty}$ is the sup–norm of a function over its domain.
+Formally, for a domain $\mathbb{D}\subset\mathbb{R}^{r}$, the Hölder ball is defined as
+$\mathcal{H}^{\alpha}_{r}(\mathbb{D},M)=\Big\{g:\mathbb{D}\to\mathbb{R}:\sum_{\boldsymbol{\beta}:|\boldsymbol{\beta}|<\alpha}\|\partial^{\boldsymbol{\beta}}g\|_{\infty}+\sum_{\boldsymbol{\beta}:|\boldsymbol{\beta}|=\lfloor\alpha\rfloor}\sup_{u,v\in\mathbb{D},u\neq v}\frac{|\partial^{\boldsymbol{\beta}}g(u)-\partial^{\boldsymbol{\beta}}g(v)|}{\|u-v\|^{\alpha-\lfloor\alpha\rfloor}}\leq M\Big\},$
+where $\boldsymbol{\beta}^{\sf T}=(\beta_{1},\ldots,\beta_{r})\in\mathbb{N}^{r}$ is a multi-index $\partial^{\boldsymbol{\beta}}=\partial^{\beta_{1}}\cdots\partial^{\beta_{r}}$, $|\boldsymbol{\beta}|=\sum_{j=1}^{r}\beta_{j}$, and $\|g\|_{\infty}$ is the sup–norm of a function over its domain.
 To formalize the function class used for the nuisance component, let $\boldsymbol{\alpha}^{\sf T}=(\alpha_{0},\ldots,\alpha_{q})\in\mathbb{R}_{+}^{q+1}$ denote the layerwise smoothness parameters, $\bf d$} the dimension vector, and $\tilde{{\bf d}}$ the intrinsic dimension vector.
 The composite Hölder class is defined as
-$\mathcal{H}(q,\boldsymbol{\alpha},{\bf d},\tilde{{\bf d}},M)=\Big\{g=g_{q}\circ\cdots\circ g_{0}:g_{\ell j}\in\mathcal{H}^{\alpha_{\ell}}_{\tilde{d}_{\ell}}([a_{\ell},b_{\ell}]^{\tilde{d}_{\ell}},M),\ |a_{\ell}|,|b_{\ell}|\leq M\Big\}.$
-This class accomodates a broad range of structured functions, such as additive models, single- and multiple-index models, and hierarchical interactions, through the intrinsic dimensions $\tilde{d}_{\ell}$.
-Finally, we assume that the true nuisance function $g_{o}$ belongs to the composite class, $g_{o}\in\mathcal{H}(q,\boldsymbol{\alpha},{\bf d},\tilde{{\bf d}},M)$.
-In the classical (non-composite) Hölder class $\mathcal{H}^{\alpha}_{d}$, the minimax-optimal rate is $n^{-\alpha/(2\alpha+d)}$ (tsybakov2009).
-For multilayer composite Hölder classes, schmidthieber2020nonparametric established matching minimax lower and upper bounds (up to polylogarithmic factors), showing that the optimal rate is governed by $\gamma_{n}\;=\;\max_{\ell=0,\ldots,q}\,n^{-\tilde{\alpha}_{\ell}/(2\tilde{\alpha}_{\ell}+\tilde{d}_{\ell})}$, where the effective smoothness indices are $\tilde{\alpha}_{\ell}\;=\;\alpha_{\ell}\prod_{k=\ell+1}^{q}(\alpha_{k}\wedge 1),\ \ a\wedge b=\min\{a,b\}$.
+$\mathcal{H}(q,\boldsymbol{\alpha},{\bf d}^{*},\tilde{{\bf d}},M)=\Big\{g=g_{q}\circ\cdots\circ g_{0}:g_{\ell j}\in\mathcal{H}^{\alpha_{\ell}}_{\tilde{d}_{\ell}}([a_{\ell},b_{\ell}]^{\tilde{d}_{\ell}},M),\ |a_{\ell}|,|b_{\ell}|\leq M\Big\}.$
+This class accommodates a broad range of structured functions, such as additive models, single- and multiple-index models, and hierarchical interactions, through the intrinsic dimensions $\tilde{d}_{\ell}$.
+Finally, we assume that the true nuisance function $g_{o}$ belongs to the composite class, $g_{o}\in\mathcal{H}(q,\boldsymbol{\alpha},{\bf d}^{*},\tilde{{\bf d}},M)$.
+In the classical (non-composite) Hölder class $\mathcal{H}^{\alpha}_{d}$, the minimax-optimal rate is $n^{-\alpha/(2\alpha+d)}$ (Tsybakov, 2009).
+For multilayer composite Hölder classes, Schmidt-Hieber (2020) established matching minimax lower and upper bounds (up to polylogarithmic factors), showing that the optimal rate is governed by $\gamma_{n}\;=\;\max_{\ell=0,\ldots,q}\,n^{-\tilde{\alpha}_{\ell}/(2\tilde{\alpha}_{\ell}+\tilde{d}_{\ell})}$, where the effective smoothness indices are $\tilde{\alpha}_{\ell}\;=\;\alpha_{\ell}\prod_{k=\ell+1}^{q}(\alpha_{k}\wedge 1),\ \ a\wedge b=\min\{a,b\}$.
 Thus, each layer contributes to the overall difficulty of the estimation problem through its intrinsic dimension $\tilde{d}_{\ell}$ and effective smoothness $\tilde{\alpha}_{\ell}$.
-In Theorem 2 we show that our estimator achieves this optimal composite rate $\gamma_{n}$, up to polylogarithmic factors.
+Taken together, Theorems 1 and 2 show that $\gamma_{n}$ is minimax-optimal up to polylogarithmic factors: Theorem 1 gives the upper bound for our estimator, and Theorem 2 gives the lower bound.
 2.2 Neural Network Class
-To approximate the nuisance function $g_{o}(t,\mathbf{X})$, we use the popular class of sparse deep ReLU neural networks (schmidthieber2020nonparametric, and reference therein).
+To approximate the nuisance function $g_{o}(t,\mathbf{X})$, we use the popular class of sparse deep ReLU neural networks (Schmidt-Hieber, 2020, and references therein).
 This choice is motivated by two considerations: (i) ReLU networks are universal approximators capable of representing complex, high-dimensional functions; and (ii) their structure lends itself to sharp theoretical control through entropy bounds and sparsity constraints.
-We define a general $(K+1)$-layer feedforward ReLU network with depth $K\in\mathbb{N}_{+}$ and width vector $\boldsymbol{p}^{\top}=(p_{0},\ldots,p_{K+1})$, where $p_{0}=r$ denotes the input dimension and $p_{K+1}=1$ the scalar output.
+We define a general $(K+1)$-layer feedforward ReLU network with depth $K\in\mathbb{N}_{+}$ and width vector $\boldsymbol{p}^{\top}=(p_{0},\ldots,p_{K+1})$, where $p_{0}=r=d+1$ denotes the input dimension and $p_{K+1}=1$ the scalar output.
 Specifically, the network is defined by
@@
 $\mathcal{G}(K,\boldsymbol{p})\;=\;\Big\{\,g:\ \max_{\ell=0,\ldots,K}\big(\|\mathbf{W}_{\ell}\|_{\infty}\vee\|\mathbf{v}_{\ell}\|_{\infty}\big)\leq 1\,\Big\}\,$,
-which ensures controlled Lipschitz behavior of the class, where for a matrix $\bf A$} and a vector $\bf a$} the entrywise sup–norms is adopted, $\|\mathbf{A}\|_{\infty}=\max_{i,j}|A_{ij}|$, $\|\mathbf{a}\|_{\infty}=\max_{j}|a_{j}|$,.
-Following schmidthieber2020nonparametric we further impose sparsity.
-For sparsity level $s\in\mathbb{N}_{+}$ and range bound $D>0$, the sparse deep ReLU class is define as
+which ensures controlled Lipschitz behavior of the class, where for a matrix $\bf A$} and a vector $\bf a$} the entrywise sup–norms are adopted, $\|\mathbf{A}\|_{\infty}=\max_{i,j}|A_{ij}|$, $\|\mathbf{a}\|_{\infty}=\max_{j}|a_{j}|$.
+Following Schmidt-Hieber (2020) we further impose sparsity.
+For sparsity level $s\in\mathbb{N}_{+}$ and range bound $D>0$, the sparse deep ReLU class is defined as
 $\mathcal{G}(K,s,\boldsymbol{p},D)\;=\;\Big\{g\in\mathcal{G}(K,\boldsymbol{p})\ :\ \sum_{\ell=0}^{K}\big(\|\mathbf{W}_{\ell}\|_{0}+\|\mathbf{v}_{\ell}\|_{0}\big)\leq s,\ \ \|g\|_{\infty}\leq D\Big\},$
@@
 Sparsity serves to control network complexity and is essential for obtaining optimal convergence rates.
-In the subsequent analysis we set the depth, width, and sparsity levels to balance bias-variance tradeoff.
-The bracketing entropy (vdvwellner1996, Chs.
-2–3) of this class is then bounded using results of schmidthieber2020nonparametric, and Dudley’s entropy integral (vdvwellner1996, Thm.
+In the subsequent analysis we set the depth, width, and sparsity levels to balance the bias-variance tradeoff.
+The bracketing entropy (van der Vaart and Wellner, 1996, Chs.
+2–3) of this class is then bounded using results of Schmidt-Hieber (2020), and Dudley’s entropy integral (van der Vaart and Wellner, 1996, Thm.
 2.14.1) is used to establish the required uniform empirical-process bounds.
-3 FLEXI-Haz Procedure
-Estimation is based on maximizing the full likelihood induced by the hazard model (1).
+3 FLEXI–Haz Procedure
+3.1 Estimation of $(\mbox{${\theta}$}_{o},g_{o})$
+Estimation is based on maximizing the full likelihood induced by the hazard model (2.1).
 Under noninformative and conditionally independent censoring given $({\bf X},{\bf Z})$, the corresponding full log–likelihood is proportional to
-$\ell_{n}(\mbox{${\theta}$},g)\;=\;\frac{1}{n}\sum_{i=1}^{n}\Big[\Delta_{i}\big\{g(T_{i},\mathbf{X}_{i})+\mbox{${\theta}$}^{\top}\mathbf{Z}_{i}\big\}-\int_{0}^{\tau}Y_{i}(t)\,\exp\!\big\{g(t,\mathbf{X}_{i})+\mbox{${\theta}$}^{\top}\mathbf{Z}_{i}\big\}\,dt\Big].$ (2)
+$\ell_{n}(\mbox{${\theta}$},g)\;=\;\frac{1}{n}\sum_{i=1}^{n}\Big[\Delta_{i}\big\{g(T_{i},\mathbf{X}_{i})+\mbox{${\theta}$}^{\top}\mathbf{Z}_{i}\big\}-\int_{0}^{\tau}Y_{i}(t)\,\exp\!\big\{g(t,\mathbf{X}_{i})+\mbox{${\theta}$}^{\top}\mathbf{Z}_{i}\big\}\,dt\Big].$ (3.1)
 where $Y_{i}(t)\;=\;\mathbf{1}(T_{i}\geq t)$, $t\in[0,\tau]$, $i=1,\ldots,n$, are at-risk processes, and $\tau$ is a prespecified constant.
-Because $g$ is modeled by a DNN, the predictable cumulative hazard in (2) contains a time-dependent integral that lacks a closed-form expression and must therefore be approximated numerically during optimization.
+Because $g$ is modeled by a DNN, the predictable cumulative hazard in (3.1) contains a time-dependent integral that lacks a closed-form expression and must therefore be approximated numerically during optimization.
 We therefore introduce a grid $0=t_{0}<t_{1}<\cdots<t_{m_{n}}=\tau$ and approximate the integral by the Riemann–type sum
@@
 Defining $\Delta_{ij}=\mathbf{1}\{T_{i}\in(t_{j-1},t_{j}],\Delta_{i}=1\}$, we obtain the discretized likelihood
-$\widetilde{\ell}_{n}(\mbox{${\theta}$},g)\;=\;\frac{1}{n}\sum_{i=1}^{n}\sum_{j:\,Y_{i}(t_{j-1})=1}\Big[\Delta_{ij}\,\big\{g(t_{j},\mathbf{X}_{i})+\mbox{${\theta}$}^{\top}\mathbf{Z}_{i}\big\}-\big\{\min(T_{i},t_{j})-t_{j-1}\big\}\,\exp\!\{g(t_{j},\mathbf{X}_{i})+\mbox{${\theta}$}^{\top}\mathbf{Z}_{i}\}\Big].$ (3)
-In practice, discretizing the time axis on a fixed fine grid and augmenting it with all observed times (event and censoring times) provides an accurate and computationally efficient approximation, as detailed in Section 5.1.
+$\widetilde{\ell}_{n}(\mbox{${\theta}$},g)\;=\;\frac{1}{n}\sum_{i=1}^{n}\sum_{j:\,Y_{i}(t_{j-1})=1}\Big[\Delta_{ij}\,\big\{g(t_{j},\mathbf{X}_{i})+\mbox{${\theta}$}^{\top}\mathbf{Z}_{i}\big\}-\big\{\min(T_{i},t_{j})-t_{j-1}\big\}\,\exp\!\{g(t_{j},\mathbf{X}_{i})+\mbox{${\theta}$}^{\top}\mathbf{Z}_{i}\}\Big].$ (3.2)
+In practice, discretizing the time axis on a fine grid and augmenting it with all observed times (event and censoring times) provides an accurate and computationally efficient approximation, as detailed in Section 5.1.
 This construction automatically produces a dense grid in regions where failures occur, precisely where the integrand contributes most to the likelihood.
 We consider two estimators. The oracle maximum–likelihood estimator is defined as
-$(\widehat{}\mbox{${\theta}$},\widehat{g})\;=\;\arg\max_{\,g\in\mathcal{G}(K,s,\boldsymbol{p},D),\;\theta\in\Theta\subseteq\mathbb{R}^{p}}\;\ell_{n}(\mbox{${\theta}$},g)\,.$ (4)
+$(\widehat{}\mbox{${\theta}$},\widehat{g})\;\in\;\arg\max_{\,g\in\mathcal{G}(K,s,\boldsymbol{p},D),\;\theta\in\Theta\subseteq\mathbb{R}^{p}}\;\ell_{n}(\mbox{${\theta}$},g)\,.$ (3.3)
 The estimator that is actually implemented replaces the at-risk integral with its discretized version $\widetilde{\ell}_{n}$. We therefore define the numerically integrated likelihood estimator as
-$(\widetilde{}\mbox{${\theta}$},\widetilde{g})\;=\;\arg\max_{\,g\in\mathcal{G}(K,s,\boldsymbol{p},D),\;\theta\in\Theta}\;\widetilde{\ell}_{n}(\mbox{${\theta}$},g)\,.$ (5)
-Theorems 1–4 establish the asymptotic properties of (4), while Theorem 5 shows that under the chosen grid these properties carry over to the estimator (5) implemented in practice.
-4 Asymptotic Analysis
-4.1 Notation
+$(\widetilde{}\mbox{${\theta}$},\widetilde{g})\;\in\;\arg\max_{\,g\in\mathcal{G}(K,s,\boldsymbol{p},D),\;\theta\in\Theta\subseteq\mathbb{R}^{p}}\;\widetilde{\ell}_{n}(\mbox{${\theta}$},g)\,.$ (3.4)
+If the maximizer is not unique, owing to the non-uniqueness of neural-network representations, we take an arbitrary element of the argmax set.
+Theorems 1–4 establish the asymptotic properties of (3.3), while Theorem 5 shows that under the chosen grid these properties carry over to the estimator (3.4) implemented in practice.
+3.2 Estimation and Marginal Inference for the Survival of a New Subject
+We now turn to estimation and inference on the survival function of a new subject.
+Let $({\bf X}_{0},{\bf Z}_{0})$ be an independent draw from the covariate distribution, independent of the training sample, and define
+$H_{o}(t\mid{\bf X}_{0},{\bf Z}_{0})=\int_{0}^{t}h_{o}(u\mid{\bf X}_{0},{\bf Z}_{0})\,du\,,\qquad S_{o}(t\mid{\bf X}_{0},{\bf Z}_{0})=\exp\{-H_{o}(t\mid{\bf X}_{0},{\bf Z}_{0})\}\,.$
+[…]
+$\widehat{H}(t\mid{\bf X}_{0},{\bf Z}_{0})=\sum_{j:t_{j}\leq t}\Delta t_{j}\exp\!\bigl\{\widetilde{g}(t_{j},{\bf X}_{0})+\widetilde{\mbox{${\theta}$}}^{\sf T}{\bf Z}_{0}\bigr\}\,$,
+where $\Delta t_{j}=t_{j}-t_{j-1}$, and $\widehat{S}(t\mid{\bf X}_{0},{\bf Z}_{0})=\exp\{-\widehat{H}(t\mid{\bf X}_{0},{\bf Z}_{0})\}$.
+Although the plug-in estimator is natural, it is generally not the most convenient starting point for first-order inference.
+The reason is that the target depends on the infinite-dimensional nuisance function $g_{o}$, and the first-order effect of nuisance estimation error need not vanish automatically.
+In particular, the nuisance-rate theory established in Section 4 controls $\widetilde{g}-g_{o}$ in an integrated norm, which is sufficient for estimation of $\mbox{${\theta}$}_{o}$ but does not by itself guarantee that the plug-in survival estimator admits an asymptotically linear expansion with an $o_{p}(n^{-1/2})$ remainder.
+The purpose of the one-step construction is to remove this first-order nuisance bias (Bickel et al., 1993; Van der Vaart, 2000).
+Concretely, we will augment the plug-in estimator by an empirical average of an estimated efficient influence function (EIF).
+This correction yields an estimator whose leading error term is linear and whose remainder is second-order under the rate conditions imposed below.
+As a consequence, Wald-type inference becomes available for the cumulative hazard and, via the delta method, for the survival function.
+In practice, we implement this correction using cross-fitting (Zheng and van der Laan, 2011; Chernozhukov et al., 2018)..
+The motivation for cross-fitting is twofold.
+First, it separates nuisance estimation from influence-function evaluation, which substantially weakens empirical-process requirements.
+Second, it makes the one-step correction more robust to overfitting in flexible nuisance estimators such as DNN.
+For this reason, we formulate the estimator below in cross-fitted form from the outset.
+Two additional points are worth emphasizing.
+First, the target considered here is marginal in the sense that it is evaluated at an independent draw $({\bf X}_{0},{\bf Z}_{0})$.
+This differs from conditional pointwise inference at a fixed covariate value $({\bf x}_{0},{\bf z}_{0})$, which may be nonregular without additional smoothness or localization assumptions.
+Second, the construction below also applies to the hazard model without a separate linear component, obtained by absorbing $\bf Z$} into the nuisance function $g_{o}$.
+In that case, the parametric score and information terms disappear, and the EIF reduces to its nuisance–Riesz component alone.
+Equivalently, one may replace $\bf X$} throughout by the enlarged covariate vector $({\bf X}^{\sf T},{\bf Z}^{\sf T})^{\sf T}$.
+We first present the estimator at a high level, postponing all technical derivations to Section 5 below.
+To define the one-step estimator, let $\phi_{t}^{0}({\bf W};{\bf X}_{0},{\bf Z}_{0})$ denote the efficient influence function for $H_{o}(t\mid{\bf X}_{0},{\bf Z}_{0})$, where ${\bf W}=(T,\Delta,{\bf X},{\bf Z})$, and let $\widehat{\phi}_{t}^{(-k)}({\bf W};{\bf X}_{0},{\bf Z}_{0})$ be its estimator computed using nuisance estimators trained without fold $k$.
+Let $\{\mathcal{I}_{k}\}_{k=1}^{K}$ be a partition of $\{1,\dots,n\}$ into folds.
+For $i\in\mathcal{I}_{k}$, the observation ${\bf W}_{i}$ is evaluated using nuisance estimators fitted on the complement of $\mathcal{I}_{k}$.
+The cross-fitted one-step estimator of the cumulative hazard is
+$\widehat{H}^{1,\mathrm{cf}}(t\mid{\bf X}_{0},{\bf Z}_{0})=\widehat{H}^{\mathrm{cf}}(t\mid{\bf X}_{0},{\bf Z}_{0})+\frac{1}{n}\sum_{k=1}^{K}\sum_{i\in\mathcal{I}_{k}}\widehat{\phi}_{t}^{(-k)}({\bf W}_{i};{\bf X}_{0},{\bf Z}_{0}),$ (3.5)
+where
+$\widehat{H}^{\mathrm{cf}}(t\mid{\bf X}_{0},{\bf Z}_{0})=\sum_{k=1}^{K}\frac{|\mathcal{I}_{k}|}{n}\widehat{H}^{(-k)}(t\mid{\bf X}_{0},{\bf Z}_{0})$
+and the corresponding one-step estimator of the survival function is
+$\widehat{S}^{1,\mathrm{cf}}(t\mid{\bf X}_{0},{\bf Z}_{0})=\exp\bigl\{-\widehat{H}^{1,\mathrm{cf}}(t\mid{\bf X}_{0},{\bf Z}_{0})\bigr\}.$ (3.6)
+At an intuitive level, (3.5) augments the plug-in estimator by a correction term that removes the first-order effect of nuisance estimation.
+Under the conditions stated below, the resulting estimator is asymptotically linear.
+The practical construction therefore reduces to two ingredients: estimation of the nuisance objects entering the plug-in hazard; and estimation of the efficient influence function $\phi_{t}^{0}(\cdot;{\bf X}_{0},{\bf Z}_{0})$.
+The one-step correction is not shape-constrained.
+Hence, in finite samples, $\widehat{H}^{1,\mathrm{cf}}(t\mid{\bf X}_{0},{\bf Z}_{0})$ need not be monotone nondecreasing in $t$, and $\widehat{S}^{1,\mathrm{cf}}(t\mid{\bf X}_{0},{\bf Z}_{0})=\exp\{-\widehat{H}^{1,\mathrm{cf}}(t\mid{\bf X}_{0},{\bf Z}_{0})\}$ need not be monotone nonincreasing.
+This is not a problem for the pointwise asymptotic inference developed here, which is for each fixed $t$.
+If a monotone curve is desired for display, the estimated survival curve can be post-processed by monotone rearrangement or isotonic projection.
+The next subsection describes the resulting estimator explicitly in the finite discrete-covariate case, where the influence function admits a closed-form representation.
+The explicit construction below focuses on the case where $bX$ has finite support, since in that setting the nuisance Riesz representer admits a closed form.
+A more general high-level formulation for arbitrary covariate distributions is possible, based on an abstract $L^{2}(P_{{\bf X},{\bf Z}})$-valued pathwise differentiability argument and an associated representer condition, but this leads to a less explicit estimator and is therefore deferred to the Appendix.
+3.2.1 Explicit form under finite discrete covariates
+We now specialize to the case where $\bf X$} has finite support $\mathcal{X}=\{{\bf x}^{(1)},\dots,{\bf x}^{(J)}\}$ with ${\Pr}({\bf X}={\bf x}^{(j)})=\pi_{j}>0$, $j=1,\dots,J$.
+In this setting, the nuisance Riesz representer underlying the efficient influence function admits a closed form, and therefore the one-step estimator becomes fully explicit.
+For $u\in[0,\tau]$ and $j=1,\dots,J$, define
+$A_{j}(u)=\mathbb{E}_{0}\!\left\{Y(u)h_{o}(u\mid{\bf X},{\bf Z})\mid{\bf X}={\bf x}^{(j)}\right\}.$
+Then the nuisance part of the efficient influence function is represented by a weight function of the form
+$\psi_{t}^{0}(u,{\bf X};{\bf X}_{0},{\bf Z}_{0})=\mathbf{1}(u\leq t)\mathbf{1}({\bf X}={\bf X}_{0})\frac{h_{o}(u\mid{\bf X}_{0},{\bf Z}_{0})}{\Pr({\bf X}={\bf X}_{0})\,A_{{\bf X}_{0}}(u)}.$ (3.7)
+Accordingly, the efficient influence function takes the form
+$\phi_{t}^{0}({\bf W};{\bf X}_{0},{\bf Z}_{0})=\int_{0}^{t}\mathbf{1}({\bf X}={\bf X}_{0})\frac{h_{o}(u\mid{\bf X}_{0},{\bf Z}_{0})}{\Pr({\bf X}={\bf X}_{0})\,A_{{\bf X}_{0}}(u)}\,dM(u)+c_{0,t}({\bf X}_{0},{\bf Z}_{0})^{\sf T}{\bf I}(\mbox{${\theta}$}_{0})^{-1}\ell_{\theta_{0}}^{\ast}({\bf W}),$ (3.8)
+where $M(t)=N(t)-\int_{0}^{t}Y(u)h_{o}(u\mid{\bf X},{\bf Z})\,du$ is the counting-process martingale under $\mathbb{P}_{0}$ with $N(t)=\Delta\mathbf{1}(T\leq t)$.
+The quantities $c_{0,t}({\bf X}_{0},{\bf Z}_{0})$, $I(\mbox{${\theta}$}_{o})$, and $\ell^{*}_{\theta_{o}}$ are defined formally in Section 5; for now, it suffices to note that they represent the sensitivity term, efficient information, and efficient score, respectively.
+Replacing the unknown quantities in (3.8) by their estimators yields an explicit estimator $\widehat{\phi}_{t}^{(-k)}({\bf W}_{i};{\bf X}_{0},{\bf Z}_{0})$ and therefore an explicit cross-fitted one-step estimator through (3.5)–(3.6).
+This discrete-support case is especially convenient in practice because the Riesz representer is explicit.
+Hence the one-step correction can be implemented directly, without introducing an additional approximation step for the representer.
+Let $\widehat{\sigma}_{H,t}^{2}({\bf X}_{0},{\bf Z}_{0})$ denote a consistent estimator of the asymptotic variance of $\widehat{H}^{1,\mathrm{cf}}(t\mid{\bf X}_{0},{\bf Z}_{0})$, for example the empirical variance of the cross-fitted influence-function estimates is given by
+$\widehat{\sigma}_{H,t}^{2}({\bf X}_{0},{\bf Z}_{0})=\frac{1}{n}\sum_{k=1}^{K}\sum_{i\in\mathcal{I}_{k}}\left\{\widehat{\phi}_{t}^{(-k)}({\bf W}_{i};{\bf X}_{0},{\bf Z}_{0})\right\}^{2}.$ (3.9)
+An asymptotic $(1-\alpha)$ confidence interval for $H_{o}(t\mid{\bf X}_{0},{\bf Z}_{0})$ is therefore
+$\widehat{H}^{1,\mathrm{cf}}(t\mid{\bf X}_{0},{\bf Z}_{0})\pm z_{1-\alpha/2}\frac{\widehat{\sigma}_{H,t}({\bf X}_{0},{\bf Z}_{0})}{\sqrt{n}},$ (3.10)
+and, by the delta method, an asymptotic $(1-\alpha)$ confidence interval for $S_{o}(t\mid{\bf X}_{0},{\bf Z}_{0})$ is
+$\widehat{S}^{1,\mathrm{cf}}(t\mid{\bf X}_{0},{\bf Z}_{0})\pm z_{1-\alpha/2}\frac{\widehat{S}^{1,\mathrm{cf}}(t\mid{\bf X}_{0},{\bf Z}_{0})\widehat{\sigma}_{H,t}({\bf X}_{0},{\bf Z}_{0})}{\sqrt{n}}.$ (3.11)
+The asymptotic validity of these intervals is established in Section 5.
+[…]
+4.1 Assumptions
 Some additional notation will be used throughout.
@@
 For a column vector $\bf a$}, ${\bf a}^{\otimes 2}={\bf a}{\bf a}^{\sf T}$.
-Let $P_{0}$ denote the distribution of a generic observation $(T,\Delta,{\bf X},{\bf Z})$.
-For any measurable function $h=h(T,{\bf X},{\bf Z})$ we define the unweighted $L^{2}_{P_{0}}$ norm $\|h\|^{2}_{L^{2}({P_{0}})}=\mathbb{E}_{0}\{h(T,{\bf X},{\bf Z})^{2}\}$ and the event–weighted norm $\|h\|^{2}_{L^{2}(P_{0},\Delta)}=\mathbb{E}_{0}\{\Delta\,h(T,{\bf X},{\bf Z})^{2}\big\}$.
+Let $\mathbb{P}_{0}$ denote the true distribution of a generic observation $(T,\Delta,{\bf X},{\bf Z})$.
+For any measurable function $\mu=\mu(T,{\bf X},{\bf Z})$, define the unweighted $L^{2}_{\mathbb{P}_{0}}$ norm $\|\mu\|^{2}_{L^{2}({\mathbb{P}_{0}})}=\mathbb{E}_{0}\{\mu(T,{\bf X},{\bf Z})^{2}\}$ and the event–weighted norm $\|\mu\|^{2}_{L^{2}(\mathbb{P}_{0},\Delta)}=\mathbb{E}_{0}\{\Delta\,\mu(T,{\bf X},{\bf Z})^{2}\big\}$.
 Let $\eta=(\mbox{${\theta}$},g)$, write
@@
 Finally, the population loss is defined as
-$\ell_{0}(\eta)\;=\;\mathbb{E}_{0}\Big[\Delta\,\chi_{\eta}(T,{\bf X},{\bf Z})-\int_{0}^{\tau}Y(t)\,\exp\!\big\{\chi_{\eta}(t,{\bf X},{\bf Z})\big\}\,dt\Big]\,.$ (6)
-4.2 Assumptions
+$\ell_{0}(\eta)\;=\;\mathbb{E}_{0}\Big[\Delta\,\chi_{\eta}(T,{\bf X},{\bf Z})-\int_{0}^{\tau}Y(t)\,\exp\!\big\{\chi_{\eta}(t,{\bf X},{\bf Z})\big\}\,dt\Big]\,.$ (4.1)
 A1.
 The failure time $U$ and censoring time $C$ are conditionally independent given $({\bf X}^{\top},{\bf Z}^{\top})$, and the censoring mechanism is noninformative in the sense that the censoring process is noninformative for $(\mbox{${\theta}$},g)$.
-The true nuisance function satisfies $g_{o}\in\mathcal{H}(q,\boldsymbol{\alpha},{\bf d},\tilde{{\bf d}},M)$.
+The true nuisance function satisfies $g_{o}\in\mathcal{H}(q,\boldsymbol{\alpha},{\bf d}^{*},\tilde{{\bf d}},M)$.
 A2.
-The covariate vector $({\bf X},\mathbf{Z})$ takes values in a bounded subset of $\mathbb{R}^{p+q}$, denoted $\mathcal{X}\times\mathcal{Z}$, where both sets are compact.
+The covariate vector $({\bf X},\mathbf{Z})$ takes values in a bounded subset of $\mathbb{R}^{d+p}$, denoted $\mathcal{X}\times\mathcal{Z}$, where both sets are compact.
 Without loss of generality, assume $\mathcal{X}\subset[0,1]^{d}$ and that $\|\mathbf{Z}\|_{2}\leq B_{{\bf Z}}$ almost surely.
@@
 A3.
-The network architecture satisfies $K=O(\log n)$, $s=O\!\big(n\gamma_{n}^{2}\log n\big)$, and $n\gamma_{n}^{2}\;\min_{0\leq\ell\leq K}p_{\ell}\lesssim\max_{0\leq\ell\leq K}p_{\ell}\lesssim n$.
+The network architecture satisfies $K=O(\log n)$, $s=O\!\big(n\gamma_{n}^{2}\log n\big)$, and $n\gamma_{n}^{2}\lesssim\;\min_{0\leq\ell\leq K}p_{\ell}\lesssim\max_{0\leq\ell\leq K}p_{\ell}\lesssim n$.
 In addition, the sieve is uniformly bounded, $\|g\|_{\infty}\leq D_{n}$ for a known sieve radius $D_{n}=O(1)$.
 A4.
-There exist $\delta_{1},\delta_{2}\in(0,1]$ such that, almost surely, $\Pr(\Delta=1\mid T,{\bf X},{\bf Z})\geq\delta_{1}$ and $\Pr(U\geq\tau\mid{\bf X},{\bf Z})\geq\delta_{2}$.
+There exist $\delta_{1},\delta_{2},\delta_{3}\in(0,1]$ such that, almost surely, $\Pr(\Delta=1\mid T,{\bf X},{\bf Z})\geq\delta_{1}$, $\Pr(U\geq\tau\mid{\bf X},{\bf Z})\geq\delta_{2}$, and $\Pr(Y(t)=1\mid{\bf X},{\bf Z})\geq\delta_{3}$ for all $t\in[0,\tau]$.
 A5.
-The covariate vector $\mathbf{Z}$ does not include a constant (intercept) term. Furthermore, the residual covariance matrix of $\mathbf{Z}$ given $\mathbf{X}$ is positive definite, that is,
-$\Sigma_{\mathbf{Z}}=\mathbb{E}\!\left[\big\{\mathbf{Z}-\mathbb{E}(\mathbf{Z}\mid\mathbf{X})\big\}\big\{\mathbf{Z}-\mathbb{E}(\mathbf{Z}\mid\mathbf{X})\big\}^{\top}\right]\succ 0\,.$
-Assumptions A1A3, A5 used also in zhong2022deep under the partial linear Cox regression model.
-A3 determines the structure of the neural network and provides a trade-off between the approximation error and estimation error (schmidthieber2020nonparametric).
+The covariate vector $\mathbf{Z}$ does not include a constant (intercept) term.
+Furthermore, the residual covariance matrix of $\mathbf{Z}$ given $\mathbf{T,X}$ among failures, is positive definite, that is,
+$\Sigma_{\mathbf{Z}}=\mathbb{E}\!\left[\big\{\mathbf{Z}-\mathbb{E}(\mathbf{Z}\mid T,\mathbf{X},\Delta=1)\big\}\big\{\mathbf{Z}-\mathbb{E}(\mathbf{Z}\mid T,\mathbf{X},\Delta=1)\big\}^{\top}\right]\succ 0\,.$
+Assumptions A1A3 are also used in Zhong et al.
+(2022) under the partial linear PH regression model.
+A3 determines the structure of the neural network and provides a trade-off between the approximation error and estimation error (Schmidt-Hieber, 2020).
 Assumption A2 is a standard assumption for semi- or non-parametric regression.
 Moreover, A1 and A2 imply that there exist constants $0<c<C<\infty$ (that depend on $D_{n},B_{{\bf Z}}$ and $M_{\Theta}$) such that for all $t\in(0,\tau]$ and $({\bf X},{\bf Z})\in\mathcal{X}\times\mathcal{Z}$, $c_{h}\leq h_{o}(t|{\bf X},{\bf Z})\leq C_{h}$.
-Assumption A5 is a standard idenfiability assumption in survival analysis.
-Assumption A4 is required since in the hazard function (1) time is one of the components in the function $g_{o}$ estimated by the DNN and therefore the effects of the covariates $\bf X$}} and time are not disatengled, in contrast to the partial linear Cox model of zhong2022deep.
-Hence we need a non-vanishing chance to observe events over the entire study horizon for each covariate pattern.
+Assumption A4 is required since in the hazard function (2.1) time is one of the components in the function $g_{o}$ estimated by the DNN and therefore the effects of the covariates $\bf X$}} and time are not disentangled, in contrast to the partial linear Cox model of Zhong et al.
+(2022).
+Hence we need a non-vanishing chance to observe events over the entire study horizon for all $({\bf x},{\bf Z})$ in the support.
+While this assumption is stronger than in standard Cox regression, it is natural in baseline-free hazard models and is comparable to positivity conditions commonly imposed in semiparametric and causal models with continuous time.
+Assumption A5 is the non-PH analogue of the standard partial-linear identifiability condition.
+Because the nuisance component is allowed to depend on $(t,{\bf X})$, the relevant residual variation for the linear covariates must be measured after conditioning on $(T,{\bf X})$ among failures, rather than on $\bf X$} alone.
 4.2 Asymptotic Results
-The proofs of all lemmas and theorems in this section appear in the appendix.
+The proofs of all lemmas and theorems in this section appear in Appendix A.
 We begin by establishing several basic properties of the model.
-A key step is to understand how the counting–process framework allows us to translate expressions involving at-risk process into simpler quantities expressed directly in terms of the event indicator $\Delta$.
+A key step is to understand how the counting–process framework allows us to translate expressions involving the at-risk processes into simpler quantities expressed directly in terms of the event indicator $\Delta$.
 The next lemma formalizes this relationship, showing that these two representations are equivalent up to a multiplicative constant.
@@
 Lemma 1.
-Under Assumptions A1A2, for any nonnegative, predictable function $f(t,{\bf X},{\bf Z})$ with $\mathbb{E}\!\left[\int_{0}^{\tau}Y(t)\,f(t,{\bf X},{\bf Z})\,dt\right]<\infty$,
+Under A1A2, for any nonnegative, predictable function $f(t,{\bf X},{\bf Z})$ with $\mathbb{E}\!\left[\int_{0}^{\tau}Y(t)\,f(t,{\bf X},{\bf Z})\,dt\right]<\infty$,
 $\mathbb{E}\!\left\{\int_{0}^{\tau}Y(t)\,f(t,{\bf X},{\bf Z})\,dt\right\}\;\asymp\;\mathbb{E}\!\left\{\Delta\,f(T,{\bf X},{\bf Z})\right\}.$
@@
 Under Assumptions A1, A2 and A4, there exist finite constants $0<c_{1}\leq c_{2}<\infty$ and $0<c_{3}\leq c_{4}<\infty$, depending only on $(\tau,\delta_{1},\delta_{2},c_{h},C_{h})$, such that
-$c_{1}\,\mathbb{E}\!\big\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\big\}$ $\;\leq\;\mathbb{E}\!\big\{\phi(T,{\bf X},{\bf Z})^{2}\big\}\;\leq\;c_{2}\,\mathbb{E}\!\big\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\big\}\,$, (7)
-$c_{3}\,\mathbb{E}\!\big\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\big\}$ $\;\leq\;\mathbb{E}_{{\bf X},{\bf Z}}\!\Big\{\int_{0}^{\tau}\phi(t,{\bf X},{\bf Z})^{2}\,dt\Big\}\;\leq\;c_{4}\,\mathbb{E}\!\big\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\big\}\,.$ (8)
-In particular, if $\mathbb{E}\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\}=0$, then $\phi(T,{\bf X},{\bf Z})=0$ $P_{0}$–a.s.
+$c_{1}\,\mathbb{E}\!\big\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\big\}$ $\;\leq\;\mathbb{E}\!\big\{\phi(T,{\bf X},{\bf Z})^{2}\big\}\;\leq\;c_{2}\,\mathbb{E}\!\big\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\big\}\,$, (4.2)
+$c_{3}\,\mathbb{E}\!\big\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\big\}$ $\;\leq\;\mathbb{E}_{{\bf X},{\bf Z}}\!\Big\{\int_{0}^{\tau}\phi(t,{\bf X},{\bf Z})^{2}\,dt\Big\}\;\leq\;c_{4}\,\mathbb{E}\!\big\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\big\}\,.$ (4.3)
+In particular, if $\mathbb{E}\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\}=0$, then $\phi(T,{\bf X},{\bf Z})=0$ $\mathbb{P}_{0}$–a.s.
 and $\phi(t,{\bf X},{\bf Z})=0$ for Lebesgue-a.e.
-$t\in[0,\tau]$ and $P_{{\bf X},{\bf Z}}$–a.e.
+$t\in[0,\tau]$ and $\mathbb{P}_{{\bf X},{\bf Z}}$–a.e.
 $({\bf X},{\bf Z})$, i.e.
-$\chi_{\eta_{1}}(t,{\bf X},{\bf Z})=\chi_{\eta_{2}}(t,{\bf X},{\bf Z})$ for $dt\otimes P_{{\bf X},{\bf Z}}$–almost all $(t,{\bf X},{\bf Z})$.
-Lemma 2 shows that the distance $d(\cdot,\cdot)$ is a valid metric on the parameter space: $d(\eta_{1},\eta_{2})=0$ if and only if the induced linear score coincide almost everywhere.
+$\chi_{\eta_{1}}(t,{\bf X},{\bf Z})=\chi_{\eta_{2}}(t,{\bf X},{\bf Z})$ for $dt\otimes\mathbb{P}_{{\bf X},{\bf Z}}$–almost all $(t,{\bf X},{\bf Z})$.
+Lemma 2 shows that the distance $d(\cdot,\cdot)$ is a valid metric on the parameter space: $d(\eta_{1},\eta_{2})=0$ if and only if the induced linear score coincides almost everywhere.
 In addition, the lemma shows that the event-weighted norm and the corresponding unweighted norms are equivalent up to multiplicative constants.
-As a result, any bound or convergence statement expressed in terms of the metric $d$ can be translated into an equivalent statement under the usual $L^{2}$ norm and vice versa.
+As a result, any bound or convergence statement expressed in terms of the metric $d$ can be translated into an equivalent statement under the unweighted $L^{2}(dt\otimes P_{{\bf X},{\bf Z}})$ norm and vice versa.
 This equivalence is used repeatedly throughout the theoretical development.
-Next we study the shape of the population loss around $\eta_{o}=(g_{o},\mbox{${\theta}$}_{o})$.
+Next we study the shape of the population loss around $\eta_{o}=(\mbox{${\theta}$}_{o},g_{o})$.
 Lemma 3.
-Assume A1A2. For any $\eta\in\Theta\times\mathcal{H}(q,\boldsymbol{\alpha},{\bf d},\tilde{{\bf d}},M)$,
+Assume A1A2. For any $\eta\in\Theta\times\mathcal{H}(q,\boldsymbol{\alpha},{\bf d}^{*},\tilde{{\bf d}},M)$,
 $\ell_{0}(\eta)-\ell_{0}(\eta_{o})\;\asymp\;-\,d(\eta,\eta_{o})^{2}.$
@@
 We begin by quantifying how well the estimator $\widehat{g}$ recovers the unknown function $g_{o}$.
-The next theorem gives a nonparametric convergence rate for $\widehat{g}$ in the event–weighted $L^{2}_{P_{0}}$ norm that naturally appears in the loss and in the distance function $d$.
+The next theorem gives a nonparametric convergence rate for $\widehat{g}$ in the event–weighted $L^{2}_{\mathbb{P}_{0}}$ norm that naturally appears in the loss and in the distance function $d(\eta,\eta_{o})$.
 The rate is expressed in terms of the complexity parameter $\gamma_{n}$ and will later be used to control remainder terms in the expansion of the efficient score.
@@
 […]
-$\left(\mathbb{E}\!\Bigl[\Delta\bigl\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\bigr\}^{2}\Bigr]\right)^{1/2}=O_{p}(\tau_{n})$ (9)
-where the expectation is taken with respect to an independent draw $(T,\Delta,{{\bf X}},{{\bf Z}})\sim P_{0}$ from the true distribution, and $\tau_{n}=\gamma_{n}\log^{2}n$.
-The quantity in (9) is random through its dependence on the training sample used to construct $\widehat{g}$.
-The conclusion of Theorem 1 is strengthened by Lemma 2, which shows that $d(\eta,\eta_{0})^{2}$ is equivalent to the unweighted norm $\|\chi_{\eta}-\chi_{\eta_{0}}\|^{2}_{L^{2}(dt\otimes P_{\mathbf{X},\mathbf{Z}})}$.
+$\left(\mathbb{E}\!\Bigl[\Delta\bigl\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\bigr\}^{2}\Bigr]\right)^{1/2}=O_{p}(\tau_{n})$ (4.4)
+where the expectation is taken with respect to an independent draw $(T,\Delta,{{\bf X}},{{\bf Z}})\sim\mathbb{P}_{0}$ from the true distribution, and $\tau_{n}=\gamma_{n}\log^{2}n$.
+The quantity in (4.4) is random through its dependence on the training sample used to construct $\widehat{g}$.
+The conclusion of Theorem 1 is strengthened by Lemma 2, which shows that $d(\eta,\eta_{o})^{2}$ is equivalent to the unweighted norm $\|\chi_{\eta}-\chi_{\eta_{0}}\|^{2}_{L^{2}(dt\otimes\mathbb{P}_{\mathbf{X},\mathbf{Z}})}$.
 Consequently, the estimator converges in $L^{2}$ over the entire time domain $(0,\tau)$ with respect to Lebesgue measure, and not merely at observed event times or in a way that depends on the censoring pattern.
+In Theorem 1, both Zhong et al.
+(2022) and our work rely on Schmidt-Hieber (2020) exclusively for deterministic approximation results of sparse ReLU neural networks over composite Hölder function classes, including the network size and sparsity scaling required to achieve optimal approximation rates.
+These results are purely analytic and do not involve any distributional assumptions.
+All stochastic components of the proof, such as control of the estimation error, empirical process bounds, and derivation of convergence rates, are fundamentally different and are derived from the survival-analysis framework.
+In Zhong et al.
+(2022), these arguments rely on the Cox partial likelihood and counting-process martingale theory, while in our setting they are based on the full likelihood and properties of the hazard function.
+Consequently, unlike Schmidt-Hieber (2020), no assumptions on an additive error distribution, including normality, are required or used at any stage of the analysis.
 The next result shows that the rate obtained in Theorem 1 is essentially optimal, up to logarithmic factors.
-It establishes a minimax lower bound for estimation of $g_{o}$ under the same event–weighted $L^{2}_{P_{0}}$ loss, taken uniformly over the parameter space $\Theta$ and the nuisance function class $\mathcal{H}(q,\boldsymbol{\alpha},{\bf d},\tilde{{\bf d}},M)$.
+It establishes a minimax lower bound for estimation of $g_{o}$ under the same event–weighted $L^{2}_{\mathbb{P}_{0}}$ loss, taken uniformly over the parameter space $\Theta$ and the nuisance function class $\mathcal{H}(q,\boldsymbol{\alpha},{\bf d}^{*},\tilde{{\bf d}},M)$.
 Theorem 2.
-Under A1A2 and A5 there exists $c>0$ such that
-$\inf_{\widehat{g}}\sup_{\theta_{o}\in\Theta,\,g_{o}\in\mathcal{H}(q,\boldsymbol{\alpha},{\bf d},\tilde{{\bf d}},M)}\mathbb{E}_{P_{g_{o},\theta_{o}}}\!\left[\Delta\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\}^{2}\right]\;\;\geq\;c\,\gamma_{n}^{2}\,.$
-Next, we turn to the parametric component $\mbox{${\theta}$}_{o}$ and characterize the efficient score and information bound in our semiparametric model.
+Under A1A2 and A4 there exists $c>0$ such that
+$\inf_{\widehat{g}}\sup_{\theta_{o}\in\Theta,\,g_{o}\in\mathcal{H}(q,\boldsymbol{\alpha},{\bf d}^{*},\tilde{{\bf d}},M)}\mathbb{E}_{\mathbb{P}_{g_{o},\theta_{o}}}\!\left[\Delta\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\}^{2}\right]\;\;\geq\;c\,\gamma_{n}^{2}\,.$
+We now shift from nuisance-rate optimality to semiparametric efficiency for the finite-dimensional parameter $\mbox{${\theta}$}_{o}$ and characterize the efficient score and information bound in our semiparametric model.
 To characterize the nuisance tangent space associated with $g_{o}$, we consider differentiable one-dimensional submodels that pass through $g_{o}$.
-Let $\mathcal{H}_{g_{o}}$ denote the collection of all functions $g\in L^{2}([0,\tau]\times[0,1]^{r})$ for which there exist a submodel $\{g_{b}:b\in(-1,1)\}\subset\mathcal{H}(q,\boldsymbol{\alpha},{\bf d},\tilde{{\bf d}},M)$ satisfying
+Let $\mathcal{H}_{g_{o}}$ denote the collection of all functions $g\in L^{2}([0,\tau]\times[0,1]^{r})$ for which there exist a submodel $\{g_{b}:b\in(-1,1)\}\subset\mathcal{H}(q,\boldsymbol{\alpha},{\bf d}^{*},\tilde{{\bf d}},M)$ satisfying
 $\lim_{b\to 0}\left\|b^{-1}(g_{b}-g_{o})-g\right\|_{L^{2}([0,\tau]\times[0,1]^{r})}=0\,.$
 Here, the scalar $b$ is a perturbation parameter indexing the submodel. The set of all attainable first–order directions at $g_{o}$ is then defined as
-$\mathcal{T}_{g_{o}}=\Big\{g\in L^{2}([0,\tau]\times[0,1]^{r}):\exists\,\{g_{b}\}\in\mathcal{H}_{g_{o}}\text{ s.t. }\lim_{b\to 0}\left\|b^{-1}(g_{b}-g_{o})-g\right\|=0\Big\}\,.$
-Finally, semiparametric efficiency theory requires a closed linear subspace, so we take the closure of the linear span of $\mathcal{T}{g_{o}}$, denoted $\overline{\mathcal{T}}{g_{o}}$, which forms the nuisance tangent space.
+$\mathcal{T}_{g_{o}}=\Big\{g\in L^{2}([0,\tau]\times[0,1]^{r}):\exists\,\{g_{b}\}\in\mathcal{H}_{g_{o}}\text{ s.t.
+}\lim_{b\to 0}\left\|b^{-1}(g_{b}-g_{o})-g\right\|_{L^{2}([0,\tau]\times[0,1]^{r})}=0\Big\}\,.$
 […]
-$\mathbb{E}\!\left\{\Delta\,\|{\bf Z}-{\bf g}^{*}(T,{\bf X})\|_{c}^{2}\right\},\quad\|v\|_{c}^{2}=(v_{1}^{2},\ldots,v_{p}^{2})^{\top}\,$,
-so ${{\bf g}}^{*}(t,{\bf X})$ projects the covariates $\bf Z$} onto the tangent space of the nuisance function.
-Finally, under the counting-process representation of the model (andersen1982cox), the counting process $N(t)=\Delta I(T\leq t)$ admits the Doob-Meyer decomposition where the associated martingale is
+$g_{j}^{*}\in\arg\min_{g\in\overline{\mathcal{T}}_{g_{o}}}\mathbb{E}\!\left[\Delta\{Z_{j}-g(T,{\bf X})\}^{2}\right],\qquad j=1,\ldots,p\,.$
+Equivalently,
+${\bf g}^{*}\in\arg\min_{{\bf g}\in(\overline{\mathcal{T}}_{g_{o}})^{p}}\mathbb{E}\!\left[\Delta\|{\bf Z}-{\bf g}(T,{\bf X})\|_{2}^{2}\right].$
+Since $(\overline{\mathcal{T}}_{g_{o}})^{p}$ is a Cartesian product and the criterion separates across coordinates, this minimization is componentwise.
+Thus, ${\bf g}^{*}(t,{\bf X})$ is the $\Delta$-weighted $L_{2}$-projection of $\bf Z$} onto $(\overline{\mathcal{T}}_{g_{o}})^{p}$.
+Finally, under the counting-process representation of the model (Andersen and Gill, 1982), the counting process $N(t)=\Delta I(T\leq t)$ has compensator $\int_{0}^{t}Y(s)\exp\{\mbox{${\theta}$}_{o}^{\top}{\bf Z}+g_{o}(s,{\bf X})\}\,ds$ so the associated martingale is
 $M(t)=N(t)-\int_{0}^{t}Y(s)\exp\{\mbox{${\theta}$}_{o}^{\sf T}{{\bf Z}}+g_{o}(s,{{\bf X}})\}\,ds\,.$
 Theorem 3.
-Under Assumptions A1A2 and A5, the efficient score for $\mbox{${\theta}$}_{o}$ in the semiparametric model (1) is
+Under Assumptions A1A5, the efficient score for $\mbox{${\theta}$}_{o}$ in the semiparametric model (2.1) is
 $\ell_{\theta_{o}}^{*}(T,{\bf X},{\bf Z},\Delta)\;=\;\int_{0}^{\tau}\big\{{\bf Z}-{\bf g}^{*}(t,{\bf X})\big\}\,dM(t)\,$,
-where ${\bf g}^{*}(t,{\bf X})$ is the projection of $\bf Z$} onto the nuisance tangent space and $M(t)$ is the associated counting–process martingale.
+where ${\bf g}^{*}(t,{\bf X})$ is the $\Delta$-weighted $L_{2}$-projection of $\bf Z$} onto $(\overline{\mathcal{T}}_{g_{o}})^{p}$, and $M(t)$ is the associated counting-process martingale.
 The semiparametric information bound for $\mbox{${\theta}$}_{0}$ is
@@
 The key technical observation is that, in the Taylor expansion of the population score around $g_{o}$, the first–order derivative in $\widehat{g}-g_{o}$ direction vanishes because the efficient score is orthogonal to the nuisance tangent space.
-The remaining second–order term is of smaller order, specifically $o_{p}(n^{-1/2})$, provided that $n\gamma_{n}^{4}\to 0$ as $n\to\infty$.
+The remaining second–order term is $o_{p}(n^{-1/2})$ provided that $n\gamma_{n}^{4}\to 0$ as $n\to\infty$.
 Finally, we justify the numerical approximation used to compute the estimator in practice.
@@
 […]
-$\delta_{n}\;=\;\max_{1\leq j\leq m_{n}}(t_{j}-t_{j-1})$
-denote the mesh size of this grid. Assume that $\delta_{n}=o(\tau_{n}^{2})$ and $\sqrt{n}\,\delta_{n}\;\rightarrow\;0$, as $n\rightarrow\infty$, and that the conditions of Theorems 1 and 4 hold. Then,
+$\omega_{n}(\delta)=\sup_{\eta\in\mathcal{N}_{n}}\sup_{{\bf X}\in\mathcal{X}}\sup_{\begin{subarray}{c}s,t\in[0,\tau]\\ |t-s|\leq\delta\end{subarray}}|g(t,{\bf X})-g(s,{\bf X})|\,,\qquad\mathcal{N}_{n}=\{\eta:d(\eta,\eta_{o})\leq M\tau_{n}\}\,$,
+for some fixed $M>0$. Assume that $\omega_{n}(\delta_{n})=o(\tau_{n}^{2})$, $\sqrt{n}\,\omega_{n}(\delta_{n})\to 0$, and that the conditions of Theorems 1 and 4 hold. Then,
   1. 1.
@@
 Consequently,
-[…]
+$\sqrt{n}(\widetilde{}\mbox{${\theta}$}-\mbox{${\theta}$}_{o})=\sqrt{n}(\widehat{}\mbox{${\theta}$}-\mbox{${\theta}$}_{o})+o_{p}(1)\;\xrightarrow{\mathcal{D}}\;N\bigl(\mathbf{0},\,I(\mbox{${\theta}$}_{o})^{-1}\bigr)\,$,
 where $I(\mbox{${\theta}$}_{o})$ is the efficient information matrix from Theorem 3.
 Theorem 5 shows that the difference between the exact and the numerically integrated likelihood estimators is $o_{p}(n^{-1/2})$, so that both share the same asymptotic normal distribution.
-Part 2 verifies that the required error bounds hold for a simple equally spaced grid with $n$ time points, where the per–observation quadrature error is of order $1/n$.
+The second conclusion verifies that the required error bounds hold for the grid with $n$ time points, where the per–observation quadrature error is of order $1/n$.
 This ensures that the asymptotic efficiency result for $\widehat{}\mbox{${\theta}$}$ carries over directly to the implementable estimator $\widetilde{}\mbox{${\theta}$}$ based on numerical integration.
-5 Pratical Implementation
+5 Semiparametric Representation and Asymptotic Theory for $\widehat{S}^{1,\sf cf}(t\mid{\bf X}_{0},{\bf Z}_{0})$
+After Section 4 established estimation and efficiency results for the linear parameter $\mbox{${\theta}$}_{o}$, this section turns to the harder problem of inference for $H_{o}(t\mid{\bf X}_{0},{\bf Z}_{0})$ and $S_{o}(t\mid{\bf X}_{0},{\bf Z}_{0})$.
+Our development follows the semiparametric approach based on pathwise differentiability and efficient influence functions (Bickel et al., 1993; Van der Vaart, 2000; Tsiatis, 2006).
+We first introduce the additional assumptions needed for this target, then derive the efficient influence function and the corresponding cross-fitted one-step estimator.
+We next show how this construction simplifies in the finite-support case for $\bf X$}, where the Riesz representer becomes explicit, and finally use these results to obtain asymptotic normality and pointwise confidence intervals for the cumulative hazard and survival function in the finite-support case for $\bf X$}.
+The proofs of all lemmas and theorems in this section appear in Appendix B.
+5.1 Additional assumptions for survival inference
+Fix $t\in[0,\tau]$, and let $({\bf X}_{0},{\bf Z}_{0})$ be an independent draw from the covariate distribution, independent of the training sample.
+The assumptions below are used to establish pathwise differentiability of the target, derive its efficient influence function, and justify the cross-fitted one-step estimator introduced in Section 3.2.
+To state these assumptions, we equip nuisance directions with the predictable-variation inner product
+$\langle a,b\rangle_{0}=\mathbb{E}_{0}\!\left\{\int_{0}^{\tau}a(u,{\bf X})b(u,{\bf X})\,Y(u)h_{o}(u\mid{\bf X},{\bf Z})\,du\right\}\,,\qquad\|b\|_{0}^{2}=\langle b,b\rangle_{0}\,.$ (5.1)
+A6.
+For the fixed time $t$, $\mathbb{E}_{0}\!\int_{0}^{t}h_{o}(u\mid{\bf X}_{0},{\bf Z}_{0})\,du<\infty$, and there exists $\kappa_{t}>0$ such that $\mathbb{P}_{0}(Y(u)=1\mid{\bf X}_{0},{\bf Z}_{0})\geq\kappa_{t}$ a.s.
+for a.e.
+$u\in[0,t]$.
+A7.
+Define the linear operator $Q:\overline{\mathcal{T}}_{g_{o}}\to L^{2}(P_{{\bf X},{\bf Z}})$,
+$(Qb)({\bf X}_{0},{\bf Z}_{0})=\int_{0}^{t}h_{o}(u\mid{\bf X}_{0},{\bf Z}_{0})\,b(u,{\bf X}_{0})\,du,\qquad b\in\overline{\mathcal{T}}_{g_{o}}\,.$
+Assume that $Q$ is bounded, i.e., there exists a constant $C_{t}<\infty$ such that
+$\|Qb\|_{L^{2}(P_{{\bf X},{\bf Z}})}^{2}=E_{{\bf X}_{0},{\bf Z}_{0}}\!\left\{(Qb)({\bf X}_{0},{\bf Z}_{0})^{2}\right\}\leq C_{t}\|b\|_{0}^{2},\qquad\forall\,b\in\overline{\mathcal{T}}_{g_{o}}.$
+Assume further that there exists a random element $\psi_{t}^{0}(\cdot,\cdot;{\bf X}_{0},{\bf Z}_{0})\in L^{2}(P_{{\bf X},{\bf Z}};\overline{\mathcal{T}}_{g_{o}})$, such that for every $b\in\overline{\mathcal{T}}_{g_{o}}$,
+$(Qb)({\bf X}_{0},{\bf Z}_{0})=\left\langle\psi_{t}^{0}(\cdot,\cdot;{\bf X}_{0},{\bf Z}_{0}),\,b\right\rangle_{0}\quad\text{in }L^{2}(P_{{\bf X},{\bf Z}}).$ (5.2)
+A8.
+The nuisance covariate $\bf X$} takes values in a finite set $\mathcal{X}=\{{\bf x}^{(1)},\dots,{\bf x}^{(J)}\}$, with $\Pr({\bf X}={\bf x}^{(j)})=\pi_{j}>0$, $j=1,\dots,J$.
+When pointwise inference is stated in explicit closed form below, we work under this finite-support assumption.
+A9.
+Let $\{\mathcal{I}_{k}\}_{k=1}^{K}$ be a fixed K-fold partition of $\{1,\dots,n\}$, and for each $k$ let $(\widetilde{}\mbox{${\theta}$}^{(-k)},\widetilde{g}^{(-k)})$ and $\widehat{\bf g}^{*,(-k)}$ be estimated using only the training sample $\mathcal{I}_{k}^{c}$.
+Assume these estimators are measurable with respect to the training sigma-field and, conditional on the training sample, are independent of the validation observations $\{{\bf W}_{i}:i\in\mathcal{I}_{k}\}$.
+Assume moreover that, uniformly in $k$, $\|\widetilde{g}^{(-k)}-g_{o}\|_{0}=O_{p}(\tau_{n})$, $\|\widehat{\bf g}^{*,(-k)}-{\bf g}^{*}\|_{0}=O_{p}(\tau_{n})$, $\|\widetilde{}\mbox{${\theta}$}^{(-k)}-\mbox{${\theta}$}_{o}\|=O_{p}(n^{-1/2})$, where $\tau_{n}=\gamma_{n}\log^{2}n$, and $\sqrt{n}\,\tau_{n}^{2}\to 0$.
+A10.
+The efficient information matrix $I(\mbox{${\theta}$}_{o})$ is nonsingular and the smallest eigenvalue satisfies $\lambda_{min}{I(\theta_{o})}\geq c_{0}$ for some constant $c_{0}>0$.
+In addition, $E_{0}\bigl\|\ell^{*}_{\theta_{o}}({\bf W})\bigr\|^{2}<\infty$, and $E_{0}\!\left[\bigl\{\phi^{0}_{t}({\bf W};{\bf X}_{0},{\bf Z}_{0})\bigr\}^{2}\right]<\infty$.
+A11.
+Let $0=t_{0}<t_{1}<\cdots<t_{m_{n}}=\tau$ be the numerical grid used in the estimation procedure, with mesh size $\delta_{n}\;=\;\max_{1\leq j\leq m_{n}}(t_{j}-t_{j-1})$.
+Assume the grid contains all observed event and censoring times, that $\sqrt{n}\,\delta_{n}\to 0$, and that the corresponding Riemann-approximation error is $O(\delta_{n})$ uniformly over a neighborhood of $(\mbox{${\theta}$}_{o},g_{o})$.
+Under A6A11, the efficient influence function and the cross-fitted one-step estimator admit a tractable representation.
+Assumptions A6A7 guarantee pathwise differentiability of the target functional $H_{o}(t\mid{\bf X}_{0},{\bf Z}_{0})$, A8 yields an explicit closed-form Riesz representer, A9 provides the foldwise nuisance-rate conditions needed for replacing oracle quantities by estimated ones, A10 ensures regularity of the efficient score and information, and A11 controls the numerical integration error.
+Under A6 and A8, the operator $Q$ admits the explicit Riesz representer given in Lemma 6; hence A7 holds in the finite-support case.
+Assumption A8 is used only to obtain an explicit closed-form Riesz representer, and hence an explicit one-step correction, for pointwise survival-curve inference.
+It is not required for estimation of $g_{o}$, for semiparametric efficiency of $\mbox{${\theta}$}_{o}$, or for inference on $\mbox{${\theta}$}_{o}$.
+Moreover, A8 concerns only the nuisance covariates entering the flexible component $g(t,{\bf X})$, not the low-dimensional covariates $\bf Z$} of primary scientific interest.
+In applications with continuous nuisance covariates, one practical route is to replace them by a sufficiently fine grouped version when implementing the explicit one-step correction.
+Because this grouping acts only on the nuisance inputs of a flexible DNN, it is often not expected to materially affect the fitted hazard or resulting survival predictions, although this should be assessed empirically and may become less convenient when the induced joint support is very large.
+5.2 Efficient influence function and one-step representation
+We now provide the semiparametric ingredients underlying the estimator introduced in Section 3.2.
+Consider nuisance perturbations of the form $g_{\varepsilon}=g_{o}+\varepsilon b$, with predictable $b=b(u,{\bf X})$.
+The corresponding nuisance score equals
+$S_{g,b}({\bf W})=\left.\frac{\partial}{\partial\varepsilon}\ell_{0}(\mbox{${\theta}$}_{o},g_{\varepsilon})\right|_{\varepsilon=0}=\int_{0}^{\tau}b(u,{\bf X})\,dM(u)\,.$ (5.3)
+Since $\overline{\mathcal{T}}_{g_{o}}$ denote the closure of predictable directions under $\|\cdot\|_{0}$, elements of $\overline{\mathcal{T}}_{g_{o}}$ are effectively functions on $0,\tau]\times\mathcal{X}$.
+As earlier, ${\bf g}^{*}(u,X)=(g_{1}^{*},\dots,g_{p}^{*})^{\top}\in(\overline{\mathcal{T}}_{g_{o}})^{p}$ denote the componentwise $L_{0}^{2}$-projection of $\bf Z$} onto $(\overline{\mathcal{T}}_{g_{o}})^{p}$, characterized by
+$\mathbb{E}_{0}\!\left\{\int_{0}^{\tau}\{{\bf Z}-{\bf g}^{*}(u,{\bf X})\}\,b(u,{\bf X})\,Y(u)h_{o}(u\mid{\bf X},{\bf Z})\,du\right\}=0\,,\qquad\forall\,\,b\in\mathcal{\overline{}}{\mathcal{T}}_{g_{o}}\,.$ (5.4)
+The efficient score for $\mbox{${\theta}$}_{o}$ is therefore
+$\ell_{\theta_{o}}^{*}({\bf W})=\int_{0}^{\tau}\{{\bf Z}-{\bf g}^{*}(u,{\bf X})\}\,dM(u)\,$,
+and $I(\mbox{${\theta}$}_{o})=\mathbb{E}_{0}\!\left\{\ell_{\theta_{o}}^{*}({\bf W})^{\otimes 2}\right\}$.
+For ${\bf a}\in\mathbb{R}^{p}$, consider the parametric path $\mbox{${\theta}$}_{\varepsilon}=\mbox{${\theta}$}_{o}+\varepsilon{\bf a}$.
+To construct a least favorable submodel, we simultaneously perturb the nuisance function as
+$g_{\varepsilon,{\bf a}}(u,{\bf X})=g_{o}(u,{\bf X})-\varepsilon\,{\bf a}^{\top}{\bf g}^{*}(u,{\bf X})\,.$ (5.5)
+More generally, allowing an additional nuisance direction $b\in\overline{\mathcal{T}}_{g_{o}}$, we consider
+$g_{\varepsilon}(u,{\bf X})=g_{o}(u,{\bf X})-\varepsilon\,{\bf a}^{\top}{\bf g}^{*}(u,{\bf X})+\varepsilon b(u,{\bf X})\,$,
+while keeping $\mbox{${\theta}$}_{\varepsilon}=\mbox{${\theta}$}_{o}+\varepsilon{\bf a}$. The score decomposes as
+$S({\bf W};{\bf a},b)={\bf a}^{\top}\ell_{\theta_{o}}^{*}({\bf W})+\int_{0}^{\tau}b(u,{\bf X})\,dM(u)\,.$ (5.6)
+Fix $t\in(0,\tau]$ and define the linear functional
+$L_{t}(b)=\int_{0}^{t}h_{o}(u\mid{\bf X}_{0},{\bf Z}_{0})\,b(u,{\bf X}_{0})\,du\,,\qquad b\in\overline{\mathcal{T}}_{g_{o}}\,.$ (5.7)
+When applied to a vector-valued function ${\bf\mu}(u,{\bf X})=(\mu_{1},\dots,\mu_{p})^{\top}$, we define $L_{t}({\bf\mu})=\big(L_{t}(\mu_{1}),\dots,L_{t}(\mu_{p})\big)^{\top}$. Define
+$c_{0,t}({\bf X}_{0},{\bf Z}_{0})=\int_{0}^{t}h_{o}(u\mid{\bf X}_{0},{\bf Z}_{0})\{{\bf Z}_{0}-{\bf g}^{*}(u,{\bf X}_{0})\}\,du={\bf Z}_{0}\,H_{o}(t\mid{\bf X}_{0},{\bf Z}_{0})-L_{t}({\bf g}^{*})\,.$ (5.8)
+The following lemma computes the pathwise derivative of the target cumulative hazard $H_{\varepsilon}(t\mid{\bf X}_{0},{\bf Z}_{0})$ along a regular submodel.
+It shows the derivative splits cleanly into a nuisance part $L_{t}(b)$ (how perturbations of the nuisance component move the target) plus a parametric part ${\bf a}^{\top}c_{0,t}({\bf X}_{0},{\bf Z}_{0})$ (how perturbations of $\theta$} move the target).
+This is the basic Gateaux derivative input needed to identify the EIF.
+Lemma 4.
+Let $P_{\varepsilon}$ be a regular submodel with score (5.6). Then
+$\left.\frac{d}{d\varepsilon}H_{\varepsilon}(t\mid{\bf X}_{0},{\bf Z}_{0})\right|_{\varepsilon=0}=L_{t}(b)+{\bf a}^{\top}c_{0,t}({\bf X}_{0},{\bf Z}_{0})\,.$ (5.9)
+The following lemma gives an explicit EIF decomposition: a nuisance correction term plus a parametric correction term.
+Lemma 5.
+Fix $t\in[0,\tau]$ and let $({\bf X}_{0},{\bf Z}_{0})$ be an independent draw from the covariate distribution, independent of the training sample.
+Assume A6A7.
+Then $\psi_{t}^{0}$ is unique in $L^{2}(P_{{\bf X},{\bf Z}};\overline{\mathcal{T}}_{g_{o}})$ and the map $P\mapsto H_{P}(t\mid{\bf X}_{0},{\bf Z}_{0})$ is pathwise differentiable at $\mathbb{P}_{0}$ as an element of $L^{2}(P_{{\bf X},{\bf Z}})$; that is, for every regular submodel $\mathbb{P}_{\varepsilon}$ through $\mathbb{P}_{0}$ with score $S({\bf W};{\bf a},b)$,
+$\left.\frac{d}{d\varepsilon}H_{\varepsilon}(t\mid{\bf X}_{0},{\bf Z}_{0})\right|_{\varepsilon=0}=\mathbb{E}_{0}\!\big\{\phi_{t}^{0}({\bf W};{\bf X}_{0},{\bf Z}_{0})\,S({\bf W};{\bf a},b)\ \big|\ {\bf X}_{0},{\bf Z}_{0}\big\}\quad\text{in }L^{2}(P_{{\bf X},{\bf Z}}).$
+Equivalently, (5.2) can be stated as a unique (up to $P_{{\bf X},{\bf Z}}$-null sets) random element $\psi^{0}_{t}(\cdot,\cdot;{\bf X}_{0},{\bf Z}_{0})\in L^{2}\big(P_{{\bf X},{\bf Z}};\overline{\mathcal{T}}_{g_{o}}\big)$ such that for all $b\in\overline{\mathcal{T}}_{g_{o}}$,
+$\mathbb{E}_{({\bf X}_{0},{\bf Z}_{0})}\!\Big[\big\{\langle\psi^{0}_{t}(\cdot,\cdot;{\bf X}_{0},{\bf Z}_{0}),\,b\rangle_{0}-L_{t}(b)\big\}^{2}\Big]=0\,.$ (5.10)
+Finally, an EIF for $H_{o}(t\mid{\bf X}_{0},{\bf Z}_{0})$ is
+$\phi_{t}^{0}({\bf W};{\bf X}_{0},{\bf Z}_{0})=\int_{0}^{\tau}\psi^{0}_{t}(u,{\bf X};{\bf X}_{0},{\bf Z}_{0})\,dM(u)+c_{0,t}({\bf X}_{0},{\bf Z}_{0})^{\top}I(\mbox{${\theta}$}_{o})^{-1}\ell_{\theta_{o}}^{*}({\bf W})\,.$ (5.11)
+We emphasize that the finite-support assumption on $\bf X$} is introduced only to obtain an explicit and implementable one-step estimator.
+More generally, survival inference for a new subject can also be developed under arbitrary covariate distributions via a high-level semiparametric characterization of the efficient influence function.
+Since that approach is more technical and does not yield an equally explicit estimator without additional strong approximation arguments, we present it in Appendix C.
+5.3 Explicit form under finite-support $\bf X$} and asymptotic theory for the cross-fitted estimator
+In the following lemma we show that under discrete finite-support $\bf X$}, the abstract Riesz representer from Lemma 5 becomes an explicit weight function.
+Therefore, the EIF and one-step estimator become fully explicit.
+Lemma 6.
+Under Assumptions A6A8, the Riesz representer $\psi^{0}_{t}(\cdot,\cdot;{\bf X}_{0},{\bf Z}_{0})$ satisfying (5.10) admits the closed form
+$\psi^{0}_{t}(u,{\bf X};{\bf X}_{0},{\bf Z}_{0})=\mathbf{1}(u\leq t)\mathbf{1}({\bf X}={\bf X}_{0})\frac{h_{o}(u\mid{\bf X}_{0},{\bf Z}_{0})}{\Pr({\bf X}={\bf X}_{0})\,A_{{\bf X}_{0}}(u)}.$ (5.12)
+The following is a cross-fitted one-step estimator.
+Let $\{\mathcal{I}_{1},\dots,\mathcal{I}_{K}\}$ be a fixed $K$-fold partition of $\{1,\dots,n\}$ and let $\mathcal{I}_{k}^{c}$ denote the complement of $\mathcal{I}_{k}$.
+For each fold $k$, let $(\widetilde{}\mbox{${\theta}$}^{(-k)},\widetilde{g}^{(-k)})$ be the estimators trained on $\mathcal{I}_{k}^{c}$.
+Define the fold-specific hazard estimator
+$\widehat{h}^{(-k)}(u\mid{\bf X},{\bf Z})=\exp\!\left\{\widetilde{g}^{(-k)}(u,{\bf X})+(\widetilde{}\mbox{${\theta}$}^{(-k)})^{\top}{\bf Z}\right\}.$
+For $t\in[0,\tau]$, define the plug-in cumulative hazard at $({\bf X}_{0},{\bf Z}_{0})$ by
+$\widehat{H}^{(-k)}(t\mid{\bf X}_{0},{\bf Z}_{0})=\sum_{j:t_{j}\leq t}(t_{j}-t_{j-1})\widehat{h}^{(-k)}(t_{j}\mid{\bf X}_{0},{\bf Z}_{0}).$
+For $i\in\mathcal{I}_{k}$, define $\Delta N_{ij}=1\{T_{i}\in(t_{j-1},t_{j}],\ \Delta_{i}=1\}$, $Y_{ij}=Y_{i}(t_{j-1})=1\{T_{i}\geq t_{j-1}\},$ and $\Delta A_{ij}^{(-k)}=(t_{j}-t_{j-1})Y_{ij}\widehat{h}^{(-k)}(t_{j}\mid{\bf X}_{i},{\bf Z}_{i})$.
+The grid-based martingale increment is $\Delta\widehat{M}_{ij}^{(-k)}=\Delta N_{ij}-\Delta A_{ij}^{(-k)}$.
+Let $\widehat{\psi}_{t}(\cdot,\cdot;{\bf X}_{0},{\bf Z}_{0})^{(-k)}$ denote the fold-specific estimator of the Riesz representer. Under discrete finite support of $\bf X$},
+$\widehat{\psi}_{t}^{(-k)}(t_{j},{\bf X}_{i};{\bf X}_{0},{\bf Z}_{0})=\mathbf{1}(t_{j}\leq t)\,\mathbf{1}({\bf X}_{i}={\bf X}_{0})\,\frac{\widehat{h}^{(-k)}(t_{j}\mid{\bf X}_{0},{\bf Z}_{0})}{\widehat{\pi}_{j_{0}}^{(-k)}\widehat{A}_{j_{0}}^{(-k)}(t_{j})},$
+where $j_{0}$ satisfies ${\bf X}_{0}={\bf x}^{(j_{0})}$, $\widehat{\pi}_{j_{0}}^{(-k)}=|\mathcal{I}_{k}^{c}|^{-1}\sum_{i\in\mathcal{I}_{k}^{c}}\mathbf{1}({\bf X}_{i}={\bf x}^{(j_{0})})$, and
+$\widehat{A}_{j_{0}}^{(-k)}(t_{j})=\frac{\sum_{i\in\mathcal{I}_{k}^{c}}Y_{ij}\widehat{h}^{(-k)}(t_{j}\mid{\bf X}_{i},{\bf Z}_{i})1({\bf X}_{i}={\bf x}^{(j_{0})})}{\sum_{i\in\mathcal{I}_{k}^{c}}\mathbf{1}({\bf X}_{i}={\bf x}^{(j_{0})})}.$
+The nuisance influence contribution is
+$\widehat{\phi}^{(-k)}_{t,g}({\bf W}_{i};{\bf X}_{0},{\bf Z}_{0})=\sum_{j=1}^{m_{n}}\widehat{\psi}_{t}^{(-k)}(t_{j},{\bf X}_{i};{\bf X}_{0},{\bf Z}_{0})\Delta\widehat{M}_{ij}^{(-k)}\,.$
+Let $\widehat{\bf g}^{*,(-k)}$ denote the estimator of the projection $\bf g$$^{*}$ and define
+$\widehat{c}_{t}^{(-k)}({\bf X}_{0},{\bf Z}_{0})=\sum_{j:t_{j}\leq t}(t_{j}-t_{j-1})\widehat{h}^{(-k)}(t_{j}\mid{\bf X}_{0},{\bf Z}_{0})\Big\{{\bf Z}_{0}-\widehat{\bf g}^{*,(-k)}(t_{j},{\bf X}_{0})\Big\}.$
+Let $\widehat{I}^{(-k)}$ and $\widehat{\ell}_{\theta}^{*,(-k)}$ be the out-of-fold efficient information and score estimators. Define
+$\widehat{\phi}^{(-k)}_{t,\theta}({\bf W}_{i};{\bf X}_{0},{\bf Z}_{0})=\widehat{c}_{t}^{(-k)}({\bf X}_{0},{\bf Z}_{0})^{\top}\big\{\widehat{I}^{(-k)}\big\}^{-1}\widehat{\ell}_{\theta}^{*,(-k)}({\bf W}_{i})\,.$
+The fold-specific influence estimator is
+$\widehat{\phi}_{t}^{(-k)}=\widehat{\phi}^{(-k)}_{t,g}+\widehat{\phi}^{(-k)}_{t,\theta}\,$.
+The cross-fitted one-step estimator is
+$\widehat{H}^{1,{\sf cf}}(t\mid{\bf X}_{0},{\bf Z}_{0})=\sum_{k=1}^{K}\frac{|\mathcal{I}_{k}|}{n}\widehat{H}^{(-k)}(t\mid{\bf X}_{0},{\bf Z}_{0})+\frac{1}{n}\sum_{k=1}^{K}\sum_{i\in\mathcal{I}_{k}}\widehat{\phi}_{t}^{(-k)}({\bf W}_{i};{\bf X}_{0},{\bf Z}_{0})\,.$
+Finally,
+$\widehat{S}^{1,{\sf cf}}(t\mid{\bf X}_{0},{\bf Z}_{0})=\exp\!\big\{-\widehat{H}^{1,{\sf cf}}(t\mid{\bf X}_{0},{\bf Z}_{0})\big\}.$
+The following lemma shows that in empirical sums, the oracle efficient score $\ell_{\theta_{o}}^{*}$ can be replaced by the cross-fitted plug-in score $\hat{\ell}_{\theta_{o}}^{*,(-k)}$ without changing the $n^{1/2}$-limit, provided the nuisance estimates converge at a suitable rate and cross-fitting ensures conditional independence.
+This is the key technical step that lets us use estimated nuisances inside the EIF expansion.
+Lemma 7.
+For $i\in\mathcal{I}_{k}$ define the plug-in martingale
+$\widehat{M}^{(-k)}_{i}(u)=N_{i}(u)-\int_{0}^{u}Y_{i}(s)\widehat{h}^{(-k)}(s\mid{\bf X}_{i},{\bf Z}_{i})\,ds\,$,
+and the plug-in efficient score
+$\widehat{\ell}^{*,(-k)}_{\theta_{o}}({\bf W}_{i})=\int_{0}^{\tau}\{{\bf Z}_{i}-\widehat{{\bf g}}^{*,(-k)}(u,{\bf X}_{i})\}\,d\widehat{M}^{(-k)}_{i}(u)\,.$ (5.13)
+Assume A1A7 and A9. Then
+$\frac{1}{\sqrt{n}}\sum_{k=1}^{K}\sum_{i\in\mathcal{I}_{k}}\Big\{\widehat{\ell}^{*,(-k)}_{\theta_{o}}({\bf W}_{i})-\ell^{*}_{\theta_{o}}({\bf W}_{i})\Big\}=o_{p}(1).$ (5.14)
+Consequently, in any root-$n$ expansion in which the oracle score $\ell^{*}_{\theta_{o}}$ appears inside an empirical sum, it may be replaced by the cross-fitted plug-in score $\widehat{\ell}^{*,(-k)}_{\theta_{o}}$ without affecting the first-order limit.
+The following lemma proves the out-of-fold estimator $\widehat{I}^{(-k)}$ of the efficient information converges to $I(\mbox{${\theta}$}_{o})$, and that the inverse $\{\widehat{I}^{(-k)}\}^{-1}$ is well-behaved when $I(\mbox{${\theta}$}_{o})$ is nonsingular.
+This justifies replacing $I(\mbox{${\theta}$}_{o})^{-1}$ in the EIF by a cross-fitted plug-in inverse.
+Lemma 8.
+Define the out-of-fold (efficient) information estimator
+$\widehat{I}^{(-k)}=\frac{1}{|\mathcal{I}_{k}^{c}|}\sum_{j\in\mathcal{I}_{k}^{c}}\Delta_{j}\Big\{{\bf Z}_{j}-\widehat{\bf g}^{*,(-k)}(T_{j},{\bf X}_{j})\Big\}^{\otimes 2},\qquad k=1,\dots,K.$
+Assume A1A7 and A9A10. Then,
+  1. (a)
+    $\max_{1\leq k\leq K}\|\widehat{I}^{(-k)}-I(\mbox{${\theta}$}_{o})\|=o_{p}(1)$.
+  2. (b)
+    With probability tending to one, each $\widehat{I}^{(-k)}$ is invertible and
+    $\max_{1\leq k\leq K}\big\|\{\widehat{I}^{(-k)}\}^{-1}-I(\mbox{${\theta}$}_{o})^{-1}\big\|=o_{p}(1).$
+  3. (c)
+    If, in addition, $E_{0}\{\|\ell^{*}_{\theta_{o}}({\bf W})\|^{2}\}<\infty$, then
+    $\frac{1}{\sqrt{n}}\sum_{k=1}^{K}\sum_{i\in\mathcal{I}_{k}}\Big(\{\widehat{I}^{(-k)}\}^{-1}-I(\mbox{${\theta}$}_{o})^{-1}\Big)\,\ell^{*}_{\theta_{o}}({\bf W}_{i})=o_{p}(1).$
+Because the estimators use a time grid, the following lemma compares the true EIF $\phi_{t}^{0}$, its grid version $\phi_{t,n}^{0}$, and the fully plug-in cross-fitted estimate $\widehat{\phi}_{t}^{(-k)}$.
+It shows $\widehat{\phi}_{t}^{(-k)}$ is $L_{2}(\mathbb{P}_{0})$-close to the grid EIF uniformly over folds, and since the grid error vanishes when $\sqrt{n}\,\delta_{n}\to 0$, this yields $L_{2}$-consistency for the true EIF as well.
+Lemma 9.
+Assume A1A11.
+Let $\phi^{0}_{t,n}({\bf W};{\bf X}_{0},{\bf Z}_{0})$ denote the grid version of (5.11) obtained by replacing $\int_{0}^{\tau}\psi_{t}^{0}(u,{\bf X};{\bf X}_{0},{\bf Z}_{0})\,dM(u)$ with $\sum_{j=1}^{m_{n}}\psi_{t}^{0}(t_{j},{\bf X};{\bf X}_{0},{\bf Z}_{0})\Delta M_{j}$ and $c_{0,t}({\bf X}_{0},{\bf Z}_{0})$ with its grid analogue.
+Then, uniformly in $k=1,\dots,K$,
+$\mathbb{E}_{0}\!\left[\Big(\widehat{\phi}_{t}^{(-k)}({\bf W};{\bf X}_{0},{\bf Z}_{0})-\phi^{0}_{t,n}({\bf W};{\bf X}_{0},{\bf Z}_{0})\Big)^{2}\right]=o_{p}(1),$
+and therefore (since $\sqrt{n}\delta_{n}\to 0$ implies $\phi^{0}_{t,n}$ and $\phi_{t}^{0}$ are asymptotically equivalent),
+$\mathbb{E}_{0}\!\left[\Big(\widehat{\phi}_{t}^{(-k)}({\bf W};{\bf X}_{0},{\bf Z}_{0})-\phi^{0}_{t}({\bf W};{\bf X}_{0},{\bf Z}_{0})\Big)^{2}\right]=o_{p}(1),\qquad\text{uniformly in }k.$
+Theorem 6.
+Fix $t\in[0,\tau]$ and let $({\bf X}_{0},{\bf Z}_{0})$ be an independent draw from the covariate distribution. Under Assumptions A1A11, and assuming moreover that for each fold $k$,
+$\|\widetilde{g}^{(-k)}-g_{o}\|_{0}=O_{p}(\gamma_{n}\log^{2}n),\qquad\gamma_{n}\log^{2}n=o(n^{-1/4}),$
+uniformly in $k$, and that $\widetilde{}\mbox{${\theta}$}^{(-k)}$ satisfies the same first-order expansion as established in Theorem 4, namely,
+$\sqrt{|\mathcal{I}_{k}^{c}|}\,(\widetilde{}\mbox{${\theta}$}^{(-k)}-\mbox{${\theta}$}_{o})=I(\mbox{${\theta}$}_{o})^{-1}\frac{1}{\sqrt{|\mathcal{I}_{k}^{c}|}}\sum_{i\in\mathcal{I}_{k}^{c}}\ell_{\theta_{o}}^{*}({\bf W}_{i})+o_{p}(1),$
+uniformly in k. Then, conditional on the independent draw $({\bf X}_{0},{\bf Z}_{0})$,
+$\sqrt{n}\Big\{\widehat{H}^{\,1,{\sf cf}}(t\mid{\bf X}_{0},{\bf Z}_{0})-H_{o}(t\mid{\bf X}_{0},{\bf Z}_{0})\Big\}\mid({\bf X}_{0},{\bf Z}_{0})\rightarrow^{\mathcal{D}}\mathcal{N}\!\big(0,\sigma_{H}^{2}(t;{\bf X}_{0},{\bf Z}_{0})\big),$
+where $\sigma_{H}^{2}(t;{\bf X}_{0},{\bf Z}_{0})={\mathbb{V}}_{0}\!\Big(\phi_{t}^{0}({\bf W};{\bf X}_{0},{\bf Z}_{0})\ \big|\ {\bf X}_{0},{\bf Z}_{0}\Big).$ Moreover, with
+$\widehat{\sigma}_{H}^{2}(t;{\bf X}_{0},{\bf Z}_{0})=\frac{1}{n}\sum_{k=1}^{K}\ \sum_{i\in\mathcal{I}_{k}}\Big\{\widehat{\phi}_{t}^{(-k)}({\bf W}_{i};{\bf X}_{0},{\bf Z}_{0})\Big\}^{2},$
+we have $\widehat{\sigma}_{H}^{2}(t;{\bf X}_{0},{\bf Z}_{0})\xrightarrow{p}\sigma_{H}^{2}(t;{\bf X}_{0},{\bf Z}_{0})$. Finally,
+$\sqrt{n}\Big\{\widehat{S}^{\,1,{\sf cf}}(t\mid{\bf X}_{0},{\bf Z}_{0})-S_{o}(t\mid{\bf X}_{0},{\bf Z}_{0})\Big\}\mid({\bf X}_{0},{\bf Z}_{0})\rightarrow{{}^{\mathcal{D}}}\mathcal{N}\!\Big(0,\sigma_{S}^{2}(t;{\bf X}_{0},{\bf Z}_{0})\Big)$
+where $\sigma_{S}^{2}(t;{\bf X}_{0},{\bf Z}_{0})=S_{o}(t\mid{\bf X}_{0},{\bf Z}_{0})^{2}\,\sigma_{H}^{2}(t;{\bf X}_{0},{\bf Z}_{0})$ and a consistent variance estimator is
+$\widehat{\sigma}_{S}^{2}(t;{\bf X}_{0},{\bf Z}_{0})=\big\{\widehat{S}^{\,1,{\sf cf}}(t\mid{\bf X}_{0},{\bf Z}_{0})\big\}^{2}\,\widehat{\sigma}_{H}^{2}(t;{\bf X}_{0},{\bf Z}_{0})\,.$
+Theorem 6 yields a Wald-type pointwise confidence interval for the survival of a new subject at any fixed time $t\in[0,\tau]$, that is
+$\widehat{S}^{1,\mathrm{cf}}(t\mid X_{0},Z_{0})\pm z_{1-\alpha/2}{\widehat{\sigma}_{S}(t;X_{0},Z_{0})}/{\sqrt{n}}\,.$
+Theorem 6 is proved conditionally on the random draw $({\bf X}_{0},{\bf Z}_{0})$; it is not a uniform statement over all fixed covariate values $({\bf x}_{0},{\bf z}_{0})$.
+Hence, the associated coverage statement is first understood conditionally,
+$\Pr\!\left(S_{o}(t\mid{\bf X}_{0},{\bf Z}_{0})\in\left[\widehat{S}^{1,\mathrm{cf}}(t\mid{\bf X}_{0},{\bf Z}_{0})\pm z_{1-\alpha/2}\frac{\widehat{\sigma}_{S}(t;{\bf X}_{0},{\bf Z}_{0})}{\sqrt{n}}\right]\Bigm|{\bf X}_{0},{\bf Z}_{0}\right)\to 1-\alpha\,.$
+Let
+$A_{n}(t)=\left\{S_{o}(t\mid{\bf X}_{0},{\bf Z}_{0})\in\left[\widehat{S}^{1,\mathrm{cf}}(t\mid{\bf X}_{0},{\bf Z}_{0})\pm z_{1-\alpha/2}\frac{\widehat{\sigma}_{S}(t;{\bf X}_{0},{\bf Z}_{0})}{\sqrt{n}}\right]\right\}\,$,
+for which Theorem 6 yields $\Pr\!\big(A_{n}(t)\mid{\bf X}_{0},{\bf Z}_{0}\big)\to 1-\alpha$, in probability.
+Then, since $0\leq\Pr(A_{n}(t)\mid{\bf X}_{0},{\bf Z}_{0})\leq 1$, bounded convergence in probability implies convergence in $L^{1}$ and therefore $\mathbb{E}\!\left\{\Pr\!\big(A_{n}(t)\mid{\bf X}_{0},{\bf Z}_{0}\big)\right\}\to 1-\alpha$.
+Applying the law of iterated expectations gives $\Pr\!\big(A_{n}(t)\big)=\mathbb{E}\!\left\{\Pr\!\big(A_{n}(t)\mid{\bf X}_{0},{\bf Z}_{0}\big)\right\}\to 1-\alpha\,.$ Thus the pointwise (in $t$) Wald interval has asymptotic marginal coverage under the joint law of the training data and the independent draw $({\bf X}_{0},{\bf Z}_{0})$, namely,
+$\Pr\!\left(S_{o}(t\mid{\bf X}_{0},{\bf Z}_{0})\in\left[\widehat{S}^{1,\mathrm{cf}}(t\mid{\bf X}_{0},{\bf Z}_{0})\pm z_{1-\alpha/2}\frac{\widehat{\sigma}_{S}(t;{\bf X}_{0},{\bf Z}_{0})}{\sqrt{n}}\right]\right)\to 1-\alpha\,.$
+In our partially linear hazard model, the nuisance function $g_{o}(t,{\bf X})$ is assumed to belong to a composite Hölder class, $g_{o}\in\mathcal{H}(q,\alpha,{\bf d}^{*},\tilde{\bf d},M)$.
+Under this assumption, sparse ReLU networks can achieve the (near-)minimax rate $\|\widehat{g}-g_{o}\|=O_{p}\!\big(\tau_{n}\big)$, where the complexity exponent $\gamma_{n}$ is governed by the hardest layer of the composition.
+The condition on $\gamma_{n}$ imposed in the theorem above is therefore a standard second-order remainder requirement for one-step inference, and it ensures $\sqrt{n}\,\|\widetilde{g}-g_{o}\|_{0}^{2}=o_{p}(1)$, so that nuisance-estimation error enters only through $o_{p}(n^{-1/2})$ terms.
+Interpreting this condition in terms of the composite Hölder parameters, it is sufficient that the slowest layer satisfies
+$\frac{\tilde{\alpha}_{\ell}}{2\tilde{\alpha}_{\ell}+\tilde{d}_{\ell}}>\frac{1}{4}\quad\text{for the maximizing }\ell,\qquad\text{equivalently}\qquad 2\tilde{\alpha}_{\ell}>\tilde{d}_{\ell}$,
+up to the polylogarithmic factor $\log^{2}n$.
+Thus, the assumption is mild when the intrinsic dimensions $\tilde{d}_{\ell}$ are small (structural sparsity or low-dimensional composition) and the effective smoothness levels $\tilde{\alpha}_{\ell}$ are not too low.
+It can fail only in very rough or high-intrinsic-dimension regimes where the best achievable $\gamma_{n}$ is slower than $n^{-1/4}$, in which case Wald-type inference would require additional structure (e.g., stronger smoothness, further dimension reduction, or alternative debiasing or regularization).
+6 Practical Implementation
 6.1 Applying DNN
 In practice, we approximate $g_{o}$ by a fully connected ReLU network and model the linear term $\mbox{${\theta}$}^{\top}{\bf Z}$ using a parallel linear layer.
-The network computes $\widetilde{g}(t,{\bf X})$ and $\widetilde{\mbox{${\theta}$}}^{\top}{\bf Z}$ and their sum is inserted into the loss based on (3).
+The network computes $\widetilde{g}(t,{\bf X})$ and $\widetilde{\mbox{${\theta}$}}^{\top}{\bf Z}$ and their sum is inserted into the loss based on (3.2).
 This architecture jointly learns the nuisance function $g$ and a finite–dimensional linear effect of $\bf Z$} in a way that mirrors the model structure.
-To evaluate the approximated log-likelihood (3) in practice, we discretize the time axis on a fixed grid (e.g., $n$) and augment it with all observed times, yielding $0=t_{0}<t_{1}<\cdots<t_{m_{n}}=\tau$.
+To evaluate the approximated log-likelihood (3.2) in practice, we discretize the time axis on a grid (e.g., $n$) and augment it with all observed times, yielding $0=t_{0}<t_{1}<\cdots<t_{m_{n}}=\tau$.
 The dataset is then expanded in the usual counting‐process manner.
@@
 All dummy-observations inherit $(\mathbf{X}_{i},\mathbf{Z}_{i})$; the event indicator $\Delta_{ik}=\mathbf{1}\{T_{i}\in(t_{k-1},t_{k}],\Delta=1\}$ is set to 0 except for the final interval containing $T_{i}$, where it takes the original value $\Delta_{i}$.
-This choice of grid and data expansion enabling efficient computation within standard deep-learning frameworks.
-The construction parallels the standard expanded-data representation used for Cox models with time-dependent covariates (therneau2000cox; hu2023conditional).
+This choice of grid and data expansion enables efficient computation within standard deep-learning frameworks.
+The construction parallels the standard expanded-data representation used for Cox models with time-dependent covariates (Therneau and Grambsch, 2000; Hu and Nan, 2023).
 In theory we allow for sparsity, but in practice we do not explicitly enforce it.
@@
 6.2 Estimation of the Covariance Matrix $I(\mbox{${\theta}$}_{o})^{-1}$
-Inference on $\mbox{${\theta}$}_{0}$ relies an estimating the efficient information matrix $I(\mbox{${\theta}$}_{o})$ from Theorem 4, whose inverse provides the asymptotic covariance of the estimator $\widetilde{\mbox{${\theta}$}}$.
+Inference on $\mbox{${\theta}$}_{0}$ relies on estimating the efficient information matrix $I(\mbox{${\theta}$}_{o})$, whose inverse provides the asymptotic covariance of the estimator $\widetilde{\mbox{${\theta}$}}$.
 Recall that the efficient information is
@@
 Specifically, we approximate $\bf g$$^{\ast}$ coordinatewise.
-For each $j=1,\ldots,p$ we fit a separate fully connected network $\widehat{g}^{\ast}_{j}(T,{\bf X})$ by least squares using only uncensored observations, that is by
-$\min\sum_{i:\,\Delta_{i}=1}\{Z_{ij}-\widehat{g}^{\ast}_{j}(T_{i},{{\bf X}}_{i})\}^{2}\,.$
+For each $j=1,\ldots,p$ we fit a separate fully connected network $\widehat{g}^{\ast}_{j}(T,{\bf X})$ by least squares using only uncensored observations, that is by $\min\sum_{i:\,\Delta_{i}=1}\{Z_{ij}-\widehat{g}^{\ast}_{j}(T_{i},{{\bf X}}_{i})\}^{2}$.
 Each network $j$, $j=1,\ldots,p$, takes $(T_{i},\mathbf{X}_{i})$ as input, $Z_{ij}$ as target, and uses the same architecture and hyperparameters as the main model; empirically, these choices produce stable estimates and are not sensitive to tuning.
@@
 The data are split into five folds, for each fold and each coordinate $j$, the corresponding network is trained on the remaining four folds (restricting to $\Delta=1$) and predictions $\widehat{g}^{\ast}_{j}(T_{i},X_{i})$ are obtained for the held-out fold.
-for all $i$ in the held–out fold.
 Aggregating across folds yields, for each observation $i$, the estimated projection vector
 $\widehat{\bf g}^{\ast}(T_{i},{{\bf X}}_{i})=\bigl(\widehat{g}^{\ast}_{1}(T_{i},{{\bf X}}_{i}),\ldots,\widehat{g}^{\ast}_{p}(T_{i},{{\bf X}}_{i})\bigr)^{\top}\,.$
-We then compute the residuals ${\bf R}_{i}={{\bf Z}}_{i}-\hat{\bf g}^{\ast}(T_{i},{{\bf X}}_{i})\in\mathbb{R}^{p}$, $i=1,\ldots,n$, and stack them into an $n\times p$ matrix $\mathcal{R}$.
+We then compute the residuals ${\bf R}_{i}=\sqrt{\Delta}_{i}\{{{\bf Z}}_{i}-\widehat{\bf g}^{\ast}(T_{i},{{\bf X}}_{i})\}\in\mathbb{R}^{p}$, $i=1,\ldots,n$, and stack them into a matrix $\mathcal{R}$.
 Our estimator of the efficient information is $\widehat{I}=\mathcal{R}^{\top}\mathcal{R}/n$, and the asymptotic covariance is $\widehat{I}^{-1}$.
 In practice this procedure is numerically stable, requires no tuning beyond the main network fit, and produces reliable covariance estimates for $\widetilde{\mbox{${\theta}$}}$.
-6 Simulation Study
-In this section, we compare the finite-sample performance of the proposed FLEXI-Haz method with the partially linear neural PH estimator of zhong2022deep.
-Our evaluation focuses on the estimation accuracy of the finite-dimensional parameter $\theta$} and on the coverage properties of the corresponding Wald confidence intervals.
-We consider a setting with both nuisance and primary covariates.
-Let ${{\bf X}}=(X_{1},X_{2},X_{3})^{\top}$ and ${{\bf Z}}=(Z_{1},Z_{2})^{\top}$, where each covariate is independently sampled from a uniform distributed over [1,1][-1,1].
-The true parameter of interest is $\mbox{${\theta}$}_{o}=(b_{1},b_{2})^{\top}=(2,-1)^{\top}$.
-For the nuisance component, we define $f({\bf X})=0.2(X_{1}+X_{2})+0.5X_{1}X_{2}+X_{3}^{2}$, which induces nonlinear and interaction effects and interactions among the nuisance covariates.
-The event time $U$ is generated from a hazard function of the form
-$h_{o}(t\mid{\bf X},{\bf Z})=0.1\exp\!\left\{(0.1+f({\bf X})^{2})t+\mbox{${\theta}$}_{o}^{\top}{\bf Z}\right\},\qquad t\in[0,\tau]\,.$
-with $\tau=30$ chosen as the fixed study horizon.
-Right censoring is introduced through an independent censoring time $C$ sampled from an exponential distribution with expectation $1/30$, combined with administration censoring at time 30.
-The parameters of the data-generating mechanism are chosen so that the censoring rate is approximately 30%.
-For both FLEXI-Haz and the estimator of zhong2022deep, the nuisance function is estimated using a fully connected feedforward network with ReLU activations.
-The only and difference between the two estimators lies in the underlying hazard model, and the training pipeline and tuning strategy are identical.
-Hyperparameters are selected once, using only one simulated dataset.
-We construct the expanded dataset using a fixed time grid of $m_{n}=n$ points augmented with all observed times, and perform 5-fold cross-validation over the following search space:
-  • number of hidden layers: $\{1,2,3,4,5,6\}$
-  • width (nodes per hidden layer): $\{20,40,80,160,320\}$
-  • learning rate: $\{10^{-3},\,5\times 10^{-3},\,10^{-2},\,5\times 10^{-2}\}$
-Training uses early stopping with a patience of 35 epochs and a batch size of $10^{5}$ on the expanded dataset.
-For each hyperparameter combination we train the network once and record the average validation loss across the five folds; the configuration achieving the smallest average loss is selected.
-Cross-validation selects, for FLEXI-Haz, a network with five hidden layers, width 20, and learning rate $10^{-3}$.
-For the estimator of zhong2022deep, the selected architecture has four hidden layers, width 80, and the same learning rate $10^{-3}$.
-Table 1 presents the results for $n=8000$ based on 200 replications.
-The proposed estimator, FLEXI-Haz, is essentially unbiased and its empirical variability closely matches the theoretical standard errors, yielding coverage probabilities near their nominal levels.
-In contrast, the partially linear PH estimator of zhong2022deep exhibits substantial bias in both components of $\boldsymbol{\theta}_{o}$, leading to near-zero empirical coverage.
+7 Simulation Study and Real-data application
+7.1 Simulation Study
+We compare the finite-sample performance of the proposed FLEXI–Haz method with the partially linear neural PH estimator of Zhong et al.
+(2022).
+Our focus is on estimation of the finite-dimensional parameter $\theta$} and on the coverage of the associated Wald confidence intervals.
+We consider nuisance and primary covariates ${{\bf X}}=(X_{1},X_{2},X_{3})^{\top}$ and ${{\bf Z}}=(Z_{1},Z_{2})^{\top}$, where all covariates are independently distributed as $\mathrm{Unif}[-1,1]$.
+The true parameter is $\mbox{${\theta}$}_{o}=(2,-1)^{\top}$.
+The nuisance component includes $f({\bf X})=0.2(X_{1}+X_{2})+0.5X_{1}X_{2}+X_{3}^{2}$, which induces nonlinear and interaction effects.
+Event times are generated from
+$h_{o}(t\mid{\bf X},{\bf Z})=0.1\exp\!\left\{(0.1+f({\bf X})^{2})t+\mbox{${\theta}$}_{o}^{\top}{\bf Z}\right\},\qquad t\in[0,\tau],$
+with fixed study horizon $\tau=30$.
+Right censoring is introduced through an independent exponential censoring time with mean $1/30$, together with administrative censoring at time 30.
+The resulting censoring rate is approximately 30%.
+For both FLEXI–Haz and the model of Zhong et al.
+(2022), the nuisance function is estimated using a fully connected feedforward ReLU network.
+The two methods differ only in the underlying hazard model; the training pipeline and tuning strategy are otherwise identical.
+Hyperparameters are selected once using a single simulated dataset.
+We construct the expanded dataset on a grid of $m_{n}=n$ time points augmented with all observed times, and perform 5-fold cross-validation over the number of hidden layers $\{1,2,3,4,5,6\}$, width $\{20,40,80,160,320\}$, and learning rate $\{10^{-3},\,5\times 10^{-3},\,10^{-2},\,5\times 10^{-2}\}$.
+Training uses early stopping with patience 35 and batch size $10^{5}$ on the expanded dataset.
+For each hyperparameter combination, the network is trained once and the average validation loss across the five folds is recorded; the configuration with the smallest average loss is selected.
+Cross-validation selects, for FLEXI–Haz, five hidden layers, width 20, and learning rate $10^{-3}$, and for Zhong et al.
+(2022), four hidden layers, width 80, and learning rate $10^{-3}$.
+Table 1 reports results for $n=8000$ based on 200 replications.
+FLEXI–Haz is essentially unbiased, and its empirical standard deviation closely matches the average estimated standard error, yielding coverage probabilities close to the nominal levels.
+In contrast, the partially linear PH estimator of Zhong et al.
+(2022) exhibits substantial bias in both components of $\mbox{${\theta}$}_{o}$, resulting in near-zero empirical coverage.
+To assess the finite-sample performance of the proposed pointwise survival-curve inference, we conducted an additional simulation experiment under the same setup as above, except that each nuisance covariate was generated independently from a Poisson(2)(2) distribution truncated at $6$.
+In each of $200$ simulation runs, we generated a training sample of size $n=8000$ and an additional evaluation sample of size $200$.
+For each target level $s\in\{0.1,0.2,\ldots,0.9\}$, we selected a fixed time point $t_{s}$ such that the mean estimated survival in the training sample was approximately $s$, and evaluated the empirical coverage of the pointwise confidence interval for the survival probability at time $t_{s}$ based on the evaluation samples.
+Table 2 demonstrates that the proposed pointwise inference procedure is well calibrated in finite samples.
+Across the selected time points, the empirical coverage remains close to the nominal level for both the 90% and 95% confidence intervals, supporting the practical reliability of the method in the discrete-covariate setting considered here.
+The slight overcoverage at $t=6.7126$ may be due to boundary effects, since this is a late time point where the true survival probability is close to zero for some values of $({\bf X}_{0},{\bf Z}_{0})$.
 Table 1: Simulation results for $n=8000$ observations, based on 200 replications.
 The true parameter is $\mbox{${\theta}$}_{o}=(2,-1)^{\top}$.
-“Est” denotes the mean of the estimated coefficients; “Emp SD” the empirical standard deviation; “Est SE” the estimated theoretical standard error; and “90% , 95%” the empirical coverage rates of the corresponding Wald-type confidence intervals.
-FLEXI-Haz zhong2022deep
-Est Emp SD Est SE 90% 95% Est Emp SD Est SE 90% 95%
+“Est” denotes the mean of the estimated coefficients; “Emp SD” the empirical standard deviation; “Est SE” the estimated theoretical standard error; and “90%CI , 95%CI” the empirical coverage rates of the corresponding Wald-type confidence intervals.
+FLEXI–Haz Zhong et al. (2022)
+Est Emp SD Est SE 90%CI 95%CI Est Emp SD Est SE 90%CI 95%CI
 $\theta_{o1}$ 1.997 0.040 0.035 0.878 0.935 1.764 0.039 0.034 0.000 0.000
 $\theta_{o2}$ -0.994 0.027 0.029 0.937 0.959 -0.886 0.030 0.028 0.015 0.035
+Table 2: Simulation results of empirical pointwise coverage of the 90% and 95% confidence intervals for the survival curve, based on 200200 replications with training sample size $n=8000$ and testing sample size of $200$.
+For each target survival time $t_{s}$, we evaluated whether the confidence interval covered the true survival value at $t_{s}$.
+Marginal survival $s$ Target $t_{s}$ 90%CI SD of coverage 95%CI SD of coverage
+0.1 6.7126 0.9377 0.0741 0.9676 0.0531
+0.2 4.2934 0.9120 0.0861 0.9548 0.0638
+0.3 3.0192 0.9167 0.0769 0.9565 0.0595
+0.4 2.2003 0.9121 0.0755 0.9595 0.0523
+0.5 1.6035 0.8903 0.0889 0.9488 0.0589
+0.6 1.2330 0.8997 0.0912 0.9520 0.0611
+0.7 0.9476 0.8995 0.0824 0.9480 0.0614
+0.8 0.6812 0.9123 0.0815 0.9577 0.0559
+0.9 0.3900 0.9044 0.0987 0.9567 0.0638
+7.2 Real-data application: Rotterdam breast cancer study
+To illustrate the practical performance of FLEXI–Haz, we analyzed the Rotterdam Breast Cancer data set, available in the survival package in R.
+These data were studied by Foekens et al.
+(2000) to assess clinical factors associated with time to breast cancer recurrence or death, where the outcome is defined as the number of days from primary surgery until the earlier of recurrence or death.
+The data set contains $n=2982$ patients, with an overall censoring proportion of $57.35\%$, and includes nine baseline covariates: age, progesterone receptor level, estrogen receptor level, number of positive lymph nodes, menopausal status, tumor size, tumor grade, chemotherapy, and hormonal therapy.
+Among these variables, the first four are continuous and the remaining five are discrete.
+Our primary scientific goal in this analysis is to assess the effects of chemotherapy and hormonal therapy.
+Accordingly, we included these two treatment indicators in the linear component $\bf Z$}, while modeling the remaining covariates nonparametrically through the DNN component.
+We used five-fold cross-validation.
+Within each fold, the training portion was further split so that $20\%$ of the training data served as a validation set for tuning hyperparameters, after which model performance was evaluated on the held-out test set.
+Table 3 summarizes the estimated treatment coefficients and corresponding $95\%$ confidence intervals for the proposed method, together with the comparator methods.
+Table 3: Estimated treatment effects and $95\%$ confidence intervals for the Rotterdam breast cancer data.
+Treatment Method Estimate (SE) 95% CI
+Chemotherapy Cox PH 0.127(0.071)-0.127\ (0.071) (0.267, 0.013)(-0.267,\ 0.013)
+Zhong et al. (2022) 0.257(0.095)-0.257\ (0.095) (0.443,0.071)(-0.443,\ -0.071)
+FLEXI–Haz 0.296(0.094)-0.296\ (0.094) (0.480,0.112)(-0.480,\ -0.112)
+Hormonal therapy Cox PH 0.103(0.083)-0.103\ (0.083) (0.266, 0.061)(-0.266,\ 0.061)
+Zhong et al. (2022) 0.372(0.080)-0.372\ (0.080) (0.529,0.215)(-0.529,\ -0.215)
+FLEXI–Haz 0.365(0.080)-0.365\ (0.080) (0.522,0.209)(-0.522,\ -0.209)
+Table 3 shows negative estimated effects of both chemotherapy and hormonal therapy, with 95% confidence intervals excluding zero under FLEXI–Haz.
+The corresponding estimates are qualitatively similar to those of Zhong et al.
+(2022), whereas the Cox PH model yields confidence intervals that include zero for both treatments.
+These findings indicate that FLEXI–Haz recovers interpretable treatment effects while accommodating flexible nonproportional nuisance effects.
+Figure 1 displays estimated survival curves for 16 randomly selected subjects.
+The plug-in estimator is visually smooth because it is evaluated on a dense time grid, whereas the one-step estimator is piecewise constant, reflecting that its bias-correction term is constructed from subjects sharing the same discrete nuisance-covariate profile ${\bf X}_{0}$.
+Figure 1: Estimated survival curves for 16 randomly selected subjects in the Rotterdam breast cancer data.
+The blue curve is the plug-in FLEXI–Haz estimator, evaluated on a dense grid.
+The red step function is the cross-fitted one-step estimator, and the shaded band gives the corresponding pointwise confidence interval.
+The visible steps in the one-step curves arise from the finite-support correction, which uses subjects with the same nuisance-covariate profile ${\bf X}_{0}$.
 8 Discussion
-In this work we proposed FLEXI–Haz, a deep semiparametric survival model that combines a fully non-PH formulation with a partially linear structure.
-By modeling the log-hazard as $\boldsymbol{\theta}_{o}^{\top}\mathbf{Z}+g_{o}(t,\mathbf{X})$ and estimating $g_{o}$ with a deep ReLU network, we retain interpretability and valid inference for the low-dimensional parameter $\boldsymbol{\theta}_{o}$ while allowing highly flexible, time-varying effects of the nuisance covariates.
-Our main theoretical result is that FLEXI–Haz enjoys the same key properties that zhong2022deep established for partially linear Cox models under PH: the neural estimator of $g_{o}$ attains the optimal composite minimax rate, and the estimator of $\boldsymbol{\theta}_{o}$ is $\sqrt{n}$–consistent, asymptotically normal, and semiparametrically efficient.
-Crucially, these properties continue to hold in our more general, baseline-free non-PH setting, despite the presence of time in the nuisance component and the use of the full likelihood rather than a partial likelihood.
-The simulation study supports these theoretical findings and illustrates their practical relevance.
-In a setting with nonlinear time–covariate interactions, FLEXI–Haz produces essentially unbiased estimates of $\boldsymbol{\theta}_{o}$, with empirical standard deviations that align closely with the estimated standard errors and nominal coverage for Wald-type confidence intervals.
-In contrast, the partially linear PH estimator of zhong2022deep, which forces the nuisance component to depend only on $\mathbf{X}$, exhibits substantial bias and severe undercoverage when the true hazard is non-proportional.
-This behavior is exactly what the theory predicts: when time-varying effects are incorrectly constrained to be time-invariant, the nuisance approximation error leaks into the finite-dimensional parameter.
-Allowing $g_{o}$ to depend on both $t$ and $\mathbf{X}$ removes this source of misspecification, and the empirical results confirm that the resulting estimator behaves in accordance with the asymptotic theory.
-Although FLEXI–Haz provides consistent estimators of the hazard and enables estimation of the conditional survival function $S(t\mid\mathbf{X},\mathbf{Z})$, the present work does not derive the asymptotic distribution of the resulting survival-curve estimator.
-Nevertheless, valid pointwise confidence intervals and simultaneous confidence bands can still be constructed by applying the ensemble–bootstrap procedure developed by arie2024confidence.
-That method is fully nonparametric, accommodates right censoring, and can be used as a post–hoc procedure on top of FLEXI–Haz without modifying the training algorithm.
-Our implementation focuses on fully connected ReLU networks, which fit naturally into the composite Hölder framework and for which sharp approximation and entropy bounds are available.
-This choice enables us to obtain clean minimax rates and empirical-process control.
-At the same time, fully connected architectures are not always the most natural representation for complex structured data such as images.
-Convolutional neural networks (CNNs) are often preferable in imaging applications, and several deep survival methods in the literature rely on convolutional encoders.
-While the FLEXI–Haz likelihood and estimation strategy can be used with convolutional backbones in practice, our current theory does not cover convolutional architectures, and the existing results on sparse ReLU networks cannot be directly transferred to CNNs.
-Extending the approximation and efficiency analysis to convolutional or other structured architectures is therefore an important direction for future work.
-A further practical advantage of our formulation is that it meshes naturally with the standard “pseudo-observation” representation used to handle time-dependent covariates.
-The full-likelihood derivation already requires us to expand each subject into multiple pseudo-observations over risk intervals.
-This is precisely the data-expansion scheme commonly used to encode time-dependent covariates in Cox-type models.
-As a result, extending FLEXI–Haz to time-dependent covariates is conceptually straightforward: one augments $\mathbf{X}$ to include time-varying components and constructs the expanded dataset in the usual way.
-The theoretical development is also compatible with this extension, as the key arguments are expressed in terms of predictable processes and integrals over the at-risk time, and do not rely on time-independence of the covariates beyond notational convenience.
-The main current limitation of FLEXI–Haz is computational rather than statistical.
-Because we work with the full likelihood and approximate the cumulative hazard via a Riemann-type sum on a fine time grid, the expanded dataset contains $O(n\,m_{n})$ pseudo-observations, where $m_{n}$ is the number of grid points.
-In the default choice $m_{n}\asymp n$, this leads to an effective $O(n^{2})$ cost per epoch, which can become prohibitive for very large studies or dense event processes.
-Our experiments show that this is manageable for moderate $n$, but improving scalability is crucial for applications with hundreds of thousands of subjects or very fine time resolution.
-Several strategies could mitigate this computational burden.
-One natural approach is to sub-sample pseudo-observations corresponding to non-events in the risk sets, e.g., by keeping all intervals containing failures and a carefully chosen random subset of non-event intervals, with importance weights to preserve unbiasedness.
-Since the information about $\boldsymbol{\theta}_{o}$ and $g_{o}$ is concentrated near failures and in regions with high hazard, aggressive sub-sampling of “uninformative” background intervals should reduce cost with limited loss in efficiency (Keret02102023).
-Another line of work is to exploit structured quadrature rules or other powerful approximation methods.
-Our theoretical treatment of numerical integration in Theorem 5 is deliberately abstract and suggests a broad class of admissible approximations.
-The only requirement is that the approximation error in the empirical loss is $O_{p}(n^{-1})$ uniformly over the parameter space, and that the corresponding error in the $\boldsymbol{\theta}$-score is $O_{p}(n^{-1/2})$.
-Any numerical scheme—whether based on fewer grid points, higher-order quadrature, adaptive time meshes, or stochastic approximations—that satisfies these two error bounds yields the same asymptotic distribution of $\widetilde{\mbox{${\theta}$}}$.
-In particular, this shows that our choice of a simple Riemann sum with $m_{n}\asymp n$ is sufficient but not necessary: one can potentially use substantially fewer time points while preserving the validity of the asymptotic theory, provided the approximation quality is high enough.
-8 Appendix - Proofs
+We proposed FLEXI–Haz, a deep semiparametric survival model that combines a partially linear structure with a flexible non-proportional hazards formulation.
+By modeling the log-hazard as $\mbox{${\theta}$}_{o}^{\top}{\bf Z}+g_{o}(t,{\bf X})$, the method preserves interpretability for the low-dimensional parameter $\mbox{${\theta}$}_{o}$ while allowing the nuisance component to vary nonlinearly with both time and covariates.
+Our main theoretical results show that the DNN estimator of $g_{o}$ attains the optimal composite minimax rate, that the estimator of $\mbox{${\theta}$}_{o}$ is $\sqrt{n}$-consistent, asymptotically normal, and semiparametrically efficient, and that pointwise frequentist asymptotic inference is available for the survival curve of a new subject through a cross-fitted one-step procedure.
+The simulation and real-data analyses support these conclusions and illustrate the practical advantages of FLEXI–Haz when proportional hazards may be too restrictive.
+The main limitation of the current survival-inference theory is the additional assumption that the nuisance covariate vector $\bf X$} has discrete finite support, which is used to obtain an explicit Riesz representer and a tractable one-step correction.
+Although this restriction is often reasonable in applications with categorical covariates or grouped continuous features, extending the pointwise survival-inference theory to more general covariate distributions is an important direction for future work.
+Another natural extension is to develop analogous theory for richer network architectures beyond fully connected ReLU networks.
+Appendix A Proofs for Section 4
 See 1
 Proof of Lemma 1.
-Let $N(t)=\Delta\,\mathbf{1}(T\leq t)$ be the counting process associated with the event time, and let $M(t)=N(t)-\int_{0}^{t}Y(s)\,h_{o}(s\,|\,{{\bf X}},{{\bf Z}})\,ds$ denote its martingale decomposition under the true hazard (1).
+Let $N(t)=\Delta\,\mathbf{1}(T\leq t)$ be the counting process associated with the event time, and let $M(t)=N(t)-\int_{0}^{t}Y(s)\,h_{o}(s\,|\,{{\bf X}},{{\bf Z}})\,ds$ denote its martingale decomposition under the true hazard (2.1).
 Because $f(t,{\bf X},{\bf Z})$ is predictable, the stochastic integral $\int_{0}^{\tau}f(t,{{\bf X}},{{\bf Z}})\,dM(t)$ has expectation zero.
 Consequently,
-$\mathbb{E}\!\left\{\int_{0}^{\tau}f(t,{\bf X},{\bf Z})\,dN(t)\right\}=\mathbb{E}\!\left\{\int_{0}^{\tau}f(t,{\bf X},{\bf Z})\,Y(t)\,h_{0}(t\,|\,{\bf X},{\bf Z})\,dt\right\}.$
+$\mathbb{E}\!\left\{\int_{0}^{\tau}f(t,{\bf X},{\bf Z})\,dN(t)\right\}=\mathbb{E}\!\left\{\int_{0}^{\tau}f(t,{\bf X},{\bf Z})\,Y(t)\,h_{o}(t\,|\,{\bf X},{\bf Z})\,dt\right\}.$
 Since $\int_{0}^{\tau}f(t,{\bf X},{\bf Z})\,dN(t)=\Delta\,f(T,{\bf X},{\bf Z})$, we obtain the identity
-$\mathbb{E}\{\Delta\,f(T,{\bf X},{\bf Z})\}=\mathbb{E}\!\left\{\int_{0}^{\tau}f(t,{\bf X},{\bf Z})\,Y(t)\,h_{0}(t\,|\,{\bf X},{\bf Z})\,dt\right\}\,.$ (10)
+$\mathbb{E}\{\Delta\,f(T,{\bf X},{\bf Z})\}=\mathbb{E}\!\left\{\int_{0}^{\tau}f(t,{\bf X},{\bf Z})\,Y(t)\,h_{o}(t\,|\,{\bf X},{\bf Z})\,dt\right\}\,.$ (A.1)
 Under Assumptions A1–A2, the hazard is uniformly bounded and bounded away from zero,
-$0<c_{\lambda}\;\leq\;h_{0}(t\,|\,{\bf X},{\bf Z})\;\leq\;C_{\lambda}<\infty.$
-Applying these bounds to the right-hand side of (10) yields
-$c_{\lambda}\,\mathbb{E}\!\left\{\int_{0}^{\tau}f(t,{\bf X},{\bf Z})\,Y(t)\,dt\right\}\;\leq\;\mathbb{E}\{\Delta\,f(T,{\bf X},{\bf Z})\}\;\leq\;C_{\lambda}\,\mathbb{E}\!\left\{\int_{0}^{\tau}f(t,{\bf X},{\bf Z})\,Y(t)\,dt\right\}\,.$
+$0<c_{h}\;\leq\;h_{o}(t\,|\,{\bf X},{\bf Z})\;\leq\;C_{h}<\infty.$
+Applying these bounds to the right-hand side of (A.1) yields
+$c_{h}\,\mathbb{E}\!\left\{\int_{0}^{\tau}f(t,{\bf X},{\bf Z})\,Y(t)\,dt\right\}\;\leq\;\mathbb{E}\{\Delta\,f(T,{\bf X},{\bf Z})\}\;\leq\;C_{h}\,\mathbb{E}\!\left\{\int_{0}^{\tau}f(t,{\bf X},{\bf Z})\,Y(t)\,dt\right\}\,.$
 This implies the desired equivalence:
@@
 $\mathbb{E}\!\big\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\big\}\;\leq\;\mathbb{E}\!\big\{\phi(T,{\bf X},{\bf Z})^{2}\big\}\;\leq\;\frac{1}{\delta_{1}}\,\mathbb{E}\!\big\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\big\}\,.$
-Thus (7) holds with $c_{1}\leq 1$ and $c_{2}=1/\delta_{1}$.
+Thus (4.2) holds with $c_{1}\leq 1$ and $c_{2}=1/\delta_{1}$.
 Next, we relate the event–weighted quantity $\mathbb{E}\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\}$ to the time–integrated $L^{2}$ norm of $\phi$. Write,
-$\mathbb{E}\!\big\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\big\}=\mathbb{E}_{{\bf X},{\bf Z}}\!\Big[\mathbb{E}\!\big\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\,\big|\,{\bf X},{\bf Z}\big\}\Big]\,$, (11)
+$\mathbb{E}\!\big\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\big\}=\mathbb{E}_{{\bf X},{\bf Z}}\!\Big[\mathbb{E}\!\big\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\,\big|\,{\bf X},{\bf Z}\big\}\Big]\,$, (A.2)
 and let $f_{T,\Delta=1\mid{\bf X},{\bf Z}}(t\mid{\bf X},{\bf Z})$ denote the conditional density of $T$ given $({\bf X},{\bf Z})$ on the event $\{\Delta=1\}$. Then,
 $\mathbb{E}\!\big\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\,\big|\,{\bf X},{\bf Z}\big\}=\int_{0}^{\tau}\phi(t,{\bf X},{\bf Z})^{2}\,f_{T,\Delta=1\mid{\bf X},{\bf Z}}(t\mid{\bf X},{\bf Z})\,dt\,.$
-Under Assumptions A1 and A2, the event-time hazard $h_{0}(t,{\bf X},{\bf Z})$ is bounded away from zero and infinity.
+Under Assumptions A1 and A2, the event-time hazard $h_{o}(t,{\bf X},{\bf Z})$ is bounded away from zero and infinity.
 Similarly, Assumption A4 implies the censoring mechanism is well behaved.
@@
 $c_{f}\;\leq\;f_{T,\Delta=1\mid{\bf X},{\bf Z}}(t\mid{\bf X},{\bf Z})\;\leq\;C_{f},\qquad t\in[0,\tau]\,.$
-Using these bounds inside the conditional expectation (11), we obtain, that given $({\bf X},{\bf Z})$,
+Using these bounds inside the conditional expectation (A.2), we obtain, that given $({\bf X},{\bf Z})$,
 $c_{f}\int_{0}^{\tau}\phi(t,{\bf X},{\bf Z})^{2}\,dt\;\leq\;\mathbb{E}\!\big\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\,\big|\,{\bf X},{\bf Z}\big\}\;\leq\;C_{f}\int_{0}^{\tau}\phi(t,{\bf X},{\bf Z})^{2}\,dt\,.$
@@
 $c_{f}\,\mathbb{E}_{{\bf X},{\bf Z}}\!\Big\{\int_{0}^{\tau}\phi(t,{\bf X},{\bf Z})^{2}\,dt\Big\}\;\leq\;\mathbb{E}\!\big\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\big\}\;\leq\;C_{f}\,\mathbb{E}_{{\bf X},{\bf Z}}\!\Big\{\int_{0}^{\tau}\phi(t,{\bf X},{\bf Z})^{2}\,dt\Big\}.$
-Thus (8) holds with $c_{3}=1/C_{f}$ and $c_{4}=1/c_{f}$ (or equivalently with $c_{3},c_{4}$ chosen directly as above).
+Thus (4.3) holds with $c_{3}=1/C_{f}$ and $c_{4}=1/c_{f}$ (or equivalently with $c_{3},c_{4}$ chosen directly as above).
 Finally, suppose $\mathbb{E}\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\}=0$.
-Since $\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\geq 0$, this implies $\Delta\,\phi(T,{\bf X},{\bf Z})^{2}=0$ almost surely, hence $\phi(T,{\bf X},{\bf Z})=0$ $P_{0}$–a.s.
+Since $\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\geq 0$, this implies $\Delta\,\phi(T,{\bf X},{\bf Z})^{2}=0$ almost surely, hence $\phi(T,{\bf X},{\bf Z})=0$ $\mathbb{P}_{0}$–a.s.
 on $\{\Delta=1\}$.
-From (7) we also have
+From (4.2) we also have
 $0\;\leq\;\mathbb{E}\!\big\{\phi(T,{\bf X},{\bf Z})^{2}\big\}\;\leq\;c_{2}\,\mathbb{E}\!\big\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\big\}=0\,$,
-so $\mathbb{E}\{\phi(T,{\bf X},{\bf Z})^{2}\}=0$, which forces $\phi(T,{\bf X},{\bf Z})=0$ $P_{0}$–almost surely. Likewise, from (8),
+so $\mathbb{E}\{\phi(T,{\bf X},{\bf Z})^{2}\}=0$, which forces $\phi(T,{\bf X},{\bf Z})=0$ $\mathbb{P}_{0}$–almost surely. Likewise, from (4.3),
 $0\;\leq\;\mathbb{E}_{{\bf X},{\bf Z}}\!\Big\{\int_{0}^{\tau}\phi(t,{\bf X},{\bf Z})^{2}\,dt\Big\}\;\leq\;c_{4}\,\mathbb{E}\!\big\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\big\}=0\,$,
@@
 Since the integrand $\phi(t,{\bf X},{\bf Z})^{2}$ is nonnegative, this implies
-$\int_{0}^{\tau}\phi(t,{\bf X},{\bf Z})^{2}\,dt=0\quad\text{for $P_{{\bf X},{\bf Z}}$–a.e. }({\bf X},{\bf Z}),$
+$\int_{0}^{\tau}\phi(t,{\bf X},{\bf Z})^{2}\,dt=0\quad\text{for $\mathbb{P}_{{\bf X},{\bf Z}}$–a.e. }({\bf X},{\bf Z}),$
 and hence $\phi(t,{\bf X},{\bf Z})=0$ for Lebesgue–a.e. $t\in[0,\tau]$ and $P_{{\bf X},{\bf Z}}$–a.e. $({\bf X},{\bf Z})$. Equivalently,
-$\chi_{\eta_{1}}(t,{\bf X},{\bf Z})=\chi_{\eta_{2}}(t,{\bf X},{\bf Z})\quad\text{for }dt\otimes P_{{\bf X},{\bf Z}}\text{–almost all }(t,{\bf X},{\bf Z})\,.$
+$\chi_{\eta_{1}}(t,{\bf X},{\bf Z})=\chi_{\eta_{2}}(t,{\bf X},{\bf Z})\quad\text{for }dt\otimes\mathbb{P}_{{\bf X},{\bf Z}}\text{–almost all }(t,{\bf X},{\bf Z})\,.$
 See 3
@@
 where $\chi_{\eta}$ is the linear predictor and $h_{\eta}(t,{\bf X},{\bf Z})$ is the corresponding instantaneous hazard process.
-Let $\eta_{0}=(g_{o},\theta_{o})$ denote the true parameter, with true hazard $h_{o}(t,{\bf X},{\bf Z})=h_{\eta_{0}}(t,{\bf X},{\bf Z})$.
-We study the behavior of $\ell_{0}$ along a one–dimensional path through $\eta_{0}$. Fix an arbitrary direction $w=(w_{g},w_{\theta})$ and define
-$\eta_{\varepsilon}=\eta_{0}+\varepsilon w\,,\qquad\varepsilon\in\mathbb{R}\,.$
-Assume that $\varepsilon$ is restricted to a small neighbourhood of 0 so that $\eta_{\varepsilon}\in\mathcal{\eta}\in\Theta\times\mathcal{H}(q,\boldsymbol{\alpha},{\bf d},\tilde{{\bf d}},M)$.
+Let $\eta_{o}=(g_{o},\theta_{o})$ denote the true parameter, with true hazard $h_{o}(t,{\bf X},{\bf Z})=h_{\eta_{0}}(t,{\bf X},{\bf Z})$.
+We study the behavior of $\ell_{0}$ along a one–dimensional path through $\eta_{o}$. Fix an arbitrary direction $w=(w_{g},w_{\theta})$ and define
+$\eta_{\varepsilon}=\eta_{o}+\varepsilon w\,,\qquad\varepsilon\in\mathbb{R}\,.$
+Assume that $\varepsilon$ is restricted to a small neighbourhood of 0 so that $\eta_{\varepsilon}\in\mathcal{\eta}\in\Theta\times\mathcal{H}(q,\boldsymbol{\alpha},{\bf d}^{*},\tilde{{\bf d}},M)$.
 Because $\chi_{\eta}$ is linear in $(g,\mbox{${\theta}$})$, the Gateaux derivative of $\chi_{\eta}$ in direction $w$ at $\eta_{0}$ is simply
@@
 and we can write
-$\chi_{\eta_{\varepsilon}}(t,{\bf X},{\bf Z})=\chi_{\eta_{0}}(t,{\bf X},{\bf Z})+\varepsilon\,\dot{\chi}_{w}(t,{\bf X},{\bf Z})\,.$
+$\chi_{\eta_{\varepsilon}}(t,{\bf X},{\bf Z})=\chi_{\eta_{o}}(t,{\bf X},{\bf Z})+\varepsilon\,\dot{\chi}_{w}(t,{\bf X},{\bf Z})\,.$
 For notational simplicity, when we fix a particular $\eta$, we also write
@@
 Next, note that
-$h_{\eta_{\varepsilon}}(t,{\bf X},{\bf Z})=h_{o}(t,{\bf X},{\bf Z})\,\exp\{\chi_{\eta_{\varepsilon}}(t,{\bf X},{\bf Z})-\chi_{\eta_{0}}(t,{\bf X},{\bf Z})\}=h_{o}(t,{\bf X},{\bf Z})\,\exp\{\varepsilon\,\dot{\chi}_{w}(t,{\bf X},{\bf Z})\}\,$,
-so at $\varepsilon=0$ we have $h_{\eta_{0}}=h_{o}$ and
-$\frac{d}{d\varepsilon}h_{\eta_{\varepsilon}}(t,{\bf X},{\bf Z})\Big|_{\varepsilon=0}=h_{o}(t,{\bf X},{\bf Z})\,\dot{\chi}_{w}(t,{\bf X},{\bf Z})\,$,
-$\frac{d^{2}}{d\varepsilon^{2}}h_{\eta_{\varepsilon}}(t,{\bf X},{\bf Z})\Big|_{\varepsilon=0}=h_{o}(t,{\bf X},{\bf Z})\,\dot{\chi}_{w}(t,{\bf X},{\bf Z})^{2}\,.$
+$h_{\eta_{\varepsilon}}(t\mid{\bf X},{\bf Z})=h_{o}(t\mid{\bf X},{\bf Z})\,\exp\{\chi_{\eta_{\varepsilon}}(t,{\bf X},{\bf Z})-\chi_{\eta_{o}}(t,{\bf X},{\bf Z})\}=h_{o}(t\mid{\bf X},{\bf Z})\,\exp\{\varepsilon\,\dot{\chi}_{w}(t,{\bf X},{\bf Z})\}\,$,
+so at $\varepsilon=0$ we have $h_{\eta_{o}}=h_{o}$ and
+$\frac{d}{d\varepsilon}h_{\eta_{\varepsilon}}(t,{\bf X},{\bf Z})\Big|_{\varepsilon=0}=h_{o}(t\mid{\bf X},{\bf Z})\,\dot{\chi}_{w}(t,{\bf X},{\bf Z})\,$,
+$\frac{d^{2}}{d\varepsilon^{2}}h_{\eta_{\varepsilon}}(t,{\bf X},{\bf Z})\Big|_{\varepsilon=0}=h_{o}(t\mid{\bf X},{\bf Z})\,\dot{\chi}_{w}(t,{\bf X},{\bf Z})^{2}\,.$
 By the definition of the Gateaux derivative,
@@
 Using the explicit form of $\ell_{0}$ and differentiating under the expectation,
-$\ell_{0}^{\prime}(\eta_{0})[w]$ $=\frac{d}{d\varepsilon}\mathbb{E}\Big\{\Delta\,\chi_{\eta_{\varepsilon}}(T,{\bf X},{\bf Z})-\int_{0}^{\tau}Y(t)\,h_{\eta_{\varepsilon}}(t,{\bf X},{\bf Z})\,dt\Big\}\Big|_{\varepsilon=0}$
-$=\mathbb{E}\Big\{\Delta\,\dot{\chi}_{w}(T,{\bf X},{\bf Z})-\int_{0}^{\tau}Y(t)\,\frac{d}{d\varepsilon}h_{\eta_{\varepsilon}}(t,{\bf X},{\bf Z})\Big|_{\varepsilon=0}\,dt\Big\}$
-$=\mathbb{E}\Big\{\Delta\,\dot{\chi}_{w}(T,{\bf X},{\bf Z})-\int_{0}^{\tau}Y(t)\,h_{o}(t,{\bf X},{\bf Z})\,\dot{\chi}_{w}(t,{\bf X},{\bf Z})\,dt\Big\}\,.$
+$\ell_{0}^{\prime}(\eta_{o})[w]$ $=\frac{d}{d\varepsilon}\mathbb{E}\Big\{\Delta\,\chi_{\eta_{\varepsilon}}(T,{\bf X},{\bf Z})-\int_{0}^{\tau}Y(t)\,h_{\eta_{\varepsilon}}(t\mid{\bf X},{\bf Z})\,dt\Big\}\Big|_{\varepsilon=0}$
+$=\mathbb{E}\Big\{\Delta\,\dot{\chi}_{w}(T,{\bf X},{\bf Z})-\int_{0}^{\tau}Y(t)\,\frac{d}{d\varepsilon}h_{\eta_{\varepsilon}}(t\mid{\bf X},{\bf Z})\Big|_{\varepsilon=0}\,dt\Big\}$
+$=\mathbb{E}\Big\{\Delta\,\dot{\chi}_{w}(T,{\bf X},{\bf Z})-\int_{0}^{\tau}Y(t)\,h_{o}(t\mid{\bf X},{\bf Z})\,\dot{\chi}_{w}(t,{\bf X},{\bf Z})\,dt\Big\}\,.$
 Since
 $\Delta\,\dot{\chi}_{w}(T,{\bf X},{\bf Z})=\int_{0}^{\tau}\dot{\chi}_{w}(t,{\bf X},{\bf Z})\,dN(t)\,$,
-the Doob–Meyer decomposition for the true model (1) gives
-$dN(t)=Y(t)h_{o}(t,{\bf X},{\bf Z})\,dt+dM(t)\,$,
+the Doob–Meyer decomposition for the true model (2.1) gives
+$dN(t)=Y(t)h_{o}(t\mid{\bf X},{\bf Z})\,dt+dM(t)\,$,
 where $M(t)$ is a martingale. Therefore,
-$\Delta\,\dot{\chi}_{w}(T,{\bf X},{\bf Z})$ $=\int_{0}^{\tau}\dot{\chi}_{w}(t,{\bf X},{\bf Z})\,Y(t)h_{o}(t,{\bf X},{\bf Z})\,dt+\int_{0}^{\tau}\dot{\chi}_{w}(t,{\bf X},{\bf Z})\,dM(t)\,.$
+$\Delta\,\dot{\chi}_{w}(T,{\bf X},{\bf Z})$ $=\int_{0}^{\tau}\dot{\chi}_{w}(t,{\bf X},{\bf Z})\,Y(t)h_{o}(t\mid{\bf X},{\bf Z})\,dt+\int_{0}^{\tau}\dot{\chi}_{w}(t,{\bf X},{\bf Z})\,dM(t)\,.$
 Substituting this into the expression for $\ell_{0}^{\prime}(\eta_{o})[w]$ and cancelling the compensator term,
-$\ell_{0}^{\prime}(\eta_{0})[w]$ $=\mathbb{E}\Big\{\int_{0}^{\tau}\dot{\chi}_{h}(t,{\bf X},{\bf Z})\,dM(t)\Big\}\,.$
+$\ell_{0}^{\prime}(\eta_{o})[w]$ $=\mathbb{E}\Big\{\int_{0}^{\tau}\dot{\chi}_{h}(t,{\bf X},{\bf Z})\,dM(t)\Big\}\,.$
 The stochastic integral $\int_{0}^{\tau}\dot{\chi}_{w}(t,{\bf X},{\bf Z})\,dM(t)$ is a martingale with mean zero. Hence,
-$\ell_{0}^{\prime}(\eta_{0})[w]=0\qquad\text{for every direction }w\,.$
-In particular, when we fix $\eta$ and set $q(t,{\bf X},{\bf Z})=\chi_{\eta}(t,{\bf X},{\bf Z})-\chi_{\eta_{0}}(t,{\bf X},{\bf Z})$, we can summarize this step as
-$\ell_{0}^{\prime}(\eta_{0})[w]=\mathbb{E}\Big\{\Delta\,q(T,{\bf X},{\bf Z})-\int_{0}^{\tau}Y(t)h_{o}(t,{\bf X},{\bf Z})\,q(t,{\bf X},{\bf Z})\,dt\Big\}=0.$
-To control $\ell_{0}(\eta)-\ell_{0}(\eta_{0})$, we apply a one–dimensional Taylor expansion of $\ell_{0}(\eta_{\varepsilon})$ around $\varepsilon=0$.
-Write $\eta=\eta_{1}$ and define the line segment $\eta_{v}=\eta_{0}+v(\eta-\eta_{0})$ for $v\in[0,1]$.
+$\ell_{0}^{\prime}(\eta_{o})[w]=0\qquad\text{for every direction }w\,.$
+In particular, when we fix $\eta$ and set $q(t,{\bf X},{\bf Z})=\chi_{\eta}(t,{\bf X},{\bf Z})-\chi_{\eta_{o}}(t,{\bf X},{\bf Z})$, we can summarize this step as
+$\ell_{0}^{\prime}(\eta_{o})[w]=\mathbb{E}\Big\{\Delta\,q(T,{\bf X},{\bf Z})-\int_{0}^{\tau}Y(t)h_{o}(t\mid{\bf X},{\bf Z})\,q(t,{\bf X},{\bf Z})\,dt\Big\}=0.$
+To control $\ell_{0}(\eta)-\ell_{0}(\eta_{o})$, we apply a one–dimensional Taylor expansion of $\ell_{0}(\eta_{\varepsilon})$ around $\varepsilon=0$.
+Write $\eta=\eta_{1}$ and define the line segment $\eta_{v}=\eta_{o}+v(\eta-\eta_{o})$ for $v\in[0,1]$.
 The mean value theorem yields
-$\ell_{0}(\eta)=\ell_{0}(\eta_{0})+\ell_{0}^{\prime}(\eta_{0})[w]+\frac{1}{2}\,\ell_{0}^{\prime\prime}(\eta_{\bar{v}})[w,w]$
-for some $\bar{v}\in(0,1)$, where $w=\eta-\eta_{0}$. Since we have already shown that $\ell_{0}^{\prime}(\eta_{0})[w]=0$, it remains to analyse the second derivative term.
+$\ell_{0}(\eta)=\ell_{0}(\eta_{o})+\ell_{0}^{\prime}(\eta_{o})[w]+\frac{1}{2}\,\ell_{0}^{\prime\prime}(\eta_{\bar{v}})[w,w]$
+for some $\bar{v}\in(0,1)$, where $w=\eta-\eta_{o}$. Since we have already shown that $\ell_{0}^{\prime}(\eta_{o})[w]=0$, it remains to analyse the second derivative term.
 For a fixed $v\in[0,1]$, consider the path $\varepsilon\mapsto\eta_{v}+\varepsilon w$ and write
 $q_{v}(t,{\bf X},{\bf Z})=\frac{d}{d\varepsilon}\,\chi_{\eta_{v}+\varepsilon w}(t,{\bf X},{\bf Z})\Big|_{\varepsilon=0}\,$,
-which is again linear in $w$. Repeating the differentiation steps above but now starting from $\eta_{v}$ instead of $\eta_{0}$ gives
+which is again linear in $w$. Repeating the differentiation steps above but now starting from $\eta_{v}$ instead of $\eta_{o}$ gives
 $\ell_{0}^{\prime\prime}(\eta_{v})[w,w]$ $=\frac{d^{2}}{d\varepsilon^{2}}\,\ell_{0}(\eta_{v}+\varepsilon w)\Big|_{\varepsilon=0}$
@@
 Since $\chi_{\eta}$ is linear in $\eta$, its second derivative in $\varepsilon$ vanishes, while the second derivative of the hazard is
-$\frac{d^{2}}{d\varepsilon^{2}}h_{\eta_{v}+\varepsilon w}(t,{\bf X},{\bf Z})\Big|_{\varepsilon=0}=h_{\eta_{v}}(t,{\bf X},{\bf Z})\,q_{v}(t,{\bf X},{\bf Z})^{2}\,.$
+$\frac{d^{2}}{d\varepsilon^{2}}h_{\eta_{v}+\varepsilon w}(t\mid{\bf X},{\bf Z})\Big|_{\varepsilon=0}=h_{\eta_{v}}(t\mid{\bf X},{\bf Z})\,q_{v}(t,{\bf X},{\bf Z})^{2}\,.$
 Hence
-$\ell_{0}^{\prime\prime}(\eta_{v})[w,w]=-\mathbb{E}\Big\{\int_{0}^{\tau}Y(t)\,q_{v}(t,{\bf X},{\bf Z})^{2}\,h_{\eta_{v}}(t,{\bf X},{\bf Z})\,dt\Big\}\,.$
+$\ell_{0}^{\prime\prime}(\eta_{v})[w,w]=-\mathbb{E}\Big\{\int_{0}^{\tau}Y(t)\,q_{v}(t,{\bf X},{\bf Z})^{2}\,h_{\eta_{v}}(t\mid{\bf X},{\bf Z})\,dt\Big\}\,.$
 For notational simplicity (since we will ultimately evaluate this at $\varepsilon=1$), we drop the subscript $v$ and simply write $q(t,{\bf X},{\bf Z})=\chi_{\eta}(t,{\bf X},{\bf Z})-\chi_{\eta_{0}}(t,{\bf X},{\bf Z})$ along the line, so that
-$\ell_{0}^{\prime\prime}(\eta_{v})[w,w]=-\mathbb{E}\Big\{\int_{0}^{\tau}Y(t)\,q(t,{\bf X},{\bf Z})^{2}\,h_{\eta_{v}}(t,{\bf X},{\bf Z})\,dt\Big\}\,.$
-Next, for relating the second derivative to $d(\eta,\eta_{0})^{2}$, we apply Lemma 1 for
-$f_{v}(t,{\bf X},{\bf Z})=q(t,{\bf X},{\bf Z})^{2}\,\frac{h_{\eta_{v}}(t,{\bf X},{\bf Z})}{h_{o}(t,{\bf X},{\bf Z})}\,.$
+$\ell_{0}^{\prime\prime}(\eta_{v})[w,w]=-\mathbb{E}\Big\{\int_{0}^{\tau}Y(t)\,q(t,{\bf X},{\bf Z})^{2}\,h_{\eta_{v}}(t\mid{\bf X},{\bf Z})\,dt\Big\}\,.$
+Next, for relating the second derivative to $d(\eta,\eta_{o})^{2}$, we apply Lemma 1 for
+$f_{v}(t,{\bf X},{\bf Z})=q(t,{\bf X},{\bf Z})^{2}\,\frac{h_{\eta_{v}}(t\mid{\bf X},{\bf Z})}{h_{o}(t\mid{\bf X},{\bf Z})}\,.$
 Indeed, Lemma 1 gives
-$\mathbb{E}\Big\{\int_{0}^{\tau}Y(t)\,q(t,{\bf X},{\bf Z})^{2}\,h_{\eta_{v}}(t,{\bf X},{\bf Z})\,dt\Big\}\asymp\mathbb{E}\Big\{\Delta\,q(T,{\bf X},{\bf Z})^{2}\,\frac{h_{\eta_{v}}(T,{\bf X},{\bf Z})}{h_{o}(T,{\bf X},{\bf Z})}\Big\}\,.$
+$\mathbb{E}\Big\{\int_{0}^{\tau}Y(t)\,q(t,{\bf X},{\bf Z})^{2}\,h_{\eta_{v}}(t\mid{\bf X},{\bf Z})\,dt\Big\}\asymp\mathbb{E}\Big\{\Delta\,q(T,{\bf X},{\bf Z})^{2}\,\frac{h_{\eta_{v}}(T\mid{\bf X},{\bf Z})}{h_{o}(T\mid{\bf X},{\bf Z})}\Big\}\,.$
 Under Assumptions A1A2, the processes $\chi_{\eta}$ are uniformly bounded, and hence
-$0<c_{\star}\;\leq\;\frac{h_{\eta_{v}}(t,{\bf X},{\bf Z})}{h_{o}(t,{\bf X},{\bf Z})}\;\leq\;C_{\star}<\infty,\qquad\text{for all }(t,{\bf X},{\bf Z}),\;v\in[0,1]\,.$
+$0<c_{\star}\;\leq\;\frac{h_{\eta_{v}}(t\mid{\bf X},{\bf Z})}{h_{o}(t\mid{\bf X},{\bf Z})}\;\leq\;C_{\star}<\infty,\qquad\text{for all }(t,{\bf X},{\bf Z}),\;v\in[0,1]\,.$
 Therefore,
-$c_{\star}\,\mathbb{E}\big[\Delta\,q(T,{\bf X},{\bf Z})^{2}\big]\;\leq\;\mathbb{E}\Big\{\Delta\,q(T,{\bf X},{\bf Z})^{2}\,\frac{h_{\eta_{v}}(T,{\bf X},{\bf Z})}{h_{o}(T,{\bf X},{\bf Z})}\Big]\;\leq\;C_{\star}\,\mathbb{E}\big[\Delta\,q(T,{\bf X},{\bf Z})^{2}\big\}\,.$
+$c_{\star}\,\mathbb{E}\big[\Delta\,q(T,{\bf X},{\bf Z})^{2}\big]\;\leq\;\mathbb{E}\Big\{\Delta\,q(T,{\bf X},{\bf Z})^{2}\,\frac{h_{\eta_{v}}(T\mid{\bf X},{\bf Z})}{h_{o}(T\mid{\bf X},{\bf Z})}\Big]\;\leq\;C_{\star}\,\mathbb{E}\big[\Delta\,q(T,{\bf X},{\bf Z})^{2}\big\}\,.$
 Combining this with the expression for $\ell_{0}^{\prime\prime}(\eta_{v})[w,w]$ yields the two-sided bound
 $C_{\star}\,\mathbb{E}\big\{\Delta\,q(T,{\bf X},{\bf Z})^{2}\big\}\;\leq\;\ell_{0}^{\prime\prime}(\eta_{v})[w,w]\;\leq\;-c_{\star}\,\mathbb{E}\big\{\Delta\,q(T,{\bf X},{\bf Z})^{2}\big\}.$
-Recall that the distance $d(\eta,\eta_{0})$ is defined by
-$d(\eta,\eta_{0})^{2}=\mathbb{E}\Big\{\Delta\big\{\chi_{\eta}(T,{\bf X},{\bf Z})-\chi_{\eta_{0}}(T,{\bf X},{\bf Z})\big\}^{2}\Big\}=\mathbb{E}\big\{\Delta\,q(T,{\bf X},{\bf Z})^{2}\big\}\,.$
+Recall that the distance $d(\eta,\eta_{o})$ is defined by
+$d(\eta,\eta_{o})^{2}=\mathbb{E}\Big\{\Delta\big\{\chi_{\eta}(T,{\bf X},{\bf Z})-\chi_{\eta_{o}}(T,{\bf X},{\bf Z})\big\}^{2}\Big\}=\mathbb{E}\big\{\Delta\,q(T,{\bf X},{\bf Z})^{2}\big\}\,.$
 Hence
-$C_{\star}\,d(\eta,\eta_{0})^{2}\;\leq\;\ell_{0}^{\prime\prime}(\eta_{v})[w,w]\;\leq\;-c_{\star}\,d(\eta,\eta_{0})^{2}\,.$
+$C_{\star}\,d(\eta,\eta_{o})^{2}\;\leq\;\ell_{0}^{\prime\prime}(\eta_{v})[w,w]\;\leq\;-c_{\star}\,d(\eta,\eta_{o})^{2}\,.$
 Therefore, by the Taylor expansion with the mean value theorem,
-$\ell_{0}(\eta)-\ell_{0}(\eta_{0})=\frac{1}{2}\,\ell_{0}^{\prime\prime}(\eta_{\bar{v}})[w,w]$
-for some $\bar{v}\in(0,1)$, and we have just shown that $\ell_{0}^{\prime\prime}(\eta_{\bar{v}})[w,w]$ is comparable to $d(\eta,\eta_{0})^{2}$. Thus,
-$\ell_{0}(\eta)-\ell_{0}(\eta_{0})\;\asymp\;-\,d(\eta,\eta_{0})^{2},$
+$\ell_{0}(\eta)-\ell_{0}(\eta_{o})=\frac{1}{2}\,\ell_{0}^{\prime\prime}(\eta_{\bar{v}})[w,w]$
+for some $\bar{v}\in(0,1)$, and we have just shown that $\ell_{0}^{\prime\prime}(\eta_{\bar{v}})[w,w]$ is comparable to $d(\eta,\eta_{o})^{2}$. Thus,
+$\ell_{0}(\eta)-\ell_{0}(\eta_{o})\;\asymp\;-\,d(\eta,\eta_{o})^{2},$
 which completes the proof. ∎
@@
   • The efficient score-type term
-    $\psi_{\eta}(T,\mathbf{X},\mathbf{Z},\Delta)=\int_{0}^{T}\bigl\{\mathbf{Z}-g^{\ast}(t,\mathbf{X})\bigr\}\,dM_{\eta}(t),$
-    where $g^{\ast}$ denotes the efficient projection defined in Theorem 3.
+    $\psi_{\eta}(T,\mathbf{X},\mathbf{Z},\Delta)=\int_{0}^{T}\bigl\{\mathbf{Z}-{\bf g}^{\ast}(t,\mathbf{X})\bigr\}\,dM_{\eta}(t),$
+    where $\bf g$$^{\ast}$ denotes the efficient projection defined in Theorem 3.
 For $0<\delta\leq 1$, define the localized classes
-$\mathcal{F}_{\delta}=\{\,f_{\eta}-f_{\eta_{0}}:d(\eta,\eta_{0})\leq\delta\,\}\,,\qquad\mathcal{G}_{\delta}=\{\,\psi_{\eta}-\psi_{\eta_{0}}:d(\eta,\eta_{0})\leq\delta\,\}\,.$
+$\mathcal{F}_{\delta}=\{\,f_{\eta}-f_{\eta_{o}}:d(\eta,\eta_{o})\leq\delta\,\}\,,\qquad\mathcal{G}_{\delta}=\{\,\psi_{\eta}-\psi_{\eta_{o}}:d(\eta,\eta_{o})\leq\delta\,\}\,.$
 Under Assumptions A1A3 and A5, there exists a constant $C>0$ such that for every $\delta\in(0,1]$,
@@
 The right-hand side is exactly the squared metric $d^{2}(\eta_{1},\eta_{2})$. Thus,
-$\|f_{\eta_{1}}-f_{\eta_{2}}\|_{L^{2}(P)}\;\lesssim\;d(\eta_{1},\eta_{2})\,.$
+$\|f_{\eta_{1}}-f_{\eta_{2}}\|_{L^{2}(\mathbb{P})}\;\lesssim\;d(\eta_{1},\eta_{2})\,.$
 Repeating the argument for $\psi_{\eta}$ requires only replacing the exponential term with the stochastic integral.
@@
 The same calculation shows
-$\|\psi_{\eta_{1}}-\psi_{\eta_{2}}\|_{L^{2}(P)}\;\lesssim\;d(\eta_{1},\eta_{2})\,.$
-Hence both localized classes $\mathcal{F}_{\delta}$ and $\mathcal{G}_{\delta}$ lie in a common $L^{2}(P)$-ball of radius $O(\delta)$ and admit a common envelope of order $O(\delta)$.
+$\|\psi_{\eta_{1}}-\psi_{\eta_{2}}\|_{L^{2}(\mathbb{P})}\;\lesssim\;d(\eta_{1},\eta_{2})\,.$
+Hence both localized classes $\mathcal{F}_{\delta}$ and $\mathcal{G}_{\delta}$ lie in a common $L^{2}(\mathbb{P})$-ball of radius $O(\delta)$ and admit a common envelope of order $O(\delta)$.
 The parametrization $\eta=(\mbox{${\theta}$},g)$ is $s$-sparse in the neural-network coordinates, and Assumption A4 ensures that the weights are uniformly bounded and that the architecture has fixed depth and layer widths.
-In particular, the map $\eta\mapsto\chi_{\eta}$ belongs to the class of sparse feedforward neural networks considered in zhong2022deep.
-Moreover, the maps $\eta\longmapsto f_{\eta}$ and $\eta\longmapsto\psi_{\eta}$ are Lipschitz transformations of $\chi_{\eta}$ in the $L^{2}(P)$-metric, with Lipschitz constant bounded uniformly under Assumptions A1A3 and A5.
+In particular, the map $\eta\mapsto\chi_{\eta}$ belongs to the class of sparse feedforward neural networks considered in Zhong et al.
+(2022).
+Moreover, the maps $\eta\longmapsto f_{\eta}$ and $\eta\longmapsto\psi_{\eta}$ are Lipschitz transformations of $\chi_{\eta}$ in the $L^{2}(\mathbb{P})$-metric, with Lipschitz constant bounded uniformly under Assumptions A1A3 and A5.
 Since Lipschitz images of a function class cannot increase its covering number by more than a multiplicative constant, the metric entropy of the localized classes $\mathcal{F}_{\delta}$ and $\mathcal{G}_{\delta}$ is therefore controlled by the entropy of the underlying sparse neural-network ball
-$\mathcal{H}_{\delta}=\{\chi_{\eta}-\chi_{\eta_{0}}:d(\eta,\eta_{0})\leq\delta\}.$
-Lemma 6 of zhong2022deep establishes that for such a sparse network class,
+$\mathcal{H}_{\delta}=\{\chi_{\eta}-\chi_{\eta_{o}}:d(\eta,\eta_{o})\leq\delta\}.$
+Lemma 6 of Zhong et al. (2022) establishes that for such a sparse network class,
 $\log N_{[]}\!\left(\varepsilon,\,\mathcal{H}_{\delta},\,L^{2}(P)\right)\;\lesssim\;s\log\!\left(\frac{\mathcal{U}\delta}{\varepsilon}\right),$
 where $\mathcal{U}$ is the architectural complexity factor.
-Because $\mathcal{F}_{\delta}$ and $\mathcal{G}_{\delta}$ lie in a common $L^{2}(P)$-ball of radius $O(\delta)$ and share a common envelope of order $O(\delta)$, the same entropy bound applies to their union.
+Because $\mathcal{F}_{\delta}$ and $\mathcal{G}_{\delta}$ lie in a common $L^{2}(\mathbb{P})$-ball of radius $O(\delta)$ and share a common envelope of order $O(\delta)$, the same entropy bound applies to their union.
 Consequently,
-$\log N_{[]}\!\left(\varepsilon,\,\mathcal{F}_{\delta}\cup\mathcal{G}_{\delta},\,L^{2}(P)\right)\;\lesssim\;s\log\!\left(\frac{\mathcal{U}\delta}{\varepsilon}\right)\,.$
+$\log N_{[]}\!\left(\varepsilon,\,\mathcal{F}_{\delta}\cup\mathcal{G}_{\delta},\,L^{2}(\mathbb{P})\right)\;\lesssim\;s\log\!\left(\frac{\mathcal{U}\delta}{\varepsilon}\right)\,.$
 Next, we bound the bracketing entropy integral. Using the entropy estimate derived above, the associated Dudley integral satisfies
 $J_{[]}(\delta)=\int_{0}^{\delta}\!\sqrt{\log N_{[]}(\varepsilon)}\;d\varepsilon\;\leq\;C_{2}\,\delta\,\sqrt{s\log\!\left(\frac{\mathcal{U}}{\delta}\right)}\,.$
-Finally, applying Lemma 3.4.2 of vdvwellner1996 to the empirical process $\mathbb{G}_{n}=\sqrt{n}(\mathbb{P}_{n}-\mathbb{P})$, with envelope $F_{\delta}=O(\delta)$, gives
-$\mathbb{E}^{\ast}\!\left\{\sup_{\mu\in\mathcal{F}_{\delta}\cup\mathcal{G}_{\delta}}|\mathbb{G}_{n}\mu|\right\}\;\leq\;C_{3}\Bigl\{J_{[]}(\delta)+n^{-1/2}\,\|F_{\delta}\|_{L^{2}(P)}\,\log N_{[]}(\delta)\Bigr\}\,.$
-Substituting the bound on $J_{[]}(\delta)$ and noting that $\|F_{\delta}\|_{L^{2}(P)}=O(\delta)$ and $\log N_{[]}(\delta)\lesssim s\log(\mathcal{U}/\delta)$ yields the asserted maximal inequality. ∎
+Finally, applying Lemma 3.4.2 of van der Vaart and Wellner (1996) to the empirical process $\mathbb{G}_{n}=\sqrt{n}(\mathbb{P}_{n}-\mathbb{P})$, with envelope $F_{\delta}=O(\delta)$, gives
+$\mathbb{E}^{\ast}\!\left\{\sup_{\mu\in\mathcal{F}_{\delta}\cup\mathcal{G}_{\delta}}|\mathbb{G}_{n}\mu|\right\}\;\leq\;C_{3}\Bigl\{J_{[]}(\delta)+n^{-1/2}\,\|F_{\delta}\|_{L^{2}(\mathbb{P})}\,\log N_{[]}(\delta)\Bigr\}\,.$
+Substituting the bound on $J_{[]}(\delta)$ and noting that $\|F_{\delta}\|_{L^{2}(\mathbb{P})}=O(\delta)$ and $\log N_{[]}(\delta)\lesssim s\log(\mathcal{U}/\delta)$ yields the asserted maximal inequality.
+∎
 The following lemma establishes equivalence of expectations under two models whose hazard functions differ only through the nuisance function $g$.
@@
 Suppose Assumptions A1A2 and A4 hold.
-Let $g_{0},g_{1}\in\mathcal{H}(q,\boldsymbol{\alpha},{\bf d},\tilde{{\bf d}},M)$ and fix $\mbox{${\theta}$}\in\Theta$.
-For $j=0,1$, let $P_{g_{j}}$ denote the distribution of $(T,{\bf X},{\bf Z},\Delta)$ under event hazard $h_{g_{j}}(t\mid{\bf X},{\bf Z})=\exp\{g_{j}(t,{\bf X})+\mbox{${\theta}$}^{\top}{\bf Z}\}$ and a common censoring hazard $\mu(t\mid{\bf X},{\bf Z})$.
+Let $g_{0},g_{1}\in\mathcal{H}(q,\boldsymbol{\alpha},{\bf d}^{*},\tilde{{\bf d}},M)$ and fix $\mbox{${\theta}$}\in\Theta$.
+For $j=0,1$, let $\mathbb{P}_{g_{j}}$ denote the distribution of $(T,{\bf X},{\bf Z},\Delta)$ under event hazard $h_{g_{j}}(t\mid{\bf X},{\bf Z})=\exp\{g_{j}(t,{\bf X})+\mbox{${\theta}$}^{\top}{\bf Z}\}$ and a common censoring hazard $\mu(t\mid{\bf X},{\bf Z})$.
 As a consequence of Assumptions A1A2, there exists a constant $0<D<\infty$ such that, almost surely,
 $e^{-D}\;\leq\;h_{g_{j}}(t\mid\boldsymbol{X},\boldsymbol{Z})\;\leq\;e^{D}\,,\qquad\mbox{for all}\,\,t\in[0,\tau]\,,\ j=0,1\,.$
-Define $c=e^{-D}$ and $C=e^{D}$. Then, for any nonnegative measurable function $q:[0,\tau]\times\mathcal{X}\to[0,\infty)$ with $\mathbb{E}_{P_{g_{0}}}[q(T,{\bf X})]<\infty$,
-$K_{-}\,\mathbb{E}_{P_{g_{0}}}\!\big\{q(T,{\bf X})\big\}\;\leq\;\mathbb{E}_{P_{g_{1}}}\!\big\{q(T,{\bf X})\big\}\;\leq\;K_{+}\,\mathbb{E}_{P_{g_{0}}}\!\big\{q(T,{\bf X})\big\}\,$,
+Define $c=e^{-D}$ and $C=e^{D}$. Then, for any nonnegative measurable function $q:[0,\tau]\times\mathcal{X}\to[0,\infty)$ with $\mathbb{E}_{\mathbb{P}_{g_{0}}}[q(T,{\bf X})]<\infty$,
+$K_{-}\,\mathbb{E}_{\mathbb{P}_{g_{0}}}\!\big\{q(T,{\bf X})\big\}\;\leq\;\mathbb{E}_{\mathbb{P}_{g_{1}}}\!\big\{q(T,{\bf X})\big\}\;\leq\;K_{+}\,\mathbb{E}_{\mathbb{P}_{g_{0}}}\!\big\{q(T,{\bf X})\big\}\,$,
 where
 $K_{-}=\frac{c}{C}\,e^{-\tau(C-c)}\,,\qquad K_{+}=\frac{C}{c}\,e^{\,\tau(C-c)}\,.$
-In particular, $\mathbb{E}_{P_{g_{1}}}[q(T,{\bf X})]\asymp\mathbb{E}_{P_{g_{0}}}[q(T,{\bf X})]$ with constants depending only on $(c,C,\tau)$.
+In particular, $\mathbb{E}_{\mathbb{P}_{g_{1}}}[q(T,{\bf X})]\asymp\mathbb{E}_{\mathbb{P}_{g_{0}}}[q(T,{\bf X})]$ with constants depending only on $(c,C,\tau)$.
 Proof of Lemma 11.
@@
 Thus the Radon–Nikodým derivative satisfies
-$K_{-}\;\leq\;\frac{dP_{g_{1}}}{dP_{g_{0}}}(T,\Delta,{\bf X},{\bf Z})\;\leq\;K_{+},\qquad P_{g_{0}}\text{--a.s.},$
-where $K_{-}=\frac{c}{C}e^{-\tau(C-c)}$ and $K_{+}=\frac{C}{c}e^{\tau(C-c)}$. For any nonnegative measurable function $q$ with finite expectation under $P_{g_{0}}$,
-$\mathbb{E}_{P_{g_{1}}}\{q(T,{\bf X})\}=\mathbb{E}_{P_{g_{0}}}\!\left\{q(T,{\bf X})\,\frac{dP_{g_{1}}}{dP_{g_{0}}}(T,\Delta,{\bf X},{\bf Z})\right\}$
+$K_{-}\;\leq\;\frac{d\mathbb{P}_{g_{1}}}{d\mathbb{P}_{g_{0}}}(T,\Delta,{\bf X},{\bf Z})\;\leq\;K_{+},\qquad\mathbb{P}_{g_{0}}\text{--a.s.},$
+where $K_{-}=\frac{c}{C}e^{-\tau(C-c)}$ and $K_{+}=\frac{C}{c}e^{\tau(C-c)}$. For any nonnegative measurable function $q$ with finite expectation under $\mathbb{P}_{g_{0}}$,
+$\mathbb{E}_{\mathbb{P}_{g_{1}}}\{q(T,{\bf X})\}=\mathbb{E}_{\mathbb{P}_{g_{0}}}\!\left\{q(T,{\bf X})\,\frac{d\mathbb{P}_{g_{1}}}{d\mathbb{P}_{g_{0}}}(T,\Delta,{\bf X},{\bf Z})\right\}$
 and the bounds on the likelihood ratio give
-$K_{-}\,\mathbb{E}_{P_{g_{0}}}\{q(T,{\bf X})\}\;\leq\;\mathbb{E}_{P_{g_{1}}}\{q(T,{\bf X})\}\;\leq\;K_{+}\,\mathbb{E}_{P_{g_{0}}}\{q(T,{\bf X})\}\,.$
+$K_{-}\,\mathbb{E}_{\mathbb{P}_{g_{0}}}\{q(T,{\bf X})\}\;\leq\;\mathbb{E}_{\mathbb{P}_{g_{1}}}\{q(T,{\bf X})\}\;\leq\;K_{+}\,\mathbb{E}_{\mathbb{P}_{g_{0}}}\{q(T,{\bf X})\}\,.$
 Lemma 12.
 Assume A2.
-Let $\eta_{j}=(\mbox{${\theta}$}_{j},g_{j})$, $j=1,2$, be two parameter values in $(\Theta\times\mathcal{H}(q,\boldsymbol{\alpha},{\bf d},\tilde{{\bf d}},M))\cup(\Theta\times\mathcal{G}(K,s,\boldsymbol{p},D))$.
+Let $\eta_{j}=(\mbox{${\theta}$}_{j},g_{j})$, $j=1,2$, be two parameter values in $(\Theta\times\mathcal{H}(q,\boldsymbol{\alpha},{\bf d}^{*},\tilde{{\bf d}},M))\cup(\Theta\times\mathcal{G}(K,s,\boldsymbol{p},D))$.
 Define
@@
 Lemma 13.
-If Assumptions A1A5 hold, then there exists a constant $c_{\theta}>0$ depending only on the constants in A1A4 such that, for all parameters in a neighborhood of $\eta_{o}$,
+If Assumptions A1A5 hold, then there exists a constant $c_{\theta}>0$ such that, for all parameters $\eta_{1}=(\mbox{${\theta}$}_{1},g_{1})$ and $\eta_{2}=(\mbox{${\theta}$}_{2},g_{2})$,
 $d(\eta_{1},\eta_{2})^{2}\;\geq\;c_{\theta}\,\|\mbox{${\theta}$}_{1}-\mbox{${\theta}$}_{2}\|^{2}\,.$
+In particular, one may take
+$c_{\theta}=\lambda_{\min}(\Sigma_{{\bf Z}})>0.$
 Proof of Lemma 13.
-Let $\mbox{${\beta}$}=\mbox{${\theta}$}_{1}-\mbox{${\theta}$}_{2}$, $a(t,{\bf X})=g_{1}(t,{\bf X})-g_{2}(t,{\bf X})$, $b=\mbox{${\beta}$}^{\top}{\bf Z}$.
-At a failure time, $\chi_{\eta_{1}}-\chi_{\eta_{2}}=a+b$.
-Thus
-$d(\eta_{1},\eta_{2})^{2}=\mathbb{E}_{0}\{\Delta(a+b)^{2}\}\,.$
-To find the smallest possible value of this expression over all possible nuisance differences $g_{1}-g_{2}$, minimize $(a+b)^{2}$ over scalar functions $a(T,{\bf X})$. The minimizing choice is
-$a^{\ast}(T,{\bf X})=-\mathbb{E}_{0}\{b\mid T,{\bf X},\Delta=1\}\,.$
-Substituting this optimal choice yields
-$\inf_{g_{1}-g_{2}}d(\eta_{1},\eta_{2})^{2}=\mathbb{E}_{0}\!\Big\{\Delta\,Var(b\mid T,{\bf X},\Delta=1)\Big\}\,.$
-This is exactly the standard least–squares argument. Assumption A4 ensures
-$\Pr(\Delta=1\mid T,{\bf X})\;\geq\;c_{\Delta}\quad\text{a.s.}$
+Let $\mbox{${\beta}$}=\mbox{${\theta}$}_{1}-\mbox{${\theta}$}_{2}$, $a(T,{\bf X})=g_{1}(T,{\bf X})-g_{2}(T,{\bf X})$, $b=\mbox{${\beta}$}^{\top}{\bf Z}$. Then
+$\chi_{\eta_{1}}(T,{\bf X},{\bf Z})-\chi_{\eta_{2}}(T,{\bf X},{\bf Z})=a(T,{\bf X})+b,$
+and hence
+$d(\eta_{1},\eta_{2})^{2}=\mathbb{E}_{0}\!\left[\Delta\{a(T,{\bf X})+b\}^{2}\right].$
+Because $a(T,{\bf X})$ is measurable with respect to $(T,{\bf X})$, we have
+$\mathbb{E}_{0}\!\left[\Delta\{a+b\}^{2}\mid T,{\bf X}\right]=\Pr(\Delta=1\mid T,{\bf X})\,\mathbb{E}_{0}\!\left[\{a+b\}^{2}\mid T,{\bf X},\Delta=1\right].$
+By the conditional variance decomposition,
+$\mathbb{E}_{0}\!\left[\{a+b\}^{2}\mid T,{\bf X},\Delta=1\right]=\Big(a+\mathbb{E}_{0}[b\mid T,{\bf X},\Delta=1]\Big)^{2}+Var_{0}(b\mid T,{\bf X},\Delta=1).$
 Therefore
-$\mathbb{E}_{0}\Big[\Pr(\Delta=1\mid T,{\bf X})\,Var(b\mid T,{\bf X},\Delta=1)\Big]\;\geq\;c_{\Delta}\,\mathbb{E}_{0}\bigl[Var(b\mid T,{\bf X},\Delta=1)\bigr]\,.$
-Assumption A5 implies that for some $c_{Z}>0$,
-$Var({\bf Z}\mid T,{\bf X},\Delta=1)\succeq c_{Z}I_{p}\,$,
-thus for any $\beta$},
-$Var(b\mid T,{\bf X},\Delta=1)=Var(\mbox{${\beta}$}^{\top}{\bf Z}\mid T,{\bf X},\Delta=1)\geq c_{\Delta}c_{Z}\|\mbox{${\beta}$}\|^{2}\,.$
-Combining the above, we get
-$d(\eta_{1},\eta_{2})^{2}\geq c_{Z}\|\mbox{${\theta}$}_{1}-\mbox{${\theta}$}_{2}\|^{2}.$
-Setting $c_{\theta}=c_{\Delta}c_{Z}>0$ completes the proof. ∎
+$\mathbb{E}_{0}\!\left[\Delta\{a+b\}^{2}\right]=\mathbb{E}_{0}\!\left[\Pr(\Delta=1\mid T,{\bf X})\Big(a+\mathbb{E}_{0}[b\mid T,{\bf X},\Delta=1]\Big)^{2}\right]+\mathbb{E}_{0}\!\left[\Delta\,Var_{0}(b\mid T,{\bf X},\Delta=1)\right].$
+Since the first term is nonnegative,
+$d(\eta_{1},\eta_{2})^{2}\geq\mathbb{E}_{0}\!\left[\Delta\,Var_{0}(b\mid T,{\bf X},\Delta=1)\right].$
+Now, because $b=\mbox{${\beta}$}^{\top}{\bf Z}$,
+$Var_{0}(b\mid T,{\bf X},\Delta=1)=\mbox{${\beta}$}^{\top}Var_{0}({\bf Z}\mid T,{\bf X},\Delta=1)\mbox{${\beta}$},$
+and thus
+$\mathbb{E}_{0}\!\left[\Delta\,Var_{0}(b\mid T,{\bf X},\Delta=1)\right]=\mbox{${\beta}$}^{\top}\mathbb{E}_{0}\!\left[\Delta\,\Big\{{\bf Z}-\mathbb{E}_{0}({\bf Z}\mid T,{\bf X},\Delta=1)\Big\}^{\otimes 2}\right]\mbox{${\beta}$}=\mbox{${\beta}$}^{\top}\Sigma_{{\bf Z}}\mbox{${\beta}$}.$
+By Assumption A5,
+$\mbox{${\beta}$}^{\top}\Sigma_{{\bf Z}}\mbox{${\beta}$}\geq\lambda_{\min}(\Sigma_{\bf Z})\,\|\mbox{${\beta}$}\|^{2}.$
+Hence
+$d(\eta_{1},\eta_{2})^{2}\geq\lambda_{\min}(\Sigma_{\bf Z})\,\|\mbox{${\theta}$}_{1}-\mbox{${\theta}$}_{2}\|^{2}.$
+This proves the lemma with $c_{\theta}=\lambda_{\min}(\Sigma_{\bf Z})$. ∎
 See 1
@@
 and let $\eta_{o}=(\mbox{${\theta}$}_{o},g_{o})$ denote the true parameter. By Lemma 10, there exists a constant $C>0$ such that for every $0<\delta\leq 1$,
-$\mathbb{E}^{*}\!\sup_{d(\eta,\eta_{o})\leq\delta}\bigl||\mathbb{G}_{n}\big\{\ell(\eta;\cdot)-\ell(\eta_{o};\cdot)\big\}\bigr|\;\leq\;\Psi_{n}(\delta)\equiv\delta\,\sqrt{s\log\frac{\mathcal{U}}{\delta}}+\frac{s}{\sqrt{n}}\log\frac{\mathcal{U}}{\delta}\,$, (12)
+$\mathbb{E}^{*}\!\sup_{d(\eta,\eta_{o})\leq\delta}\bigl|\mathbb{G}_{n}\big\{\ell(\eta;\cdot)-\ell(\eta_{o};\cdot)\big\}\bigr|\;\leq\;\Psi_{n}(\delta)\equiv\delta\,\sqrt{s\log\frac{\mathcal{U}}{\delta}}+\frac{s}{\sqrt{n}}\log\frac{\mathcal{U}}{\delta}\,$, (A.3)
 and by Markov’s inequality, this implies
@@
 By Assumption A3 and the specific choice of $(K,s,\boldsymbol{p})$ (cf.
-the construction in schmidthieber2020nonparametric and zhong2022deep), we have
+the construction in Schmidt-Hieber (2020) and Zhong et al.
+(2022)), we have
 $s\log\!\Bigl(\frac{\mathcal{U}}{\tau_{n}}\Bigr)=O\!\bigl(n\,\gamma_{n}^{2}\log n\bigr)\,$,
@@
 for $n$ large enough. Hence, there exists a constant $c>0$ such that
-$\Psi_{n}(\tau_{n})\;\leq\;\frac{c}{2}\,\sqrt{n}\,\tau_{n}^{2}\quad\mbox{and}\quad\frac{1}{\sqrt{n}}\Psi_{n}(\tau_{n})\;\leq\;\frac{c}{2}\,\tau_{n}^{2}\,.$ (13)
+$\Psi_{n}(\tau_{n})\;\leq\;\frac{c}{2}\,\sqrt{n}\,\tau_{n}^{2}\quad\mbox{and}\quad\frac{1}{\sqrt{n}}\Psi_{n}(\tau_{n})\;\leq\;\frac{c}{2}\,\tau_{n}^{2}\,.$ (A.4)
 Consequently,
-$\sup_{d(\eta,\eta_{o})\leq\tau_{n}}\bigl|(\ell_{n}-\ell_{0})(\eta)-(\ell_{n}-\ell_{0})(\eta_{o})\bigr|=\sup_{d(\eta,\eta_{o})\leq\tau_{n}}\frac{1}{\sqrt{n}}\bigl|\mathbb{G}_{n}\big\{\ell(\eta;\cdot)-\ell(\eta_{o};\cdot)\big\}\bigr|=O_{p}\bigl(\tfrac{1}{\sqrt{n}}\Psi_{n}(\tau_{n})\bigr)=O_{p}(\tau_{n}^{2}),$ (14)
-where we used (12) and (13).
-By Theorem 1 of schmidthieber2020nonparametric, applied to the composite Hölder class $\mathcal{H}(q,\boldsymbol{\alpha},{\bf d},\tilde{\bf d},M)$ with the choice of $(K,s,\boldsymbol{p},D)$ given in Assumption A3, there exists $\tilde{g}_{1}\in\mathcal{G}(K,s,\boldsymbol{p},D)$ such that
+$\sup_{d(\eta,\eta_{o})\leq\tau_{n}}\bigl|(\ell_{n}-\ell_{0})(\eta)-(\ell_{n}-\ell_{0})(\eta_{o})\bigr|=\sup_{d(\eta,\eta_{o})\leq\tau_{n}}\frac{1}{\sqrt{n}}\bigl|\mathbb{G}_{n}\big\{\ell(\eta;\cdot)-\ell(\eta_{o};\cdot)\big\}\bigr|=O_{p}\bigl(\tfrac{1}{\sqrt{n}}\Psi_{n}(\tau_{n})\bigr)=O_{p}(\tau_{n}^{2}),$ (A.5)
+where we used (A.3) and (A.4).
+By Theorem 1 of Schmidt-Hieber (2020), applied to the composite Hölder class $\mathcal{H}(q,\boldsymbol{\alpha},{\bf d}^{*},\tilde{\bf d},M)$ with the choice of $(K,s,\boldsymbol{p},D)$ given in Assumption A3, there exists $\tilde{g}_{1}\in\mathcal{G}(K,s,\boldsymbol{p},D)$ such that
 $\mathbb{E}\bigl\{\tilde{g}_{1}(T,{\bf X})-g_{o}(T,{\bf X})\bigr\}^{2}\leq C_{1}\,\tau_{n}^{2}\,,\qquad\mathbb{E}\bigl\{\Delta\big(\tilde{g}_{1}(T,{\bf X})-g_{o}(T,{\bf X})\big)^{2}\bigr\}\leq C_{1}\,\tau_{n}^{2}$
@@
 since the $\theta$}–component is the same in both $\tilde{\eta}_{1}$ and $\eta_{o}$. Therefore,
-$d(\tilde{\eta}_{1},\eta_{o})\;\leq\;\sqrt{C_{1}}\,\tau_{n}\,$, (15)
+$d(\tilde{\eta}_{1},\eta_{o})\;\leq\;\sqrt{C_{1}}\,\tau_{n}\,$, (A.6)
 and in particular $\tilde{\eta}_{1}$ lies inside a ball of radius proportional to $\tau_{n}$ around $\eta_{o}$.
@@
 $c_{+}\,d(\eta,\eta_{o})^{2}\;\leq\;\ell_{0}(\eta)-\ell_{0}(\eta_{o})\;\leq\;-\,c_{-}\,d(\eta,\eta_{o})^{2}\,.$
-Applying this to $\eta=\tilde{\eta}_{1}$ and using (15), we obtain
-$\ell_{0}(\tilde{\eta}_{1})-\ell_{0}(\eta_{o})\;\geq\;-\,c_{+}\,d(\tilde{\eta}_{1},\eta_{o})^{2}\;\geq\;-\,c_{+}C_{1}\,\tau_{n}^{2}\;=\;-\,C_{2}\,\tau_{n}^{2}\,$, (16)
+Applying this to $\eta=\tilde{\eta}_{1}$ and using (A.6), we obtain
+$\ell_{0}(\tilde{\eta}_{1})-\ell_{0}(\eta_{o})\;\geq\;-\,c_{+}\,d(\tilde{\eta}_{1},\eta_{o})^{2}\;\geq\;-\,c_{+}C_{1}\,\tau_{n}^{2}\;=\;-\,C_{2}\,\tau_{n}^{2}\,$, (A.7)
 for some constant $C_{2}>0$.
-By combining (14) and (16) we get
+By combining (A.5) and (A.7) we get
 $\ell_{n}(\tilde{\eta}_{1})-\ell_{n}(\eta_{o})$ $=\bigl\{\ell_{n}(\tilde{\eta}_{1})-\ell_{0}(\tilde{\eta}_{1})\bigr\}-\bigl\{\ell_{n}(\eta_{o})-\ell_{0}(\eta_{o})\bigr\}+\bigl\{\ell_{0}(\tilde{\eta}_{1})-\ell_{0}(\eta_{o})\bigr\}$
 $\geq-\,\sup_{d(\eta,\eta_{o})\leq\tau_{n}}\bigl|(\ell_{n}-\ell_{0})(\eta)-(\ell_{n}-\ell_{0})(\eta_{o})\bigr|\;-\;C_{2}\,\tau_{n}^{2}$
-$=-\,O_{p}(\tau_{n}^{2}).$ (17)
+$=-\,O_{p}(\tau_{n}^{2}).$ (A.8)
 […]
-$\ell_{n}(\widehat{\eta})=\max_{\eta}\ell_{n}(\eta)\;\geq\;\ell_{n}(\tilde{\eta}_{1})\;\geq\;\ell_{n}(\eta_{o})-O_{p}(\tau_{n}^{2})\,.$ (18)
+$\ell_{n}(\widehat{\eta})=\max_{\eta}\ell_{n}(\eta)\;\geq\;\ell_{n}(\tilde{\eta}_{1})\;\geq\;\ell_{n}(\eta_{o})-O_{p}(\tau_{n}^{2})\,.$ (A.9)
 On the other hand, by Lemma 3, the population loss satisfies
 $\ell_{0}(\eta)-\ell_{0}(\eta_{o})\;\leq\;-\,c_{-}\,d(\eta,\eta_{o})^{2}\quad\text{for all }\eta\in\mathcal{F}_{D}\,.$
-The empirical-process bound (14) and the likelihood chain (18) verify the conditions (i)–(ii) of Theorem 3.4.1 in vdvwellner1996 with radius of order $\tau_{n}$. Therefore,
+The empirical-process bound (A.5) and the likelihood chain (A.9) verify the conditions (i)–(ii) of Theorem 3.4.1 in van der Vaart and Wellner (1996) with radius of order $\tau_{n}$. Therefore,
 $d(\widehat{\eta},\eta_{o})=O_{p}(\tau_{n})\,.$
@@
 […]
-$\mathbb{E}\Bigl[\Delta\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\}^{2}\Bigr]\;\leq\;C\Bigl(d(\widehat{\eta},\eta_{o})^{2}+\|\widehat{}\mbox{${\theta}$}-\mbox{${\theta}$}_{o}\|^{2}\Bigr)\,.$ (19)
+$\mathbb{E}\Bigl[\Delta\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\}^{2}\Bigr]\;\leq\;C\Bigl(d(\widehat{\eta},\eta_{o})^{2}+\|\widehat{}\mbox{${\theta}$}-\mbox{${\theta}$}_{o}\|^{2}\Bigr)\,.$ (A.10)
 Applying Lemma 13 with $\eta_{1}=\widehat{\eta}$ and $\eta_{2}=\eta_{o}$ yields
 $\|\widehat{}\mbox{${\theta}$}-\mbox{${\theta}$}_{o}\|^{2}\leq c_{\theta}^{-1}d(\widehat{\eta},\eta_{o})^{2}=O_{p}(\tau_{n}^{2})\,.$
-Hence, $\|\widehat{}\mbox{${\theta}$}-\mbox{${\theta}$}_{o}\|=O_{p}(\tau_{n})$, and the right-hand side of (19) is $O_{p}(\tau_{n}^{2})$, which is exactly the claim of Theorem 1.
-The expectation is taken with respect to a new observation $(T,\Delta,{\bf X},{\bf Z})\sim P_{0}$, while the $O_{p}(\cdot)$ is with respect to the randomness in the training sample used to construct $\widehat{g}$.
+Hence, $\|\widehat{}\mbox{${\theta}$}-\mbox{${\theta}$}_{o}\|=O_{p}(\tau_{n})$, and the right-hand side of (A.10) is $O_{p}(\tau_{n}^{2})$, which is exactly the claim of Theorem 1.
+The expectation is taken with respect to a new observation $(T,\Delta,{\bf X},{\bf Z})\sim\mathbb{P}_{0}$, while the $O_{p}(\cdot)$ is with respect to the randomness in the training sample used to construct $\widehat{g}$.
 ∎
@@
 Proof of Theorem 2.
-For $g\in\mathcal{H}(q,\boldsymbol{\alpha},{\bf d},\tilde{{\bf d}},M)$ and and $\mbox{${\theta}$}_{o}\in\Theta$, let $P_{g,\theta_{o}}$ denote the joint distribution of $(T,\Delta,{\bf X},{\bf Z})$ induced by the hazard $h_{g,\theta_{o}}(t\mid{\bf X},{\bf Z})=\exp\{g(t,{\bf X})+\mbox{${\theta}$}_{o}^{\top}{\bf Z}\}$ together with a censoring hazard that is held fixed across all $g$.
-By the standard Kullback-Leibler (KL) distance for bounded hazards (Assumptions A1A2), there exists $C>0$ such that for any $g_{0},g_{1}\in\mathcal{\mathcal{}}{H}(q,\boldsymbol{\alpha},{\bf d},\tilde{{\bf d}},M)$,
-$\mbox{KL}(P_{g_{1},\theta_{o}}\,\|\,P_{g_{0},\theta_{o}})\;\leq\;C\,n\,\mathbb{E}_{P_{g_{1},\theta_{o}}}\!\left[\Delta\{g_{1}(T,{\bf X})-g_{0}(T,{\bf X})\}^{2}\right]\,.$
+For $g\in\mathcal{H}(q,\boldsymbol{\alpha},{\bf d}^{*},\tilde{{\bf d}},M)$ and $\mbox{${\theta}$}_{o}\in\Theta$, let $\mathbb{P}_{g,\theta_{o}}$ denote the joint distribution of $(T,\Delta,{\bf X},{\bf Z})$ induced by the hazard $h_{g,\theta_{o}}(t\mid{\bf X},{\bf Z})=\exp\{g(t,{\bf X})+\mbox{${\theta}$}_{o}^{\top}{\bf Z}\}$ together with a censoring hazard that is held fixed across all $g$.
+By the standard Kullback-Leibler (KL) distance for bounded hazards (Assumptions A1A2), there exists $C>0$ such that for any $g_{0},g_{1}\in\mathcal{\mathcal{}}{H}(q,\boldsymbol{\alpha},{\bf d}^{*},\tilde{{\bf d}},M)$,
+$\mbox{KL}(\mathbb{P}_{g_{1},\theta_{o}}\,\|\,\mathbb{P}_{g_{0},\theta_{o}})\;\leq\;C\,n\,\mathbb{E}_{\mathbb{P}_{g_{1},\theta_{o}}}\!\left[\Delta\{g_{1}(T,{\bf X})-g_{0}(T,{\bf X})\}^{2}\right]\,.$
 Since $\Delta\leq 1$,
-$\mbox{KL}(P_{g_{1},\theta_{o}}\,\|\,P_{g_{0},\theta_{o}})\;\leq\;C\,n\,\mathbb{E}_{P_{g_{1},\theta_{o}}}\!\bigl[\{g_{1}(T,{\bf X})-g_{0}(T,{\bf X})\}^{2}\bigr].$
+$\mbox{KL}(\mathbb{P}_{g_{1},\theta_{o}}\,\|\,\mathbb{P}_{g_{0},\theta_{o}})\;\leq\;C\,n\,\mathbb{E}_{\mathbb{P}_{g_{1},\theta_{o}}}\!\bigl[\{g_{1}(T,{\bf X})-g_{0}(T,{\bf X})\}^{2}\bigr].$
 Lemma 11 yields the equivalence
-$\mbox{KL}(P_{g_{1},\theta_{o}}\,\|\,P_{g_{0},\theta_{o}})\;\leq\;C^{\prime}\,n\,\mathbb{E}_{P_{g_{0},\theta_{o}}}\!\bigl[\{g_{1}(T,{\bf X})-g_{0}(T,{\bf X})\}^{2}\bigr]$
+$\mbox{KL}(\mathbb{P}_{g_{1},\theta_{o}}\,\|\,\mathbb{P}_{g_{0},\theta_{o}})\;\leq\;C^{\prime}\,n\,\mathbb{E}_{\mathbb{P}_{g_{0},\theta_{o}}}\!\bigl[\{g_{1}(T,{\bf X})-g_{0}(T,{\bf X})\}^{2}\bigr]$
 for a constant $C^{\prime}$ depending only on A1A2.
-By the packing construction for composite Hölder networks (schmidthieber2020nonparametric), there exist functions $g^{(0)},\dots,g^{(N)}\in\mathcal{H}(q,\boldsymbol{\alpha},{\bf d},\tilde{{\bf d}},M)$ and constants $c_{1},c_{2}>0$ such that
-$\|g^{(j)}-g^{(k)}\|_{L^{2}(P_{g^{(0)},\theta_{o}})}\;\geq\;2c_{1}\,\gamma_{n}\,,\qquad j\neq k\,$, (20)
+By the packing construction for composite Hölder networks (Schmidt-Hieber, 2020), there exist functions $g^{(0)},\dots,g^{(N)}\in\mathcal{H}(q,\boldsymbol{\alpha},{\bf d}^{*},\tilde{{\bf d}},M)$ and constants $c_{1},c_{2}>0$ such that
+$\|g^{(j)}-g^{(k)}\|_{L^{2}(P_{g^{(0)},\theta_{o}})}\;\geq\;2c_{1}\,\gamma_{n}\,,\qquad j\neq k\,$, (A.11)
 and
-$\frac{1}{N}\sum_{j=1}^{N}\|g^{(j)}-g^{(0)}\|_{L^{2}(P_{g^{(0)},\theta_{o}})}^{2}\;\leq\;\frac{c_{2}\,\log N}{n}\,.$ (21)
-Applying the KL bound with $g_{0}=g^{(0)}$ and $g_{1}=g^{(j)}$ and the using (21),
-$\frac{1}{N}\sum_{j=1}^{N}\mbox{KL}\!\big(P_{g^{(j)},\theta_{o}}\,\|\,P_{g^{(0)},\theta_{o}}\big)\;\leq\;C^{\prime}\,c_{2}\,\log N\,.$
+$\frac{1}{N}\sum_{j=1}^{N}\|g^{(j)}-g^{(0)}\|_{L^{2}(P_{g^{(0)},\theta_{o}})}^{2}\;\leq\;\frac{c_{2}\,\log N}{n}\,.$ (A.12)
+Applying the KL bound with $g_{0}=g^{(0)}$ and $g_{1}=g^{(j)}$ and using (A.12),
+$\frac{1}{N}\sum_{j=1}^{N}\mbox{KL}\!\big(\mathbb{P}_{g^{(j)},\theta_{o}}\,\|\,\mathbb{P}_{g^{(0)},\theta_{o}}\big)\;\leq\;C^{\prime}\,c_{2}\,\log N\,.$
 Choose $\alpha\in(0,1/8)$ so that $C^{\prime}c_{2}\leq\alpha$. Then,
-$\frac{1}{N}\sum_{j=1}^{N}\mbox{KL}(P_{g^{(j)},\theta_{o}}\,\|\,P_{g^{(0)},\theta_{o}})\;\leq\;\alpha\log N\,.$
-Applying Tsybakov’s Fano inequality to $\{g^{(0)},\dots,g^{(N)}\}$ with separation (20), we obtain that there exists $c>0$ (depending only on $c_{1}$ and $\alpha$) such that for any estimator $\widehat{g}$,
-$\sup_{j=0,\dots,N}\ \mathbb{E}_{P_{g^{(j)},\theta_{o}}}\!\Big(\|\widehat{g}-g^{(j)}\|_{L^{2}(P_{g^{(0)},\theta_{o}})}^{2}\Big)\;\geq\;c\,\gamma_{n}^{2}\,.$
+$\frac{1}{N}\sum_{j=1}^{N}\mbox{KL}(\mathbb{P}_{g^{(j)},\theta_{o}}\,\|\,\mathbb{P}_{g^{(0)},\theta_{o}})\;\leq\;\alpha\log N\,.$
+Applying Tsybakov’s Fano inequality to $\{g^{(0)},\dots,g^{(N)}\}$ with separation (A.11), we obtain that there exists $c>0$ (depending only on $c_{1}$ and $\alpha$) such that for any estimator $\widehat{g}$,
+$\sup_{j=0,\dots,N}\ \mathbb{E}_{\mathbb{P}_{g^{(j)},\theta_{o}}}\!\Big(\|\widehat{g}-g^{(j)}\|_{L^{2}(\mathbb{P}_{g^{(0)},\theta_{o}})}^{2}\Big)\;\geq\;c\,\gamma_{n}^{2}\,.$
 Since the supremum over $g_{o}\in\mathcal{H}$ dominates this finite set,
-$\inf_{\widehat{g}}\ \sup_{g_{o}\in\mathcal{H}}\mathbb{E}_{P_{g_{o},\theta_{o}}}\!\Bigl(\|\widehat{g}-g_{o}\|_{L^{2}(P_{g^{(0)},\theta_{o}})}^{2}\Bigr)\;\geq\;c\,\gamma_{n}^{2}\,.$
+$\inf_{\widehat{g}}\ \sup_{g_{o}\in\mathcal{H}}\mathbb{E}_{\mathbb{P}_{g_{o},\theta_{o}}}\!\Bigl(\|\widehat{g}-g_{o}\|_{L^{2}(\mathbb{P}_{g^{(0)},\theta_{o}})}^{2}\Bigr)\;\geq\;c\,\gamma_{n}^{2}\,.$
 By Lemma 11, there exists a constant $C_{\mathrm{eq}}>0$ (depending only on A1A2) such that, uniformly over $g_{o}\in\mathcal{H}$ and all sieve estimators $\widehat{g}$,
-$\mathbb{E}_{P_{g^{(0)},\theta_{o}}}\!\big[\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\}^{2}\big]\;\leq\;C_{\mathrm{eq}}\,\mathbb{E}_{P_{g_{o},\theta_{o}}}\!\big[\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\}^{2}\big].$
+$\mathbb{E}_{\mathbb{P}_{g^{(0)},\theta_{o}}}\!\big[\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\}^{2}\big]\;\leq\;C_{\mathrm{eq}}\,\mathbb{E}_{\mathbb{P}_{g_{o},\theta_{o}}}\!\big[\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\}^{2}\big].$
 Combining this with the previous display yields
-$\inf_{\widehat{g}}\ \sup_{g_{o}\in\mathcal{H}}\mathbb{E}_{P_{g_{o},\theta_{o}}}\!\big[\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\}^{2}\big]\;\geq\;\frac{c}{C_{\mathrm{eq}}}\,\gamma_{n}^{2}.$
+$\inf_{\widehat{g}}\ \sup_{g_{o}\in\mathcal{H}}\mathbb{E}_{\mathbb{P}_{g_{o},\theta_{o}}}\!\big[\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\}^{2}\big]\;\geq\;\frac{c}{C_{\mathrm{eq}}}\,\gamma_{n}^{2}.$
 Finally, since the supremum in the theorem is also taken over $\mbox{${\theta}$}_{o}\in\Theta$, and the above bound holds for each fixed $\theta_{o}$, we obtain
-$\inf_{\widehat{g}}\ \sup_{\theta_{o}\in\Theta,\ g_{o}\in\mathcal{H}}\mathbb{E}_{P_{g_{o},\theta_{o}}}\!\big[\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\}^{2}\big]\;\geq\;c_{1}\,\gamma_{n}^{2}$ (22)
+$\inf_{\widehat{g}}\ \sup_{\theta_{o}\in\Theta,\ g_{o}\in\mathcal{H}}\mathbb{E}_{\mathbb{P}_{g_{o},\theta_{o}}}\!\big[\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\}^{2}\big]\;\geq\;c_{1}\,\gamma_{n}^{2}$ (A.13)
 for some $c_{1}>0$. By Lemma 2 and Assumptions A1, A2, A4, there exists a constant $C_{\Delta}>0$ such that, for all $g_{o}\in\mathcal{H}$ and all $\widehat{g}$,
-$\mathbb{E}_{P_{g_{o},\theta_{o}}}\!\big[\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\}^{2}\big]\;\leq\;C_{\Delta}\,\mathbb{E}_{P_{g_{o},\theta_{o}}}\!\big[\Delta\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\}^{2}\big]\,.$
-Combining this with (22) gives
-$\inf_{\widehat{g}}\ \sup_{\theta_{o}\in\Theta,\ g_{o}\in\mathcal{H}}\mathbb{E}_{P_{g_{o},\theta_{o}}}\!\big[\Delta\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\}^{2}\big]\;\geq\;\frac{c_{1}}{C_{\Delta}}\,\gamma_{n}^{2}\,.$
+$\mathbb{E}_{\mathbb{P}_{g_{o},\theta_{o}}}\!\big[\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\}^{2}\big]\;\leq\;C_{\Delta}\,\mathbb{E}_{\mathbb{P}_{g_{o},\theta_{o}}}\!\big[\Delta\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\}^{2}\big]\,.$
+Combining this with (A.13) gives
+$\inf_{\widehat{g}}\ \sup_{\theta_{o}\in\Theta,\ g_{o}\in\mathcal{H}}\mathbb{E}_{\mathbb{P}_{g_{o},\theta_{o}}}\!\big[\Delta\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\}^{2}\big]\;\geq\;\frac{c_{1}}{C_{\Delta}}\,\gamma_{n}^{2}\,.$
 Renaming the constant $c=c_{1}/C_{\Delta}$ completes the proof. ∎
@@
 Proof of Theorem 3.
-Let $\mbox{${\theta}$}\in\mathbb{R}^{p}$ and let $\{g_{s}:s\in(-1,1)\}\subset\mathcal{H}_{g_{o}}$ be a one–dimensional regular submodel of $g$, with
-$g_{s}=g_{o}+sg,\qquad g=\left.\frac{\partial g_{s}}{\partial s}\right|_{s=0}\in\mathcal{T}_{g_{o}}\,.$
+Let ${\bf u}\in\mathbb{R}^{p}$ and let $\{g_{s}:s\in(-1,1)\}\in\mathcal{H}_{g_{o}}$ be a one-dimensional regular submodel through $g_{o}$, with
+$\dot{g}_{0}=\left.\frac{\partial g_{s}}{\partial s}\right|_{s=0}\in\mathcal{T}_{g_{o}}.$
 Consider the semiparametric submodel
-$\left\{P_{\theta_{o}+s\theta,\;g_{s}}:s\in(-1,1)\right\}\,.$
-For a single observation $(T,\Delta,{\bf X},{\bf Z})$ the log-likelihood is
+$\{\mathbb{P}_{\theta_{o}+s{\bf u},\;g_{s}}:s\in(-1,1)\}.$
+For one observation $(T,\Delta,{\bf X},{\bf Z})$,
 $\ell(\mbox{${\theta}$},g)=\Delta\{g(T,{\bf X})+\mbox{${\theta}$}^{\top}{\bf Z}\}-\int_{0}^{\tau}Y(t)\exp\{g(t,{\bf X})+\mbox{${\theta}$}^{\top}{\bf Z}\}\,dt.$
-Differentiating $\ell(\mbox{${\theta}$}_{o}+s\mbox{${\theta}$},g_{s})$ at $s=0$ gives
-$\left.\frac{d}{ds}\ell(\mbox{${\theta}$}_{o}+s\mbox{${\theta}$},g_{s})\right|_{s=0}=\mbox{${\theta}$}^{\top}\dot{\ell}_{\theta_{o}}+\dot{\ell}_{g_{o},g},$
+Differentiating at $s=0$ gives
+$\left.\frac{d}{ds}\ell(\mbox{${\theta}$}_{o}+s{\bf u}\,,\,g_{s})\right|_{s=0}={\bf u}^{\top}\dot{\ell}_{\theta_{o}}+\dot{\ell}_{g_{o},\dot{g}_{0}},$
 where
-$\dot{\ell}_{\theta_{o}}=\Delta{\bf Z}-\int_{0}^{\tau}{\bf Z}Y(t)\exp\{g_{o}(t,{\bf X})+\mbox{${\theta}$}_{o}^{\top}{\bf Z}\}\,dt=\int_{0}^{\tau}{\bf Z}\,dM(t),$
-and
-$\dot{\ell}_{g_{o},g}=\Delta g(T,{\bf X})-\int_{0}^{\tau}g(t,{\bf X})\,Y(t)\exp\{g_{o}(t,{\bf X})+\mbox{${\theta}$}_{o}^{\top}{\bf Z}\}\,dt=\int_{0}^{\tau}g(t,{\bf X})\,dM(t)\,.$
-Thus the (naive) score for $\mbox{${\theta}$}_{o}$ is $\dot{\ell}_{\theta_{o}}$. To obtain the efficient score, we project this onto the orthocomplement of the nuisance tangent space
-$\mathcal{T}_{g_{o}}^{\mathrm{nuis}}=\left\{\dot{\ell}_{g_{o},g}:g\in\mathcal{T}_{g_{o}}\right\},$
-but the projection must be taken onto the closure $\overline{\mathcal{T}_{g_{o}}^{\mathrm{nuis}}}$ in $L^{2}(P_{0})$. Hence the efficient score is
+$\dot{\ell}_{\theta_{o}}=\int_{0}^{\tau}{\bf Z}\,dM(t),\qquad\dot{\ell}_{g_{o},g}=\int_{0}^{\tau}g(t,{\bf X})\,dM(t).$
+Hence the nuisance score space is
+$\mathcal{T}^{\mathrm{nuis}}_{g_{o}}=\{\dot{\ell}_{g_{o},g}:g\in\mathcal{T}_{g_{o}}\},$
+and the efficient score is
 $\ell^{*}_{\theta_{o}}=\dot{\ell}_{\theta_{o}}-\Pi_{g_{o}}(\dot{\ell}_{\theta_{o}}),$
-where $\Pi_{g_{o}}$ denotes the $L^{2}(P_{0})$–projection onto $\overline{\mathcal{T}_{g_{o}}^{\mathrm{nuis}}}$.
-The projection $\Pi_{g_{o}}(\dot{\ell}_{\theta_{o}})$ has the representation $\dot{\ell}_{g_{o},{\bf g}^{*}}$ for a pp–vector ${\bf g}^{*}=(g_{1}^{*},\dots,g_{p}^{*})^{\top}\in(\overline{\mathcal{T}_{g_{o}}})^{p}$ satisfying the orthogonality condition
-$\mathbb{E}\!\Big\{(\dot{\ell}_{\theta_{o}}-\dot{\ell}_{g_{o},{\bf g}^{*}})\,\dot{\ell}_{g_{o},g}\Big\}={\bf 0}\qquad\text{for all }g\in\mathcal{T}_{g_{o}}\,.$
-Using Lemma 1 of sasieni1992information, this orthogonality is equivalent to
-$\mathbb{E}\!\left[\Delta\{{\bf Z}-{\bf g}^{*}(T,{\bf X})\}\,g(T,{\bf X})\right]={\bf 0}\qquad\text{for all }g\in\mathcal{T}_{g_{o}}\,.$
-Thus, componentwise, $g_{j}^{*}$ is the (unique) $\Delta$–weighted $L^{2}(P_{0})$ projection of $Z_{j}$ onto $\overline{\mathcal{T}_{g_{o}}}$. Equivalently, $\bf g$$^{*}$ minimizes
-$\mathbb{E}\!\left(\Delta\|{\bf Z}-{\bf g}(T,{\bf X})\|_{c}^{2}\right)\,,\qquad{\bf g}\in(\overline{\mathcal{T}_{g_{o}}})^{p}\,.$
-Substituting $\bf g$$^{*}$, the efficient score becomes
+where $\Pi_{g_{o}}$ denotes projection onto $\overline{\mathcal{T}^{\mathrm{nuis}}_{g_{o}}}$ in $L_{2}(\mathbb{P}_{0})$.
+By Lemma 1 of Sasieni (1992), for any square-integrable predictable functions $\mu_{1},\mu_{2}$,
+$E_{0}\!\left[\left(\int_{0}^{\tau}\mu_{1}\,dM\right)\left(\int_{0}^{\tau}\mu_{2}\,dM\right)\right]=E_{0}\!\left[\Delta\,\mu_{1}(T,{\bf X})\mu_{2}(T,{\bf X})\right].$
+Therefore the map
+$J:h\mapsto\int_{0}^{\tau}\mu(t,{\bf X})\,dM(t)$
+is an isometry from $L_{2}(\mathbb{P}_{0},\Delta)$ onto its image.
+Since $\overline{\mathcal{T}}_{g_{o}}$ is a closed linear subspace of $L_{2}(\mathbb{P}_{0},\Delta)$ and $Z_{j}\in L_{2}(\mathbb{P}_{0},\Delta)$ by Assumption A2, the Hilbert projection theorem implies that for each $j=1,\dots,p$ there exists a unique $g_{j}^{*}\in\overline{\mathcal{T}}_{g_{o}}$ such that
+$\Pi_{g_{o}}(\dot{\ell}_{\theta_{o},j})=\dot{\ell}_{g_{o},g_{j}^{*}}.$
+Let ${\bf g}^{*}=(g_{1}^{*},\dots,g_{p}^{*})^{\top}$. Then
 $\ell^{*}_{\theta_{o}}(T,{\bf X},{\bf Z},\Delta)=\int_{0}^{\tau}\{{\bf Z}-{\bf g}^{*}(t,{\bf X})\}\,dM(t).$
-By the martingale isometry (again Lemma 1 of sasieni1992information), the efficient information matrix is
-$I(\theta_{o})=\mathbb{E}\left[\Delta\,\{{\bf Z}-{\bf g}^{*}(T,{\bf X})\}^{\otimes 2}\right].$
-Finally, under the regularity conditions in Assumptions (A1)–(A4), existence and uniqueness of the weighted projection $\bf g$$^{*}$ in $\overline{\mathcal{T}_{g_{o}}}^{p}$ follow from the approximation theorems of stone1985additive and the general semiparametric projection theory in Appendix A.4 of bickel1993efficient.
+The orthogonality condition for the projection is
+$E_{0}\!\left[\{\dot{\ell}_{\theta_{o},j}-\dot{\ell}_{g_{o},g_{j}^{*}}\}\dot{\ell}_{g_{o},g}\right]=0\qquad\forall g\in\mathcal{T}_{g_{o}},$
+which, by the isometry above, is equivalent to
+$E_{0}\!\left[\Delta\{Z_{j}-g_{j}^{*}(T,{\bf X})\}g(T,{\bf X})\right]=0\qquad\forall g\in\mathcal{T}_{g_{o}}.$
+By density, the same holds for all $g\in\overline{\mathcal{T}}_{g_{o}}$.
+Thus $g_{j}^{*}$ is the $\Delta$-weighted $L_{2}(\mathbb{P}_{0})$ projection of $Z_{j}$ onto $\overline{\mathcal{T}}_{g_{o}}$, and equivalently
+${\bf g}^{*}\in\arg\min_{{\bf g}\in(\overline{\mathcal{T}}_{g_{o}})^{p}}E_{0}\!\left[\Delta\|{\bf Z}-{\bf g}(T,{\bf X})\|_{2}^{2}\right].$
+Finally, by the martingale isometry,
+$I(\mbox{${\theta}$}_{o})=E_{0}\!\left[\ell_{\theta_{o}}^{*\otimes 2}\right]=E_{0}\!\left[\Delta\{{\bf Z}-{\bf g}^{*}(T,{\bf X})\}^{\otimes 2}\right].$
+Finally, under the regularity conditions in Assumptions (A1)–(A4), existence and uniqueness of the weighted projection $\bf g$$^{*}$ in $\overline{\mathcal{T}}_{g_{o}}^{p}$ follow from the approximation theorems of Stone (1985) and the general semiparametric projection theory in Appendix A.4 of Bickel et al.
+(1993).
 Therefore, both the efficient score and the efficient information matrix are well-defined and given by the formulas above.
@@
 $r_{\theta,g}(T,{\bf X},{\bf Z},\Delta)=\Delta\{{\bf Z}-{\bf g}^{*}(T,{\bf X})\}-\int_{0}^{\tau}\{{\bf Z}-{\bf g}^{*}(t,{\bf X})\}Y(t)\exp\{g(t,{\bf X})+\mbox{${\theta}$}^{\top}{\bf Z}\}\,dt\,$,
-where $\bf g$$^{*}$ is the pp–dimensional minimizer from Theorem 3, which depends only on $(\mbox{${\theta}$}_{o},g_{o})$.
+where ${\bf g}^{*}=(g_{1}^{*},\dots,g_{p}^{*})^{\top}$ is the pp–dimensional minimizer from Theorem 3, which depends only on $(\mbox{${\theta}$}_{o},g_{o})$.
 By Theorem 3, at the true parameter $(\mbox{${\theta}$}_{o},g_{o})$ this coincides with the efficient score,
@@
 $I(\mbox{${\theta}$}_{o})=\mathbb{P}\bigl\{r_{\theta_{o},g_{o}}^{\otimes 2}\bigr\}=\mathbb{E}\!\bigl[\Delta\{{\bf Z}-{\bf g}^{*}(T,{\bf X})\}^{\otimes 2}\bigr]\,.$
-The empirical efficient score at $(\mbox{${\theta}$},g)$ is $\mathbb{P}_{n}r_{\theta,g}$. By the first–order condition for the (sieve) maximum–likelihood estimator, the estimated parameters satisfy
-$\mathbb{P}_{n}r_{\widehat{\theta},\widehat{g}}={\bf 0}\,.$
-Write
+The empirical efficient score at $(\mbox{${\theta}$},g)$ is $\mathbb{P}_{n}r_{\theta,g}$.
+In the following it will be shown that $\mathbb{P}_{n}r_{\widehat{\theta},\widehat{g}}=\mathbf{0}$. For the $j$-th coordinate of $r_{\theta,g}$, we have
+$r_{\theta,g}(T,{\bf X},{\bf Z},\Delta;j)$ $=\Delta\{Z_{j}-g_{j}^{*}(T,{\bf X})\}-\int_{0}^{\tau}\{Z_{j}-g_{j}^{*}(t,{\bf X})\}Y(t)\exp\{g(t,{\bf X})+\mbox{${\theta}$}^{\top}{\bf Z}\}\,dt$
+$=\left[\Delta Z_{j}-\int_{0}^{\tau}Z_{j}\,Y(t)\exp\{g(t,{\bf X})+\mbox{${\theta}$}^{\top}{\bf Z}\}\,dt\right]$
+$\qquad-\left[\Delta g_{j}^{*}(T,{\bf X})-\int_{0}^{\tau}g_{j}^{*}(t,{\bf X})\,Y(t)\exp\{g(t,{\bf X})+\mbox{${\theta}$}^{\top}{\bf Z}\}\,dt\right].$
+Therefore,
+$\mathbb{P}_{n}\!\bigl[r_{\theta,g}(\cdot\,;j)\bigr]$ $=\mathbb{P}_{n}\!\left[\Delta Z_{j}-\int_{0}^{\tau}Z_{j}\,Y(t)\exp\{g(t,{\bf X})+\mbox{${\theta}$}^{\top}{\bf Z}\}\,dt\right]$
+$\qquad-\mathbb{P}_{n}\!\left[\Delta g_{j}^{*}(T,{\bf X})-\int_{0}^{\tau}g_{j}^{*}(t,{\bf X})\,Y(t)\exp\{g(t,{\bf X})+\mbox{${\theta}$}^{\top}{\bf Z}\}\,dt\right]$
+$=\dot{\ell}_{n,\theta,j}(\mbox{${\theta}$},g)-\dot{\ell}_{n,g}(\mbox{${\theta}$},g)[g_{j}^{*}],$
+where
+$\dot{\ell}_{n,\theta}(\mbox{${\theta}$},g)=\mathbb{P}_{n}\!\left[\Delta{\bf Z}-\int_{0}^{\tau}{\bf Z}\,Y(t)\exp\{g(t,{\bf X})+\mbox{${\theta}$}^{\top}{\bf Z}\}\,dt\right],$
+and, for any admissible nuisance direction $\mu\in\overline{\mathcal{T}}_{g_{o}}$, the directional Gateaux derivative score
+$\dot{\ell}_{n,g}(\mbox{${\theta}$},g)[\mu]=\mathbb{P}_{n}\!\left[\Delta\mu(T,{\bf X})-\int_{0}^{\tau}\mu(t,{\bf X})\,Y(t)\exp\{g(t,{\bf X})+\mbox{${\theta}$}^{\top}{\bf Z}\}\,dt\right].$
+[…]
+$\mathbb{P}_{n}\,r_{\widehat{\theta},\widehat{g}}(\cdot\,;j)=\dot{\ell}_{n,\theta,j}(\widehat{}\mbox{${\theta}$},\widehat{g})-\dot{\ell}_{n,g}(\widehat{}\mbox{${\theta}$},\widehat{g})[g_{j}^{*}]=0,$
+for every $j=1,\dots,p$. Hence $\mathbb{P}_{n}r_{\hat{\theta},\hat{g}}=\mathbf{0}$.
+Now, write
 $\mathbb{P}_{n}\bigl\{r_{\widehat{\theta},\widehat{g}}-r_{\theta_{o},g_{o}}\bigr\}=\mathbb{P}\bigl\{r_{\widehat{\theta},\widehat{g}}-r_{\theta_{o},g_{o}}\bigr\}+(\mathbb{P}_{n}-\mathbb{P})\bigl\{r_{\widehat{\theta},\widehat{g}}-r_{\theta_{o},g_{o}}\bigr\}\,.$
@@
 Using $\mathbb{P}_{n}r_{\widehat{\theta},\widehat{g}}={\bf 0}$, this identity can be written as
-$\mathbb{P}\bigl\{r_{\widehat{\theta},\widehat{g}}-r_{\theta_{o},g_{o}}\bigr\}=-\,P_{n}r_{\theta_{o},g_{o}}+o_{p}(n^{-1/2}).$ (23)
+$\mathbb{P}\bigl\{r_{\widehat{\theta},\widehat{g}}-r_{\theta_{o},g_{o}}\bigr\}=-\,\mathbb{P}_{n}r_{\theta_{o},g_{o}}+o_{p}(n^{-1/2}).$ (A.14)
 Consider the map
@@
 Hence the contribution of $\widehat{}\mbox{${\theta}$}-\mbox{${\theta}$}_{o}$ in the expansion is $I(\mbox{${\theta}$}_{o})(\widehat{}\mbox{${\theta}$}-\mbox{${\theta}$}_{o})$.
-Fix a deterministic direction $g$ and consider $g_{s}=g_{o}+sg$. Differentiating under the integral sign and applying the martingale compensation identity yields
-$\left.\frac{d}{ds}\right|_{s=0}m(\mbox{${\theta}$}_{o},g_{s})=-\,\mathbb{E}\bigl[\Delta\{{\bf Z}-{\bf g}^{*}(T,{\bf X})\}g(T,{\bf X})\bigr]\,.$
-By Theorem 3, $\bf g$$^{*}$ is the $\Delta$–weighted $L^{2}(P)$–projection of $\bf Z$} onto the nuisance tangent space $(\overline{\mathcal{T}}_{g_{o}})^{p}$, so
-$\mathbb{E}\bigl[\Delta\{{\bf Z}-{\bf g}^{*}(T,{\bf X})\}g(T,{\bf X})\bigr]={\bf 0}\quad\text{for all }h\in\overline{\mathcal{T}}_{g_{o}}.$
-By Assumptions A1A4, $\widehat{g}-g_{o}\in\overline{\mathcal{T}}_{g_{o}}$, and therefore
-$\dot{m}_{g}(\mbox{${\theta}$}_{o},g_{o})[\widehat{g}-g_{o}]=0.$
+Fix a deterministic scalar direction $\mu\in\overline{T}_{g_{o}}$ and consider $g_{s}=g_{o}+s\mu$. Differentiating under the integral sign and applying the martingale compensation identity yields
+$\left.\frac{d}{ds}\right|_{s=0}m(\mbox{${\theta}$}_{o},g_{s})=-\,\mathbb{E}\bigl[\Delta\{{\bf Z}-{\bf g}^{*}(T,{\bf X})\}\mu(T,{\bf X})\bigr]\,.$
+By Theorem 3, $\bf g$$^{*}$ is the coordinatewise $\Delta$–weighted $L^{2}(\mathbb{P})$–projection of $\bf Z$} onto the nuisance tangent space $(\overline{\mathcal{T}}_{g_{o}})^{p}$. Therefore,
+$\mathbb{E}\bigl[\Delta\{{\bf Z}-{\bf g}^{*}(T,{\bf X})\}\mu(T,{\bf X})\bigr]={\bf 0}\quad\text{for all }\mu\in\overline{\mathcal{T}}_{g_{o}}\,$,
+and hence $\dot{m}_{g}(\mbox{${\theta}$}_{o},g_{o})[\mu]={\bf 0}$ for every $\mu\in\overline{\mathcal{T}}_{g_{o}}$.
+Arguing as in Zhong et al.
+(2022) (Lemma 5), the sieve score equation together with the tangent-space approximation implies that
+$\dot{m}_{g}(\mbox{${\theta}$}_{o},g_{o})[\widehat{g}-g_{o}]=o_{p}(n^{-1/2})\,.$
 The map $g\mapsto m(\mbox{${\theta}$}_{o},g)$ is twice Fréchet differentiable, and its second derivative is bounded by a constant multiple of $\mathbb{E}[\Delta\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\}^{2}]$.
@@
 […]
-$\mathbb{P}\bigl\{r_{\widehat{\theta},\widehat{g}}-r_{\theta_{o},g_{o}}\bigr\}=-\,I(_{o})(\widehat{}\mbox{${\theta}$}-\mbox{${\theta}$}_{o})+o_{p}(n^{-1/2})\,.$
-Substituting this into (23) yields
+$\mathbb{P}\bigl\{r_{\widehat{\theta},\widehat{g}}-r_{\theta_{o},g_{o}}\bigr\}=-\,I(\theta_{o})(\widehat{}\mbox{${\theta}$}-\mbox{${\theta}$}_{o})+o_{p}(n^{-1/2})\,.$
+Substituting this into (A.14) yields
 $I(\mbox{${\theta}$}_{o})(\widehat{}\mbox{${\theta}$}-\mbox{${\theta}$}_{o})=-\,\mathbb{P}_{n}r_{\theta_{o},g_{o}}+o_{p}(n^{-1/2})\,$,
@@
 Lemma 14.
-Let $\delta_{n}=\max_{j}(t_{j}-t_{j-1})$ be the mesh size of the time grid.
-Under Assumptions A1A4, there exist constants $C_{1},C_{2}>0$ such that, for all $n$ and all $\eta$ in the neighborhood $\mathcal{N}_{n}=\{\eta:d(\eta,\eta_{o})\leq M\tau_{n}\}$,
-$\sup_{\eta\in\mathcal{N}_{n}}\bigl|\widetilde{\ell}_{n}(\eta)-\ell_{n}(\eta)\bigr|\leq C_{1}\,\delta_{n}$ (24)
+Let $0=t_{0}<t_{1}<\cdots<t_{m_{n}}=\tau$ be a grid with mesh size $\delta_{n}=\max_{1\leq j\leq m_{n}}(t_{j}-t_{j-1})$, and suppose that the grid contains all observed times $T_{1},\dots,T_{n}$.
+For $\eta=(\mbox{${\theta}$},g)$, let $\mathcal{N}_{n}=\{\eta:d(\eta,\eta_{o})\leq M\tau_{n}\}$ for some fixed $M>0$.
+Assume A2A3.
+Also define the uniform time-modulus
+$\omega_{n}(\delta)=\sup_{\eta\in\mathcal{N}_{n}}\sup_{{\bf X}\in\mathcal{X}}\sup_{\begin{subarray}{c}s,t\in[0,\tau]\\ |t-s|\leq\delta\end{subarray}}|g(t,{\bf X})-g(s,{\bf X})|.$
+Then there exist constants $C_{1},C_{2}>0$, depending only on the bounds in A2A3, such that
+$\sup_{\eta\in\mathcal{N}_{n}}\bigl|\tilde{\ell}_{n}(\eta)-\ell_{n}(\eta)\bigr|\leq C_{1}\,\omega_{n}(\delta_{n}),$
 and
-$\sup_{\eta\in\mathcal{N}_{n}}\Bigl\|\nabla_{\theta}\widetilde{\ell}_{n}(\eta)-\nabla_{\theta}\ell_{n}(\eta)\Bigr\|\leq C_{2}\,\delta_{n}\,$, (25)
-where $\nabla_{\theta}\widetilde{\ell}_{n}(\eta)$ and $\nabla_{\theta}\ell_{n}(\eta)$ are the $\theta$}-score functions of $\widetilde{\ell}_{n}(\eta)$ and $\ell_{n}(\eta)$, respectively.
-The proof of Lemma 8 follows directly from classical approximation results for Riemann sums.
-Under Assumptions A1–A4, $g(t,{\bf x})$ is uniformly Hölder continuous in $t$, uniformly over $\bf x$} and over $\eta\in\mathcal{N}_{n}$.
-Because the hazard is Lipschitz in $g(t,{\bf x})$, the integrand $Y_{i}(t)\exp\{g(t,{\bf X}_{i})+\mbox{${\theta}$}^{\top}{\bf Z}_{i}\}$ is Lipschitz on each interval where $Y_{i}(t)=1$.
-Therefore the Riemann approximation error is $O(\delta_{n})$.
-Applying this bound uniformly over all observations and all $\eta\in\mathcal{N}_{n}$ yields inequalities (24) and (25).
-Because the argument involves no stochastic elements and is fully deterministic, we omit the proof.
+$\sup_{\eta\in\mathcal{N}_{n}}\bigl\|\nabla_{\theta}\tilde{\ell}_{n}(\eta)-\nabla_{\theta}\ell_{n}(\eta)\bigr\|\leq C_{2}\,\omega_{n}(\delta_{n}).$
+In particular, if $\omega_{n}(\delta_{n})=O(\delta_{n})$, then both differences are $O(\delta_{n})$ uniformly on $\mathcal{N}_{n}$.
+Proof of Lemma 14.
+Fix $\eta=(\mbox{${\theta}$},g)\in\mathcal{N}_{n}$. Recall
+$\ell_{n}(\eta)=\frac{1}{n}\sum_{i=1}^{n}\left[\Delta_{i}\chi_{\eta}(T_{i},{\bf X}_{i},{\bf Z}_{i})-\int_{0}^{\tau}Y_{i}(t)e^{\chi_{\eta}(t,{\bf X}_{i},{\bf Z}_{i})}\,dt\right],$
+and
+$\tilde{\ell}_{n}(\eta)=\frac{1}{n}\sum_{i=1}^{n}\sum_{j:Y_{i}(t_{j-1})=1}\left[\Delta_{ij}\chi_{\eta}(t_{j},{\bf X}_{i},{\bf Z}_{i})-\bigl\{\min(T_{i},t_{j})-t_{j-1}\bigr\}e^{\chi_{\eta}(t_{j},{\bf X}_{i},{\bf Z}_{i})}\right]\,$,
+where $\chi_{\eta}(t,{\bf X},{\bf Z})=g(t,{\bf X})+\mbox{${\theta}$}^{\top}{\bf Z}$.
+Because the grid contains all observed times, for each $i$ there exists an index $j(i)$ such that $T_{i}=t_{j(i)}$.
+Hence
+$\sum_{j:Y_{i}(t_{j-1})=1}\Delta_{ij}\chi_{\eta}(t_{j},{\bf X}_{i},{\bf Z}_{i})=\Delta_{i}\chi_{\eta}(T_{i},{\bf X}_{i},{\bf Z}_{i}),$
+so the event contribution is represented exactly. Therefore
+$\tilde{\ell}_{n}(\eta)-\ell_{n}(\eta)=-\frac{1}{n}\sum_{i=1}^{n}R_{i}(\eta),$
+where
+$R_{i}(\eta)=\sum_{j:Y_{i}(t_{j-1})=1}\bigl\{\min(T_{i},t_{j})-t_{j-1}\bigr\}e^{\chi_{\eta}(t_{j},{\bf X}_{i},{\bf Z}_{i})}-\int_{0}^{\tau}Y_{i}(t)e^{\chi_{\eta}(t,{\bf X}_{i},{\bf Z}_{i})}\,dt.$
+For each $i$, define $\psi_{i,\eta}(t)=Y_{i}(t)e^{\chi_{\eta}(t,{\bf X}_{i},{\bf Z}_{i})}$, $t\in[0,\tau]$.
+On any interval $(t_{j-1},t_{j}]$ such that $Y_{i}(t_{j-1})=1$, we have $Y_{i}(t)=1$ for $t\in[t_{j-1},\min(T_{i},t_{j})]$, and thus
+$\int_{t_{j-1}}^{\min(T_{i},t_{j})}\psi_{i,\eta}(t)\,dt=\int_{t_{j-1}}^{\min(T_{i},t_{j})}e^{\chi_{\eta}(t,{\bf X}_{i},{\bf Z}_{i})}\,dt.$
+Since $\Theta$ is compact, $\|{\bf Z}\|\leq B_{{\bf Z}}$ a.s.
+by A2, and $\|g\|_{\infty}\leq D_{n}=O(1)$ by A3, there exists a constant $C<\infty$ such that uniformly over $i$, $\eta\in\mathcal{N}_{n}$, and $t\in[0,\tau]$,
+$e^{\chi_{\eta}(t,{\bf X}_{i},{\bf Z}_{i})}\leq C.$
+Moreover, by the mean-value theorem,
+$\bigl|e^{\chi_{\eta}(t,{\bf X}_{i},{\bf Z}_{i})}-e^{\chi_{\eta}(s,{\bf X}_{i},{\bf Z}_{i})}\bigr|\leq C\,|g(t,{\bf X}_{i})-g(s,{\bf X}_{i})|\leq C\,\omega_{n}(\delta_{n})$
+whenever $|t-s|\leq\delta_{n}$. Therefore, for each interval,
+$\left|\int_{t_{j-1}}^{\min(T_{i},t_{j})}e^{\chi_{\eta}(t,{\bf X}_{i},{\bf Z}_{i})}\,dt-\bigl\{\min(T_{i},t_{j})-t_{j-1}\bigr\}e^{\chi_{\eta}(t_{j},{\bf X}_{i},{\bf Z}_{i})}\right|\leq(t_{j}-t_{j-1})\,C\,\omega_{n}(\delta_{n}).$
+Summing over $j$ and using $\sum_{j}(t_{j}-t_{j-1})=\tau$, we obtain
+$|R_{i}(\eta)|\leq C\tau\,\omega_{n}(\delta_{n})$
+uniformly in $i$ and $\eta\in\mathcal{N}_{n}$. Hence
+$\sup_{\eta\in\mathcal{N}_{n}}|\tilde{\ell}_{n}(\eta)-\ell_{n}(\eta)|\leq\frac{1}{n}\sum_{i=1}^{n}\sup_{\eta\in\mathcal{N}_{n}}|R_{i}(\eta)|\leq C_{1}\,\omega_{n}(\delta_{n})$
+for some constant $C_{1}>0$.
+Next consider the $\theta$}-scores. Since $g$ does not depend on $\theta$,
+$\nabla_{\theta}\ell_{n}(\eta)=\frac{1}{n}\sum_{i=1}^{n}\left[\Delta_{i}{\bf Z}_{i}-\int_{0}^{\tau}{\bf Z}_{i}Y_{i}(t)e^{\chi_{\eta}(t,{\bf X}_{i},{\bf Z}_{i})}\,dt\right],$
+and
+$\nabla_{\theta}\tilde{\ell}_{n}(\eta)=\frac{1}{n}\sum_{i=1}^{n}\left[\Delta_{i}{\bf Z}_{i}-\sum_{j:Y_{i}(t_{j-1})=1}\bigl\{\min(T_{i},t_{j})-t_{j-1}\bigr\}{\bf Z}_{i}e^{\chi_{\eta}(t_{j},{\bf X}_{i},{\bf Z}_{i})}\right].$
+Therefore,
+$\nabla_{\theta}\tilde{\ell}_{n}(\eta)-\nabla_{\theta}\ell_{n}(\eta)=-\frac{1}{n}\sum_{i=1}^{n}{\bf Z}_{i}R_{i}(\eta).$
+Using $\|{\bf Z}_{i}\|\leq B_{Z}$ a.s. from A2 together with the bound on $R_{i}(\eta)$,
+$\sup_{\eta\in\mathcal{N}_{n}}\bigl\|\nabla_{\theta}\tilde{\ell}_{n}(\eta)-\nabla_{\theta}\ell_{n}(\eta)\bigr\|\leq\frac{1}{n}\sum_{i=1}^{n}\|{\bf Z}_{i}\|\,\sup_{\eta\in\mathcal{N}_{n}}|R_{i}(\eta)|\leq B_{{\bf Z}}C_{1}\,\omega_{n}(\delta_{n}).$
+This proves the second bound with $C_{2}=B_{{\bf Z}}C_{1}$. ∎
 Lemma 15.
-Let $S_{n}(\eta)=\nabla_{\theta}\ell_{n}(\eta)$, and $\dot{S}_{n}(\eta)=\nabla_{\theta}S_{n}(\eta)$ is the Hessian of the empirical log–likelihood with respect to $\theta$}.
-For a fixed $M>0$, define the shrinking neighbourhood $\mathcal{N}_{n}=\bigl\{\eta:d(\eta,\eta_{o})\leq M\tau_{n}\bigr\}$, $\tau_{n}=\gamma_{n}\log^{2}n$.
+Let $S_{n}(\eta)=\nabla_{\theta}\ell_{n}(\eta)$ and $\dot{S}_{n}(\eta)=\nabla_{\theta}S_{n}(\eta)$ be the Hessian of the empirical log–likelihood with respect to $\theta$}.
+For a fixed $M>0$, define the shrinking neighbourhood $\mathcal{N}_{n}=\{\eta:d(\eta,\eta_{o})\leq M\tau_{n}\}$, where $\tau_{n}=\gamma_{n}\log^{2}n$.
 Assume A1A4.
 Then,
-$\sup_{\eta\in\mathcal{N}_{n}}\bigl\|\dot{S}_{n}(\eta)+I(\mbox{${\theta}$}_{o})\bigr\|\;\xrightarrow{P}\;0\,$,
-where $I(\mbox{${\theta}$}_{o})$ is the efficient information matrix defined in Theorem 3, and here $\|\cdot\|$ may be any fixed matrix norm; for concreteness, and because it provides the appropriate control on the inverse Hessian, we take it to be the operator norm induced by the Euclidean norm.
+$\sup_{\eta\in\mathcal{N}_{n}}\bigl\|\dot{S}_{n}(\eta)+I_{0}(\mbox{${\theta}$}_{o})\bigr\|\;\xrightarrow{P}\;0,$
+where
+$I_{0}(\mbox{${\theta}$}_{o})=\mathbb{E}_{0}\!\left[\int_{0}^{\tau}{\bf Z}{\bf Z}^{\top}Y(t)\exp\{\mbox{${\theta}$}_{o}^{\top}{\bf Z}+g_{o}(t,{\bf X})\}\,dt\right]=\mathbb{E}_{0}\!\left[\Delta\,{\bf Z}{\bf Z}^{\top}\right].$
+If, in addition, $I_{0}(\mbox{${\theta}$}_{o})$ is nonsingular, then
+$\sup_{\eta\in\mathcal{N}_{n}}\bigl\|\dot{S}_{n}(\eta)^{-1}\bigr\|=O_{p}(1).$
 Proof of Lemma 15.
@@
 Therefore,
-$\dot{S}_{n}(\eta)=\mathbb{P}_{n}\dot{s}(\mbox{${\theta}$},g)$
-and
-$\mathbb{P}\,\dot{s}(\mbox{${\theta}$},g)=E_{0}\!\left[-\int_{0}^{\tau}{\bf Z}{\bf Z}^{\top}Y(t)\exp\{\mbox{${\theta}$}^{\top}{\bf Z}+g(t,{\bf X})\}\,dt\right],$
-where $\mathbb{P}_{n}$ is the empirical measure and $\mathbb{P}=P_{0}$ denotes expectation under the true law.
-Define the matrix-valued per-observation information integrand by
-$J(\eta;{\bf W})=\int_{0}^{\tau}{\bf Z}{\bf Z}^{\top}Y(t)\exp\{\mbox{${\theta}$}^{\top}{\bf Z}+g(t,{\bf X})\}\,dt,$
-where ${\bf W}=(T,\Delta,{\bf X},{\bf Z})$ so that $\dot{S}_{n}(\eta)=-\,\mathbb{P}_{n}J(\eta;\cdot)$ and $\mathbb{P}\dot{s}(\mbox{${\theta}$},g)=-\,\mathbb{P}J(\eta;\cdot)$.
-By Assumptions A1A4, the covariates are bounded, $\|{\bf Z}\|\leq B_{{\bf Z}}$ a.s., and the hazards are uniformly bounded. Then, there exist $0<c_{h}<C_{h}<\infty$ such that
-$c_{h}\;\leq\;\exp\{\mbox{${\theta}$}^{\top}{\bf Z}+g(t,{\bf X})\}\;\leq\;C_{h}$
-for all $\eta$ in the sieve and all $(t,{\bf X},{\bf Z})$. Hence the integrand in $J(\eta;{\bf W})$ is uniformly bounded by
-$\|{\bf Z}{\bf Z}^{\top}Y(t)\exp\{\mbox{${\theta}$}^{\top}{\bf Z}+g(t,{\bf X})\}\|\;\leq\;B_{{\bf Z}}^{2}C_{h}$
-for all $\eta\in\mathcal{N}_{n}$ and all $\bf W$}.
-Thus the class $\mathcal{J}_{n}=\bigl\{J(\eta;\cdot):\eta\in\mathcal{N}_{n}\bigr\}$ has a finite envelope and inherits the same entropy bounds as the likelihood class used in Lemma 4.
-In particular, $\mathcal{J}_{n}$ is $\mathbb{P}$–Glivenko–Cantelli, so
-$\sup_{\eta\in\mathcal{N}_{n}}\bigl\|\mathbb{P}_{n}J(\eta;\cdot)-\mathbb{P}J(\eta;\cdot)\bigr\|\;\xrightarrow{P}\;0\,.$ (26)
-Fix $\eta_{1}=(\mbox{${\theta}$}_{1},g_{1})$ and $\eta_{2}=(\mbox{${\theta}$}_{2},g_{2})$ in $\mathcal{N}_{n}$. For each observation $\bf W$},
-$\bigl\|J(\eta_{1};{\bf W})-J(\eta_{2};{\bf W})\bigr\|$ $=\Biggl\|\int_{0}^{\tau}{\bf Z}{\bf Z}^{\top}Y(t)\Bigl[\exp\{\chi_{\eta_{1}}(t,{\bf X},{\bf Z})\}-\exp\{\chi_{\eta_{2}}(t,{\bf X},{\bf Z})\}\Bigr]\,dt\Biggr\|$
-$\leq\int_{0}^{\tau}\|{\bf Z}{\bf Z}^{\top}\|\,Y(t)\Bigl|\exp\{\chi_{\eta_{1}}(t,{\bf X},{\bf Z})\}-\exp\{\chi_{\eta_{2}}(t,{\bf X},{\bf Z})\}\Bigr|\,dt\,.$
-Using the mean-value theorem for the exponential and boundedness of $\bf Z$} and $g$, there exists a constant $C>0$, depending only on model constants, such that
-$\Bigl|\exp\{\mbox{${\theta}$}_{1}^{\top}{\bf Z}+g_{1}(t,{\bf X})\}-\exp\{\mbox{${\theta}$}_{2}^{\top}{\bf Z}+g_{2}(t,{\bf X})\}\Bigr|\;\leq\;C\Bigl\{\|\mbox{${\theta}$}_{1}-\mbox{${\theta}$}_{2}\|+|g_{1}(t,{\bf X})-g_{2}(t,{\bf X})|\Bigr\}\,.$
-Therefore,
-$\bigl\|J(\eta_{1};{\bf W})-J(\eta_{2};{\bf W})\bigr\|\;\leq\;C^{\prime}\int_{0}^{\tau}Y(t)\Bigl\{\|\mbox{${\theta}$}_{1}-\mbox{${\theta}$}_{2}\|+|g_{1}(t,{\bf X})-g_{2}(t,{\bf X})|\Bigr\}\,dt$
-for some $C^{\prime}>0$. Taking expectations and using $\int_{0}^{\tau}Y(t)\,dt\leq\tau$,
-$\bigl\|\mathbb{P}J(\eta_{1};\cdot)-\mathbb{P}J(\eta_{2};\cdot)\bigr\|$ $\leq C^{\prime}\tau\|\mbox{${\theta}$}_{1}-\mbox{${\theta}$}_{2}\|+C^{\prime}\,E_{0}\!\left\{\int_{0}^{\tau}Y(t)\,|g_{1}(t,{\bf X})-g_{2}(t,{\bf X})|\,dt\right\}.$
-By the Hölder regularity and boundedness assumptions on $g$, together with the fact that $d(\eta_{1},\eta_{2})$ controls the $L^{2}$–distance between $\chi_{\eta_{1}}$ and $\chi_{\eta_{2}}$, one can bound the right-hand side by $C^{\prime\prime}d(\eta_{1},\eta_{2})$ for some $C^{\prime\prime}>0$.
-In particular, for all $\eta\in\mathcal{N}_{n}$,
-$\bigl\|\mathbb{P}J(\eta;\cdot)-\mathbb{P}J(\eta_{o};\cdot)\bigr\|\;\leq\;C^{\prime\prime}d(\eta,\eta_{o})\;\leq\;C^{\prime\prime}M\tau_{n}\;\longrightarrow\;0\,$,
-since $\tau_{n}\to 0$. Hence
-$\sup_{\eta\in\mathcal{N}_{n}}\bigl\|\mathbb{P}J(\eta;\cdot)-\mathbb{P}J(\eta_{o};\cdot)\bigr\|\;\longrightarrow\;0\,.$ (27)
-By Theorem 3, the efficient score for $\mbox{${\theta}$}_{o}$ is
-$\ell^{*}_{\theta_{o}}(T,{\bf X},{\bf Z},\Delta)=\int_{0}^{\tau}\{{\bf Z}-{\bf g}^{*}(t,{\bf X})\}\,dM(t)\,$,
-and the efficient information matrix is
-$I(\mbox{${\theta}$}_{o})=E_{0}\bigl[\Delta\{{\bf Z}-{\bf g}^{*}(T,{\bf X})\}^{\otimes 2}\bigr]\,.$
-By semiparametric information theory, the efficient score satisfies the usual information identity along the least favorable submodel.
-In particular, the efficient Fisher information equals minus the expectation of the derivative of the efficient score,
-$\mathbb{P}\dot{s}(\mbox{${\theta}$}_{o},g_{o})=I(\mbox{${\theta}$}_{o})\,.$
-Recalling that $\dot{s}(\mbox{${\theta}$}_{o},g_{o})=-J(\eta_{o};{\bf W})$, this gives
-$\mathbb{P}J(\eta_{o};\cdot)=I(\mbox{${\theta}$}_{o})\,.$ (28)
-Finally, from (26), (27), and (28),
-$\sup_{\eta\in\mathcal{N}_{n}}\bigl\|\dot{S}_{n}(\eta)+I(\mbox{${\theta}$}_{o})\bigr\|$ $=\sup_{\eta\in\mathcal{N}_{n}}\bigl\|-\mathbb{P}_{n}J(\eta;\cdot)+I(\mbox{${\theta}$}_{o})\bigr\|$
+$\dot{S}_{n}(\eta)=\mathbb{P}_{n}\dot{s}(\mbox{${\theta}$},g).$
+Define the matrix-valued per-observation integrand
+$J(\eta;{\bf W})=\int_{0}^{\tau}{\bf Z}{\bf Z}^{\top}Y(t)\exp\{\mbox{${\theta}$}^{\top}{\bf Z}+g(t,{\bf X})\}\,dt,\qquad{\bf W}=(T,\Delta,{\bf X},{\bf Z}),$
+so that
+$\dot{S}_{n}(\eta)=-\,\mathbb{P}_{n}J(\eta;\cdot),\qquad\mathbb{P}\dot{s}(\mbox{${\theta}$},g)=-\,\mathbb{P}J(\eta;\cdot).$
+By Assumptions A2A3, the covariates are bounded and the sieve is uniformly bounded, so there exists a constant $C<\infty$ such that
+$\sup_{\eta\in\mathcal{N}_{n}}\|J(\eta;{\bf W})\|\leq C\qquad\text{a.s.}$
+Since $p$ is fixed, it suffices to control each matrix entry separately. The class
+$\mathcal{J}_{n}=\{J(\eta;\cdot):\eta\in\mathcal{N}_{n}\}$
+inherits the same bracketing entropy bound as the localized likelihood/score classes used in Lemma 10. Hence it is $\mathbb{P}$–Glivenko–Cantelli, and
+$\sup_{\eta\in\mathcal{N}_{n}}\bigl\|\mathbb{P}_{n}J(\eta;\cdot)-\mathbb{P}J(\eta;\cdot)\bigr\|\;\xrightarrow{P}\;0.$ (A.15)
+Next we show continuity of $\eta\mapsto\mathbb{P}J(\eta;\cdot)$ in the metric $d$.
+Let $\eta_{1}=(\mbox{${\theta}$}_{1},g_{1})$ and $\eta_{2}=(\mbox{${\theta}$}_{2},g_{2})$, and write $\phi(t,{\bf X},{\bf Z})=\chi_{\eta_{1}}(t,{\bf X},{\bf Z})-\chi_{\eta_{2}}(t,{\bf X},{\bf Z})$.
+Then, by boundedness of $\bf Z$} and the mean-value theorem for the exponential, there exists a constant $C_{1}>0$ such that
+$\bigl\|J(\eta_{1};{\bf W})-J(\eta_{2};{\bf W})\bigr\|$ $\leq C_{1}\int_{0}^{\tau}Y(t)\,|\phi(t,{\bf X},{\bf Z})|\,dt.$
+Taking expectations and applying Cauchy–Schwarz,
+$\bigl\|\mathbb{P}J(\eta_{1};\cdot)-\mathbb{P}J(\eta_{2};\cdot)\bigr\|$ $\leq C_{1}\mathbb{E}_{0}\!\left[\int_{0}^{\tau}Y(t)\,|\phi(t,{\bf X},{\bf Z})|\,dt\right]$
+$\leq C_{1}\tau^{1/2}\left\{\mathbb{E}_{0}\!\left[\int_{0}^{\tau}Y(t)\phi(t,{\bf X},{\bf Z})^{2}\,dt\right]\right\}^{1/2}.$
+Under Assumptions A1A4, the true hazard $h_{o}(t\mid{\bf X},{\bf Z})=\exp\{\mbox{${\theta}$}_{o}^{\top}{\bf Z}+g_{o}(t,{\bf X})\}$ is bounded away from zero, say $h_{o}\geq c_{h}>0$. Therefore,
+$\mathbb{E}_{0}\!\left[\int_{0}^{\tau}Y(t)\phi(t,{\bf X},{\bf Z})^{2}\,dt\right]$ $\leq c_{h}^{-1}\mathbb{E}_{0}\!\left[\int_{0}^{\tau}Y(t)h_{o}(t\mid{\bf X},{\bf Z})\phi(t,{\bf X},{\bf Z})^{2}\,dt\right]$
+$=c_{h}^{-1}\mathbb{E}_{0}\!\left[\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\right]=c_{h}^{-1}d(\eta_{1},\eta_{2})^{2},$
+where the equality follows from the compensator identity. Hence
+$\bigl\|\mathbb{P}J(\eta_{1};\cdot)-\mathbb{P}J(\eta_{2};\cdot)\bigr\|\leq C_{2}\,d(\eta_{1},\eta_{2})$
+for some constant $C_{2}>0$. In particular,
+$\sup_{\eta\in\mathcal{N}_{n}}\bigl\|\mathbb{P}J(\eta;\cdot)-\mathbb{P}J(\eta_{o};\cdot)\bigr\|\leq C_{2}M\tau_{n}\longrightarrow 0.$ (A.16)
+At the true parameter,
+$\mathbb{P}J(\eta_{o};\cdot)$ $=\mathbb{E}_{0}\!\left[\int_{0}^{\tau}{\bf Z}{\bf Z}^{\top}Y(t)\exp\{\mbox{${\theta}$}_{o}^{\top}{\bf Z}+g_{o}(t,{\bf X})\}\,dt\right]$
+$=\mathbb{E}_{0}\!\left[\Delta\,{\bf Z}{\bf Z}^{\top}\right]=I_{0}(\mbox{${\theta}$}_{o}),$
+again by the compensator identity. Therefore, using (A.15) and (A.16),
+$\sup_{\eta\in\mathcal{N}_{n}}\bigl\|\dot{S}_{n}(\eta)+I_{0}(\mbox{${\theta}$}_{o})\bigr\|$ $=\sup_{\eta\in\mathcal{N}_{n}}\bigl\|-\mathbb{P}_{n}J(\eta;\cdot)+I_{0}(\mbox{${\theta}$}_{o})\bigr\|$
 $\leq\sup_{\eta\in\mathcal{N}_{n}}\bigl\|\mathbb{P}_{n}J(\eta;\cdot)-\mathbb{P}J(\eta;\cdot)\bigr\|+\sup_{\eta\in\mathcal{N}_{n}}\bigl\|\mathbb{P}J(\eta;\cdot)-\mathbb{P}J(\eta_{o};\cdot)\bigr\|$
-$\xrightarrow{P}0+0\;=\;0.$
-This proves the lemma. ∎
+$\xrightarrow{P}0.$
+If $I_{0}(\mbox{${\theta}$}_{o})$ is nonsingular, let $\lambda_{0}=\lambda_{\min}(I_{0}(\mbox{${\theta}$}_{o}))>0$. On the event
+$\sup_{\eta\in\mathcal{N}_{n}}\bigl\|\dot{S}_{n}(\eta)+I_{0}(\mbox{${\theta}$}_{o})\bigr\|<\lambda_{0}/2,$
+all matrices $\dot{S}_{n}(\eta)$ are nonsingular and
+$\sup_{\eta\in\mathcal{N}_{n}}\bigl\|\dot{S}_{n}(\eta)^{-1}\bigr\|\leq 2/\lambda_{0}.$
+Since this event has probability tending to one, we conclude
+$\sup_{\eta\in\mathcal{N}_{n}}\bigl\|\dot{S}_{n}(\eta)^{-1}\bigr\|=O_{p}(1).$
+This completes the proof. ∎
 See 5
@@
 By Lemma 14, for all $\eta$ in the shrinking neighborhood $\mathcal{N}_{n}=\{\eta:\ d(\eta,\eta_{o})\leq M\tau_{n}\}$ for some fixed $M\geq 1$, we have the uniform bounds
-$\sup_{\eta\in\mathcal{N}_{n}}|\widetilde{\ell}_{n}(\eta)-\ell_{n}(\eta)|\;\leq\;C_{1}\delta_{n}\,$,
+$\sup_{\eta\in\mathcal{N}_{n}}|\widetilde{\ell}_{n}(\eta)-\ell_{n}(\eta)|\;\leq\;C_{1}\omega_{n}(\delta_{n})\,$,
 and
-$\sup_{\eta\in\mathcal{N}_{n}}\|\widetilde{S}_{n}(\eta)-S_{n}(\eta)\|\;\leq\;C_{2}\delta_{n}\,.$
+$\sup_{\eta\in\mathcal{N}_{n}}\|\widetilde{S}_{n}(\eta)-S_{n}(\eta)\|\;\leq\;C_{2}\omega_{n}(\delta_{n})\,.$
 Since $\eta_{o}\in\mathcal{N}_{n}$ for all $n$ large enough, subtracting the values at $\eta_{o}$ gives
-$\sup_{d(\eta,\eta_{o})\leq M\tau_{n}}\bigl|\ell_{n}(\eta)-\widetilde{\ell}_{n}(\eta)-\{\ell_{n}(\eta_{o})-\widetilde{\ell}_{n}(\eta_{o})\}\bigr|=O_{p}(\delta_{n})\,$, (29)
+$\sup_{d(\eta,\eta_{o})\leq M\tau_{n}}\bigl|\ell_{n}(\eta)-\widetilde{\ell}_{n}(\eta)-\{\ell_{n}(\eta_{o})-\widetilde{\ell}_{n}(\eta_{o})\}\bigr|=O_{p}\!\bigl(\omega_{n}(\delta_{n})\bigr)\,$, (A.17)
 and
-$\sup_{d(\eta,\eta_{o})\leq M\tau_{n}}\bigl\|\widetilde{S}_{n}(\eta)-S_{n}(\eta)\bigr\|=O_{p}(\delta_{n})\,.$ (30)
-These bounds reflect that the per-subject approximation error is $O(\delta_{n})$ and the network outputs are uniformly bounded because the sieve $\mathcal{G}(K,s,\boldsymbol{p},D)$ is constructed with bounded weights.
-By (29) and Theorem 1, on the ball $\{d(\eta,\eta_{o})\leq\tau_{n}\}$ we have
+$\sup_{d(\eta,\eta_{o})\leq M\tau_{n}}\bigl\|\widetilde{S}_{n}(\eta)-S_{n}(\eta)\bigr\|=O_{p}\!\bigl(\omega_{n}(\delta_{n})\bigr)\,.$ (A.18)
+These bounds reflect that the per-subject approximation error is controlled by the local time-modulus $\omega_{n}(\delta_{n})$ from Lemma 14, while the network outputs are uniformly bounded because the sieve $\mathcal{G}(K,s,\boldsymbol{p},D)$ is constructed with bounded weights.
+By (A.17) and Theorem 1, on the ball $\{d(\eta,\eta_{o})\leq\tau_{n}\}$ we have
 $\sup_{d(\eta,\eta_{o})\leq\tau_{n}}\bigl|(\widetilde{\ell}_{n}-\ell_{0})(\eta)-(\widetilde{\ell}_{n}-\ell_{0})(\eta_{o})\bigr|$ $\leq\sup_{d(\eta,\eta_{o})\leq\tau_{n}}\bigl|(\ell_{n}-\ell_{0})(\eta)-(\ell_{n}-\ell_{0})(\eta_{o})\bigr|$
 $\qquad+\sup_{d(\eta,\eta_{o})\leq\tau_{n}}\bigl|\widetilde{\ell}_{n}(\eta)-\ell_{n}(\eta)-\{\widetilde{\ell}_{n}(\eta_{o})-\ell_{n}(\eta_{o})\}\bigr|$
-$=O_{p}(\tau_{n}^{2})+O_{p}(\delta_{n})=O_{p}(\tau_{n}^{2}),$
-[…]
+$=O_{p}(\tau_{n}^{2})+O_{p}\!\bigl(\omega_{n}(\delta_{n})\bigr)=O_{p}(\tau_{n}^{2}),$
+because $\omega_{n}(\delta_{n})=o(\tau_{n}^{2})$ by assumption.
+The population loss satisfies $\ell_{0}(\eta)-\ell_{0}(\eta_{o})\asymp-d(\eta,\eta_{o})^{2}$ by Lemma 3.
+Therefore, the same M-estimation argument used in the proof of Theorem 1 for the oracle estimator (via Theorem 3.4.1 in van der Vaart and Wellner, 1996) applies to $\widetilde{\ell}_{n}$, yielding $d(\widetilde{\eta},\eta_{o})=O_{p}(\tau_{n})$ and $\|\widetilde{}\mbox{${\theta}$}-\mbox{${\theta}$}_{o}\|=O_{p}(\tau_{n})$, which proves part 1.
 By the score equations, $S_{n}(\widehat{\eta})={\bf 0}$ and $\widetilde{S}_{n}(\widetilde{\eta})={\bf 0}$. Therefore,
@@
 In addition,
-$S_{n}(\widetilde{\eta})-S_{n}(\widehat{\eta})=\underbrace{S_{n}(\widetilde{g},\widetilde{}\mbox{${\theta}$})-S_{n}(\widetilde{g},\widehat{}\mbox{${\theta}$})}_{A}+\underbrace{S_{n}(\widetilde{g},\widehat{}\mbox{${\theta}$})-S_{n}(\widehat{g},\widehat{}\mbox{${\theta}$})}_{B}.$
-and (31) becomes
+$S_{n}(\widetilde{\eta})-S_{n}(\widehat{\eta})=\underbrace{S_{n}(\widetilde{}\mbox{${\theta}$},\widetilde{g})-S_{n}(\widehat{}\mbox{${\theta}$},\widetilde{g})}_{A}+\underbrace{S_{n}(\widehat{}\mbox{${\theta}$},\widetilde{g})-S_{n}(\widehat{}\mbox{${\theta}$},\widehat{g})}_{B}\,.$
 […]
-By the mean value theorem in $\theta$}, there exists $\mbox{${\theta}$}^{\dagger}$ on the segment between $\widehat{}\mbox{${\theta}$}$ and $\widetilde{}\mbox{${\theta}$}$ such that
-$A=\dot{S}_{n}(\widetilde{g},\mbox{${\theta}$}^{\dagger})\,(\widetilde{}\mbox{${\theta}$}-\widehat{}\mbox{${\theta}$})\,.$
-By Lemma 9, $\dot{S}_{n}(\widetilde{g},\mbox{${\theta}$}^{\dagger})\xrightarrow{P}-I(\mbox{${\theta}$}_{o})$ and $\|\dot{S}_{n}(\widetilde{g},\mbox{${\theta}$}^{\dagger})^{-1}\|=O_{p}(1)$.
-Substituting this into (31), we obtain
+$A=\dot{S}_{n}(\mbox{${\theta}$}^{\dagger},\widetilde{g})\,(\widetilde{}\mbox{${\theta}$}-\widehat{}\mbox{${\theta}$})\,.$
+By Lemma 15, $\dot{S}_{n}(\mbox{${\theta}$}^{\dagger},\widetilde{g})\xrightarrow{P}-I_{0}(\mbox{${\theta}$}_{o})$ and $\|\dot{S}_{n}(\mbox{${\theta}$}^{\dagger},\widetilde{g})^{-1}\|=O_{p}(1)$.
+Substituting this into (A.19), we obtain
 […]
@@
 Lemma 14 gives uniformly over $\mathcal{N}_{n}$,
-$\|R_{n}(\eta)\|\leq C\delta_{n},$
+$\|R_{n}(\eta)\|\leq C\omega_{n}(\delta_{n})\,$,
 so
-$\sqrt{n}\,R_{n}(\widetilde{\eta})=O_{p}(\sqrt{n}\delta_{n})=o_{p}(1),$
-because $\sqrt{n}\delta_{n}\to 0$ by assumption.
+$\sqrt{n}\,R_{n}(\widetilde{\eta})=O_{p}\!\bigl(\sqrt{n}\,\omega_{n}(\delta_{n})\bigr)=o_{p}(1),$
+because $\sqrt{n}\,\omega_{n}(\delta_{n})\to 0$ by assumption.
 We decompose
-$B=S_{n}(\widetilde{g},\widehat{}\mbox{${\theta}$})-S_{n}(\widehat{g},\widehat{}\mbox{${\theta}$})=\{\mathbb{P}S(\widetilde{g},\widehat{}\mbox{${\theta}$})-\mathbb{P}S(\widehat{g},\widehat{}\mbox{${\theta}$})\}+\{(\mathbb{P}_{n}-\mathbb{P})S(\widetilde{g},\widehat{}\mbox{${\theta}$})-(\mathbb{P}_{n}-\mathbb{P})S(\widehat{g},\widehat{}\mbox{${\theta}$})\},$
+[…]
 where $S(\eta;{\bf W})$ is the per–observation score in $\theta$} and $S_{n}(\eta)=\mathbb{P}_{n}S(\eta;\cdot)$. The second–order expansion of the population score (see the proof of Theorem 4) yields
-$\|\mathbb{P}S(\widetilde{g},\widehat{}\mbox{${\theta}$})-\mathbb{P}S(\widehat{g},\widehat{}\mbox{${\theta}$})\|\lesssim\|\widetilde{g}-g_{o}\|_{L^{2}(\Delta)}^{2}+\|\widehat{g}-g_{o}\|_{L^{2}(\Delta)}^{2}=O_{p}(\tau_{n}^{2})\,.$
-Moreover, Lemma 4 applied to the localized score class $\{S(g,\widehat{}\mbox{${\theta}$};\cdot):d((g,\widehat{}\mbox{${\theta}$}),\eta_{o})\leq C\tau_{n}\}$ implies
-$\sup_{d((g,\widehat{\theta}),\eta_{o})\leq C\tau_{n}}|(\mathbb{P}_{n}-\mathbb{P})S(g,\widehat{}\mbox{${\theta}$};\cdot)|=O_{p}(\tau_{n}^{2})\,$,
-and since both $(\widetilde{g},\widehat{}\mbox{${\theta}$})$ and $(\widehat{g},\widehat{}\mbox{${\theta}$})$ lie in this ball with probability 1o(1)1-o(1), we obtain
+$\|\mathbb{P}S(\widehat{}\mbox{${\theta}$},\widetilde{g})-\mathbb{P}S(\widehat{}\mbox{${\theta}$},\widehat{g})\|\lesssim\|\widetilde{g}-g_{o}\|_{L^{2}(\Delta)}^{2}+\|\widehat{g}-g_{o}\|_{L^{2}(\Delta)}^{2}=O_{p}(\tau_{n}^{2})\,.$
+Moreover, Lemma 10 applied to the localized score class $\{S(\widehat{}\mbox{${\theta}$},g;\cdot):d((\widehat{}\mbox{${\theta}$},g),\eta_{o})\leq C\tau_{n}\}$ implies
+$\sup_{d((\widehat{\theta},g),\eta_{o})\leq C\tau_{n}}|(\mathbb{P}_{n}-\mathbb{P})S(\widehat{}\mbox{${\theta}$},g;\cdot)|=O_{p}(\tau_{n}^{2})\,$,
+[…]
 $\|B\|=O_{p}(\tau_{n}^{2}),\qquad\sqrt{n}B=o_{p}(1),$
 because $\sqrt{n}\tau_{n}^{2}\to 0$.
-Using (32) together with $\dot{S}_{n}(\widetilde{g},\mbox{${\theta}$}^{\dagger})^{-1}=O_{p}(1)$, $\sqrt{n}B=o_{p}(1)$, and $\sqrt{n}R_{n}(\widetilde{\eta})=o_{p}(1)$, we obtain
+Using (A.20) together with $\dot{S}_{n}(\mbox{${\theta}$}^{\dagger},\widetilde{g})^{-1}=O_{p}(1)$, $\sqrt{n}B=o_{p}(1)$, and $\sqrt{n}R_{n}(\widetilde{\eta})=o_{p}(1)$, we obtain
 $\sqrt{n}(\widetilde{}\mbox{${\theta}$}-\widehat{}\mbox{${\theta}$})=o_{p}(1)\,.$
@@
 as claimed. ∎
+Appendix B Proofs of Section 5
+See 4
+Proof of Lemma 4.
+Along the joint submodel,
+$h_{\varepsilon}(u\mid{\bf X}_{0},{\bf Z}_{0})=h_{o}(u\mid{\bf X}_{0},{\bf Z}_{0})\exp\!\Big\{\varepsilon\,{\bf a}^{\top}({\bf Z}_{0}-{\bf g}^{*}(u,{\bf X}_{0}))+\varepsilon\,b(u,{\bf X}_{0})\Big\}\,.$
+Differentiating at $\varepsilon=0$ gives
+$\left.\frac{d}{d\varepsilon}h_{\varepsilon}(u\mid{\bf X}_{0},{\bf Z}_{0})\right|_{\varepsilon=0}=h_{o}(u\mid{\bf X}_{0},{\bf Z}_{0})\Big\{{\bf a}^{\top}({\bf Z}_{0}-{\bf g}^{*}(u,{\bf X}_{0}))+b(u,{\bf X}_{0})\Big\}\,.$
+Integration over $[0,t]$ yields the stated result. ∎
+See 5
+Proof of Lemma 5.
+By A7, $Q:\overline{\mathcal{T}}_{g_{o}}\to L^{2}(P_{{\bf X},{\bf Z}})$ is a bounded linear operator.
+Next, we invoke assumption A7, there exists $\psi_{t}^{0}(\cdot,\cdot;{\bf X}_{0},{\bf Z}_{0})\in L^{2}(P_{{\bf X},{\bf Z}};\overline{\mathcal{T}}_{g_{o}})$ such that for all $b\in\mathcal{\overline{}}{\mathcal{T}}_{g_{o}}$,
+$(Qb)({\bf X}_{0},{\bf Z}_{0})=\langle\psi_{t}^{0}(\cdot,\cdot;{\bf X}_{0},{\bf Z}_{0}),\,b\rangle_{0}\quad\text{in }L^{2}(P_{{\bf X},{\bf Z}})\,.$ (B.1)
+Uniqueness follows from the Riesz representation theorem in the Hilbert space $\overline{\mathcal{T}}_{g_{o}}$: if $\psi,\tilde{\psi}\in L^{2}(P_{{\bf X},{\bf Z}};\overline{\mathcal{T}}_{g_{o}})$ both satisfy (B.1), then $\langle\psi-\tilde{\psi},b\rangle_{0}=0$ in $L^{2}(P_{{\bf X},{\bf Z}})$ for all $b$, hence $\psi=\tilde{\psi}$ in $L^{2}(P_{{\bf X},{\bf Z}};\overline{\mathcal{T}}_{g_{o}})$.
+For each $b\in\overline{\mathcal{T}}_{g_{o}}$, the martingale isometry associated with (5.1) yields
+$\langle\psi_{t}^{0}(\cdot,\cdot;{\bf X}_{0},{\bf Z}_{0}),\,b\rangle_{0}=\mathbb{E}_{0}\!\left[\left(\int_{0}^{\tau}\psi_{t}^{0}(u,{\bf X};{\bf X}_{0},{\bf Z}_{0})\,dM(u)\right)\left(\int_{0}^{\tau}b(u,{\bf X})\,dM(u)\right)\ \middle|\ {\bf X}_{0},{\bf Z}_{0}\right],$
+and since $\langle\psi_{t}^{0},b\rangle_{0}=L_{t}(b)$ in $L^{2}(P_{{\bf X},{\bf Z}})$, the nuisance contribution to the canonical gradient in the $L^{2}(P_{{\bf X},{\bf Z}})$ sense is $\int_{0}^{\tau}\psi_{t}^{0}(u,{\bf X};{\bf X}_{0},{\bf Z}_{0})\,dM(u)$.
+The parametric derivative term in (5.9) equals ${\bf a}^{\top}c_{0,t}({\bf X}_{0},{\bf Z}_{0})$.
+Since $\ell_{\theta_{o}}^{*}$ is the efficient score and $I(\mbox{${\theta}$}_{o})=\mathbb{E}_{0}\{\ell_{\theta_{o}}^{*}({\bf W})^{\otimes 2}\}$, the corresponding efficient influence function for $\mbox{${\theta}$}_{o}$ is $I(\mbox{${\theta}$}_{o})^{-1}\ell_{\theta_{o}}^{*}({\bf W})$.
+Thus the parametric contribution to the EIF for $H_{o}(t\mid{\bf X}_{0},{\bf Z}_{0})$ is
+$c_{0,t}({\bf X}_{0},{\bf Z}_{0})^{\top}I(\mbox{${\theta}$}_{o})^{-1}\ell_{\theta_{o}}^{*}({\bf W})\,.$
+Combining nuisance and parametric pieces gives (5.11). Finally, using (5.9), (5.10), and the score decomposition (5.6), one checks that
+$\left.\frac{d}{d\varepsilon}H_{\varepsilon}(t\mid{\bf X}_{0},{\bf Z}_{0})\right|_{\varepsilon=0}-\mathbb{E}_{0}\!\big[\phi_{t}^{0}({\bf W};{\bf X}_{0},{\bf Z}_{0})\,S({\bf W};{\bf a},b)\mid{\bf X}_{0},{\bf Z}_{0}\big]=0\quad\text{in }L^{2}(P_{{\bf X},{\bf Z}}),$
+which proves pathwise differentiability of $H_{P}(t\mid{\bf X}_{0},{\bf Z}_{0})$ as an $L^{2}(P_{{\bf X},{\bf Z}})$-valued functional. ∎
+See 6
+Proof of Lemma 6.
+Since $\bf X$} takes values in the finite set $\{{\bf x}^{(1)},\ldots,{\bf x}^{(J)}\}$, every $b\in\overline{\mathcal{T}}_{g_{o}}$ admits the representation
+$b(u,{\bf X})=\sum_{j=1}^{J}b_{j}(u)\,\mathbf{1}\{{\bf X}={\bf x}^{(j)}\},\qquad b_{j}\in L_{2}([0,\tau]).$
+Let $j_{0}$ be such that ${\bf X}_{0}={\bf x}^{(j_{0})}$. Using the inner product (5.1), for any $\psi\in\overline{\mathcal{T}}_{g_{o}}$ we obtain
+$\langle\psi,b\rangle_{0}$ $=\mathbb{E}_{0}\!\left[\int_{0}^{\tau}\psi(u,{\bf X};{\bf X}_{0},{\bf Z}_{0})\,b(u,{\bf X})\,Y(u)h_{o}(u\mid{\bf X},{\bf Z})\,du\right]$
+$=\sum_{j=1}^{J}\mathbb{E}_{0}\!\left[\mathbf{1}\{{\bf X}={\bf x}^{(j)}\}\int_{0}^{\tau}\psi(u,{\bf x}^{(j)};{\bf X}_{0},{\bf Z}_{0})\,b_{j}(u)\,Y(u)h_{o}(u\mid{\bf X},{\bf Z})\,du\right]$
+$=\sum_{j=1}^{J}\mathbb{P}_{0}({\bf X}={\bf x}^{(j)})\int_{0}^{\tau}\psi(u,{\bf x}^{(j)};{\bf X}_{0},{\bf Z}_{0})\,b_{j}(u)\,\mathbb{E}_{0}\!\left[Y(u)h_{o}(u\mid{\bf X},{\bf Z})\mid{\bf X}={\bf x}^{(j)}\right]du$
+$=\sum_{j=1}^{J}\mathbb{P}_{0}({\bf X}={\bf x}^{(j)})\int_{0}^{\tau}\psi(u,{\bf x}^{(j)};{\bf X}_{0},{\bf Z}_{0})\,b_{j}(u)\,A_{j}(u)\,du.$
+On the other hand, by definition of the evaluation functional $L_{t}$ (as in (5.7)),
+$L_{t}(b)=\int_{0}^{t}h_{o}(u\mid{\bf X}_{0},{\bf Z}_{0})\,b(u,{\bf X}_{0})\,du=\int_{0}^{t}h_{o}(u\mid{\bf X}_{0},{\bf Z}_{0})\,b_{j_{0}}(u)\,du.$
+Therefore, the Riesz equation $\langle\psi_{t}^{0},b\rangle_{0}=L_{t}(b)$ for all $b\in\overline{\mathcal{T}}_{g_{o}}$ implies that $\psi_{t}^{0}(u,{\bf x}^{(j)};{\bf X}_{0},{\bf Z}_{0})=0$ for $j\neq j_{0}$ (a.e.
+$u$), and for $j=j_{0}$,
+$\Pr({\bf X}={\bf X}_{0})\int_{0}^{\tau}\psi_{t}^{0}(u,{\bf X}_{0};{\bf X}_{0},{\bf Z}_{0})\,b_{j_{0}}(u)\,A_{{\bf X}_{0}}(u)\,du=\int_{0}^{t}h_{o}(u\mid{\bf X}_{0},{\bf Z}_{0})\,b_{j_{0}}(u)\,du,\quad\forall\,b_{j_{0}}.$
+Since this holds for all $b_{j_{0}}\in L_{2}([0,\tau])$, it follows that for a.e. $u\in[0,\tau]$,
+$\Pr({\bf X}={\bf X}_{0})\,A_{{\bf X}_{0}}(u)\,\psi_{t}^{0}(u,{\bf X}_{0};{\bf X}_{0},{\bf Z}_{0})=\mathbf{1}(u\leq t)\,h_{o}(u\mid{\bf X}_{0},{\bf Z}_{0})\,.$
+Thus,
+$\psi_{t}^{0}(u,{\bf X}_{0};{\bf X}_{0},{\bf Z}_{0})=\mathbf{1}(u\leq t)\,\frac{h_{o}(u\mid{\bf X}_{0},{\bf Z}_{0})}{\Pr({\bf X}={\bf X}_{0})\,A_{{\bf X}_{0}}(u)}\quad\text{for a.e. }u.$
+Combining this with the fact that $\psi_{t}^{0}(u,{\bf x}^{(j)};{\bf X}_{0},{\bf Z}_{0})=0$ for all $j\neq j_{0}$ (a.e. $u$), we obtain for general $\bf X$},
+$\psi_{t}^{0}(u,{\bf X};{\bf X}_{0},{\bf Z}_{0})=\mathbf{1}({\bf X}={\bf X}_{0})\,\psi_{t}^{0}(u,{\bf X}_{0};{\bf X}_{0},{\bf Z}_{0})=\mathbf{1}(u\leq t)\mathbf{1}({\bf X}={\bf X}_{0})\frac{h_{o}(u\mid{\bf X}_{0},{\bf Z}_{0})}{\Pr({\bf X}={\bf X}_{0})\,A_{{\bf X}_{0}}(u)},$
+which is (5.12). ∎
+See 7
+Proof of Lemma 7.
+Fix $k\in\{1,\dots,K\}$ and write $\mathcal{F}_{-k}$ for the sigma-field generated by the training sample $\{W_{j}:j\in\mathcal{I}_{k}^{c}\}$.
+By Assumption A9, conditional on $\mathcal{F}_{-k}$ the functions $\widetilde{g}^{(-k)}$, $\widetilde{}\mbox{${\theta}$}^{(-k)}$ and $\widehat{\bf g}^{*,(-k)}$ are fixed, and the observations $\{{\bf W}_{i}:i\in\mathcal{I}_{k}\}$ are i.i.d.
+from $\mathbb{P}_{0}$.
+For $i\in\mathcal{I}_{k}$ set
+$R_{i,k}=\widehat{\ell}^{*,(-k)}_{\theta_{o}}({\bf W}_{i})-\ell^{*}_{\theta_{o}}({\bf W}_{i}).$
+Insert and subtract the intermediate integral $\int_{0}^{\tau}\{{\bf Z}_{i}-\widehat{\bf g}^{*,(-k)}(u,{\bf X}_{i})\}\,dM_{i}(u)$ to obtain
+$R_{i,k}=\int_{0}^{\tau}\big\{\widehat{\bf g}^{*,(-k)}(u,{\bf X}_{i})-{\bf g}^{*}(u,{\bf X}_{i})\big\}\,dM_{i}(u)+\int_{0}^{\tau}\{{\bf Z}_{i}-\widehat{\bf g}^{*,(-k)}(u,{\bf X}_{i})\}\,d\!\big(\widehat{M}^{(-k)}_{i}-M_{i}\big)(u).$
+Denote the first and second terms by $R^{(1)}_{i,k}$ and $R^{(2)}_{i,k}$, respectively.
+Consider $R^{(1)}_{i,k}$.
+Conditional on $\mathcal{F}_{-k}$, the integrand $\widehat{\bf g}^{*,(-k)}-{\bf g}^{*}$ is predictable, and $M_{i}$ is a square-integrable martingale.
+Therefore, the martingale isometry yields
+$E_{0}\!\left[\big(R^{(1)}_{i,k}\big)^{2}\mid\mathcal{F}_{-k}\right]$ $=$ $E_{0}\!\left[\int_{0}^{\tau}\big\{\widehat{\bf g}^{*,(-k)}(u,{\bf X})-{\bf g}^{*}(u,{\bf X})\big\}^{2}Y(u)h_{o}(u\mid{\bf X},{\bf Z})\,du\ \Big|\ \mathcal{F}_{-k}\right]$
+$=$ $\|\widehat{\bf g}^{*,(-k)}-{\bf g}^{*}\|_{0}^{2}.$
+By Assumption A9, the right-hand side is $O_{p}(\tau_{n}^{2})$ uniformly in $k$. Since $\{R^{(1)}_{i,k}:i\in\mathcal{I}_{k}\}$ are i.i.d. conditional on $\mathcal{F}_{-k}$, we have
+${\mathbb{V}}_{0}\!\left(\frac{1}{\sqrt{n}}\sum_{i\in\mathcal{I}_{k}}R^{(1)}_{i,k}\ \Big|\ \mathcal{F}_{-k}\right)=\frac{|\mathcal{I}_{k}|}{n}{\mathbb{V}}_{0}\!\left(R^{(1)}_{1,k}\mid\mathcal{F}_{-k}\right)\leq\frac{|\mathcal{I}_{k}|}{n}E_{0}\!\left[\big(R^{(1)}_{1,k}\big)^{2}\mid\mathcal{F}_{-k}\right]=O_{p}(\tau_{n}^{2}).$
+Chebyshev’s inequality implies
+$\frac{1}{\sqrt{n}}\sum_{i\in\mathcal{I}_{k}}R^{(1)}_{i,k}=O_{p}(\tau_{n})=o_{p}(1),$
+because $\tau_{n}\to 0$. Next consider $R^{(2)}_{i,k}$. Since
+$\widehat{M}^{(-k)}_{i}(u)-M_{i}(u)=-\int_{0}^{u}Y_{i}(s)\Big\{\widehat{h}^{(-k)}(s\mid{\bf X}_{i},{\bf Z}_{i})-h_{o}(s\mid{\bf X}_{i},{\bf Z}_{i})\Big\}\,ds,$
+we have
+$R^{(2)}_{i,k}=-\int_{0}^{\tau}\{{\bf Z}_{i}-\widehat{\bf g}^{*,(-k)}(u,{\bf X}_{i})\}\,Y_{i}(u)\Big\{\widehat{h}^{(-k)}(u\mid{\bf X}_{i},{\bf Z}_{i})-h_{o}(u\mid{\bf X}_{i},{\bf Z}_{i})\Big\}\,du.$ (B.2)
+Write
+$\widehat{h}^{(-k)}(u\mid{\bf X},{\bf Z})-h_{o}(u\mid{\bf X},{\bf Z})$ $=h_{o}(u\mid{\bf X},{\bf Z})$ (B.3)
+[…]
+By the mean value theorem, for each $(u,{\bf X},{\bf Z})$ there exists a random $\xi^{(-k)}(u,{\bf X},{\bf Z})$ between 0 and $\delta^{(-k)}(u,{\bf X},{\bf Z})=\widetilde{g}^{(-k)}(u,{\bf X})-g_{o}(u,{\bf X})+{\bf Z}^{\top}(\widetilde{}\mbox{${\theta}$}^{(-k)}-\mbox{${\theta}$}_{o})$ such that
+$\widehat{h}^{(-k)}(u\mid{\bf X},{\bf Z})-h_{o}(u\mid{\bf X},{\bf Z})=h_{o}(u\mid{\bf X},{\bf Z})\exp\{\xi^{(-k)}(u,{\bf X},{\bf Z})\}\delta^{(-k)}(u,{\bf X},{\bf Z}).$
+By the bounded-sieve condition for $\widehat{g}^{(-k)}$, boundedness of $g_{o}$, boundedness of $\bf Z$}, and boundedness of the parameter space for $\theta$}, there exists a deterministic constant $C<\infty$ such that
+$\sup_{u\leq\tau,\,{\bf X},\,{\bf Z}}\left|\delta^{(-k)}(u,{\bf X},{\bf Z})\right|\leq C$
+with probability tending to one. Therefore, uniformly on this event, the exponential map admits the second-order expansion
+$e^{x}=1+x+R_{C}(x),\qquad|R_{C}(x)|\leq C_{1}x^{2},\quad|x|\leq C,$
+for some finite constant $C_{1}$. Hence
+$\widehat{h}^{(-k)}(u\mid{\bf X},{\bf Z})-h_{o}(u\mid{\bf X},{\bf Z})=h_{o}(u\mid{\bf X},{\bf Z})\,\delta^{(-k)}(u,{\bf X},{\bf Z})+h_{o}(u\mid{\bf X},{\bf Z})\,R^{(-k)}(u,{\bf X},{\bf Z}),$
+where $R^{(-k)}(u,{\bf X},{\bf Z})=e^{\delta^{(-k)}(u,{\bf X},{\bf Z})}-1-\delta^{(-k)}(u,{\bf X},{\bf Z})$ and
+$|R^{(-k)}(u,{\bf X},{\bf Z})|\leq C_{1}\bigl(\delta^{(-k)}(u,{\bf X},{\bf Z})\bigr)^{2}.$
+Substituting into (B.2) gives
+$R^{(2)}_{i,k}$ $=-\int_{0}^{\tau}\{{\bf Z}_{i}-\widehat{\bf g}^{*,(-k)}(u,{\bf X}_{i})\}\,Y_{i}(u)h_{o}(u\mid{\bf X}_{i},{\bf Z}_{i})\delta^{(-k)}(u,{\bf X}_{i},{\bf Z}_{i}),du+\;\widetilde{R}_{i,k},$
+where the remainder term $\widetilde{R}_{i,k}$ satisfies
+$E_{0}\!\left[|\widetilde{R}_{i,k}|\mid\mathcal{F}_{-k}\right]\lesssim\|\widetilde{g}^{(-k)}-g_{o}\|_{0}^{2}+\|\widetilde{}\mbox{${\theta}$}^{(-k)}-\mbox{${\theta}$}_{o}\|^{2}\,.$ (B.4)
+To verify (B.4), note first that
+$|\widetilde{R}_{i,k}|\lesssim\int_{0}^{\tau}|{\bf Z}_{i}-\widehat{\bf g}^{*,(-k)}(u,{\bf X}_{i})|\,Y_{i}(u)h_{o}(u\mid{\bf X}_{i},{\bf Z}_{i})\,|\delta^{(-k)}(u,{\bf X}_{i},{\bf Z}_{i})|^{2}\,du\,.$
+By boundedness of ${\bf Z}_{i}-\widehat{\bf g}^{*,(-k)}(u,{\bf X}_{i})$ and assumptions A2 and A6, the factor
+$|{\bf Z}_{i}-\widehat{\bf g}^{*,(-k)}(u,{\bf X}_{i})|\,Y_{i}(u)h_{o}(u\mid{\bf X}_{i},{\bf Z}_{i})$
+is bounded by a random quantity with finite expectation, uniformly over the relevant indices. Therefore, conditional on $\mathcal{F}_{-k}$,
+$E_{0}\!\left[|\widetilde{R}_{i,k}|\mid\mathcal{F}_{-k}\right]\lesssim E_{0}\!\left[\int_{0}^{\tau}\bigl|\delta^{(-k)}(u,{\bf X}_{i},{\bf Z}_{i})\bigr|^{2}\,du\,\middle|\,\mathcal{F}_{-k}\right].$
+Expanding $\delta^{(-k)}$ and using $(a+b)^{2}\leq 2a^{2}+2b^{2}$, we obtain (B.4).
+Now treat the leading linear term in $R^{(2)}_{i,k}$.
+Conditional on $\mathcal{F}_{-k}$, Cauchy–Schwarz and Assumption A9 imply that the conditional second moment of the integrand is bounded by a constant multiple of $\|\widetilde{g}^{(-k)}-g_{o}\|_{0}^{2}+\|\widetilde{}\mbox{${\theta}$}^{(-k)}-\mbox{${\theta}$}_{o}\|^{2}$, and hence
+$E_{0}\!\left[\big(R^{(2)}_{i,k}-\widetilde{R}_{i,k}\big)^{2}\mid\mathcal{F}_{-k}\right]\lesssim\|\widetilde{g}^{(-k)}-g_{o}\|_{0}^{2}+\|\widetilde{}\mbox{${\theta}$}^{(-k)}-\mbox{${\theta}$}_{o}\|^{2}=O_{p}(\tau_{n}^{2}+n^{-1}).$
+Since $\{R^{(2)}_{i,k}:i\in\mathcal{I}_{k}\}$ are i.i.d. conditional on $\mathcal{F}_{-k}$, it follows that
+${\mathbb{V}}_{0}\!\left(\frac{1}{\sqrt{n}}\sum_{i\in\mathcal{I}_{k}}(R^{(2)}_{i,k}-\widetilde{R}_{i,k})\ \Big|\ \mathcal{F}_{-k}\right)=\frac{|\mathcal{I}_{k}|}{n}{\mathbb{V}}_{0}(R^{(2)}_{1,k}-\widetilde{R}_{1,k}\mid\mathcal{F}_{-k})\lesssim\tau_{n}^{2}+n^{-1},$
+and therefore
+$\frac{1}{\sqrt{n}}\sum_{i\in\mathcal{I}_{k}}(R^{(2)}_{i,k}-\widetilde{R}_{i,k})=O_{p}(\tau_{n}+n^{-1/2})=o_{p}(1).$
+It remains to control the remainder contribution coming from $\widetilde{R}_{i,k}$.
+By Assumption A9, $\|\widetilde{g}^{(-k)}-g_{o}\|_{0}=O_{p}(\tau_{n})$ and $\|\widetilde{}\mbox{${\theta}$}^{(-k)}-\mbox{${\theta}$}_{o}\|=O_{p}(n^{-1/2})$, uniformly in $k$.
+Plugging these into (B.4) yields
+$E_{0}\!\left[|\widetilde{R}_{i,k}|\mid\mathcal{F}_{-k}\right]=O_{p}(\tau_{n}^{2}+n^{-1}),$
+so
+$E_{0}\!\left[\left|\frac{1}{\sqrt{n}}\sum_{i\in\mathcal{I}_{k}}\widetilde{R}_{i,k}\right|\ \Big|\ \mathcal{F}_{-k}\right]\leq\frac{|\mathcal{I}_{k}|}{\sqrt{n}}\,E_{0}\!\left[|\widetilde{R}_{1,k}|\mid\mathcal{F}_{-k}\right]=O_{p}(\sqrt{n}\,\tau_{n}^{2})+O_{p}(n^{-1/2})=o_{p}(1),$
+because $\sqrt{n}\,\tau_{n}^{2}\to 0$. Combining the bounds for $R^{(1)}_{i,k}$ and $R^{(2)}_{i,k}$ yields
+$\frac{1}{\sqrt{n}}\sum_{i\in\mathcal{I}_{k}}R_{i,k}=o_{p}(1)\qquad\text{for each }k\,.$
+Since $K$ is fixed, summing over $k=1,\dots,K$ proves (5.14) and completes the proof. ∎
+See 8
+Proof of Lemma 8.
+Write $\mathbb{P}_{n,-k}$ for the empirical measure on the training subsample $\mathcal{I}_{k}^{c}$ and define
+$\Psi_{h}({\bf W})=\Delta\,\{{\bf Z}-h(T,{\bf X})\}^{\otimes 2}.$
+Then $\widehat{I}^{(-k)}=\mathbb{P}_{n,-k}\Psi_{\widehat{\bf g}^{*,(-k)}}$ and $I(\theta_{o})=\mathbb{P}_{0}\Psi_{g^{*}}$, hence
+$\widehat{I}^{(-k)}-I(\mbox{${\theta}$}_{o})=(\mathbb{P}_{n,-k}-\mathbb{P}_{0})\Psi_{\widehat{\bf g}^{*,(-k)}}+\mathbb{P}_{0}\big(\Psi_{\widehat{\bf g}^{*,(-k)}}-\Psi_{{\bf g}^{*}}\big).$
+We first control the second (bias) term. Let $r^{*}(T,{\bf X})={\bf Z}-g^{*}(T,{\bf X})$ and $\widehat{r}^{(-k)}(T,{\bf X})={\bf Z}-\widehat{g}^{*,(-k)}(T,{\bf X})$. Expanding,
+$\Psi_{\widehat{\bf g}^{*,(-k)}}({\bf W})-\Psi_{{\bf g}^{*}}({\bf W})=\Delta\Big[(\widehat{r}^{(-k)}-r^{*})\otimes(\widehat{r}^{(-k)}-r^{*})+(\widehat{r}^{(-k)}-r^{*})\otimes r^{*}+r^{*}\otimes(\widehat{r}^{(-k)}-r^{*})\Big].$
+Under A2 and A6, $\|r^{*}\|$ is bounded almost surely and $\|\widehat{r}^{(-k)}-r^{*}\|=\|\widehat{\bf g}^{*,(-k)}-{\bf g}^{*}\|$. Therefore, for a constant $C_{1}<\infty$,
+$\big\|\mathbb{P}_{0}(\Psi_{\widehat{\bf g}^{*,(-k)}}-\Psi_{{\bf g}^{*}})\big\|\leq C_{1}\,\mathbb{P}_{0}\|\widehat{\bf g}^{*,(-k)}(T,{\bf X})-{\bf g}^{*}(T,{\bf X})\|+C_{1}\,\mathbb{P}_{0}\|\widehat{\bf g}^{*,(-k)}(T,{\bf X})-{\bf g}^{*}(T,{\bf X})\|^{2}.$
+By Cauchy–Schwarz and the definition of $\|\cdot\|_{0}$, each of the two expectations above is $O_{p}(\|\widehat{\bf g}^{*,(-k)}-{\bf g}^{*}\|_{0})$ and $O_{p}(\|\widehat{\bf g}^{*,(-k)}-{\bf g}^{*}\|_{0}^{2})$, respectively.
+Using A9, the bias term is $O_{p}(\tau_{n})+O_{p}(\tau_{n}^{2})=o_{p}(1)$, uniformly in $k$.
+For the empirical term $(\mathbb{P}_{n,-k}-\mathbb{P}_{0})\Psi_{\widehat{\bf g}^{*,(-k)}}$, condition on the training sigma-field in A9.
+Conditional on it, $\widehat{\bf g}^{*,(-k)}$ is fixed and the summands $\Psi_{\widehat{\bf g}^{*,(-k)}}({\bf W}_{j})$, $j\in\mathcal{I}_{k}^{c}$, are i.i.d.
+Moreover, under A2, A6 and boundedness of $\widehat{\bf g}^{*,(-k)}$ induced by the model setup, $\|\Psi_{\widehat{\bf g}^{*,(-k)}}({\bf W})\|$ is uniformly bounded by a constant.
+Hence each entry of $(\mathbb{P}_{n,-k}-\mathbb{P}_{0})\Psi_{\widehat{\bf g}^{*,(-k)}}$ has conditional variance of order $|\mathcal{I}_{k}^{c}|^{-1}$, and a conditional Chebyshev (or LLN) yields
+$(\mathbb{P}_{n,-k}-\mathbb{P}_{0})\Psi_{\widehat{\bf g}^{*,(-k)}}=o_{p}(1),$
+uniformly in $k$. Combining with the bias bound proves (a).
+For (b), fix $k$ and write $A=\widehat{I}^{(-k)}$ and $B=I(\theta_{o})$.
+By (a), $\|A-B\|=o_{p}(1)$ uniformly in $k$, and by A10, $\|B^{-1}\|\leq c_{0}^{-1}$.
+On the event $\|A-B\|\leq(2\|B^{-1}\|)^{-1}$, the matrix $A$ is invertible and $\|A^{-1}\|\leq 2\|B^{-1}\|$.
+Using the resolvent identity,
+$A^{-1}-B^{-1}=A^{-1}(B-A)B^{-1},$
+we obtain
+$\|A^{-1}-B^{-1}\|\leq\|A^{-1}\|\,\|A-B\|\,\|B^{-1}\|\leq 2\|B^{-1}\|^{2}\,\|A-B\|.$
+Taking maxima over $k$ and using $\max_{k}\|A-B\|=o_{p}(1)$ gives (b).
+For (c), write
+$R_{n}=\frac{1}{\sqrt{n}}\sum_{k=1}^{K}\sum_{i\in\mathcal{I}_{k}}\Big(\{\widehat{I}^{(-k)}\}^{-1}-I(\mbox{${\theta}$}_{o})^{-1}\Big)\,\ell^{*}_{\theta_{o}}({\bf W}_{i}).$
+Then
+$\|R_{n}\|\leq\Big(\max_{1\leq k\leq K}\big\|\{\widehat{I}^{(-k)}\}^{-1}-I(\mbox{${\theta}$}_{o})^{-1}\big\|\Big)\,\Big\|\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\ell^{*}_{\theta_{o}}({\bf W}_{i})\Big\|.$
+The first factor is $o_{p}(1)$ by (b).
+The second factor is $O_{p}(1)$ by the multivariate CLT under $E_{0}\|\ell^{*}_{\theta_{o}}({\bf W})\|^{2}<\infty$.
+Therefore $\|R_{n}\|=o_{p}(1)$, proving (c).
+∎
+See 9
+Proof of lemma 9.
+Write
+$\widehat{\phi}_{t}^{(-k)}-\phi^{0}_{t,n}=\Big(\widehat{\phi}^{(-k)}_{t,g}-\phi^{0}_{t,g,n}\Big)+\Big(\widehat{\phi}^{(-k)}_{t,\theta}-\phi^{0}_{t,\theta,n}\Big),$
+and control each term in $L^{2}(P_{0})$. We start with the nuisance Riesz part. For brevity write
+$\widehat{\psi}_{j}^{(-k)}=\widehat{\psi}^{(-k)}_{t}(t_{j},{\bf X};{\bf X}_{0},{\bf Z}_{0})\,,\qquad\psi^{0}_{j}=\psi^{0}_{t}(t_{j},{\bf X};{\bf X}_{0},{\bf Z}_{0}).$
+Also write $\Delta\widehat{M}^{(-k)}_{j}=\Delta N_{j}-\Delta A^{(-k)}_{j}$ and $\Delta M^{0}_{j}=\Delta N_{j}-\Delta A^{0}_{j}$, where
+$\Delta A^{(-k)}_{j}=(t_{j}-t_{j-1})Y_{j}\widehat{h}^{(-k)}(t_{j}\mid{\bf X},{\bf Z})$
+and
+$\Delta A^{0}_{j}=(t_{j}-t_{j-1})Y_{j}h_{o}(t_{j}\mid{\bf X},{\bf Z})\,.$
+Then
+$\widehat{\phi}^{(-k)}_{t,g}-\phi^{0}_{t,g,n}=\sum_{j=1}^{m_{n}}\big(\widehat{\psi}_{j}^{(-k)}-\psi^{0}_{j}\big)\Delta M^{0}_{j}+\sum_{j=1}^{m_{n}}\widehat{\psi}_{j}^{(-k)}\big(\Delta\widehat{M}^{(-k)}_{j}-\Delta M^{0}_{j}\big).$ (B.5)
+The first sum is a martingale transform with predictable integrand $\widehat{\psi}_{j}^{(-k)}-\psi^{0}_{j}$, hence by the discrete-time martingale isometry,
+$\mathbb{E}_{0}\!\left[\left(\sum_{j=1}^{m_{n}}(\widehat{\psi}_{j}^{(-k)}-\psi^{0}_{j})\Delta M^{0}_{j}\right)^{2}\Bigg|\mathcal{F}_{-k}\right]=\mathbb{E}_{0}\!\left[\sum_{j=1}^{m_{n}}(\widehat{\psi}_{j}^{(-k)}-\psi^{0}_{j})^{2}\,\Delta A^{0}_{j}\Bigg|\mathcal{F}_{-k}\right].$
+In the discrete finite-support case, $\psi_{t}^{0}$ has the closed form in (5.12), and $\widehat{\psi}_{t}^{(-k)}$ is obtained by replacing $h_{o}$ and the nuisance quantities appearing in (5.12) by their foldwise plug-in versions.
+Write
+$\pi_{0}=\Pr({\bf X}={\bf X}_{0}),\qquad\widehat{\pi}^{(-k)}=\widehat{\mathbb{P}}^{(-k)}({\bf X}={\bf X}_{0}),$
+and
+$A_{{\bf X}_{0}}(u)=E_{0}\!\left[Y(u)h_{o}(u\mid{\bf X},{\bf Z})\mid{\bf X}={\bf X}_{0}\right],\qquad\widehat{A}^{(-k)}_{{\bf X}_{0}}(u)=\widehat{E}_{-k}\!\left[Y(u)\widehat{h}^{(-k)}(u\mid{\bf X},{\bf Z})\mid{\bf X}={\bf X}_{0}\right].$
+Because $\Pr(X=X_{0})=\pi_{0}>0$ and $X$ has finite support, we have
+$\inf_{k}\widehat{\pi}^{(-k)}\geq c_{\pi}>0$
+with probability tending to one.
+Next, by the bounded-sieve restriction for $\widehat{g}^{(-k)}$, boundedness of $\bf Z$}, and boundedness of the parameter space for $\theta$}, there exists a deterministic constant $c_{\lambda}>0$ such that
+$\inf_{k}\inf_{u\leq t}\inf_{{\bf x},{\bf z}}\widehat{h}^{(-k)}(u\mid{\bf x},{\bf z})\geq c_{\lambda}$
+with probability tending to one. Hence, for $u\leq t$,
+$\widehat{A}^{(-k)}_{{\bf X}_{0}}(u)=\widehat{E}_{-k}\!\left[Y(u)\widehat{h}^{(-k)}(u\mid{\bf X},{\bf Z})\mid{\bf X}={\bf X}_{0}\right]\geq c_{\lambda}\,\widehat{E}_{-k}\!\left[Y(u)\mid{\bf X}={\bf X}_{0}\right].$
+By the positivity assumptions, there exists $c_{Y}>0$ such that
+$\inf_{u\leq t}E_{0}\!\left[Y(u)\mid{\bf X}={\bf X}_{0}\right]\geq c_{Y}.$
+Moreover, since $Y(u)$ is monotone, the class $\{Y(u):0\leq u\leq t\}$ is Glivenko–Cantelli, and therefore
+$\sup_{k}\sup_{u\leq t}\left|\widehat{E}_{-k}\!\left[Y(u)\mid{\bf X}={\bf X}_{0}\right]-E_{0}\!\left[Y(u)\mid{\bf X}={\bf X}_{0}\right]\right|=o_{p}(1).$
+It follows that
+$\inf_{k}\min_{j:t_{j}\leq t}\widehat{A}^{(-k)}_{{\bf X}_{0}}(t_{j})\geq c_{0}$
+with probability tending to one for some constant $c_{0}>0$.
+Because $\Pr({\bf X}={\bf X}_{0})\geq\pi_{0}>0$ and $\bf X$} has finite support, $\widehat{\pi}^{(-k)}\to\pi_{0}$ in probability uniformly in $k$. Define the event
+$\mathcal{E}_{n}=\left\{\inf_{k}\widehat{\pi}^{(-k)}\geq c_{0}\ \text{ and }\ \inf_{k}\min_{j:t_{j}\leq t}\widehat{A}^{(-k)}_{{\bf X}_{0}}(t_{j})\geq c_{0}\right\}.$
+By the above arguments, there exists $c_{0}>0$ such that
+$\mathbb{P}(\mathcal{E}_{n})\to 1.$
+On $\mathcal{E}_{n}$, all denominator terms appearing in $\widehat{\psi}_{t}^{(-k)}$ are uniformly bounded away from zero.
+Now
+$\widehat{\psi}^{(-k)}_{t}(u,{\bf X};{\bf X}_{0},{\bf Z}_{0})=\mathbf{1}(u\leq t)\mathbf{1}({\bf X}={\bf X}_{0})\,\frac{\widehat{h}^{(-k)}(u\mid{\bf X}_{0},{\bf Z}_{0})}{\widehat{\pi}^{(-k)}\,\widehat{A}^{(-k)}_{{\bf X}_{0}}(u)},$
+while
+$\psi^{0}_{t}(u,{\bf X};{\bf X}_{0},{\bf Z}_{0})=\mathbf{1}(u\leq t)\mathbf{1}({\bf X}={\bf X}_{0})\,\frac{h_{o}(u\mid{\bf X}_{0},{\bf Z}_{0})}{\pi_{0}\,A_{{\bf X}_{0}}(u)}.$
+For readability, suppress the common arguments $(u\mid{\bf X}_{0},{\bf Z}_{0})$ and write
+$\Delta_{h,j}^{(-k)}=\widehat{h}^{(-k)}(t_{j}\mid{\bf X}_{0},{\bf Z}_{0})-h_{o}(t_{j}\mid{\bf X}_{0},{\bf Z}_{0}),$
+$\Delta_{\pi}^{(-k)}=\widehat{\pi}^{(-k)}-\pi_{0},\qquad\Delta_{A,j}^{(-k)}=\widehat{A}^{(-k)}_{{\bf X}_{0}}(t_{j})-A_{{\bf X}_{0}}(t_{j}).$
+On $\mathcal{E}_{n}$, a first-order ratio expansion gives, for each $j$,
+$\widehat{\psi}_{j}^{(-k)}-\psi_{j}^{0}=\mathbf{1}(t_{j}\leq t)\mathbf{1}({\bf X}={\bf X}_{0})\Big\{L^{(-k)}_{h,j}+L^{(-k)}_{\pi,j}+L^{(-k)}_{A,j}+R^{(-k)}_{\psi,j}\Big\}$,
+where
+$L^{(-k)}_{h,j}=\frac{\Delta_{h,j}^{(-k)}}{\pi_{0}\,A_{{\bf X}_{0}}(t_{j})},\qquad L^{(-k)}_{\pi,j}=\frac{h_{o}(t_{j}\mid{\bf X}_{0},{\bf Z}_{0})}{A_{{\bf X}_{0}}(t_{j})}\Big(\frac{1}{\widehat{\pi}^{(-k)}}-\frac{1}{\pi_{0}}\Big),$
+$L^{(-k)}_{A,j}=\frac{h_{o}(t_{j}\mid{\bf X}_{0},{\bf Z}_{0})}{\widehat{\pi}^{(-k)}}\Big(\frac{1}{\widehat{A}^{(-k)}_{{\bf X}_{0}}(t_{j})}-\frac{1}{A_{{\bf X}_{0}}(t_{j})}\Big),$
+and $R^{(-k)}_{\psi,j}$ is a finite sum of products of estimation errors, hence second order. Using $(a+b+c+d)^{2}\lesssim a^{2}+b^{2}+c^{2}+d^{2}$, we obtain on $\mathcal{E}_{n}$,
+$(\widehat{\psi}_{j}^{(-k)}-\psi_{j}^{0})^{2}\lesssim\mathbf{1}({\bf X}={\bf X}_{0})\Big\{|L^{(-k)}_{h,j}|^{2}+|L^{(-k)}_{\pi,j}|^{2}+|L^{(-k)}_{A,j}|^{2}+|R^{(-k)}_{\psi,j}|^{2}\Big\}.$
+Therefore
+$\mathbb{E}_{0}\!\left[\sum_{j=1}^{m_{n}}(\widehat{\psi}_{j}^{(-k)}-\psi_{j}^{0})^{2}\,\Delta A^{0}_{j}\right]$ $\lesssim T^{(-k)}_{h}+T^{(-k)}_{\pi}+T^{(-k)}_{A}+T^{(-k)}_{R}$,
+where
+$T^{(-k)}_{h}=\mathbb{E}_{0}\!\left[\sum_{j=1}^{m_{n}}\mathbf{1}({\bf X}={\bf X}_{0})\,|\Delta_{h,j}^{(-k)}|^{2}\,\Delta A^{0}_{j}\right],$
+$T^{(-k)}_{\pi}=|\Delta_{\pi}^{(-k)}|^{2}\,\mathbb{E}_{0}\!\left[\sum_{j=1}^{m_{n}}\mathbf{1}({\bf X}={\bf X}_{0})\,\Delta A^{0}_{j}\right],$
+$T^{(-k)}_{A}=\mathbb{E}_{0}\!\left[\sum_{j=1}^{m_{n}}\mathbf{1}({\bf X}={\bf X}_{0})\,|\Delta_{A,j}^{(-k)}|^{2}\,\Delta A^{0}_{j}\right],$
+and $T_{R}^{(-k)}$ is the analogous contribution of the second-order remainder $R^{(-k)}_{\psi,j}$.
+We bound these four terms separately.
+For $T^{(-k)}_{h}$, note that
+$\Delta A^{0}_{j}=(t_{j}-t_{j-1})Y(t_{j-1})h_{o}(t_{j}\mid{\bf X},{\bf Z}),$
+so
+$T^{(-k)}_{h}=\mathbb{E}_{0}\!\left[\sum_{j=1}^{m_{n}}(t_{j}-t_{j-1})\,\mathbf{1}({\bf X}={\bf X}_{0})\,Y(t_{j-1})h_{o}(t_{j}\mid{\bf X},{\bf Z})\,|\Delta_{h,j}^{(-k)}|^{2}\right].$
+By the same exponential-link expansion used in Lemma 3, $\widehat{h}^{(-k)}-h_{o}$ is linear in $\widetilde{g}^{(-k)}-g_{o}$ and $\widetilde{}\mbox{${\theta}$}^{(-k)}-\mbox{${\theta}$}_{o}$ plus a quadratic remainder.
+Therefore, exactly as in Lemma 3,
+$T^{(-k)}_{h}\lesssim\|\widetilde{g}^{(-k)}-g_{o}\|_{0}^{2}+\|\widetilde{}\mbox{${\theta}$}^{(-k)}-\mbox{${\theta}$}_{o}\|^{2}+o_{p}(1)=O_{p}(\tau_{n}^{2})+O_{p}(n^{-1})+o_{p}(1)=o_{p}(1),$
+uniformly in $k$, by Assumption A9.
+For $T_{A}^{(-k)}$, write
+$\Delta_{A,j}^{(-k)}=B_{A,j}^{(-k)}+D_{A,j}^{(-k)},$
+where
+$B_{A,j}^{(-k)}=\widehat{E}_{-k}\!\left\{Y(t_{j-1})\widehat{h}^{(-k)}(t_{j}\mid{\bf X},{\bf Z})\mid{\bf X}={\bf X}_{0}\right\}-E_{0}\!\left\{Y(t_{j-1})\widehat{h}^{(-k)}(t_{j}\mid{\bf X},{\bf Z})\mid{\bf X}={\bf X}_{0}\right\},$
+and
+$D_{A,j}^{(-k)}=E_{0}\!\left\{Y(t_{j-1})\big(\widehat{h}^{(-k)}(t_{j}\mid{\bf X},{\bf Z})-h_{o}(t_{j}\mid{\bf X},{\bf Z})\big)\mid{\bf X}={\bf X}_{0}\right\}.$
+Hence,
+$T_{A}^{(-k)}\lesssim T_{A,1}^{(-k)}+T_{A,2}^{(-k)},$
+where
+$T_{A,1}^{(-k)}=E_{0}\!\left[\sum_{j=1}^{m_{n}}\mathbf{1}({\bf X}={\bf X}_{0})\,|B_{A,j}^{(-k)}|^{2}\,\Delta A_{j}^{0}\right],$
+and
+$T_{A,2}^{(-k)}=E_{0}\!\left[\sum_{j=1}^{m_{n}}\mathbf{1}({\bf X}={\bf X}_{0})\,|D_{A,j}^{(-k)}|^{2}\,\Delta A_{j}^{0}\right].$
+We first bound $T_{A,1}^{(-k)}$. Because $\bf X$} has finite support and $\widehat{h}^{(-k)}$ is uniformly bounded with probability tending to one, the variables
+$Y(t_{j-1})\widehat{h}^{(-k)}(t_{j}\mid{\bf X},{\bf Z})\mathbf{1}({\bf X}={\bf X}_{0})$
+are uniformly bounded over $j$ and $k$. Therefore, conditional on $\mathcal{F}_{-k}$,
+$E_{0}\!\left[\big(B_{A,j}^{(-k)}\big)^{2}\mid\mathcal{F}_{-k}\right]\lesssim\frac{1}{n}\qquad\text{uniformly in }j,k.$
+Thus
+$E_{0}\!\left[T_{A,1}^{(-k)}\mid\mathcal{F}_{-k}\right]$ $\lesssim\frac{1}{n}\,E_{0}\!\left[\sum_{j=1}^{m_{n}}\mathbf{1}({\bf X}={\bf X}_{0})\,\Delta A_{j}^{0}\,\middle|\,\mathcal{F}_{-k}\right]$
+$\lesssim\frac{1}{n}\,E_{0}\!\left[\sum_{j=1}^{m_{n}}\Delta A_{j}^{0}\right]=O\!\left(\frac{1}{n}\right),$
+since
+$\sum_{j=1}^{m_{n}}\Delta A_{j}^{0}=\sum_{j=1}^{m_{n}}(t_{j}-t_{j-1})Y(t_{j-1})h_{o}(t_{j}\mid{\bf X},{\bf Z})=O_{p}(1).$
+Hence
+$T_{A,1}^{(-k)}=O_{p}(n^{-1})=o_{p}(1)$
+uniformly in $k$.
+Next, for $T_{A,2}^{(-k)}$, Jensen’s inequality gives
+$|D_{A,j}^{(-k)}|^{2}\leq E_{0}\!\left[Y(t_{j-1})\big(\widehat{h}^{(-k)}(t_{j}\mid{\bf X},{\bf Z})-h_{o}(t_{j}\mid{\bf X},{\bf Z})\big)^{2}\mid{\bf X}={\bf X}_{0}\right].$
+Therefore,
+$T_{A,2}^{(-k)}$ $\lesssim E_{0}\!\left[\sum_{j=1}^{m_{n}}\mathbf{1}({\bf X}={\bf X}_{0})\,E_{0}\!\left\{Y(t_{j-1})\big(\widehat{h}^{(-k)}(t_{j}\mid{\bf X},{\bf Z})-h_{o}(t_{j}\mid{\bf X},{\bf Z})\big)^{2}\mid{\bf X}={\bf X}_{0}\right\}\Delta A_{j}^{0}\right].$
+Since $\bf X$} has finite support, this is bounded by a constant multiple of
+$E_{0}\!\left[\sum_{j=1}^{m_{n}}(t_{j}-t_{j-1})\,Y(t_{j-1})h_{o}(t_{j}\mid{\bf X},{\bf Z})\,\big(\widehat{h}^{(-k)}(t_{j}\mid{\bf X},{\bf Z})-h_{o}(t_{j}\mid{\bf X},{\bf Z})\big)^{2}\right],$
+which is exactly the same integrated hazard-error quantity already controlled in the bound for $T_{h}^{(-k)}$.
+Hence, by the exponential-link expansion together with A9 and the same quadratic remainder control as in Lemma 3,
+$T_{A,2}^{(-k)}=o_{p}(1)$
+uniformly in $k$. Combining the two bounds yields
+$T_{A}^{(-k)}=o_{p}(1)\qquad\text{uniformly in }k.$
+Finally, $T_{R}^{(-k)}=o_{p}(1)$ uniformly in $k$, because each term in $R^{(-k)}_{\psi,j}$ is a product of first-order estimation errors and hence contributes only second-order terms after summation with weight $\Delta A_{j}^{0}$.
+Combining the four bounds yields
+$\mathbb{E}_{0}\!\left[\sum_{j=1}^{m_{n}}(\widehat{\psi}_{j}^{(-k)}-\psi_{j}^{0})^{2}\,\Delta A^{0}_{j}\right]=o_{p}(1),\qquad\text{uniformly in }k.$
+Consequently, the first sum is $o_{p}(1)$ in $L^{2}(P_{0})$ uniformly in $k$.
+For the second sum of (B.5), note that
+$\Delta\widehat{M}^{(-k)}_{j}-\Delta M^{0}_{j}=-(\Delta A^{(-k)}_{j}-\Delta A^{0}_{j}),$
+so we set
+$S_{2,k}=\sum_{j=1}^{m_{n}}\widehat{\psi}_{j}^{(-k)}\big(\Delta\widehat{M}_{j}^{(-k)}-\Delta M_{j}^{0}\big)=-\sum_{j=1}^{m_{n}}\widehat{\psi}_{j}^{(-k)}(t_{j}-t_{j-1})Y_{j}\big\{\widehat{h}^{(-k)}(t_{j}\mid{\bf X},{\bf Z})-h_{o}(t_{j}\mid{\bf X},{\bf Z})\big\}.$
+On the event $\mathcal{E}_{n}$, the discrete closed form of $\widehat{\psi}_{j}^{(-k)}$ and boundedness of the denominators imply
+$\sup_{k}\sup_{j:t_{j}\leq t}\big|\widehat{\psi}_{j}^{(-k)}\big|=O_{p}(1).$
+Hence, by Cauchy–Schwarz,
+$|S_{2,k}|^{2}$ $\lesssim\left(\sum_{j=1}^{m_{n}}(t_{j}-t_{j-1})Y_{j}\big|\widehat{h}^{(-k)}(t_{j}\mid{\bf X},{\bf Z})-h_{o}(t_{j}\mid{\bf X},{\bf Z})\big|\right)^{2}$
+$\leq\left(\sum_{j=1}^{m_{n}}\Delta A_{j}^{0}\right)\left(\sum_{j=1}^{m_{n}}(t_{j}-t_{j-1})Y_{j}\frac{\big(\widehat{h}^{(-k)}(t_{j}\mid{\bf X},{\bf Z})-h_{o}(t_{j}\mid{\bf X},{\bf Z})\big)^{2}}{h_{o}(t_{j}\mid{\bf X},{\bf Z})}\right).$
+Since $h_{o}$ is bounded away from zero and infinity, and
+$\sum_{j=1}^{m_{n}}\Delta A_{j}^{0}=\sum_{j=1}^{m_{n}}(t_{j}-t_{j-1})Y_{j}h_{o}(t_{j}\mid{\bf X},{\bf Z})=O_{P_{0}}(1),$
+it follows that
+$\mathbb{E}_{0}\!\big[|S_{2,k}|^{2}\big]\lesssim\mathbb{E}_{0}\!\left[\sum_{j=1}^{m_{n}}(t_{j}-t_{j-1})Y_{j}\big(\widehat{h}^{(-k)}(t_{j}\mid{\bf X},{\bf Z})-h_{o}(t_{j}\mid{\bf X},{\bf Z})\big)^{2}\right].$
+The right-hand side is exactly the same weighted integrated hazard-error quantity already controlled above in the bound for $T_{h}^{(-k)}$.
+Therefore, by the exponential-link expansion, Assumption A9, and the same quadratic remainder control as in Lemma 3,
+$\mathbb{E}_{0}\!\big[|S_{2,k}|^{2}\big]=o_{p}(1)\qquad\text{uniformly in }k.$
+Thus the second sum of (B.5) is $o_{p}(1)$ in $L^{2}(P_{0})$, uniformly in $k$. Therefore,
+$\mathbb{E}_{0}\Big[\big(\widehat{\phi}^{(-k)}_{t,g}-\phi^{0}_{t,g,n}\big)^{2}\Big]=o_{p}(1)\quad\text{uniformly in }k.$
+For the parametric part, we write
+$\widehat{\phi}^{(-k)}_{t,\theta}-\phi^{0}_{t,\theta,n}=\Big(\widehat{c}_{t}^{(-k)}-c_{0,t,n}\Big)^{\top}I(\mbox{${\theta}$}_{o})^{-1}\ell^{*}_{\theta_{o}}+\widehat{c}_{t}^{(-k)\top}\Big(\{\widehat{I}^{(-k)}\}^{-1}-I(\mbox{${\theta}$}_{o})^{-1}\Big)\ell^{*}_{\theta_{o}}+\widehat{c}_{t}^{(-k)\top}\{\widehat{I}^{(-k)}\}^{-1}\Big(\widehat{\ell}^{*,(-k)}_{\theta_{o}}-\ell^{*}_{\theta_{o}}\Big),$
+where all quantities are evaluated at $({\bf X}_{0},{\bf Z}_{0})$ and $\bf W$} (we suppress these arguments). In the discrete case,
+$\widehat{c}_{t}^{(-k)}({\bf X}_{0},{\bf Z}_{0})=\sum_{j:t_{j}\leq t}(t_{j}-t_{j-1})\widehat{h}^{(-k)}(t_{j}\mid{\bf X}_{0},{\bf Z}_{0})\Big\{{\bf Z}_{0}-\widehat{\bf g}^{*,(-k)}(t_{j},{\bf X}_{0})\Big\},$
+so by boundedness and the rate conditions in Assumption A9 $\|\widehat{c}_{t}^{(-k)}-c_{0,t,n}\|=o_{p}(1)$ uniformly in $k$.
+Since $\mathbb{E}_{0}\|\ell^{*}_{\theta_{o}}({\bf W})\|^{2}<\infty$ under the standing integrability assumptions, the first term is $o_{p}(1)$ in $L^{2}(P_{0})$.
+For the second term, Lemma 8 yields
+$\max_{k}\|(\widehat{I}^{(-k)})^{-1}-I(\mbox{${\theta}$}_{o})^{-1}\|=o_{p}(1),$
+while $\widehat{c}_{t}^{(-k)}$ is bounded in probability and $\ell^{*}_{\theta_{o}}\in L^{2}(P_{0})$, so the product is $o_{p}(1)$ in $L^{2}(P_{0})$ uniformly in $k$.
+By the decomposition in the proof of Lemma 3, we have
+$E_{0}\!\left[\big(\widehat{\ell}_{\theta_{o}}^{*,(-k)}-\ell_{\theta_{o}}^{*}\big)^{2}\mid F_{-k}\right]=O_{p}(\tau_{n}^{2}+n^{-1}),$
+uniformly in $k$. Therefore
+$\widehat{\ell}_{\theta_{o}}^{*,(-k)}-\ell_{\theta_{o}}^{*}=O_{p}(\tau_{n})\quad\text{in }L^{2}(\mathbb{P}_{0}),$
+and the third term is $o_{p}(1)$ in $L^{2}(\mathbb{P}_{0})$.
+Putting the three bounds together gives
+$\mathbb{E}_{0}\Big[\big(\widehat{\phi}^{(-k)}_{t,\theta}-\phi^{0}_{t,\theta,n}\big)^{2}\Big]=o_{p}(1)\quad\text{uniformly in }k.$
+Combining the nuisance and parametric parts proves
+$\mathbb{E}_{0}\Big[(\widehat{\phi}^{(-k)}_{t}-\phi^{0}_{t,n})^{2}\Big]=o_{p}(1)\quad\text{uniformly in }k.$
+Finally, $\sqrt{n}\,\delta_{n}\to 0$ implies that the grid approximation error between $\phi^{0}_{t,n}$ and $\phi_{t}^{0}$ is $o(1)$ in $L^{2}(P_{0})$, so the same conclusion holds with $\phi_{t}^{0}$ in place of $\phi^{0}_{t,n}$.
+∎
+See 6
+Proof of Theorem 6.
+We suppress $(t,{\bf X}_{0},{\bf Z}_{0})$ from the notation when no confusion arises.
+Let $\{\mathcal{I}_{k}\}_{k=1}^{K}$ be the fold partition of $\{1,\dots,n\}$ and, for each $k$, let $\mathcal{F}_{-k}$ be the $\sigma$–field generated by the training subsample $\{{\bf W}_{i}:i\in\mathcal{I}_{k}^{c}\}$, the fold split, and $({\bf X}_{0},{\bf Z}_{0})$.
+By construction, $\{{\bf W}_{i}:i\in\mathcal{I}_{k}\}$ is independent of $\mathcal{F}_{-k}$.
+For each fold $k$, write
+$\delta g^{(-k)}(u,{\bf X})=\widetilde{g}^{(-k)}(u,{\bf X})-g_{o}(u,{\bf X}),\qquad\delta\mbox{${\theta}$}^{(-k)}=\widetilde{}\mbox{${\theta}$}^{(-k)}-\mbox{${\theta}$}_{o},$
+and write $\Delta t_{j}=t_{j}-t_{j-1}$. The cross-fitted one-step estimator is
+$\widehat{H}^{\,1,{\sf cf}}=\widehat{H}^{\,{\sf cf}}+\frac{1}{n}\sum_{k=1}^{K}\sum_{i\in\mathcal{I}_{k}}\widehat{\phi}_{t}^{(-k)}({\bf W}_{i}),\qquad\widehat{H}^{\,{\sf cf}}=\sum_{k=1}^{K}\frac{|\mathcal{I}_{k}|}{n}\widehat{H}^{(-k)},$
+where the out-of-fold plug-in estimator uses the Riemann sum
+$\widehat{H}^{(-k)}(t,{\bf X}_{0},{\bf Z}_{0})=\sum_{{j:t_{j}\leq t}}\widehat{h}^{(-k)}(t_{j}\mid{\bf X}_{0},{\bf Z}_{0})\,\Delta t_{j}\,.$
+Introduce also the Riemann approximation of the true cumulative hazard,
+$\widetilde{H}_{o}(t,{\bf X}_{0},{\bf Z}_{0})=\sum_{{j:t_{j}\leq t}}h_{o}(t_{j}\mid{\bf X}_{0},{\bf Z}_{0})\,\Delta t_{j},\qquad H_{o}(t,{\bf X}_{0},{\bf Z}_{0})=\int_{0}^{t}h_{o}(u\mid{\bf X}_{0},{\bf Z}_{0})\,du.$
+Under the time-discretization condition A11, we have
+$\big|\widetilde{H}_{o}-H_{o}\big|\leq C\,\delta_{n},$ (B.6)
+for a finite constant $C$ (possibly random but tight under the joint law), and thus $\sqrt{n}(\widetilde{H}_{o}-H_{o})=o(1)$ whenever $\sqrt{n}\,\delta_{n}\to 0$.
+Decompose
+$\widehat{H}^{\,1,{\sf cf}}-H_{o}=\Big(\widehat{H}^{\,{\sf cf}}-\widetilde{H}_{o}\Big)+\Big(\widetilde{H}_{o}-H_{o}\Big)+\frac{1}{n}\sum_{k=1}^{K}\sum_{i\in\mathcal{I}_{k}}\widehat{\phi}_{t}^{(-k)}({\bf W}_{i}).$
+The second term is negligible by (B.6), so it suffices to analyze
+$A_{n}=\widehat{H}^{\,{\sf cf}}-\widetilde{H}_{o},\qquad B_{n}=\frac{1}{n}\sum_{k=1}^{K}\sum_{i\in\mathcal{I}_{k}}\widehat{\phi}_{t}^{(-k)}({\bf W}_{i}).$
+Write $A_{n}=\sum_{k=1}^{K}\frac{|\mathcal{I}_{k}|}{n}(\widehat{H}^{(-k)}-\widetilde{H}_{o})$. For each $k$,
+$\widehat{H}^{(-k)}-\widetilde{H}_{o}=\sum_{j:t_{j}\leq t}\Big\{\widehat{h}^{(-k)}(t_{j}\mid{\bf X}_{0},{\bf Z}_{0})-h_{o}(t_{j}\mid{\bf X}_{0},{\bf Z}_{0})\Big\}\Delta t_{j}.$
+Define the fold-wise log-perturbation on the grid
+$r_{j}^{(-k)}=\delta g^{(-k)}(t_{j},{\bf X}_{0})+(\delta\mbox{${\theta}$}^{(-k)})^{\top}{\bf Z}_{0},\qquad j=1,\dots,m_{n}.$
+Then
+$\widehat{h}^{(-k)}(t_{j}\mid{\bf X}_{0},{\bf Z}_{0})=h_{o}(t_{j}\mid{\bf X}_{0},{\bf Z}_{0})\exp\{r_{j}^{(-k)}\}.$
+[…]
+$\widehat{H}^{(-k)}-\widetilde{H}_{o}=\sum_{j:t_{j}\leq t}h_{o}(t_{j}\mid{\bf X}_{0},{\bf Z}_{0})\,r_{j}^{(-k)}\,\Delta t_{j}+R_{n,1}^{(-k)},$ (B.7)
+with remainder bounded by
+$\big|R_{n,1}^{(-k)}\big|\leq\frac{1}{2}\sum_{j:t_{j}\leq t}h_{o}(t_{j}\mid{\bf X}_{0},{\bf Z}_{0})\,(r_{j}^{(-k)})^{2}\exp\{|r_{j}^{(-k)}|\}\,\Delta t_{j}.$ (B.8)
+The linear term in (B.7) decomposes into a $g$-part and a $\theta$}-part:
+$\sum_{j:t_{j}\leq t}h_{o}(t_{j}\mid{\bf X}_{0},{\bf Z}_{0})\,r_{j}^{(-k)}\,\Delta t_{j}=L_{t,n}\!\big(\delta g^{(-k)}\big)+\big({\bf Z}_{0}\,\widetilde{H}_{o}\big)^{\top}\delta\mbox{${\theta}$}^{(-k)},$
+where the discrete functional is
+$L_{t,n}(b)=\sum_{j:t_{j}\leq t}h_{o}(t_{j}\mid{\bf X}_{0},{\bf Z}_{0})\,b(t_{j},{\bf X}_{0})\,\Delta t_{j}.$
+Since $\widetilde{H}_{o}=H_{o}+O(\delta_{n})$, we may replace ${\bf Z}_{0}\widetilde{H}_{o}$ by ${\bf Z}_{0}H_{o}$ at cost $O(\|{\bf Z}_{0}\|\,\delta_{n}\,\|\delta\mbox{${\theta}$}^{(-k)}\|)$, which is $o_{p}(n^{-1/2})$ under $\sqrt{n}\,\delta_{n}\to 0$ and $\|\delta\mbox{${\theta}$}^{(-k)}\|=O_{p}(n^{-1/2})$.
+Next decompose the one-step correction into conditional mean plus fluctuation. Define
+$B_{n,0}=\frac{1}{n}\sum_{k=1}^{K}\sum_{i\in\mathcal{I}_{k}}\Big\{\widehat{\phi}_{t}^{(-k)}({\bf W}_{i})-\mathbb{E}_{0}\big[\widehat{\phi}_{t}^{(-k)}({\bf W})\mid\mathcal{F}_{-k}\big]\Big\},\qquad B_{n,1}=\sum_{k=1}^{K}\frac{|\mathcal{I}_{k}|}{n}\mathbb{E}_{0}\big[\widehat{\phi}_{t}^{(-k)}({\bf W})\mid\mathcal{F}_{-k}\big],$
+so that $B_{n}=B_{n,0}+B_{n,1}$.
+In the discrete finite-support setting, the Riesz representer has a closed form and the foldwise estimator $\widehat{\psi}_{t}^{(-k)}$ is $\mathcal{F}_{-k}$-measurable.
+For each fold $k$, conditional on $\mathcal{F}_{-k}$ the estimators $\widetilde{g}^{(-k)}$, $\widetilde{}\mbox{${\theta}$}^{(-k)}$, and $\widehat{\psi}_{t}^{(-k)}$ are fixed functions, and the observations $\{{\bf W}_{i}:i\in\mathcal{I}_{k}\}$ are independent of $\mathcal{F}_{-k}$.
+Since, for each $i$, the process
+$M_{0,i}(t)=N_{i}(t)-\int_{0}^{t}Y_{i}(u)h_{o}(u\mid{\bf X}_{i},{\bf Z}_{i})\,du$
+is a square-integrable martingale with respect to its natural filtration, it follows that for any $\mathcal{F}_{-k}$-measurable predictable process $h(u,{\bf X}_{i})$,
+$\mathbb{E}_{0}\!\left[\int_{0}^{\tau}h(u,{\bf X}_{i})\,dM_{0,i}(u)\ \middle|\ \mathcal{F}_{-k}\right]=0.$
+This identity justifies the foldwise conditional mean calculations below. In particular, the foldwise conditional mean satisfies
+$\mathbb{E}_{0}\big[\widehat{\phi}_{t}^{(-k)}({\bf W})\mid\mathcal{F}_{-k}\big]=-\,L_{t,n}\!\big(\delta g^{(-k)}\big)-\big({\bf Z}_{0}H_{o}(t\mid{\bf X}_{0},{\bf Z}_{0})\big)^{\top}\delta\mbox{${\theta}$}^{(-k)}+R_{n,2}^{(-k)},$ (B.9)
+where $R_{n,2}^{(-k)}$ is second order and satisfies
+$\big|R_{n,2}^{(-k)}\big|\lesssim\|\widetilde{g}^{(-k)}-g_{o}\|_{0}^{2}+\|\widetilde{}\mbox{${\theta}$}^{(-k)}-\mbox{${\theta}$}_{o}\|^{2},$ (B.10)
+under the conditions from Lemma 5 (and the corresponding foldwise analogues), together with the uniform boundedness of $h_{o}$ ensured by Assumptions A1–A4.
+Combine (B.7) with (B.9) and the foldwise expansions and use the replacement $\widetilde{H}_{o}=H_{o}+O(\delta_{n})$ discussed above to obtain
+$A_{n}+B_{n,1}=\sum_{k=1}^{K}\frac{|\mathcal{I}_{k}|}{n}\Big(R_{n,1}^{(-k)}+R_{n,2}^{(-k)}\Big)+o_{p}(n^{-1/2})$
+under the joint law.
+It remains to bound the remainders.
+The bound (B.8) together with $\exp\{|r|\}\leq 1+|r|\exp\{|r|\}$ implies that, on events where $\max_{j\leq m_{n}}|r_{j}^{(-k)}|$ is bounded (which holds with probability tending to one under $\|\delta g^{(-k)}\|_{0}=o_{p}(1)$ and $\|\delta\mbox{${\theta}$}^{(-k)}\|=o_{p}(1)$ plus the uniform boundedness assumptions),
+$|R_{n,1}^{(-k)}|\lesssim\sum_{j:t_{j}\leq t}h_{o}(t_{j}\mid{\bf X}_{0},{\bf Z}_{0})\,(r_{j}^{(-k)})^{2}\,\Delta t_{j}\lesssim\big(L_{t,n}(|\delta g^{(-k)}|)\big)^{2}+\|\delta\mbox{${\theta}$}^{(-k)}\|^{2}\,.$
+We now bound the discrete linear functional $L_{t,n}(b)$. By Cauchy–Schwarz and boundedness of $h_{o}$,
+$|L_{t,n}(b)|\lesssim\left(\sum_{j:t_{j}\leq t}h_{o}(t_{j}\mid{\bf X}_{0},{\bf Z}_{0})b(t_{j},{\bf X}_{0})^{2}\Delta t_{j}\right)^{1/2}.$
+The discrete sum is a Riemann approximation of $\int_{0}^{t}h_{o}(u\mid{\bf X}_{0},{\bf Z}_{0})b(u,{\bf X}_{0})^{2}du$ with error $O(\delta_{n}\|b\|_{\infty}^{2})$.
+Under the boundedness of the sieve and $\sqrt{n}\,\delta_{n}\to 0$, this error is negligible.
+Because $\bf X$} takes values in a finite set and $\mathbb{P}({\bf X}=x_{0})\geq c_{0}>0$, we may write
+$\int_{0}^{t}h_{o}(u\mid{\bf X}_{0},{\bf Z}_{0})b(u,{\bf X}_{0})^{2}du=\frac{1}{\mathbb{P}({\bf X}=x_{0})}\mathbb{E}_{0}\!\left[\int_{0}^{t}h_{o}(u\mid{\bf X},{\bf Z})b(u,{\bf X})^{2}du\,\mathbf{1}\{{\bf X}=x_{0}\}\right].$
+Under the overlap condition $\mathbb{E}_{0}\{Y(u)\mid{\bf X},{\bf Z}\}\geq\kappa>0$ and boundedness of $h_{o}$, the right-hand side is bounded by a constant multiple of
+$\mathbb{E}_{0}\!\left[\int_{0}^{\tau}Y(u)h_{o}(u\mid{\bf X},{\bf Z})b(u,{\bf X})^{2}du\right]=\|b\|_{0}^{2}.$
+Therefore
+$|L_{t,n}(b)|\lesssim\|b\|_{0}+o(n^{-1/2}).$
+Consequently, $|R_{n,1}^{(-k)}|\lesssim\|\delta g^{(-k)}\|_{0}^{2}+\|\delta\mbox{${\theta}$}^{(-k)}\|^{2}$ in probability, uniformly in $k$. Together with (B.10) we obtain
+$|A_{n}+B_{n,1}|\lesssim\max_{k}\|\widetilde{g}^{(-k)}-g_{o}\|_{0}^{2}+\max_{k}\|\widetilde{}\mbox{${\theta}$}^{(-k)}-\mbox{${\theta}$}_{o}\|^{2}+o_{p}(n^{-1/2}).$ (B.11)
+Under the assumption $\max_{k}\|\widetilde{g}^{(-k)}-g_{o}\|_{0}=O_{p}(\gamma_{n}\log^{2}n)$ with $\gamma_{n}\log^{2}n=o(n^{-1/4})$, we have $\sqrt{n}\,\max_{k}\|\widetilde{g}^{(-k)}-g_{o}\|_{0}^{2}=o_{p}(1)$.
+Under $\max_{k}\|\widetilde{}\mbox{${\theta}$}^{(-k)}-\mbox{${\theta}$}_{o}\|=O_{p}(n^{-1/2})$, we also have
+$\sqrt{n}\,\max_{k}\|\widetilde{}\mbox{${\theta}$}^{(-k)}-\mbox{${\theta}$}_{o}\|^{2}=o_{p}(1)\,.$
+Therefore (B.11) implies
+$\sqrt{n}\,(A_{n}+B_{n,1})=o_{p}(1).$
+We now analyze $B_{n,0}$ fold by fold. For each $k$, define
+$B_{n,0}^{(k)}=\frac{1}{|\mathcal{I}_{k}|}\sum_{i\in\mathcal{I}_{k}}\Big\{\widehat{\phi}_{t}^{(-k)}({\bf W}_{i})-\mathbb{E}_{0}\big[\widehat{\phi}_{t}^{(-k)}({\bf W})\mid\mathcal{F}_{-k}\big]\Big\},$
+so that $B_{n,0}=\sum_{k=1}^{K}\frac{|\mathcal{I}_{k}|}{n}\,B_{n,0}^{(k)}$.
+Fix $k$.
+Conditional on $\mathcal{F}_{-k}$, the variables $\{{\bf W}_{i}:i\in\mathcal{I}_{k}\}$ are i.i.d., and $\widehat{\phi}_{t}^{(-k)}$ is fixed.
+Hence the summands in $B_{n,0}^{(k)}$ are conditionally i.i.d.
+and mean zero.
+Now insert and subtract the true EIF.
+For $i\in\mathcal{I}_{k}$, write
+$\Delta_{i,k}=\widehat{\phi}_{t}^{(-k)}({\bf W}_{i};{\bf X}_{0},{\bf Z}_{0})-\phi_{t}^{0}({\bf W}_{i};{\bf X}_{0},{\bf Z}_{0}),\qquad\Delta_{k}({\bf W})=\widehat{\phi}_{t}^{(-k)}({\bf W};{\bf X}_{0},{\bf Z}_{0})-\phi_{t}^{0}({\bf W};{\bf X}_{0},{\bf Z}_{0}).$
+Then
+$B_{n,0}^{(k)}$ $=\frac{1}{|\mathcal{I}_{k}|}\sum_{i\in\mathcal{I}_{k}}\Big\{\phi_{t}^{0}({\bf W}_{i};{\bf X}_{0},{\bf Z}_{0})-\mathbb{E}_{0}[\phi_{t}^{0}({\bf W};{\bf X}_{0},{\bf Z}_{0})\mid\mathcal{F}_{-k}]\Big\}$
+$\qquad+\frac{1}{|\mathcal{I}_{k}|}\sum_{i\in\mathcal{I}_{k}}\Big\{\Delta_{i,k}-\mathbb{E}_{0}[\Delta_{k}({\bf W})\mid\mathcal{F}_{-k}]\Big\}.$
+Because $\bf W$} is independent of the training subsample generating $\mathcal{F}_{-k}$, and $({\bf X}_{0},{\bf Z}_{0})\in\mathcal{F}_{-k}$, we have
+$\mathbb{E}_{0}[\phi_{t}^{0}({\bf W};{\bf X}_{0},{\bf Z}_{0})\mid\mathcal{F}_{-k}]=\mathbb{E}_{0}[\phi_{t}^{0}({\bf W};{\bf X}_{0},{\bf Z}_{0})\mid{\bf X}_{0},{\bf Z}_{0}].$
+Since $\phi_{t}^{0}$ is the EIF,
+$\mathbb{E}_{0}[\phi_{t}^{0}({\bf W};{\bf X}_{0},{\bf Z}_{0})\mid{\bf X}_{0},{\bf Z}_{0}]=0\,$,
+and
+$\mathbb{E}_{0}[\phi_{t}^{0}({\bf W};{\bf X}_{0},{\bf Z}_{0})\mid\mathcal{F}_{-k}]=0.$
+Therefore
+$B_{n,0}^{(k)}=\frac{1}{|\mathcal{I}_{k}|}\sum_{i\in\mathcal{I}_{k}}\phi_{t}^{0}({\bf W}_{i};{\bf X}_{0},{\bf Z}_{0})+R_{n,3}^{(k)},$
+where
+$R_{n,3}^{(k)}=\frac{1}{|\mathcal{I}_{k}|}\sum_{i\in\mathcal{I}_{k}}\Big\{\Delta_{i,k}-\mathbb{E}_{0}[\Delta_{k}({\bf W})\mid\mathcal{F}_{-k}]\Big\}.$
+We next show that the remainder is negligible. Conditional on $\mathcal{F}_{-k}$, the summands in $R_{n,3}^{(k)}$ are centered and independent, so
+$\mathbb{E}_{0}\!\left[\left(R_{n,3}^{(k)}\right)^{2}\middle|\mathcal{F}_{-k}\right]\leq\frac{1}{|\mathcal{I}_{k}|}\mathbb{E}_{0}\!\left[\Delta_{k}({\bf W})^{2}\middle|\mathcal{F}_{-k}\right].$
+Based on Lemma 9,
+$\max_{1\leq k\leq K}\mathbb{E}_{0}\!\left[\Delta_{k}({\bf W})^{2}\right]=o_{p}(1),$
+and therefore, by taking expectations,
+$\mathbb{E}_{0}\!\left[\left(R_{n,3}^{(k)}\right)^{2}\right]\leq\frac{1}{|\mathcal{I}_{k}|}\mathbb{E}_{0}\!\left[\Delta_{k}({\bf W})^{2}\right]=o\!\left(\frac{1}{n}\right),$
+uniformly in $k$, because $|\mathcal{I}_{k}|\asymp n$ and $K$ is fixed. Hence
+$\sum_{k=1}^{K}\frac{|\mathcal{I}_{k}|}{n}R_{n,3}^{(k)}=o_{p}(n^{-1/2}).$
+Consequently,
+$B_{n,0}=\sum_{k=1}^{K}\frac{|\mathcal{I}_{k}|}{n}\left\{\frac{1}{|\mathcal{I}_{k}|}\sum_{i\in\mathcal{I}_{k}}\phi_{t}^{0}({\bf W}_{i};{\bf X}_{0},{\bf Z}_{0})+R_{n,3}^{(k)}\right\}=\frac{1}{n}\sum_{i=1}^{n}\phi_{t}^{0}({\bf W}_{i};{\bf X}_{0},{\bf Z}_{0})+o_{p}(n^{-1/2}),$
+and hence
+$\sqrt{n}\,B_{n,0}=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\phi_{t}^{0}({\bf W}_{i};{\bf X}_{0},{\bf Z}_{0})+o_{p}(1).$
+Combining this with $\sqrt{n}(A_{n}+B_{n,1})=o_{p}(1)$ and $\sqrt{n}(\widetilde{H}_{o}-H_{o})=o(1)$ yields
+$\sqrt{n}\big(\widehat{H}^{\,1,{\sf cf}}-H_{o}\big)=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\phi_{t}^{0}({\bf W}_{i};{\bf X}_{0},{\bf Z}_{0})+o_{p}(1).$
+Conditional on $({\bf X}_{0},{\bf Z}_{0})$, the summands $\phi_{t}^{0}({\bf W}_{i};{\bf X}_{0},{\bf Z}_{0})$ are i.i.d.
+with mean 0 and variance $\sigma_{H}^{2}(t;{\bf X}_{0},{\bf Z}_{0})={\mathbb{V}}(\phi_{t}^{0}({\bf W};{\bf X}_{0},{\bf Z}_{0})\mid{\bf X}_{0},{\bf Z}_{0})$, and have finite second moment by Lemma 5.
+The (conditional) Lindeberg–Feller CLT therefore gives
+$\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\phi_{t}^{0}({\bf W}_{i};{\bf X}_{0},{\bf Z}_{0})\Rightarrow\mathcal{N}\!\big(0,\sigma_{H}^{2}(t;{\bf X}_{0},{\bf Z}_{0})\big),$
+proving conditional asymptotic normality of $\widehat{H}^{\,1,{\sf cf}}(t\mid{\bf X}_{0},{\bf Z}_{0})$ given $({\bf X}_{0},{\bf Z}_{0})$.
+For $\widehat{S}^{\,1,{\sf cf}}=\exp\{-\widehat{H}^{\,1,{\sf cf}}\}$, the delta method applies since $x\mapsto e^{-x}$ is continuously differentiable, giving
+$\sqrt{n}\big(\widehat{S}^{\,1,{\sf cf}}-S_{o}\big)=-\,S_{o}\,\sqrt{n}\big(\widehat{H}^{\,1,{\sf cf}}-H_{o}\big)+o_{p}(1),$
+and thus
+$\sqrt{n}\big(\widehat{S}^{\,1,{\sf cf}}-S_{o}\big)\rightarrow^{\mathcal{D}}\mathcal{N}\!\big(0,S_{o}(t\mid{\bf X}_{0},{\bf Z}_{0})^{2}\,\sigma_{H}^{2}(t;{\bf X}_{0},{\bf Z}_{0})\big).$
+Finally, consider the variance estimator
+$\widehat{\sigma}_{H}^{2}(t;{\bf X}_{0},{\bf Z}_{0})=\frac{1}{n}\sum_{k=1}^{K}\sum_{i\in\mathcal{I}_{k}}\Big\{\widehat{\phi}_{t}^{(-k)}({\bf W}_{i};{\bf X}_{0},{\bf Z}_{0})\Big\}^{2}.$
+For each $i$, let $k(i)$ denote the fold index such that $i\in\mathcal{I}_{k(i)}$, and write
+$\widehat{\phi}_{i}=\widehat{\phi}_{t}^{(-k(i))}({\bf W}_{i};{\bf X}_{0},{\bf Z}_{0}),\qquad\phi_{i}=\phi_{t}^{0}({\bf W}_{i};{\bf X}_{0},{\bf Z}_{0}),\qquad\mathbb{P}_{n}f=\frac{1}{n}\sum_{i=1}^{n}f({\bf W}_{i}).$
+Then
+$\widehat{\sigma}_{H}^{2}(t;{\bf X}_{0},{\bf Z}_{0})-\sigma_{H}^{2}(t;{\bf X}_{0},{\bf Z}_{0})=\big(\mathbb{P}_{n}-\mathbb{P}_{0}\big)\phi^{2}+\mathbb{P}_{n}\big(\widehat{\phi}^{2}-\phi^{2}\big).$
+Conditionally on $({\bf X}_{0},{\bf Z}_{0})$, $\{\phi_{i}^{2}\}_{i=1}^{n}$ are i.i.d.
+with $\mathbb{E}_{0}(\phi^{2}\mid{\bf X}_{0},{\bf Z}_{0})<\infty$, hence the law of large numbers yields $\big(\mathbb{P}_{n}-\mathbb{P}_{0}\big)\phi^{2}=o_{p}(1)$.
+For the second term, use $\widehat{\phi}^{2}-\phi^{2}=(\widehat{\phi}-\phi)(\widehat{\phi}+\phi)$ and Cauchy–Schwarz, and get
+$\Big|\mathbb{P}_{n}(\widehat{\phi}^{2}-\phi^{2})\Big|\leq\Big\{\mathbb{P}_{n}(\widehat{\phi}-\phi)^{2}\Big\}^{1/2}\Big\{\mathbb{P}_{n}(\widehat{\phi}+\phi)^{2}\Big\}^{1/2}.$
+By the $L^{2}(P_{0})$-consistency established above, $\mathbb{E}_{0}[(\widehat{\phi}-\phi)^{2}]\to 0$ (uniformly in folds), which implies $\mathbb{P}_{n}(\widehat{\phi}-\phi)^{2}=o_{p}(1)$. Moreover,
+$\mathbb{P}_{n}(\widehat{\phi}+\phi)^{2}\leq 2\mathbb{P}_{n}\widehat{\phi}^{2}+2\mathbb{P}_{n}\phi^{2}=O_{p}(1),$
+since $\sup_{n}\mathbb{E}_{0}(\widehat{\phi}^{2})<\infty$ and $\mathbb{E}_{0}(\phi^{2})<\infty$. Therefore $\mathbb{P}_{n}(\widehat{\phi}^{2}-\phi^{2})=o_{p}(1)$, and hence
+$\widehat{\sigma}_{H}^{2}(t;{\bf X}_{0},{\bf Z}_{0})\xrightarrow{p}\sigma_{H}^{2}(t;{\bf X}_{0},{\bf Z}_{0}).$
+Finally, by the delta method,
+$\sqrt{n}\Big\{\widehat{S}^{\,1,{\sf cf}}(t\mid{\bf X}_{0},{\bf Z}_{0})-S_{o}(t\mid{\bf X}_{0},{\bf Z}_{0})\Big\}\Rightarrow\mathcal{N}\!\left(0,\ \sigma_{S}^{2}(t;{\bf X}_{0},{\bf Z}_{0})\right)\,$,
+and consistency of $\widehat{\sigma}_{S}^{2}(t;{\bf X}_{0},{\bf Z}_{0})=\{\widehat{S}^{\,1,{\sf cf}}(t\mid{\bf X}_{0},{\bf Z}_{0})\}^{2}\widehat{\sigma}_{H}^{2}(t;{\bf X}_{0},{\bf Z}_{0})$ follows by Slutsky’s theorem together with $\widehat{S}^{\,1,{\sf cf}}(t\mid{\bf X}_{0},{\bf Z}_{0})\xrightarrow{p}S_{o}(t\mid{\bf X}_{0},{\bf Z}_{0})$.
+∎
+Appendix C General covariate distributions of X: sieve approximation of the Riesz representer
+When $\bf X$} has a general distribution, no closed-form expression is available and $\psi^{0}_{t}(\cdot,\cdot;{\bf X}_{0},{\bf Z}_{0})$ must be approximated in a finite-dimensional sieve.
+In this section we construct a spline-based approximation $\widehat{\psi}^{(-k)}_{t}(\cdot,\cdot;{\bf X}_{0},{\bf Z}_{0})$ and define a cross-fitted one-step estimator based on the same Riemann grid $\{t_{j}\}_{j=0}^{m_{n}}$ used throughout (with $\Delta t_{j}=t_{j}-t_{j-1}$), and we state the corresponding $\sqrt{n}$ asymptotic normality under additional sieve conditions.
+Fix $t\in[0,\tau]$ and a random evaluation point $({\bf X}_{0},{\bf Z}_{0})$ drawn independently from the target covariate distribution.
+By conditions, $L_{t}$ is bounded on $(\mathcal{H},\|\cdot\|_{0})$, $\mathcal{H}=\overline{\mathcal{T}}_{g_{o}}$, and there exists a unique $\psi^{0}_{t}(\cdot,\cdot;{\bf X}_{0},{\bf Z}_{0})\in\mathcal{H}$ solving the Riesz equation
+$\langle\psi^{0}_{t}(\cdot,\cdot;{\bf X}_{0},{\bf Z}_{0}),\,b\rangle_{0}=L_{t}(b),\qquad\forall\,b\in\mathcal{H}.$
+Let $\{\mu_{r}(u,{\bf X}):r=1,\ldots,d_{n}\}$ be a spline basis on $0,\tau]\times\mathcal{X}$ (e.g., tensor-product $B$-splines in $u$ and each coordinate of $\bf X$} after rescaling $\bf X$} to $0,1]^{r}$), and define the sieve space
+$\mathcal{H}_{d_{n}}=\mathrm{span}\{\mu_{1},\ldots,\mu_{d_{n}}\}\subset\mathcal{H}.$
+Define the population sieve-Riesz projection $\psi^{0,(d_{n})}_{t}(\cdot,\cdot;{\bf X}_{0},{\bf Z}_{0})\in\mathcal{H}_{d_{n}}$ as the unique solution of
+$\langle\psi^{0,(d_{n})}_{t}(\cdot,\cdot;{\bf X}_{0},{\bf Z}_{0}),\,b\rangle_{0}=L_{t}(b),\qquad\forall\,b\in\mathcal{H}_{d_{n}}.$
+The key additional requirement for $\sqrt{n}$ inference is that the sieve approximation bias is negligible, namely,
+$\big\|\psi^{0,(d_{n})}_{t}(\cdot,\cdot;{\bf X}_{0},{\bf Z}_{0})-\psi^{0}_{t}(\cdot,\cdot;{\bf X}_{0},{\bf Z}_{0})\big\|_{0}=o(n^{-1/2})\,.$ (C.1)
+It might be tempting to assume that for each $t\in[0,\tau]$, the Riesz representer $\psi^{0}_{t}(\cdot,\cdot;{\bf X}_{0},{\bf Z}_{0})$ (viewed as a function of $(u,{\bf X})\in[0,\tau]\times\mathcal{X}$) belongs to a Sobolev (or Hölder) ball of smoothness $s$, uniformly over $t\in[0,\tau]$ (and uniformly over $({\bf X}_{0},{\bf Z}_{0})$ in its support), so that the spline sieve $\mathcal{H}_{d_{n}}$ satisfies the approximation property (C.1).
+Clearly, this condition is separate from the composite Hölder assumption on $g_{o}$, which is used to guarantee $\|\widehat{g}-g_{o}\|_{0}=O_{p}(\gamma_{n}\log^{2}n)$.
+In the absence of additional structural restrictions, approximation rates in the norm $\|\cdot\|_{0}$ deteriorate rapidly with the dimension of $\bf X$} (curse of dimensionality), so that achieving $o(n^{-1/2})$ accuracy typically requires sieve dimension $d_{n}$ that grows so quickly with $n$ that the required $n$ for a given accuracy becomes impractically large.
+Section 1.3 provides two alternative assumption sets that ensure (C.1).
+Throughout the remainder of this section we assume that (C.1) holds.
+Now, we provide an estimator of $\psi^{0,(d_{n})}_{t}(\cdot,\cdot;{\bf X}_{0},{\bf Z}_{0})$ under the cross-fitted setting.
+For each fold $k$ and functions $f,h\in\mathcal{H}$, define the fold-wise empirical inner product (computed on $\mathcal{I}_{k}^{c}$) by the Riemann sum
+$\langle f,h\rangle_{n,-k}=\frac{1}{|\mathcal{I}_{k}^{c}|}\sum_{i\in\mathcal{I}_{k}^{c}}\sum_{j=1}^{m_{n}}f(t_{j},{\bf X}_{i})\,h(t_{j},{\bf X}_{i})\,Y_{i}(t_{j-1})\,\widehat{h}^{(-k)}(t_{j}\mid{\bf X}_{i},{\bf Z}_{i})\,\Delta t_{j}.$ (C.2)
+Similarly, define the fold-wise empirical linear functional by
+$\widehat{L}^{(-k)}_{t}(b)=\sum_{j:\,t_{j}\leq t}\widehat{h}^{(-k)}(t_{j}\mid{\bf X}_{0},{\bf Z}_{0})\,b(t_{j},{\bf X}_{0})\,\Delta t_{j}.$ (C.3)
+We estimate the Riesz representer in the sieve space $\mathcal{H}_{d_{n}}$ by solving the ridge-stabilized empirical Riesz equation.
+Namely, we find $\widehat{\psi}^{(-k)}_{t}(\cdot,\cdot;{\bf X}_{0},{\bf Z}_{0})\in\mathcal{H}_{d_{n}}$ such that for all $b\in\mathcal{H}_{d_{n}}$,
+$\langle\widehat{\psi}^{(-k)}_{t}(\cdot,\cdot;{\bf X}_{0},{\bf Z}_{0}),\,b\rangle_{n,-k}+\rho_{n}\,\langle\widehat{\psi}^{(-k)}_{t}(\cdot,\cdot;{\bf X}_{0},{\bf Z}_{0}),\,b\rangle_{\mathrm{coef}}=\widehat{L}^{(-k)}_{t}(b),$ (C.4)
+where $\rho_{n}\downarrow 0$, $\sqrt{n}\rho_{n}\to 0$, and $\langle\cdot,\cdot\rangle_{\mathrm{coef}}$ is the Euclidean inner product of spline coefficients.
+Here $\rho_{n}>0$ is a vanishing ridge parameter used only to stabilize the inversion of the empirical Gram matrix.
+Specifically, write $\widehat{\psi}^{(-k)}_{t}(u,{\bf X};{\bf X}_{0},{\bf Z}_{0})=\sum_{r=1}^{d_{n}}\widehat{\alpha}^{(-k)}_{r}\,\mu_{r}(u,{\bf X})$ and $b(u,{\bf X})=\sum_{s=1}^{d_{n}}\beta_{s}\,\mu_{s}(u,{\bf X})$.
+Let $\widehat{G}_{-k}\in\mathbb{R}^{d_{n}\times d_{n}}$ and $\widehat{\ell}_{-k}\in\mathbb{R}^{d_{n}}$ be defined such that their components are given by
+$(\widehat{G}_{-k})_{rs}=\langle\mu_{r},\mu_{s}\rangle_{n,-k},\qquad r,s=1,\ldots,d_{n}.$
+$(\widehat{\ell}_{-k})_{s}=\widehat{L}_{t}^{(-k)}(\mu_{s})=\sum_{j:\,t_{j}\leq t}\widehat{h}^{(-k)}(t_{j}\mid{\bf X}_{0},{\bf Z}_{0})\,\mu_{s}(t_{j},{\bf X}_{0})\,\Delta t_{j},\qquad s=1,\ldots,d_{n}.$
+Let $\widehat{\alpha}^{(-k)}=\big(\widehat{\alpha}^{(-k)}_{1},\ldots,\widehat{\alpha}^{(-k)}_{d_{n}}\big)^{\top}\in\mathbb{R}^{d_{n}}$. Then (C.4) is equivalent to
+$(\widehat{G}_{-k}+\rho_{n}I_{d_{n}})\,\widehat{\alpha}^{(-k)}=\widehat{\ell}_{-k}\,$,
+so that $\widehat{\alpha}^{(-k)}=(\widehat{G}_{-k}+\rho_{n}I_{d_{n}})^{-1}\widehat{\ell}_{-k}$, where $I_{d_{n}}$ denotes the $d_{n}\times d_{n}$ identity matrix.
+Finally, we are ready to define the cross-fitted one-step estimator based on the above spline approximation.
+Using the fold-wise spline weight $\widehat{\psi}^{(-k)}_{t}(\cdot,\cdot;{\bf X}_{0},{\bf Z}_{0})$, define the fold-wise nuisance influence contribution in discrete (Riemann-sum) form
+$\widehat{\phi}^{(-k)}_{t,g}(W_{i};{\bf X}_{0},{\bf Z}_{0})=\sum_{j=1}^{m_{n}}\widehat{\psi}^{(-k)}_{t}(t_{j},{\bf X}_{i};{\bf X}_{0},{\bf Z}_{0})\,\Delta\widehat{M}^{(-k)}_{i}(t_{j}).$
+Let $\widehat{\phi}^{(-k)}_{t,\theta}(W_{i};{\bf X}_{0},{\bf Z}_{0})$ denote the parametric EIF piece, constructed exactly as in the discrete-support section.
+Set $\widehat{\phi}^{(-k)}_{t}=\widehat{\phi}^{(-k)}_{t,g}+\widehat{\phi}^{(-k)}_{t,\theta}$.
+The cross-fitted one-step estimator is then
+$\widehat{H}^{1,\mathrm{cf}}(t\mid{\bf X}_{0},{\bf Z}_{0})=\widehat{H}^{\mathrm{cf}}(t\mid{\bf X}_{0},{\bf Z}_{0})+\frac{1}{n}\sum_{k=1}^{K}\sum_{i\in\mathcal{I}_{k}}\widehat{\phi}^{(-k)}_{t}(W_{i};{\bf X}_{0},{\bf Z}_{0}),$
+and $\widehat{S}^{1,\mathrm{cf}}(t\mid{\bf X}_{0},{\bf Z}_{0})=\exp\{-\widehat{H}^{1,\mathrm{cf}}(t\mid{\bf X}_{0},{\bf Z}_{0})\}$.
+Relative to the discrete case, the general-$\bf X$} setting introduces three additional sources of error that must be controlled to obtain a $\sqrt{n}$ expansion:
+  1. 1.
+    Sieve approximation (bias): the population sieve-Riesz representer $\psi^{0,(d_{n})}_{t}\in\mathcal{H}_{d_{n}}$ must approximate $\psi_{t}^{0}$ in the $\|\cdot\|_{0}$ norm, i.e.
+$\|\psi^{0,(d_{n})}_{t}-\psi_{t}^{0}\|_{0}$ (cf.
+(C.1)).
+  2. 2.
+    Estimation of the sieve-Riesz weights: the data-driven solution $\widehat{\psi}_{t}$ must be close to $\psi^{0,(d_{n})}_{t}$, i.e. $\|\widehat{\psi}_{t}-\psi^{0,(d_{n})}_{t}\|_{0}=o_{p}(1)$.
+  3. 3.
+    Stability: the (regularized) Gram matrix defining $\widehat{\psi}_{t}$ must be well-conditioned, and the effective sieve complexity must grow slowly enough with $n$ so that the associated empirical-process remainder terms are $o_{p}(n^{-1/2})$.
+The general-$\bf X$} setting is substantially more delicate than the finite-support case.
+In particular, a fully rigorous analysis of the sieve-Riesz step would require additional assumptions on the covariate distribution and on the regularity and approximability of the population Riesz representer, together with quantitative control of the sieve approximation bias, the ridge-stabilized Gram system, and the empirical-process terms arising from estimation of the Riesz coefficients.
+A complete investigation of these issues for the specific sieve-Riesz implementation considered here would take us beyond the intended scope of the present paper.
+Accordingly, the following theorem is best viewed as a high-level conditional asymptotic result for the general-$\bf X$} setting.
+Even in this form, however, the result remains valuable, as it helps clarify the ingredients that would be needed to extend the one-step asymptotic theory beyond the finite-support case.
+From a practical point of view, Theorem 7 is already highly relevant, since when $\bf X$} is handled through the DNN nuisance-estimation step, discretization or grouping of covariates is often not expected to materially compromise prediction accuracy.
+Theorem 7.
+Fix $t\in[0,\tau]$ and let $({\bf X}_{0},{\bf Z}_{0})$ be an independent draw from the covariate distribution, independent of the training sample.
+Assume Assumptions A1–A4 hold, so that the hazard is uniformly bounded.
+Let $\{t_{j}\}_{j=0}^{m_{n}}$ be a deterministic grid with $0=t_{0}<\cdots<t_{m_{n}}=t$ and mesh $\delta_{n}=\max_{j}(t_{j}-t_{j-1})$.
+Assume $\sqrt{n}\,\delta_{n}\to 0$ and the time-grid Riemann approximation error is $O(\delta_{n})$ uniformly over a neighborhood of $\eta_{o}=(\mbox{${\theta}$}_{o},g_{o})$.
+Suppose the conditions of Lemma 5 hold and $\mathbb{E}_{0}\{\phi_{t}^{0}({\bf W};{\bf X}_{0},{\bf Z}_{0})^{2}\}<\infty$.
+Suppose moreover that for each fold $k$,
+  1. (i)
+    $\|\widetilde{g}^{(-k)}-g_{o}\|_{0}=o_{p}(n^{-1/4})$ (up to polylogarithmic factors) and, uniformly in $k$,
+    $\sqrt{|\mathcal{I}_{k}^{c}|}\,(\widetilde{}\mbox{${\theta}$}^{(-k)}-\mbox{${\theta}$}_{o})=I(\mbox{${\theta}$}_{o})^{-1}\frac{1}{\sqrt{|\mathcal{I}_{k}^{c}|}}\sum_{i\in\mathcal{I}_{k}^{c}}\ell_{\theta_{o}}^{*}({\bf W}_{i})+o_{p}(1);$
+  2. (ii)
+    the sieve bias condition (C.1) holds;
+  3. (iii)
+    $\|\widehat{\psi}^{(-k)}_{t}(\cdot,\cdot;{\bf X}_{0},{\bf Z}_{0})-\psi^{0,(d_{n})}_{t}(\cdot,\cdot;{\bf X}_{0},{\bf Z}_{0})\|_{0}=o_{p}(1)$;
+  4. (iv)
+    the ridge-stabilized Gram matrix $\widehat{G}_{-k}+\rho_{n}I_{d_{n}}$ is well-conditioned and $d_{n}$ grows slowly enough so that the additional empirical-process terms due to estimating $\widehat{\psi}$ are $o_{p}(n^{-1/2})$.
+Then, conditional on the independent draw $({\bf X}_{0},{\bf Z}_{0})$,
+$\sqrt{n}\Big\{\widehat{H}^{1,\mathrm{cf}}(t\mid{\bf X}_{0},{\bf Z}_{0})-H_{0}(t\mid{\bf X}_{0},{\bf Z}_{0})\Big\}\mid({\bf X}_{0},{\bf Z}_{0})\Rightarrow N\!\Big(0,\sigma_{H}^{2}(t;{\bf X}_{0},{\bf Z}_{0})\Big),$
+where $\sigma_{H}^{2}(t;{\bf X}_{0},{\bf Z}_{0})={\mathbb{V}}_{0}\!\big(\phi_{t}^{0}(W;{\bf X}_{0},{\bf Z}_{0})\mid{\bf X}_{0},{\bf Z}_{0}\big)$. Consequently,
+$\sqrt{n}\Big\{\widehat{S}^{1,\mathrm{cf}}(t\mid{\bf X}_{0},{\bf Z}_{0})-S_{0}(t\mid{\bf X}_{0},{\bf Z}_{0})\Big\}\mid({\bf X}_{0},{\bf Z}_{0})\Rightarrow N\!\Big(0,\sigma_{S}^{2}(t;{\bf X}_{0},{\bf Z}_{0})\Big),$
+and $\sigma_{S}^{2}(t;{\bf X}_{0},{\bf Z}_{0})=S_{0}(t\mid{\bf X}_{0},{\bf Z}_{0})^{2}\,\sigma_{H}^{2}(t;{\bf X}_{0},{\bf Z}_{0})$.
+Theorem 7 should be read as a conditional asymptotic-distribution result under assumptions (i)–(iv), where assumption (iv) collects the high-level control needed for the sieve-Riesz estimation step, including the associated empirical-process and projection-approximation errors.
+Proof of Theorem 7.
+We suppress $(t,{\bf X}_{0},{\bf Z}_{0})$ from the notation when no confusion arises.
+Let $\{\mathcal{I}_{k}\}_{k=1}^{K}$ be the fixed KK–fold partition and let $\mathcal{F}_{-k}$ denote the $\sigma$–field generated by the training subsample $\{{\bf W}_{j}:j\in\mathcal{I}_{k}^{c}\}$, the fold split, and $({\bf X}_{0},{\bf Z}_{0})$.
+By construction, conditional on $\mathcal{F}_{-k}$, the observations $\{{\bf W}_{i}:i\in\mathcal{I}_{k}\}$ are i.i.d.
+from $\mathbb{P}_{0}$.
+Throughout this proof, $L_{t,n}$ and $c_{0,t,n}$ denote the grid-based analogues of $L_{t}$ and $c_{0,t}$, respectively.
+Under $\sqrt{n}\,\delta_{n}\to 0$, the discrepancy between the continuous and grid-based versions is $o(n^{-1/2})$ and is absorbed into the remainder notation.
+Let $\widetilde{H}_{0}$ be the Riemann approximation of $H_{0}$ on the grid, so that $\|\widetilde{H}_{0}-H_{0}\|\lesssim\delta_{n}$ and hence $\sqrt{n}(\widetilde{H}_{0}-H_{0})=o(1)$ whenever $\sqrt{n}\,\delta_{n}\to 0$.
+Write
+$\widehat{H}^{1,{\sf cf}}-H_{0}=\underbrace{\big(\widehat{H}^{{\sf cf}}-\widetilde{H}_{0}\big)}_{A_{n}}+\underbrace{\big(\widetilde{H}_{0}-H_{0}\big)}_{o(n^{-1/2})}+\underbrace{\mathbb{P}_{n}\widehat{\phi}_{t}}_{B_{n}},$
+where $\widehat{H}^{{\sf cf}}=\sum_{k=1}^{K}\frac{|\mathcal{I}_{k}|}{n}\widehat{H}^{(-k)}$ and $\mathbb{P}_{n}\widehat{\phi}_{t}=n^{-1}\sum_{k=1}^{K}\sum_{i\in\mathcal{I}_{k}}\widehat{\phi}_{t}^{(-k)}({\bf W}_{i})$.
+As in Theorem 6, a first-order exponential expansion yields
+$A_{n}=\sum_{k=1}^{K}\frac{|\mathcal{I}_{k}|}{n}\sum_{j:t_{j}\leq t}\Delta t_{j}\,h_{o}(t_{j}\mid{\bf X}_{0},{\bf Z}_{0})\Big\{\delta g^{(-k)}(t_{j},{\bf X}_{0})+(\delta\mbox{${\theta}$}^{(-k)})^{\top}{\bf Z}_{0}\Big\}+R_{n,1},$
+with $\sqrt{n}\,R_{n,1}=o_{p}(1)$ under assumption (i), the same second-order remainder control as in Theorem 6.
+Let $\phi_{t}^{0}({\bf W})$ denote the population EIF from Lemma 5.
+Since $\phi_{t}^{0}$ is an influence function, $\mathbb{P}_{0}\!\big(\phi_{t}^{0}({\bf W})\mid{\bf X}_{0},{\bf Z}_{0}\big)=0$, hence $\mathbb{P}_{n}\phi_{t}^{0}=(\mathbb{P}_{n}-\mathbb{P}_{0})\phi_{t}^{0}$.
+Insert and subtract $\mathbb{P}_{n}\phi_{t}^{0}$ to obtain
+$\sqrt{n}\big(\widehat{H}^{1,{\sf cf}}-H_{0}\big)=\underbrace{\sqrt{n}\,(\mathbb{P}_{n}-\mathbb{P}_{0})\phi_{t}^{0}}_{\text{CLT term}}+\sqrt{n}\Big\{A_{n}+\mathbb{P}_{n}\big(\widehat{\phi}_{t}-\phi_{t}^{0}\big)\Big\}+o_{p}(1).$
+By conditional Lindeberg–Feller (given $({\bf X}_{0},{\bf Z}_{0})$) and $\mathbb{E}_{0}[(\phi_{t}^{0})^{2}\mid{\bf X}_{0},{\bf Z}_{0}]<\infty$,
+$\sqrt{n}\,(\mathbb{P}_{n}-\mathbb{P}_{0})\phi_{t}^{0}\xrightarrow{d}N\!\left(0,\sigma_{H}^{2}(t;{\bf X}_{0},{\bf Z}_{0})\right),\qquad\sigma_{H}^{2}(t;{\bf X}_{0},{\bf Z}_{0})={\mathbb{V}}_{0}\!\big(\phi_{t}^{0}({\bf W})\mid{\bf X}_{0},{\bf Z}_{0}\big).$
+Thus it remains to show
+$\sqrt{n}\Big\{A_{n}+\mathbb{P}_{n}\big(\widehat{\phi}_{t}-\phi_{t}^{0}\big)\Big\}=o_{p}(1).$
+Decomposing $\widehat{\phi}_{t}-\phi_{t}^{0}$ into the nuisance-Riesz part and the $\theta$}-part gives
+$\widehat{\phi}_{t}^{(-k)}-\phi_{t}^{0}=\big(\widehat{\phi}_{t,g}^{(-k)}-\phi_{t,g}^{0}\big)+\big(\widehat{\phi}_{t,\theta}^{(-k)}-\phi_{t,\theta}^{0}\big).$
+By the same first-order expansion used in the discrete setting for the parametric EIF component,
+$\sum_{k=1}^{K}\frac{|\mathcal{I}_{k}|}{n}\Big\{c_{0,t,n}({\bf X}_{0},{\bf Z}_{0})^{\top}\delta\mbox{${\theta}$}^{(-k)}-T_{n,1,\theta}^{(-k)}\Big\}+\mathbb{P}_{n}\!\big(\hat{\phi}_{t,\theta}-\phi_{t,\theta}^{0}\big)=o_{p}(n^{-1/2}),$
+by Lemmas 78 and assumption (i).
+Now we focus on the nuisance-Riesz part, where the new ingredient is the sieve estimation of $\psi_{t}^{0}$. Write
+$\widehat{\phi}_{t,g}^{(-k)}-\phi_{t,g}^{0}=\sum_{j=1}^{m_{n}}\big(\widehat{\psi}_{j}^{(-k)}-\psi_{j}^{0}\big)\,\Delta M_{0,j}\;+\;\sum_{j=1}^{m_{n}}\widehat{\psi}_{j}^{(-k)}\big(\Delta\widehat{M}_{j}^{(-k)}-\Delta M_{0,j}\big)\,.$ (C.5)
+We start with the second sum of the right-hand side of (C.5). Set
+$B_{n,1}^{(g)}=\frac{1}{n}\sum_{k=1}^{K}\sum_{i\in\mathcal{I}_{k}}\sum_{j=1}^{m_{n}}\widehat{\psi}_{t}^{(-k)}(t_{j},{\bf X}_{i};{\bf X}_{0},{\bf Z}_{0})\,\big(\Delta\widehat{M}_{ij}^{(-k)}-\Delta M_{0,ij}\big),$
+so that
+$\mathbb{P}_{n}\big(\widehat{\phi}_{t,g}-\phi_{t,g}^{0}\big)=B_{n,0}^{(g)}+B_{n,1}^{(g)},$
+where
+$B_{n,0}^{(g)}=\sum_{k=1}^{K}\frac{|\mathcal{I}_{k}|}{n}\left\{\frac{1}{|\mathcal{I}_{k}|}\sum_{i\in\mathcal{I}_{k}}\sum_{j=1}^{m_{n}}\big(\widehat{\psi}_{t}^{(-k)}(t_{j},{\bf X}_{i};{\bf X}_{0},{\bf Z}_{0})-\psi_{t}^{0}(t_{j},{\bf X}_{i};{\bf X}_{0},{\bf Z}_{0})\big)\,\Delta M_{0,ij}\right\}.$
+We first analyze $B_{n,1}^{(g)}$. Since $\Delta\widehat{M}_{ij}^{(-k)}-\Delta M_{0,ij}=-\big(\Delta\widehat{A}_{ij}^{(-k)}-\Delta A_{0,ij}\big)$, we have
+$B_{n,1}^{(g)}=-\frac{1}{n}\sum_{k=1}^{K}\sum_{i\in\mathcal{I}_{k}}\sum_{j=1}^{m_{n}}\widehat{\psi}_{t}^{(-k)}(t_{j},{\bf X}_{i};{\bf X}_{0},{\bf Z}_{0})\,Y_{i}(t_{j-1})\big\{\widehat{h}^{(-k)}(t_{j}\mid{\bf X}_{i},{\bf Z}_{i})-h_{o}(t_{j}\mid{\bf X}_{i},{\bf Z}_{i})\big\}\,\Delta t_{j},$
+and $B_{n,1}^{(g)}=\sum_{k=1}^{K}\frac{|\mathcal{I}_{k}|}{n}\ B_{n,1}^{(g,k)}$ where
+$B_{n,1}^{(g,k)}=\frac{1}{|\mathcal{I}_{k}|}\sum_{i\in\mathcal{I}_{k}}\sum_{j=1}^{m_{n}}\widehat{\psi}_{t}^{(-k)}(t_{j},{\bf X}_{i};{\bf X}_{0},{\bf Z}_{0})\,\big(\Delta\widehat{M}_{ij}^{(-k)}-\Delta M_{0,ij}\big)\,.$
+Conditional on $\mathcal{F}_{-k}$, the observations $\{{\bf W}_{i}:i\in\mathcal{I}_{k}\}$ are i.i.d., hence
+$\mathbb{E}_{0}\!\big(B_{n,1}^{(g,k)}\mid\mathcal{F}_{-k}\big)=-\sum_{j=1}^{m_{n}}\mathbb{E}_{0}\!\left[\widehat{\psi}_{t}^{(-k)}(t_{j},{\bf X};{\bf X}_{0},{\bf Z}_{0})\,Y(t_{j-1})\big\{\widehat{\lambda}^{(-k)}(t_{j}\mid{\bf X},{\bf Z})-h_{o}(t_{j}\mid{\bf X},{\bf Z})\big\}\Bigm|\mathcal{F}_{-k}\right]\Delta t_{j}.$
+Now write
+$\widehat{h}^{(-k)}(t_{j}\mid{\bf X},{\bf Z})-h_{o}(t_{j}\mid{\bf X},{\bf Z})=h_{o}(t_{j}\mid{\bf X},{\bf Z})\,r_{j}^{(-k)}({\bf X},{\bf Z})+R_{\lambda,j}^{(-k)}({\bf X},{\bf Z}),$
+where
+$r_{j}^{(-k)}({\bf X},{\bf Z})=\delta g^{(-k)}(t_{j},{\bf X})+(\delta\mbox{${\theta}$}^{(-k)})^{\top}{\bf Z},$
+and the exponential-link remainder satisfies
+$|R_{\lambda,j}^{(-k)}({\bf X},{\bf Z})|\lesssim h_{o}(t_{j}\mid{\bf X},{\bf Z})\,\big(r_{j}^{(-k)}({\bf X},{\bf Z})\big)^{2}$
+with probability tending to one, uniformly in $j$ and $k$. Let ${\bf g}^{*}(u,{\bf X})$ denote the $L_{0}^{2}$-projection of $\bf Z$} onto $\mathcal{H}$, as in Lemma 5, and define
+$b_{j}^{(-k)}({\bf X})=\delta g^{(-k)}(t_{j},{\bf X})+(\delta\mbox{${\theta}$}^{(-k)})^{\top}{\bf g}^{*}(t_{j},{\bf X}).$
+Then
+$r_{j}^{(-k)}({\bf X},{\bf Z})=b_{j}^{(-k)}({\bf X})+(\delta\mbox{${\theta}$}^{(-k)})^{\top}\!\big\{{\bf Z}-{\bf g}^{*}(t_{j},{\bf X})\big\}.$
+Accordingly,
+$\mathbb{E}_{0}\!\big(B_{n,1}^{(g,k)}\mid\mathcal{F}_{-k}\big)=-\Big\{T_{n,1,g}^{(-k)}+T_{n,1,\theta}^{(-k)}+R_{n,2}^{(-k)}\Big\},$
+where
+$T_{n,1,g}^{(-k)}=\sum_{j=1}^{m_{n}}\mathbb{E}_{0}\!\left[\widehat{\psi}_{t}^{(-k)}(t_{j},{\bf X};{\bf X}_{0},{\bf Z}_{0})\,Y(t_{j-1})h_{o}(t_{j}\mid{\bf X},{\bf Z})\,b_{j}^{(-k)}({\bf X})\Bigm|\mathcal{F}_{-k}\right]\Delta t_{j},$
+and
+$T_{n,1,\theta}^{(-k)}=\sum_{j=1}^{m_{n}}\mathbb{E}_{0}\!\left[\widehat{\psi}_{t}^{(-k)}(t_{j},{\bf X};{\bf X}_{0},{\bf Z}_{0})\,Y(t_{j-1})h_{o}(t_{j}\mid{\bf X},{\bf Z})\,(\delta\mbox{${\theta}$}^{(-k)})^{\top}\!\big\{{\bf Z}-{\bf g}^{*}(t_{j},{\bf X})\big\}\Bigm|\mathcal{F}_{-k}\right]\Delta t_{j},$
+and
+$|R_{n,2}^{(-k)}|\lesssim\sum_{j=1}^{m_{n}}\mathbb{E}_{0}\!\left[\big|\widehat{\psi}_{t}^{(-k)}(t_{j},{\bf X};{\bf X}_{0},{\bf Z}_{0})\big|\,Y(t_{j-1})h_{o}(t_{j}\mid{\bf X},{\bf Z})\,\big(r_{j}^{(-k)}({\bf X},{\bf Z})\big)^{2}\Bigm|\mathcal{F}_{-k}\right]\Delta t_{j}.$
+The term $T_{n,1,g}^{(-k)}$ is the nuisance-space part, since $b^{(-k)}$ is a function of $(u,{\bf X})$ only.
+The term $T_{n,1,\theta}^{(-k)}$ corresponds to the residual parametric direction generated by ${\bf Z}-{\bf g}^{*}(u,{\bf X})$; this will match the first-order parametric contribution in $\mathbb{P}_{n}(\widehat{\phi}_{t,\theta}-\phi_{t,\theta}^{0})$.
+Write $\Pi_{d_{n}}b^{(-k)}$ for its sieve projection onto $\mathcal{H}_{d_{n}}$ on the grid.
+We next identify the leading nuisance term $T_{n,1,g}^{(-k)}$ with the evaluation functional applied to $b^{(-k)}$.
+Write
+$T_{n,1,g}^{(-k)}=\Big\langle\hat{\psi}_{t}^{(-k)},\,\Pi_{d_{n}}b^{(-k)}\Big\rangle_{0,n}+R_{n,3}^{(-k)},$
+where $R_{n,3}^{(-k)}=o_{p}(n^{-1/2})$ uniformly in $k$ by the projection approximation and the same Riemann discretization error bound used throughout. Next,
+$\Big\langle\hat{\psi}_{t}^{(-k)},\,\Pi_{d_{n}}b^{(-k)}\Big\rangle_{0,n}=\Big\langle\psi_{t}^{0,(d_{n})},\,\Pi_{d_{n}}b^{(-k)}\Big\rangle_{0}+R_{n,4}^{(-k)},$
+where $R_{n,4}^{(-k)}=o_{p}(n^{-1/2})$ uniformly in $k$ by assumption (iv), which absorbs the additional empirical-process terms arising from estimating $\hat{\psi}_{t}^{(-k)}$ and replacing the empirical inner product by its population version.
+Finally, by the defining property of the population sieve-Riesz representer,
+$\Big\langle\psi_{t}^{0,(d_{n})},\,\Pi_{d_{n}}b^{(-k)}\Big\rangle_{0}=L_{t,n}\!\big(\Pi_{d_{n}}b^{(-k)}\big).$
+Hence
+$T_{n,1,g}^{(-k)}=L_{t,n}\!\big(\Pi_{d_{n}}b^{(-k)}\big)+R_{n,3}^{(-k)}+R_{n,4}^{(-k)}.$
+It remains to remove the sieve projection. Assumption (iv) also absorbs the projection-removal error inside the evaluation functional, hence
+$L_{t,n}\!\big(\Pi_{d_{n}}b^{(-k)}\big)-L_{t,n}\!\big(b^{(-k)}\big)=o_{p}(n^{-1/2}),$
+uniformly in $k$. Therefore
+$T_{n,1,g}^{(-k)}=L_{t,n}\!\big(b^{(-k)}\big)+o_{p}(n^{-1/2})\,$,
+uniformly in $k$.
+On the other hand, by the definition of
+$c_{0,t}({\bf X}_{0},{\bf Z}_{0})={\bf Z}_{0}H_{0}(t\mid{\bf X}_{0},{\bf Z}_{0})-L_{t}({\bf g}^{*}),$
+and its Riemann approximation $c_{0,t,n}({\bf X}_{0},{\bf Z}_{0})$, we have
+$\sum_{j:t_{j}\leq t}h_{o}(t_{j}\mid{\bf X}_{0},{\bf Z}_{0})\,(\delta\mbox{${\theta}$}^{(-k)})^{\top}\!\big\{{\bf Z}_{0}-{\bf g}^{*}(t_{j},{\bf X}_{0})\big\}\,\Delta t_{j}=\big(c_{0,t,n}({\bf X}_{0},{\bf Z}_{0})\big)^{\top}\delta\mbox{${\theta}$}^{(-k)},$
+with $c_{0,t,n}({\bf X}_{0},{\bf Z}_{0})-c_{0,t}({\bf X}_{0},{\bf Z}_{0})=O(\delta_{n})=o(n^{-1/2})$.
+Thus the first-order expansion of $A_{n}$ can be written as a nuisance component $L_{t,n}(b^{(-k)})$ plus the parametric component $c_{0,t,n}({\bf X}_{0},{\bf Z}_{0})^{\top}\delta\mbox{${\theta}$}^{(-k)}$.
+The latter is canceled by the corresponding first-order contribution from $\mathbb{P}_{n}(\widehat{\phi}_{t,\theta}-\phi_{t,\theta}^{0})$, exactly as in the discrete setting.
+Now rewrite the first-order expansion of $A_{n}$ as
+$A_{n}=\sum_{k=1}^{K}\frac{|\mathcal{I}_{k}|}{n}\left[L_{t,n}\!\big(b^{(-k)}\big)+c_{0,t,n}({\bf X}_{0},{\bf Z}_{0})^{\top}\delta\mbox{${\theta}$}^{(-k)}\right]+R_{n,1},$
+where
+$b_{j}^{(-k)}({\bf X})=\delta g^{(-k)}(t_{j},{\bf X})+(\delta\mbox{${\theta}$}^{(-k)})^{\top}{\bf g}^{*}(t_{j},{\bf X}),$
+and
+$c_{0,t,n}({\bf X}_{0},{\bf Z}_{0})=\sum_{j:t_{j}\leq t}h_{o}(t_{j}\mid{\bf X}_{0},{\bf Z}_{0})\,\big\{{\bf Z}_{0}-{\bf g}^{*}(t_{j},{\bf X}_{0})\big\}\,\Delta t_{j},\qquad c_{0,t,n}({\bf X}_{0},{\bf Z}_{0})=c_{0,t}({\bf X}_{0},{\bf Z}_{0})+o(1).$
+Combining the previous display with
+$T_{n,1,g}^{(-k)}=L_{t,n}\!\big(b^{(-k)}\big)+o_{p}(n^{-1/2}),$
+we see that the nuisance-space contribution in
+$A_{n}=\sum_{k=1}^{K}\frac{|\mathcal{I}_{k}|}{n}\Big\{L_{t,n}\!\big(b^{(-k)}\big)+c_{0,t,n}({\bf X}_{0},{\bf Z}_{0})^{\top}\delta\mbox{${\theta}$}^{(-k)}\Big\}+R_{n,1}$
+cancels fold-by-fold with the term
+$\sum_{k=1}^{K}\frac{|\mathcal{I}_{k}|}{n}T_{n,1,g}^{(-k)}$
+appearing in $\sum_{k}\frac{|\mathcal{I}_{k}|}{n}\mathbb{E}_{0}(B_{n,1}^{(g,k)}\mid\mathcal{F}_{-k})$. Therefore,
+$A_{n}+\sum_{k=1}^{K}\frac{|\mathcal{I}_{k}|}{n}\mathbb{E}_{0}\!\big(B_{n,1}^{(g,k)}\mid\mathcal{F}_{-k}\big)=\sum_{k=1}^{K}\frac{|\mathcal{I}_{k}|}{n}\Big\{c_{0,t,n}({\bf X}_{0},{\bf Z}_{0})^{\top}\delta\mbox{${\theta}$}^{(-k)}-T_{n,1,\theta}^{(-k)}\Big\}+R_{n,1}+\sum_{k=1}^{K}\frac{|\mathcal{I}_{k}|}{n}\widetilde{R}_{n,2}^{(-k)},$
+where $\widetilde{R}_{n,2}^{(-k)}=o_{p}(n^{-1/2})$ uniformly in $k$.
+By the same first-order expansion used earlier for the parametric EIF component, the weighted sum
+$\sum_{k=1}^{K}\frac{|\mathcal{I}_{k}|}{n}\Big\{c_{0,t,n}({\bf X}_{0},{\bf Z}_{0})^{\top}\delta\mbox{${\theta}$}^{(-k)}-T_{n,1,\theta}^{(-k)}\Big\}+\mathbb{P}_{n}\big(\widehat{\phi}_{t,\theta}-\phi_{t,\theta}^{0}\big)=o_{p}(n^{-1/2})\,.$
+Consequently,
+$A_{n}+\sum_{k=1}^{K}\frac{|\mathcal{I}_{k}|}{n}\mathbb{E}_{0}\!\big(B_{n,1}^{(g,k)}\mid\mathcal{F}_{-k}\big)+\mathbb{P}_{n}\big(\widehat{\phi}_{t,\theta}-\phi_{t,\theta}^{0}\big)=o_{p}(n^{-1/2}),$
+because assumption (i) also implies
+$\sqrt{n}\,\max_{k}\|\delta g^{(-k)}\|_{0}^{2}=o_{p}(1),\qquad\sqrt{n}\,\max_{k}\|\delta\mbox{${\theta}$}^{(-k)}\|^{2}=o_{p}(1).$
+It remains to control the centered empirical fluctuation
+$\sum_{k=1}^{K}\frac{|\mathcal{I}_{k}|}{n}\left\{B_{n,1}^{(g,k)}-\mathbb{E}_{0}\!\big(B_{n,1}^{(g,k)}\mid\mathcal{F}_{-k}\big)\right\},$
+which is $o_{p}(n^{-1/2})$ by conditional independence across $i\in\mathcal{I}_{k}$, bounded conditional second moments, and assumption (iv).
+Here the dependence on $\widehat{\psi}_{t}^{(-k)}$ is handled entirely through assumption (iv), conditional on $\mathcal{F}_{-k}$.
+Hence
+$A_{n}+B_{n,1}^{(g)}=-\mathbb{P}_{n}\big(\widehat{\phi}_{t,\theta}-\phi_{t,\theta}^{0}\big)+o_{p}(n^{-1/2}).$
+It remains to control $B_{n,0}^{(g)}$, which reflects sieve-Riesz approximation and estimation error.
+Let $\psi_{t}^{0,(d_{n})}\in\mathcal{H}_{d_{n}}$ be the population sieve-Riesz representer and write
+$\widehat{\psi}_{t}^{(-k)}-\psi_{t}^{0}=\underbrace{\big(\psi_{t}^{0,(d_{n})}-\psi_{t}^{0}\big)}_{\text{sieve bias}}+\underbrace{\big(\widehat{\psi}_{t}^{(-k)}-\psi_{t}^{0,(d_{n})}\big)}_{\text{estimation error}}.$
+Using the triangle inequality,
+$\|\widehat{\psi}_{t}^{(-k)}-\psi_{t}^{0}\|_{0}\leq\|\psi_{t}^{0,(d_{n})}-\psi_{t}^{0}\|_{0}+\|\widehat{\psi}_{t}^{(-k)}-\psi_{t}^{0,(d_{n})}\|_{0}.$
+By assumptions (ii) and (iii),
+$\max_{1\leq k\leq K}\|\widehat{\psi}_{t}^{(-k)}-\psi_{t}^{0}\|_{0}=o_{p}(1).$
+Now, conditional on $\mathcal{F}_{-k}$ we can apply the discrete-time martingale isometry to obtain
+$\mathbb{E}_{0}\!\left[\left(\sum_{j=1}^{m_{n}}\big(\widehat{\psi}_{t}^{(-k)}(t_{j},{\bf X})-\psi_{t}^{0}(t_{j},{\bf X})\big)\,\Delta M_{0,j}\right)^{2}\Bigm|\mathcal{F}_{-k}\right]$
+$\qquad=\mathbb{E}_{0}\!\left[\sum_{j=1}^{m_{n}}\big(\widehat{\psi}_{t}^{(-k)}(t_{j},{\bf X})-\psi_{t}^{0}(t_{j},{\bf X})\big)^{2}\,\Delta A_{0,j}\Bigm|\mathcal{F}_{-k}\right].$
+Conditional on $\mathcal{F}_{-k}$, the summands
+$\sum_{j=1}^{m_{n}}(\widehat{\psi}_{t}^{(-k)}(t_{j},{\bf X}_{i})-\psi_{t}^{0}(t_{j},{\bf X}_{i}))\,\Delta M_{0,ij},\qquad i\in\mathcal{I}_{k},$
+are independent and centered. Hence, by the discrete-time martingale isometry,
+${\mathbb{V}}_{0}\!\left(\frac{1}{|\mathcal{I}_{k}|}\sum_{i\in\mathcal{I}_{k}}\sum_{j=1}^{m_{n}}(\widehat{\psi}_{t}^{(-k)}(t_{j},{\bf X}_{i})-\psi_{t}^{0}(t_{j},{\bf X}_{i}))\,\Delta M_{0,ij}\Bigm|\mathcal{F}_{-k}\right)\lesssim\frac{1}{|\mathcal{I}_{k}|}\|\widehat{\psi}_{t}^{(-k)}-\psi_{t}^{0}\|_{0}^{2}.$
+Therefore,
+$\frac{1}{|\mathcal{I}_{k}|}\sum_{i\in\mathcal{I}_{k}}\sum_{j=1}^{m_{n}}(\widehat{\psi}_{t}^{(-k)}(t_{j},{\bf X}_{i})-\psi_{t}^{0}(t_{j},{\bf X}_{i}))\,\Delta M_{0,ij}=O_{p}\!\left(|\mathcal{I}_{k}|^{-1/2}\|\widehat{\psi}_{t}^{(-k)}-\psi_{t}^{0}\|_{0}\right).$
+Since $K$ is fixed and $|\mathcal{I}_{k}|\asymp n$,
+$B_{n,0}^{(g)}=O_{p}\!\left(n^{-1/2}\max_{1\leq k\leq K}\|\widehat{\psi}_{t}^{(-k)}-\psi_{t}^{0}\|_{0}\right)$
+and hence
+$B_{n,0}^{(g)}=o_{p}(n^{-1/2}).$
+Combining this with the previously established bound
+$A_{n}+B_{n,1}^{(g)}+\mathbb{P}_{n}\big(\widehat{\phi}_{t,\theta}-\phi_{t,\theta}^{0}\big)=o_{p}(n^{-1/2}),$
+and with
+$B_{n,0}^{(g)}=o_{p}(n^{-1/2}),$
+we obtain
+$A_{n}+\mathbb{P}_{n}\big(\widehat{\phi}_{t}-\phi_{t}^{0}\big)=o_{p}(n^{-1/2}).$
+and therefore
+$\sqrt{n}\big(\widehat{H}^{1,{\sf cf}}(t\mid{\bf X}_{0},{\bf Z}_{0})-H_{0}(t\mid{\bf X}_{0},{\bf Z}_{0})\big)\xrightarrow{d}N\!\left(0,\sigma_{H}^{2}(t;{\bf X}_{0},{\bf Z}_{0})\right).$
+The result for $\widehat{S}^{1,{\sf cf}}=\exp(-\widehat{H}^{1,{\sf cf}})$ follows by the delta method with derivative $S_{0}$. ∎